Texinfo

Texinfo

The GNU Documentation Format for Texinfo version 4.2, 28 March 2002

Robert J. Chassell Richard M. Stallman

This manual is for GNU Texinfo (version 4.2, 28 March 2002), a documentation system that can produce both online information and a printed manual from a single source. Copyright (C) 1988, 90, 91, 92, 93, 95, 96, 97, 98, 99, 2000, 01, 02 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.1 or any later version published by the Free Software Foundation; with no Invariant Sections, with the Front-Cover texts being “A GNU Manual,” and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled “GNU Free Documentation License.”

(a) The FSF’s Back-Cover Text is: “You have freedom to copy and modify this GNU Manual, like GNU software. Copies published by the Free Software Foundation raise funds for GNU development.”

Published by the Free Software Foundation 59 Temple Place Suite 330 Boston, MA 02111-1307 USA ISBN 1-882114-67-1 Cover art by Etienne Suvasa.

Short Contents

Texinfo Copying Conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1 Overview of Texinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2 Using Texinfo Mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3 Beginning a Texinfo File. . . . . . . . . . . . . . . . . . . . . . . . . . 26 4 Ending a Texinfo File . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5 Chapter Structuring . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 6 Nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 7 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 8 Cross References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 9 Marking Words and Phrases . . . . . . . . . . . . . . . . . . . . . . . 71 10 Quotations and Examples. . . . . . . . . . . . . . . . . . . . . . . . . 81 11 Lists and Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 12 Indices. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 13 Special Insertions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 14 Making and Preventing Breaks . . . . . . . . . . . . . . . . . . . . 113 15 Definition Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . 116 16 Conditionally Visible Text . . . . . . . . . . . . . . . . . . . . . . . 129 17 Internationalization . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 18 Defining New Texinfo Commands. . . . . . . . . . . . . . . . . . . 137 19 Formatting and Printing Hardcopy. . . . . . . . . . . . . . . . . . 142 20 Creating and Installing Info Files . . . . . . . . . . . . . . . . . . . 153 A @-Command List. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 B Tips and Hints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 C Sample Texinfo Files . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 D Include Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 E Page Headings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 F Formatting Mistakes . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 G Refilling Paragraphs. . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 H @-Command Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 I How to Obtain TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 J Copying This Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 Command and Variable Index . . . . . . . . . . . . . . . . . . . . . . . . 222 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226

Table of Contents

Texinfo Copying Conditions . . . . . . . . . . . . . . . . . . . 2

1 Overview of Texinfo. . . . . . . . . . . . . . . . . . . . . . . . 3

1.1 Reporting Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Using Texinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Info files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.4 Printed Books . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.5 @-commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.6 General Syntactic Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.7 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.8 What a Texinfo File Must Have . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.9 Six Parts of a Texinfo File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.10 A Short Sample Texinfo File . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 1.11 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

2 Using Texinfo Mode . . . . . . . . . . . . . . . . . . . . . . 14

2.1 The Usual GNU Emacs Editing Commands . . . . . . . . . . . . . . 14 2.2 Inserting Frequently Used Commands. . . . . . . . . . . . . . . . . . . . 15 2.3 Showing the Section Structure of a File . . . . . . . . . . . . . . . . . . 16 2.4 Updating Nodes and Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

2.4.1 Updating Requirements . . . . . . . . . . . . . . . . . . . . . . . . 20 2.4.2 Other Updating Commands . . . . . . . . . . . . . . . . . . . . 20

2.5 Formatting for Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.6 Formatting and Printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.7 Texinfo Mode Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

3 Beginning a Texinfo File . . . . . . . . . . . . . . . . . . 26

3.1 Sample Texinfo File Beginning . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.2 Texinfo File Header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

3.2.1 The First Line of a Texinfo File. . . . . . . . . . . . . . . . . 27 3.2.2 Start of Header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.2.3 @setfilename: Set the output file name. . . . . . . . . 28 3.2.4 @settitle: Set the document title . . . . . . . . . . . . . . 29 3.2.5 End of Header. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

3.3 Document Permissions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.3.1 @copying: Declare copying permissions. . . . . . . . . . 30 3.3.2 @insertcopying: Include permissions text . . . . . . 31

3.4 Title and Copyright Pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.4.1 @titlepage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.4.2 @titlefont, @center, and @sp . . . . . . . . . . . . . . . . . 32 3.4.3 @title, @subtitle, and @author . . . . . . . . . . . . . . . 33 3.4.4 Copyright Page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 3.4.5 Heading Generation. . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 3.4.6 The @headings Command. . . . . . . . . . . . . . . . . . . . . . 35

3.5 The ‘Top’ Node and Master Menu . . . . . . . . . . . . . . . . . . . . . . . 36 3.5.1 Top Node Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 3.5.2 Parts of a Master Menu . . . . . . . . . . . . . . . . . . . . . . . . 37

3.6 Global Document Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.6.1 @documentdescription: Summary text . . . . . . . . . 38 3.6.2 @setchapternewpage:. . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.6.3 Paragraph Indenting . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 3.6.4 @exampleindent: Environment Indenting. . . . . . . . 40

3.7 Software Copying Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

4 Ending a Texinfo File . . . . . . . . . . . . . . . . . . . . . 41

4.1 Printing Indices and Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.2 Generating a Table of Contents. . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.3 @bye File Ending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

5 Chapter Structuring . . . . . . . . . . . . . . . . . . . . . . 44

5.1 Tree Structure of Sections. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.2 Structuring Command Types. . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.3 @top . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.4 @chapter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.5 @unnumbered and @appendix. . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.6 @majorheading, @chapheading. . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.7 @section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

5.8 @unnumberedsec, @appendixsec, @heading . . . . . . . . . . . . . . 47

5.9 The @subsection Command . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.10 The @subsection-like Commands . . . . . . . . . . . . . . . . . . . . . . 48 5.11 The ‘subsub’ Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

5.12 @raisesections and @lowersections . . . . . . . . . . . . . . . . . . 49

6 Nodes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

6.1 Two Paths. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 6.2 Node and Menu Illustration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 6.3 The @node Command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

6.3.1 Choosing Node and Pointer Names. . . . . . . . . . . . . . 52 6.3.2 How to Write an @node Line. . . . . . . . . . . . . . . . . . . . 53 6.3.3 @node Line Tips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 6.3.4 @node Line Requirements . . . . . . . . . . . . . . . . . . . . . . 54 6.3.5 The First Node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 6.3.6 The @top Sectioning Command . . . . . . . . . . . . . . . . . 55

6.4 Creating Pointers with makeinfo . . . . . . . . . . . . . . . . . . . . . . . . 55 6.5 @anchor: Defining Arbitrary Cross-reference Targets . . . . . . 56

7 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

7.1 Writing a Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 7.2 The Parts of a Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 7.3 Less Cluttered Menu Entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 7.4 A Menu Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 7.5 Referring to Other Info Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

8 Cross References . . . . . . . . . . . . . . . . . . . . . . . . . 61

8.1 Different Cross Reference Commands . . . . . . . . . . . . . . . . . . . . 61 8.2 Parts of a Cross Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 8.3 @xref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

8.3.1 @xref with One Argument . . . . . . . . . . . . . . . . . . . . . 63 8.3.2 @xref with Two Arguments . . . . . . . . . . . . . . . . . . . . 64 8.3.3 @xref with Three Arguments . . . . . . . . . . . . . . . . . . . 65 8.3.4 @xref with Four and Five Arguments . . . . . . . . . . . 66

8.4 Naming a ‘Top’ Node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 8.5 @ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 8.6 @pxref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 8.7 @inforef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 8.8 @uref{url[, text][, replacement]} . . . . . . . . . . . . . . . . . . . . . 69

9 Marking Words and Phrases. . . . . . . . . . . . . . . 71

9.1 Indicating Definitions, Commands, etc. . . . . . . . . . . . . . . . . . . 71 9.1.1 @code{sample-code} . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 9.1.2 @kbd{keyboard-characters} . . . . . . . . . . . . . . . . . . . . . 73 9.1.3 @key{key-name}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 9.1.4 @samp{text}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 9.1.5 @verb{<char>text<char>} . . . . . . . . . . . . . . . . . . . . . . 75 9.1.6 @var{metasyntactic-variable}. . . . . . . . . . . . . . . . . . . 75 9.1.7 @env{environment-variable} . . . . . . . . . . . . . . . . . . . . 76 9.1.8 @file{file-name} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 9.1.9 @command{command-name} . . . . . . . . . . . . . . . . . . . . 77 9.1.10 @option{option-name} . . . . . . . . . . . . . . . . . . . . . . . 77 9.1.11 @dfn{term} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 9.1.12 @cite{reference} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 9.1.13 @acronym{acronym} . . . . . . . . . . . . . . . . . . . . . . . . . . 78 9.1.14 @url{uniform-resource-locator}. . . . . . . . . . . . . . . . 78 9.1.15 @email{email-address[, displayed-text]} . . . . . . . . 78

9.2 Emphasizing Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 9.2.1 @emph{text} and @strong{text} . . . . . . . . . . . . . . . . 78 9.2.2 @sc{text}: The Small Caps Font. . . . . . . . . . . . . . . . 79 9.2.3 Fonts for Printing, Not Info . . . . . . . . . . . . . . . . . . . . 80

10 Quotations and Examples . . . . . . . . . . . . . . . . 81

10.1 Block Enclosing Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 10.2 @quotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 10.3 @example: Example Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 10.4 @verbatim: Literal Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 10.5 @verbatiminclude file: Include a File Verbatim . . . . . . . . . 84 10.6 @lisp: Marking a Lisp Example. . . . . . . . . . . . . . . . . . . . . . . . 84 10.7 @small... Block Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . 84 10.8 @display and @smalldisplay . . . . . . . . . . . . . . . . . . . . . . . . . 85 10.9 @format and @smallformat . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 10.10 @exdent: Undoing a Line’s Indentation . . . . . . . . . . . . . . . . 86 10.11 @flushleft and @flushright . . . . . . . . . . . . . . . . . . . . . . . . 86 10.12 @noindent: Omitting Indentation . . . . . . . . . . . . . . . . . . . . . 87 10.13 @cartouche: Rounded Rectangles Around Examples …. 87

11 Lists and Tables . . . . . . . . . . . . . . . . . . . . . . . . . 89

11.1 @itemize: Making an Itemized List . . . . . . . . . . . . . . . . . . . . 89 11.2 @enumerate: Making a Numbered or Lettered List. . . . . . . 91 11.3 Making a Two-column Table . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

11.3.1 @ftable and @vtable. . . . . . . . . . . . . . . . . . . . . . . . . 93 11.3.2 @itemx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

11.4 Multi-column Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 11.4.1 Multitable Column Widths . . . . . . . . . . . . . . . . . . . . 94 11.4.2 Multitable Rows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

12 Indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

12.1 Making Index Entries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 12.2 Predefined Indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 12.3 Defining the Entries of an Index . . . . . . . . . . . . . . . . . . . . . . . . 97 12.4 Combining Indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

12.4.1 @syncodeindex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 12.4.2 @synindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

12.5 Defining New Indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

13 Special Insertions. . . . . . . . . . . . . . . . . . . . . . . 101

13.1 Inserting @ and Braces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 13.1.1 Inserting ‘@’ with @@ . . . . . . . . . . . . . . . . . . . . . . . . 101 13.1.2 Inserting ‘{’ and ‘}’with @{ and @}. . . . . . . . . . . 101

13.2 Inserting Space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 13.2.1 Not Ending a Sentence. . . . . . . . . . . . . . . . . . . . . . . 102 13.2.2 Ending a Sentence . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 13.2.3 Multiple Spaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 13.2.4 @dmn{dimension}: Format a Dimension. . . . . . . . 103

13.3 Inserting Accents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 13.4 Inserting Ellipsis and Bullets . . . . . . . . . . . . . . . . . . . . . . . . . . 104

13.4.1 @dots{} (…) and @enddots{} (….) . . . . . . . . 104 13.4.2 @bullet{} (•). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 13.5 Inserting TEX and the Copyright Symbol . . . . . . . . . . . . . . 105

13.5.1 @TeX{} (TEX) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 13.5.2 @copyright\{\} (\begin{matrix}c \\ ◯ \end{matrix}) . . . . . . . . . . . . . . . . . . . . . . . . . . 105

13.6 @pounds{} (£): Pounds Sterling . . . . . . . . . . . . . . . . . . . . . . 105 13.7 @minus{} (−): Inserting a Minus Sign . . . . . . . . . . . . . . . . . 105 13.8 @math: Inserting Mathematical Expressions . . . . . . . . . . . . 106 13.9 Glyphs for Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

13.9.1 Glyphs Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 13.9.2 @result{} (⇒): Indicating Evaluation . . . . . . . . 107 13.9.3 @expansion{} (↦): Indicating an Expansion . . 107 13.9.4 @print{} ( ⊣): Indicating Printed Output. . . . . 107 13.9.5 @error{} ( error ): Indicating an Error Message

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 13.9.6 @equiv{} (≡): Indicating Equivalence . . . . . . . . 108 13.9.7 @point{} (⋆): Indicating Point in a Buffer. . . . . 108

13.10 Footnotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 13.10.1 Footnote Commands. . . . . . . . . . . . . . . . . . . . . . . . 109 13.10.2 Footnote Styles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

13.11 Inserting Images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

14 Making and Preventing Breaks . . . . . . . . . . 113

14.1 @*: Generate Line Breaks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 14.2 @- and @hyphenation: Helping TEX hyphenate . . . . . . . . . 114 14.3 @w{text}: Prevent Line Breaks . . . . . . . . . . . . . . . . . . . . . . . . 114 14.4 @sp n: Insert Blank Lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 14.5 @page: Start a New Page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 14.6 @group: Prevent Page Breaks . . . . . . . . . . . . . . . . . . . . . . . . . 115 14.7 @need mils: Prevent Page Breaks. . . . . . . . . . . . . . . . . . . . . . 115

15 Definition Commands . . . . . . . . . . . . . . . . . . . 116

15.1 The Template for a Definition . . . . . . . . . . . . . . . . . . . . . . . . . 116 15.2 Optional and Repeated Arguments . . . . . . . . . . . . . . . . . . . . 117 15.3 Two or More ‘First’ Lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 15.4 The Definition Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

15.4.1 Functions and Similar Entities . . . . . . . . . . . . . . . . 119 15.4.2 Variables and Similar Entities . . . . . . . . . . . . . . . . 120 15.4.3 Functions in Typed Languages. . . . . . . . . . . . . . . . 121 15.4.4 Variables in Typed Languages . . . . . . . . . . . . . . . . 122 15.4.5 Object-Oriented Programming. . . . . . . . . . . . . . . . 123 15.4.6 Data Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

15.5 Conventions for Writing Definitions. . . . . . . . . . . . . . . . . . . . 126 15.6 A Sample Function Definition . . . . . . . . . . . . . . . . . . . . . . . . . 126

16 Conditionally Visible Text . . . . . . . . . . . . . . 129

16.1 Conditional Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 16.2 Conditional Not Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . 130 16.3 Raw Formatter Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 16.4 @set, @clear, and @value . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

16.4.1 @set and @value . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 16.4.2 @ifset and @ifclear. . . . . . . . . . . . . . . . . . . . . . . . 132 16.4.3 @value Example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133

17 Internationalization . . . . . . . . . . . . . . . . . . . . . 135

17.1 @documentlanguage cc: Set the Document Language … 135 17.2 @documentencoding enc: Set Input Encoding . . . . . . . . . . 136

18 Defining New Texinfo Commands. . . . . . . . 137

18.1 Defining Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 18.2 Invoking Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 18.3 Macro Details. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 18.4 ‘@alias new=existing’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 18.5 ‘definfoenclose’: Customized Highlighting . . . . . . . . . . . 140

19 Formatting and Printing Hardcopy . . . . . . 142

19.1 Use TEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 19.2 Format with tex and texindex . . . . . . . . . . . . . . . . . . . . . . . 142 19.3 Format with texi2dvi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 19.4 Shell Print Using lpr -d. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 19.5 From an Emacs Shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 19.6 Formatting and Printing in Texinfo Mode . . . . . . . . . . . . . . 145 19.7 Using the Local Variables List . . . . . . . . . . . . . . . . . . . . . . . . 147 19.8 TEX Formatting Requirements Summary. . . . . . . . . . . . . . . 147 19.9 Preparing for TEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 19.10 Overfull “hboxes” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 19.11 Printing “Small” Books. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 19.12 Printing on A4 Paper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 19.13 @pagesizes [width][, height]: Custom page sizes. . . . . . . 150 19.14 Cropmarks and Magnification . . . . . . . . . . . . . . . . . . . . . . . . 151 19.15 PDF Output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152

20 Creating and Installing Info Files . . . . . . . . 153

20.1 Creating an Info File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 20.1.1 makeinfo Preferred . . . . . . . . . . . . . . . . . . . . . . . . . . 153 20.1.2 Running makeinfo from a Shell. . . . . . . . . . . . . . . 153 20.1.3 Options for makeinfo . . . . . . . . . . . . . . . . . . . . . . . . 153 20.1.4 Pointer Validation . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 20.1.5 Running makeinfo inside Emacs . . . . . . . . . . . . . . 157 20.1.6 The texinfo-format... Commands . . . . . . . . . . 158 20.1.7 Batch Formatting . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 20.1.8 Tag Files and Split Files . . . . . . . . . . . . . . . . . . . . . 159 20.1.9 Generating HTML. . . . . . . . . . . . . . . . . . . . . . . . . . . 160

20.2 Installing an Info File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 20.2.1 The Directory File ‘dir’ . . . . . . . . . . . . . . . . . . . . . 161 20.2.2 Listing a New Info File. . . . . . . . . . . . . . . . . . . . . . . 162 20.2.3 Info Files in Other Directories . . . . . . . . . . . . . . . . 162 20.2.4 Installing Info Directory Files. . . . . . . . . . . . . . . . . 164 20.2.5 Invoking install-info. . . . . . . . . . . . . . . . . . . . . . . . . . 164

Appendix A @-Command List. . . . . . . . . . . . . . 167

Appendix B Tips and Hints . . . . . . . . . . . . . . . . 186

Appendix C Sample Texinfo Files . . . . . . . . . . 191

C.1 Short Sample. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 C.2 GNU Sample Texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192

Appendix D Include Files . . . . . . . . . . . . . . . . . . 196

D.1 How to Use Include Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196

D.2 texinfo-multiple-files-update . . . . . . . . . . . . . . . . . . . . . 196

D.3 Include File Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 D.4 Sample File with @include . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 D.5 Evolution of Include Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

Appendix E Page Headings . . . . . . . . . . . . . . . . 200

E.1 Standard Heading Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 E.2 Specifying the Type of Heading . . . . . . . . . . . . . . . . . . . . . . . . 201 E.3 How to Make Your Own Headings . . . . . . . . . . . . . . . . . . . . . 202

Appendix F Formatting Mistakes. . . . . . . . . . . 204

F.1 Catching Errors with Info Formatting . . . . . . . . . . . . . . . . . . 204 F.2 Catching Errors with TEX Formatting . . . . . . . . . . . . . . . . . . 205 F.3 Using texinfo-show-structure . . . . . . . . . . . . . . . . . . . . . . . 207 F.4 Using occur . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 F.5 Finding Badly Referenced Nodes . . . . . . . . . . . . . . . . . . . . . . . 209

F.5.1 Running Info-validate. . . . . . . . . . . . . . . . . . . . . . 209 F.5.2 Creating an Unsplit File . . . . . . . . . . . . . . . . . . . . . . 209 F.5.3 Tagifying a File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 F.5.4 Splitting a File Manually . . . . . . . . . . . . . . . . . . . . . 210

Appendix G Refilling Paragraphs. . . . . . . . . . . 212

Appendix H @-Command Syntax. . . . . . . . . . . 213

Appendix I How to Obtain TEX . . . . . . . . . . . . 214

Appendix J Copying This Manual . . . . . . . . . . 215

J.1 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . 215

J.1.1 ADDENDUM: How to use this License for your documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

Command and Variable Index . . . . . . . . . . . . . . . 222

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226

Documentation is like sex: when it is good, it is very, very good; and when it is bad, it is better than nothing. —Dick Brandon

Texinfo Copying Conditions

The programs currently being distributed that relate to Texinfo include makeinfo, info, texindex, and ‘texinfo.tex’. These programs are free; this means that everyone is free to use them and free to redistribute them on a free basis. The Texinfo-related programs are not in the public domain; they are copyrighted and there are restrictions on their distribution, but these restrictions are designed to permit everything that a good cooperating citizen would want to do. What is not allowed is to try to prevent others from further sharing any version of these programs that they might get from you.

Specifically, we want to make sure that you have the right to give away copies of the programs that relate to Texinfo, that you receive source code or else can get it if you want it, that you can change these programs or use pieces of them in new free programs, and that you know you can do these things.

To make sure that everyone has such rights, we have to forbid you to deprive anyone else of these rights. For example, if you distribute copies of the Texinfo related programs, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must tell them their rights.

Also, for our own protection, we must make certain that everyone finds out that there is no warranty for the programs that relate to Texinfo. If these programs are modified by someone else and passed on, we want their recipients to know that what they have is not what we distributed, so that any problems introduced by others will not reflect on our reputation.

The precise conditions of the licenses for the programs currently being distributed that relate to Texinfo are found in the General Public Licenses that accompany them. This manual specifically is covered by the GNU Free Documentation License (see Section J.1 [GNU Free Documentation License], page 215).

1 Overview of Texinfo

Texinfo^{1} is a documentation system that uses a single source file to produce both online information and printed output. This means that instead of writing two different documents, one for the online information and the other for a printed work, you need write only one document. Therefore, when the work is revised, you need revise only that one document.

1.1 Reporting Bugs

We welcome bug reports and suggestions for any aspect of the Texinfo system, programs, documentation, installation, anything. Please email them to bug-texinfo@gnu.org. You can get the latest version of Texinfo from ftp://ftp.gnu.org/gnu/texinfo/ and its mirrors worldwide.

For bug reports, please include enough information for the maintainers to reproduce the problem. Generally speaking, that means:

• the version number of Texinfo and the program(s) or manual(s) involved.

• hardware and operating system names and versions.

• the contents of any input files necessary to reproduce the bug.

• a description of the problem and samples of any erroneous output.

• any unusual options you gave to configure.

• anything else that you think would be helpful.

When in doubt whether something is needed or not, include it. It’s better to include too much than to leave out something important.

Patches are most welcome; if possible, please make them with ‘diff -c’ (see section “Overview” in Comparing and Merging Files) and include ‘ChangeLog’ entries (see section “Change Log” in The GNU Emacs Manual).

When sending patches, if possible please do not encode or split them in any way; it’s much easier to deal with one plain text message, however large, than many small ones. GNU shar (ftp://ftp.gnu.org/gnu/sharutils/) is a convenient way of packaging multiple and/or binary files for email.

1.2 Using Texinfo

Using Texinfo, you can create a printed document with the normal features of a book, including chapters, sections, cross references, and indices. From the same Texinfo source file, you can create a menu-driven, online Info file with nodes, menus, cross references, and indices. You can also create from that same source file an HTML output file suitable for use with a web browser, or an XML file. The GNU Emacs Manual is a good example of a Texinfo file, as is this manual.

1 The first syllable of “Texinfo” is pronounced like “speck”, not “hex”. This odd pronunciation is derived

from, but is not the same as, the pronunciation of TEX. In the word TEX, the ‘X’ is actually the Greek

letter “chi” rather than the English letter “ex”. Pronounce TEX as if the ‘X’ were the last sound in the

name ‘Bach’; but pronounce Texinfo as if the ‘x’ were a ‘k’. Spell “Texinfo” with a capital “T” and

the other letters in lower case.

To make a printed document, you process a Texinfo source file with the TEX typesetting program (but the Texinfo language is very different and much stricter than TEX’s usual language, plain TEX). This creates a DVI file that you can typeset and print as a book or report (see Chapter 19 [Hardcopy], page 142).

To output an Info file, process your Texinfo source with the makeinfo utility or Emacs’s texinfo-format-buffer command. You can install the result in your Info tree (see Section 20.2 [Installing an Info File], page 161).

To output an HTML file, run makeinfo --html on your Texinfo source. You can (for example) install the result on your web site.

To output an XML file, run makeinfo --xml on your Texinfo source. To output Doc-Book (a particular form of XML), run makeinfo --docbook. If you want to convert from

Docbook to Texinfo, please see http://docbook2X.sourceforge.net/.

If you are a programmer and would like to contribute to the GNU project by implementing additional output formats for Texinfo, that would be excellent. But please do not write a separate translator texi2foo for your favorite format foo! That is the hard way to do the job, and makes extra work in subsequent maintenance, since the Texinfo language is continually being enhanced and updated. Instead, the best approach is modify makeinfo to generate the new format, as it does now for Info, plain text, HTML, XML, and DocBook.

TEX works with virtually all printers; Info works with virtually all computer terminals; the HTML output works with virtually all web browsers. Thus Texinfo can be used by almost any computer user.

A Texinfo source file is a plain ascii file containing text and @-commands (words preceded by an ‘@’) that tell the typesetting and formatting programs what to do. You may edit a Texinfo file with any text editor; but it is especially convenient to use GNU Emacs since that editor has a special mode, called Texinfo mode, that provides various Texinfo-related features. (See Chapter 2 [Texinfo Mode], page 14.)

Before writing a Texinfo source file, you should learn about nodes, menus, cross references, and the rest, for example by reading this manual.

You can use Texinfo to create both online help and printed manuals; moreover, Texinfo is freely redistributable. For these reasons, Texinfo is the official documentation format of the GNU project. More information is available at the GNU documentation web page

(http://www.gnu.org/doc/).

From time to time, proposals are made to generate traditional Unix man pages from Texinfo source. This is not likely to ever be supported, because man pages have a very strict conventional format. Merely enhancing makeinfo to output troff format would be insufficient. Generating a good man page therefore requires a completely different source than the typical Texinfo applications of writing a good user tutorial or a good reference manual. This makes generating man pages incompatible with the Texinfo design goal of not having to document the same information in different ways for different output formats. You might as well just write the man page directly.

Man pages still have their place, and if you wish to support them, the program help2man may be useful; it generates a traditional man page from the ‘--help’ output of a program. In fact, this is currently used to generate man pages for the Texinfo programs themselves. It is GNU software written by Brendan O’Dea, available from

ftp://ftp.gnu.org/gnu/help2man/.

1.3 Info files

An Info file is a Texinfo file formatted so that the Info documentation reading program can operate on it. (makeinfo and texinfo-format-buffer are two commands that convert a Texinfo file into an Info file.)

Info files are divided into pieces called nodes, each of which contains the discussion of one topic. Each node has a name, and contains both text for the user to read and pointers to other nodes, which are identified by their names. The Info program displays one node at a time, and provides commands with which the user can move to other related nodes.

Each node of an Info file may have any number of child nodes that describe subtopics of the node’s topic. The names of child nodes are listed in a menu within the parent node; this allows you to use certain Info commands to move to one of the child nodes. Generally, an Info file is organized like a book. If a node is at the logical level of a chapter, its child nodes are at the level of sections; likewise, the child nodes of sections are at the level of subsections.

All the children of any one parent are linked together in a bidirectional chain of ‘Next’ and ‘Previous’ pointers. The ‘Next’ pointer provides a link to the next section, and the ‘Previous’ pointer provides a link to the previous section. This means that all the nodes that are at the level of sections within a chapter are linked together. Normally the order in this chain is the same as the order of the children in the parent’s menu. Each child node records the parent node name as its ‘Up’ pointer. The last child has no ‘Next’ pointer, and the first child has the parent both as its ‘Previous’ and as its ‘Up’ pointer.^{2}

The book-like structuring of an Info file into nodes that correspond to chapters, sections, and the like is a matter of convention, not a requirement. The ‘Up’, ‘Previous’, and ‘Next’ pointers of a node can point to any other nodes, and a menu can contain any other nodes. Thus, the node structure can be any directed graph. But it is usually more comprehensible to follow a structure that corresponds to the structure of chapters and sections in a printed book or report.

In addition to menus and to ‘Next’, ‘Previous’, and ‘Up’ pointers, Info provides pointers of another kind, called references, that can be sprinkled throughout the text. This is usually the best way to represent links that do not fit a hierarchical structure.

Usually, you will design a document so that its nodes match the structure of chapters and sections in the printed output. But occasionally there are times when this is not right for the material being discussed. Therefore, Texinfo uses separate commands to specify the node structure for the Info file and the section structure for the printed output.

Generally, you enter an Info file through a node that by convention is named ‘Top’. This node normally contains just a brief summary of the file’s purpose, and a large menu through which the rest of the file is reached. From this node, you can either traverse the file systematically by going from node to node, or you can go to a specific node listed in the main menu, or you can search the index menus and then go directly to the node that has the information you want. Alternatively, with the standalone Info program, you can specify specific menu items on the command line (see section “Top” in Info).

2 In some documents, the first child has no ‘Previous’ pointer. Occasionally, the last child has the node

name of the next following higher level node as its ‘Next’ pointer.

If you want to read through an Info file in sequence, as if it were a printed manual, you can hit ^{⟨}\underline{\overline{SPC}}^{⟩} repeatedly, or you get the whole file with the advanced Info command g *. (See Info file ‘info’, node ‘Expert’.)

The ‘dir’ file in the ‘info’ directory serves as the departure point for the whole Info system. From it, you can reach the ‘Top’ nodes of each of the documents in a complete Info system.

If you wish to refer to an Info file in a URI, you can use the (unofficial) syntax exemplified in the following. This works with Emacs/W3, for example:

     info:///usr/info/emacs#Dissociated%20Press
     info:emacs#Dissociated%20Press
     info://localhost/usr/info/emacs#Dissociated%20Press

The info program itself does not follow URI’s of any kind.

1.4 Printed Books

A Texinfo file can be formatted and typeset as a printed book or manual. To do this, you need TEX, a powerful, sophisticated typesetting program written by Donald Knuth.^{3}

A Texinfo-based book is similar to any other typeset, printed work: it can have a title page, copyright page, table of contents, and preface, as well as chapters, numbered or unnumbered sections and subsections, page headers, cross references, footnotes, and indices.

You can use Texinfo to write a book without ever having the intention of converting it into online information. You can use Texinfo for writing a printed novel, and even to write a printed memo, although this latter application is not recommended since electronic mail is so much easier.

TEX is a general purpose typesetting program. Texinfo provides a file ‘texinfo.tex’ that contains information (definitions or macros) that TEX uses when it typesets a Texinfo file. (‘texinfo.tex’ tells TEX how to convert the Texinfo @-commands to TEX commands, which TEX can then process to create the typeset document.) ‘texinfo.tex’ contains the specifications for printing a document. You can get the latest version of ‘texinfo.tex’

from ftp://ftp.gnu.org/gnu/texinfo.tex.

In the United States, documents are most often printed on 8.5 inch by 11 inch pages (216 mm by 280 mm); this is the default size. But you can also print for 7 inch by 9.25 inch pages (178 mm by 235 mm, the @smallbook size; or on A4 or A5 size paper (@afourpaper, @afivepaper). (See Section 19.11 [Printing “Small” Books], page 150. Also, see Section 19.12 [Printing on A4 Paper], page 150.)

By changing the parameters in ‘texinfo.tex’, you can change the size of the printed document. In addition, you can change the style in which the printed document is formatted; for example, you can change the sizes and fonts used, the amount of indentation for each paragraph, the degree to which words are hyphenated, and the like. By changing the specifications, you can make a book look dignified, old and serious, or light-hearted, young and cheery.

3 You can also use the texi2roff (ftp://tug.org/texi2roff.tar.gz) program if you do not have TEX;

since Texinfo is designed for use with TEX, texi2roff is not described here. texi2roff is not part

of the standard GNU distribution and is not maintained or up-to-date with all the Texinfo features

described in this manual.

TEX is freely distributable. It is written in a superset of Pascal called WEB and can be compiled either in Pascal or (by using a conversion program that comes with the TEX distribution) in C. (See section “TEX Mode” in The GNU Emacs Manual, for information about TEX.)

TEX is very powerful and has a great many features. Because a Texinfo file must be able to present information both on a character-only terminal in Info form and in a typeset book, the formatting commands that Texinfo supports are necessarily limited.

To get a copy of TEX, see Appendix I [How to Obtain TEX], page 214.

1.5 @-commands

In a Texinfo file, the commands that tell TEX how to typeset the printed manual and tell makeinfo and texinfo-format-buffer how to create an Info file are preceded by ‘@’; they are called @-commands. For example, @node is the command to indicate a node and @chapter is the command to indicate the start of a chapter.

Please note: All the @-commands, with the exception of the @TeX{} command, must be written entirely in lower case. The Texinfo @-commands are a strictly limited set of constructs. The strict limits make

it possible for Texinfo files to be understood both by TEX and by the code that converts them into Info files. You can display Info files on any terminal that displays alphabetic and numeric characters. Similarly, you can print the output generated by TEX on a wide variety of printers.

Depending on what they do or what arguments^{4} they take, you need to write @- commands on lines of their own or as part of sentences:

• Write a command such as @noindent at the beginning of a line as the only text on the line. (@noindent prevents the beginning of the next line from being indented as the beginning of a paragraph.)

• Write a command such as @chapter at the beginning of a line followed by the command’s arguments, in this case the chapter title, on the rest of the line. (@chapter creates chapter titles.)

• Write a command such as @dots{} wherever you wish but usually within a sentence. (@dots{} creates dots …)

• Write a command such as @code{sample-code} wherever you wish (but usually within a sentence) with its argument, sample-code in this example, between the braces. (@code marks text as being code.)

4 The word argument comes from the way it is used in mathematics and does not refer to a dispute

between two people; it refers to the information presented to the command. According to the Oxford

English Dictionary, the word derives from the Latin for to make clear, prove; thus it came to mean

‘the evidence offered as proof’, which is to say, ‘the information offered’, which led to its mathematical

meaning. In its other thread of derivation, the word came to mean ‘to assert in a manner against which

others may make counter assertions’, which led to the meaning of ‘argument’ as a dispute.

• Write a command such as @example on a line of its own; write the body-text on following lines; and write the matching @end command, @end example in this case, at the on a line of its own after the body-text. (@example … @end example indents and typesets body-text as an example.) It’s usually ok to indent environment commands like this, but in complicated and hard-to-define circumstances the extra spaces cause extra space to appear in the output, so beware.

As a general rule, a command requires braces if it mingles among other text; but it does not need braces if it starts a line of its own. The non-alphabetic commands, such as @:, are exceptions to the rule; they do not need braces.

As you gain experience with Texinfo, you will rapidly learn how to write the different commands: the different ways to write commands make it easier to write and read Texinfo files than if all commands followed exactly the same syntax. (For details about @-command syntax, see Appendix H [@-Command Syntax], page 213.)

1.6 General Syntactic Conventions

This section describes the general conventions used in all Texinfo documents.

• All printable ascii characters except ‘@’, ‘{’ and ‘}’ can appear in a Texinfo file and stand for themselves. ‘@’ is the escape character which introduces commands, while ‘{’ and ‘}’ are used to surround arguments to certain commands. To put one of these special characters into the document, put an ‘@’ character in front of it, like this: ‘@@’, ‘@{’, and ‘@}’.

• It is customary in TEX to use doubled single-quote characters to begin and end quotations: ``...''. This convention should be followed in Texinfo files. TEX converts two single quotes to left- and right-hand doubled quotation marks, “like this”, and Info converts doubled single-quote characters to ascii double-quotes: ``...'' becomes

"...".

• Use three hyphens in a row, ‘---’, for a dash—like this. In TEX, a single or double hyphen produces a printed dash that is shorter than the usual typeset dash. Info reduces three hyphens to two for display on the screen.

• To prevent a paragraph from being indented in the printed manual, put the command @noindent on a line by itself before the paragraph.

• If you mark off a region of the Texinfo file with the @iftex and @end iftex commands, that region will appear only in the printed copy; in that region, you can use certain commands borrowed from plain TEX that you cannot use in Info. Conversely, text surrounded by @ifnottex and @end ifnottex will appear in all output formats except TEX. Each of the other output formats (html, info, plaintext) have an analogous pair of commands. See Chapter 16 [Conditionals], page 129.

Caution: Do not use tab characters in a Texinfo file (except in verbatim modes)! TEX uses variable-width fonts, which means that it is impractical at best to define a tab to work in all circumstances. Consequently, TEX treats tabs like single spaces, and that is not what they look like. Furthermore, makeinfo does nothing special with tabs, and thus a tab character in your input file may appear differently in the output, for example, in indented text.

To avoid this problem, Texinfo mode causes GNU Emacs to insert multiple spaces when you press the ^{⟨}\underline{\overline{TAB}}^{⟩} key. Also, you can run untabify in Emacs to convert tabs in a region to multiple spaces.

1.7 Comments

You can write comments in a Texinfo file that will not appear in either the Info file or the printed manual by using the @comment command (which may be abbreviated to @c). Such comments are for the person who revises the Texinfo file. All the text on a line that follows either @comment or @c is a comment; the rest of the line does not appear in either the Info file or the printed manual.

Often, you can write the @comment or @c in the middle of a line, and only the text that follows after the @comment or @c command does not appear; but some commands, such as @settitle and @setfilename, work on a whole line. You cannot use @comment or @c in a line beginning with such a command.

You can write long stretches of text that will not appear in either the Info file or the printed manual by using the @ignore and @end ignore commands. Write each of these commands on a line of its own, starting each command at the beginning of the line. Text between these two commands does not appear in the processed output. You can use @ignore and @end ignore for writing comments.

Text enclosed by @ignore or by failing @ifset or @ifclear conditions is ignored in the sense that it will not contribute to the formatted output. However, TEX and makeinfo must still parse the ignored text, in order to understand when to stop ignoring text from the source file; that means that you may still get error messages if you have invalid Texinfo commands within ignored text.

1.8 What a Texinfo File Must Have

By convention, the namea of a Texinfo file ends with (in order of preference) one of the extensions ‘.texinfo’, ‘.texi’, ‘.txi’, or ‘.tex’. The longer extensions are preferred since they describe more clearly to a human reader the nature of the file. The shorter extensions are for operating systems that cannot handle long file names.

In order to be made into a printed manual and an Info file, a Texinfo file must begin with lines like this:

     \input texinfo

@setfilename info-file-name @settitle name-of-manual

The contents of the file follow this beginning, and then you must end a Texinfo file with a line like this:

     @bye

Here’s an explanation:

• The ‘\input texinfo’ line tells TEX to use the ‘texinfo.tex’ file, which tells TEX how to translate the Texinfo @-commands into TEX typesetting commands. (Note the use of the backslash, ‘\’; this is correct for TEX.)

• The @setfilename line provides a name for the Info file and tells TEX to open auxiliary files. All text before @setfilename is ignored!

• The @settitle line specifies a title for the page headers (or footers) of the printed manual, and the default document description for the ‘<head>’ in HTML format. Strictly speaking, @settitle is optional—if you don’t mind your document being titled ‘Untitled’.

• The @bye line at the end of the file on a line of its own tells the formatters that the file is ended and to stop formatting.

Typically, you will not use quite such a spare format, but will include mode setting and start-of-header and end-of-header lines at the beginning of a Texinfo file, like this:

     \input texinfo   @c -*-texinfo-*-
     @c %**start of header

@setfilename info-file-name @settitle name-of-manual

     @c %**end of header

In the first line, ‘-*-texinfo-*-’ causes Emacs to switch into Texinfo mode when you edit the file.

The @c lines which surround the @setfilename and @settitle lines are optional, but you need them in order to run TEX or Info on just part of the file. (See Section 3.2.2 [Start of Header], page 28.)

Furthermore, you will usually provide a Texinfo file with a title page, indices, and the like, all of which are explained in this manual. But the minimum, which can be useful for short documents, is just the three lines at the beginning and the one line at the end.

1.9 Six Parts of a Texinfo File

Generally, a Texinfo file contains more than the minimal beginning and end described in the previous section—it usually contains the six parts listed below. These are described fully in the following sections.

1. Header The Header names the file, tells TEX which definitions file to use, and other such housekeeping tasks.

2. Summary and Copyright The Summary and Copyright segment describes the document and contains the copyright notice and copying permissions. This is done with the @copying command.

3. Title and Copyright The Title and Copyright segment contains the title and copyright pages for the printed manual. The segment must be enclosed between @titlepage and @end titlepage commands. The title and copyright page appear only in the printed manual.

4. ‘Top’ Node and Master Menu The ‘Top’ node starts off the online output; it does not appear in the printed manual. We recommend including the copying permissions here as well as the segments above. And it contains at least a top-level menu listing the chapters, and possibly a Master Menu listing all the nodes in the entire document.

5. Body The Body of the document is typically structured like a traditional book or encyclopedia, but it may be free form.

6. End The End segment contains commands for printing indices and generating the table of contents, and the @bye command on a line of its own.

1.10 A Short Sample Texinfo File

Here is a very short but complete Texinfo file, in the six conventional parts enumerated in the previous section, so you can see how Texinfo source appears in practice. The first three parts of the file, from ‘\input texinfo’ through to ‘@end titlepage’, look more intimidating than they are: most of the material is standard boilerplate; when writing a manual, you simply change the names as appropriate.

See Chapter 3 [Beginning a File], page 26, for full documentation on the commands listed here. See Section C.2 [GNU Sample Texts], page 192, for the full texts to be used in GNU manuals.

In the following, the sample text is indented; comments on it are not. The complete file, without interspersed comments, is shown in Section C.1 [Short Sample Texinfo File], page 191.

Part 1: Header

The header does not appear in either the Info file or the printed output. It sets various parameters, including the name of the Info file and the title used in the header.

     \input texinfo   @c -*-texinfo-*-
     @c %**start of header
     @setfilename sample.info
     @settitle Sample Manual 1.0
     @c %**end of header

Part 2: Summary Description and Copyright

A real manual includes more text here, according to the license under which it is distributed. See Section C.2 [GNU Sample Texts], page 192.

     @copying
     This is a short example of a complete Texinfo file, version 1.0.
     Copyright @copyright{} 2002 Free Software Foundation, Inc.
     @end copying

Part 3: Titlepage, Contents, Copyright

The titlepage segment does not appear in the online output, only in the printed manual. We use the @insertcopying command to include the permission text from the previous section, instead of writing it out again; it is output on the back of the title page. The @contents command generates a table of contents.

@titlepage
@title Sample Title
@c The following two commands start the copyright page.
@page
@vskip 0pt plus 1filll
@insertcopying
@end titlepage
@c Output the table of contents at the beginning.
@contents

Part 4: ‘Top’ Node and Master Menu

The ‘Top’ node contains the master menu for the Info file. Since a printed manual uses a table of contents rather than a menu, the master menu appears only in online output. We also include the copying text again for the benefit of online readers. And since the copying text begins with a brief description of the manual, no other text is needed.

@ifnottex
@node Top
@end ifnottex
@insertcopying
@menu
* First Chapter::    The first chapter is the
                       only chapter in this sample.
* Index::            Complete index.
@end menu

Part 5: The Body of the Document

The body segment contains all the text of the document, but not the indices or table of contents. This example illustrates a node and a chapter containing an enumerated list.

@node First Chapter
@chapter First Chapter
@cindex chapter, first
This is the first chapter.
@cindex index entry, another
Here is a numbered list.
@enumerate
@item
This is the first item.
@item
This is the second item.
@end enumerate

Part 6: The End of the Document

The end segment contains commands for generating an index in a node and unnumbered chapter of its own, and the @bye command that marks the end of the document.

     @node Index
     @unnumbered Index
     @printindex cp
     @bye

Some Results

Here is what the contents of the first chapter of the sample look like:

This is the first chapter. Here is a numbered list.

1. This is the first item.

2. This is the second item.

1.11 History

Richard M. Stallman invented the Texinfo format, wrote the initial processors, and created Edition 1.0 of this manual. Robert J. Chassell greatly revised and extended the manual, starting with Edition 1.1. Brian Fox was responsible for the standalone Texinfo distribution until version 3.8, and wrote the standalone makeinfo and info programs. Karl Berry has continued maintenance since Texinfo 3.8 (manual edition 2.22).

Our thanks go out to all who helped improve this work, particularly the indefatigable Eli Zaretskii and Andreas Schwab, who have provided patches beyond counting. François Pinard and David D. Zuhn, tirelessly recorded and reported mistakes and obscurities. Zack Weinberg did the impossible by implementing the macro syntax in ‘texinfo.tex’. Special thanks go to Melissa Weisshaus for her frequent reviews of nearly similar editions. Dozens of others have contributed patches and suggestions, they are gratefully acknowledged in the ‘ChangeLog’ file. Our mistakes are our own.

A bit of history: in the 1970’s at CMU, Brian Reid developed a program and format named Scribe to mark up documents for printing. It used the @ character to introduce commands, as Texinfo does. Much more consequentially, it strived to describe document contents rather than formatting, an idea wholeheartedly adopted by Texinfo.

Meanwhile, people at MIT developed another, not too dissimilar format called Bolio. This then was converted to using TEX as its typesetting language: BoTEX. The earliest BoTEX version seems to have been 0.02 on October 31, 1984.

BoTEX could only be used as a markup language for documents to be printed, not for online documents. Richard Stallman (RMS) worked on both Bolio and BoTEX. He also developed a nifty on-line help format called Info, and then combined BoTEX and Info to create Texinfo, a mark up language for text that is intended to be read both online and as printed hard copy.

2 Using Texinfo Mode

You may edit a Texinfo file with any text editor you choose. A Texinfo file is no different from any other ascii file. However, GNU Emacs comes with a special mode, called Texinfo mode, that provides Emacs commands and tools to help ease your work.

This chapter describes features of GNU Emacs’ Texinfo mode but not any features of the Texinfo formatting language. So if you are reading this manual straight through from the beginning, you may want to skim through this chapter briefly and come back to it after reading succeeding chapters which describe the Texinfo formatting language in detail.

Texinfo mode provides special features for working with Texinfo files. You can:

• Insert frequently used @-commands.

• Automatically create @node lines.

• Show the structure of a Texinfo source file.

• Automatically create or update the ‘Next’, ‘Previous’, and ‘Up’ pointers of a node.

• Automatically create or update menus.

• Automatically create a master menu.

• Format a part or all of a file for Info.

• Typeset and print part or all of a file.

Perhaps the two most helpful features are those for inserting frequently used @-commands and for creating node pointers and menus.

2.1 The Usual GNU Emacs Editing Commands

In most cases, the usual Text mode commands work the same in Texinfo mode as they do in Text mode. Texinfo mode adds new editing commands and tools to GNU Emacs’ general purpose editing features. The major difference concerns filling. In Texinfo mode, the paragraph separation variable and syntax table are redefined so that Texinfo commands that should be on lines of their own are not inadvertently included in paragraphs. Thus, the M-q (fill-paragraph) command will refill a paragraph but not mix an indexing command on a line adjacent to it into the paragraph.

In addition, Texinfo mode sets the page-delimiter variable to the value of texinfochapter-level-regexp; by default, this is a regular expression matching the commands for chapters and their equivalents, such as appendices. With this value for the page delimiter, you can jump from chapter title to chapter title with the C-x ] (forward-page) and C-x [ (backward-page) commands and narrow to a chapter with the C-x p (narrow-to-page) command. (See section “Pages” in The GNU Emacs Manual, for details about the page commands.)

You may name a Texinfo file however you wish, but the convention is to end a Texinfo file name with one of the extensions ‘.texinfo’, ‘.texi’, ‘.txi’, or ‘.tex’. A longer extension is preferred, since it is explicit, but a shorter extension may be necessary for operating systems that limit the length of file names. GNU Emacs automatically enters Texinfo mode when you visit a file with a ‘.texinfo’, ‘.texi’ or ‘.txi’ extension. Also, Emacs switches to Texinfo mode when you visit a file that has ‘-*-texinfo-*-’ in its first line. If ever you are in another mode and wish to switch to Texinfo mode, type M-x texinfo-mode.

Like all other Emacs features, you can customize or enhance Texinfo mode as you wish. In particular, the keybindings are very easy to change. The keybindings described here are the default or standard ones.

2.2 Inserting Frequently Used Commands

Texinfo mode provides commands to insert various frequently used @-commands into the buffer. You can use these commands to save keystrokes.

The insert commands are invoked by typing C-c twice and then the first letter of the @-command:

C-c C-c c
M-x texinfo-insert-@code

Insert @code{} and put the cursor between the braces.

C-c C-c d
M-x texinfo-insert-@dfn

Insert @dfn{} and put the cursor between the braces.

C-c C-c e
M-x texinfo-insert-@end

Insert @end and attempt to insert the correct following word, such as ‘example’ or ‘table’. (This command does not handle nested lists correctly, but inserts the word appropriate to the immediately preceding list.)

C-c C-c i
M-x texinfo-insert-@item

Insert @item and put the cursor at the beginning of the next line.

C-c C-c k
M-x texinfo-insert-@kbd

Insert @kbd{} and put the cursor between the braces.

C-c C-c n
M-x texinfo-insert-@node

Insert @node and a comment line listing the sequence for the ‘Next’, ‘Previous’, and ‘Up’ nodes. Leave point after the @node.

C-c C-c o
M-x texinfo-insert-@noindent

Insert @noindent and put the cursor at the beginning of the next line.

C-c C-c s
M-x texinfo-insert-@samp

Insert @samp{} and put the cursor between the braces.

C-c C-c t
M-x texinfo-insert-@table

Insert @table followed by a ^{⟨}\underline{\overline{SPC}}^{⟩} and leave the cursor after the ^{⟨}\underline{\overline{SPC}}^{⟩}.

C-c C-c v
M-x texinfo-insert-@var

Insert @var{} and put the cursor between the braces.

C-c C-c x
M-x texinfo-insert-@example

Insert @example and put the cursor at the beginning of the next line.

C-c C-c {
M-x texinfo-insert-braces

Insert {} and put the cursor between the braces.

C-c C-c }
C-c C-c ]
M-x up-list

Move from between a pair of braces forward past the closing brace. Typing C-c C-c ] is easier than typing C-c C-c }, which is, however, more mnemonic; hence the two keybindings. (Also, you can move out from between braces by typing C-f.)

To put a command such as @code{...} around an existing word, position the cursor in front of the word and type C-u 1 C-c C-c c. This makes it easy to edit existing plain text. The value of the prefix argument tells Emacs how many words following point to include between braces—‘1’ for one word, ‘2’ for two words, and so on. Use a negative argument to enclose the previous word or words. If you do not specify a prefix argument, Emacs inserts the @-command string and positions the cursor between the braces. This feature works only for those @-commands that operate on a word or words within one line, such as @kbd and

@var.

This set of insert commands was created after analyzing the frequency with which different @-commands are used in the GNU Emacs Manual and the GDB Manual. If you wish to add your own insert commands, you can bind a keyboard macro to a key, use abbreviations, or extend the code in ‘texinfo.el’.

    C-c C-c C-d (texinfo-start-menu-description) is an insert command that works

differently from the other insert commands. It inserts a node’s section or chapter title in the space for the description in a menu entry line. (A menu entry has three parts, the entry name, the node name, and the description. Only the node name is required, but a description helps explain what the node is about. See Section 7.2 [The Parts of a Menu], page 58.)

To use texinfo-start-menu-description, position point in a menu entry line and type C-c C-c C-d. The command looks for and copies the title that goes with the node name, and inserts the title as a description; it positions point at beginning of the inserted text so you can edit it. The function does not insert the title if the menu entry line already contains a description.

This command is only an aid to writing descriptions; it does not do the whole job. You must edit the inserted text since a title tends to use the same words as a node name but a useful description uses different words.

2.3 Showing the Section Structure of a File

You can show the section structure of a Texinfo file by using the C-c C-s command (texinfo-show-structure). This command shows the section structure of a Texinfo file by listing the lines that begin with the @-commands for @chapter, @section, and the like.

It constructs what amounts to a table of contents. These lines are displayed in another buffer called the ‘*Occur*’ buffer. In that buffer, you can position the cursor over one of the lines and use the C-c C-c command (occur-mode-goto-occurrence), to jump to the corresponding spot in the Texinfo file.

C-c C-s
M-x texinfo-show-structure

Show the @chapter, @section, and such lines of a Texinfo file.

C-c C-c
M-x occur-mode-goto-occurrence

Go to the line in the Texinfo file corresponding to the line under the cursor in the ‘*Occur*’ buffer.

If you call texinfo-show-structure with a prefix argument by typing C-u C-c C-s, it will list not only those lines with the @-commands for @chapter, @section, and the like, but also the @node lines. You can use texinfo-show-structure with a prefix argument to check whether the ‘Next’, ‘Previous’, and ‘Up’ pointers of an @node line are correct.

Often, when you are working on a manual, you will be interested only in the structure of the current chapter. In this case, you can mark off the region of the buffer that you are interested in by using the C-x n n (narrow-to-region) command and texinfo-show-structure will work on only that region. To see the whole buffer again, use C-x n w (widen). (See section “Narrowing” in The GNU Emacs Manual, for more information about the narrowing commands.)

In addition to providing the texinfo-show-structure command, Texinfo mode sets the value of the page delimiter variable to match the chapter-level @-commands. This enables

you to use the C-x ] (forward-page) and C-x [ (backward-page) commands to move

forward and backward by chapter, and to use the C-x p (narrow-to-page) command to narrow to a chapter. See section “Pages” in The GNU Emacs Manual, for more information about the page commands.

2.4 Updating Nodes and Menus

Texinfo mode provides commands for automatically creating or updating menus and node pointers. The commands are called “update” commands because their most frequent use is for updating a Texinfo file after you have worked on it; but you can use them to insert the ‘Next’, ‘Previous’, and ‘Up’ pointers into an @node line that has none and to create menus in a file that has none.

If you do not use the updating commands, you need to write menus and node pointers by hand, which is a tedious task.

You can use the updating commands to:

• insert or update the ‘Next’, ‘Previous’, and ‘Up’ pointers of a node,

• insert or update the menu for a section, and

• create a master menu for a Texinfo source file.

You can also use the commands to update all the nodes and menus in a region or in a whole Texinfo file.

The updating commands work only with conventional Texinfo files, which are structured hierarchically like books. In such files, a structuring command line must follow closely after each @node line, except for the ‘Top’ @node line. (A structuring command line is a line beginning with @chapter, @section, or other similar command.)

You can write the structuring command line on the line that follows immediately after an @node line or else on the line that follows after a single @comment line or a single @ifinfo line. You cannot interpose more than one line between the @node line and the structuring command line; and you may interpose only an @comment line or an @ifinfo line.

Commands which work on a whole buffer require that the ‘Top’ node be followed by a node with an @chapter or equivalent-level command. The menu updating commands will not create a main or master menu for a Texinfo file that has only @chapter-level nodes! The menu updating commands only create menus within nodes for lower level nodes. To create a menu of chapters, you must provide a ‘Top’ node.

The menu updating commands remove menu entries that refer to other Info files since they do not refer to nodes within the current buffer. This is a deficiency. Rather than use menu entries, you can use cross references to refer to other Info files. None of the updating commands affect cross references.

Texinfo mode has five updating commands that are used most often: two are for updating the node pointers or menu of a single node (or a region); two are for updating every node pointer and menu in a file; and one, the texinfo-master-menu command, is for creating a master menu for a complete file, and optionally, for updating every node and menu in the whole Texinfo file.

The texinfo-master-menu command is the primary command:

C-c C-u m
M-x texinfo-master-menu

Create or update a master menu that includes all the other menus (incorporating the descriptions from pre-existing menus, if any).

With an argument (prefix argument, C-u, if interactive), first create or update all the nodes and all the regular menus in the buffer before constructing the master menu. (See Section 3.5 [The Top Node and Master Menu], page 36, for more about a master menu.)

For texinfo-master-menu to work, the Texinfo file must have a ‘Top’ node and at least one subsequent node.

After extensively editing a Texinfo file, you can type the following:

               C-u M-x texinfo-master-menu

or

               C-u C-c C-u m

This updates all the nodes and menus completely and all at once.

The other major updating commands do smaller jobs and are designed for the person who updates nodes and menus as he or she writes a Texinfo file.

The commands are:

C-c C-u C-n
M-x texinfo-update-node

Insert the ‘Next’, ‘Previous’, and ‘Up’ pointers for the node that point is within (i.e., for the @node line preceding point). If the @node line has pre-existing ‘Next’, ‘Previous’, or ‘Up’ pointers in it, the old pointers are removed and new ones inserted. With an argument (prefix argument, C-u, if interactive), this command updates all @node lines in the region (which is the text between point and mark).

C-c C-u C-m
M-x texinfo-make-menu

Create or update the menu in the node that point is within. With an argument (C-u as prefix argument, if interactive), the command makes or updates menus for the nodes which are either within or a part of the region. Whenever texinfo-make-menu updates an existing menu, the descriptions from that menu are incorporated into the new menu. This is done by copying descriptions from the existing menu to the entries in the new menu that have the same node names. If the node names are different, the descriptions are not copied to the new menu.

C-c C-u C-e
M-x texinfo-every-node-update

Insert or update the ‘Next’, ‘Previous’, and ‘Up’ pointers for every node in the buffer.

C-c C-u C-a
M-x texinfo-all-menus-update

Create or update all the menus in the buffer. With an argument (C-u as prefix argument, if interactive), first insert or update all the node pointers before working on the menus. If a master menu exists, the texinfo-all-menus-update command updates it; but the command does not create a new master menu if none already exists. (Use the texinfo-master-menu command for that.) When working on a document that does not merit a master menu, you can type the following:

     C-u C-c C-u C-a

or

     C-u M-x texinfo-all-menus-update

This updates all the nodes and menus.

The texinfo-column-for-description variable specifies the column to which menu descriptions are indented. By default, the value is 32 although it is often useful to reduce it to as low as 24. You can set the variable with the M-x edit-options command (see section “Editing Variable Values” in The GNU Emacs Manual) or with the M-x set-variable command (see section “Examining and Setting Variables” in The GNU Emacs Manual).

Also, the texinfo-indent-menu-description command may be used to indent existing menu descriptions to a specified column. Finally, if you wish, you can use the texinfoinsert-node-lines command to insert missing @node lines into a file. (See Section 2.4.2 [Other Updating Commands], page 20, for more information.)

2.4.1 Updating Requirements

To use the updating commands, you must organize the Texinfo file hierarchically with chapters, sections, subsections, and the like. When you construct the hierarchy of the manual, do not ‘jump down’ more than one level at a time: you can follow the ‘Top’ node with a chapter, but not with a section; you can follow a chapter with a section, but not with a subsection. However, you may ‘jump up’ any number of levels at one time—for example, from a subsection to a chapter.

Each @node line, with the exception of the line for the ‘Top’ node, must be followed by a line with a structuring command such as @chapter, @section, or @unnumberedsubsec.

Each @node line/structuring-command line combination must look either like this:

     @node     Comments,  Minimum, Conventions, Overview
     @comment  node-name, next,    previous,    up
     @section Comments

or like this (without the @comment line):

     @node Comments, Minimum, Conventions, Overview
     @section Comments

or like this (without the explicit node pointers):

     @node Comments
     @section Comments

In this example, ‘Comments’ is the name of both the node and the section. The next node is called ‘Minimum’ and the previous node is called ‘Conventions’. The ‘Comments’ section is within the ‘Overview’ node, which is specified by the ‘Up’ pointer. (Instead of an @comment line, you may also write an @ifinfo line.)

If a file has a ‘Top’ node, it must be called ‘top’ or ‘Top’ and be the first node in the file.

The menu updating commands create a menu of sections within a chapter, a menu of subsections within a section, and so on. This means that you must have a ‘Top’ node if you want a menu of chapters.

Incidentally, the makeinfo command will create an Info file for a hierarchically organized Texinfo file that lacks ‘Next’, ‘Previous’ and ‘Up’ pointers. Thus, if you can be sure that your Texinfo file will be formatted with makeinfo, you have no need for the update node commands. (See Section 20.1 [Creating an Info File], page 153, for more information about makeinfo.) However, both makeinfo and the texinfo-format-... commands require that you insert menus in the file.

2.4.2 Other Updating Commands

In addition to the five major updating commands, Texinfo mode possesses several less frequently used updating commands:

M-x texinfo-insert-node-lines

Insert @node lines before the @chapter, @section, and other sectioning commands wherever they are missing throughout a region in a Texinfo file.

With an argument (C-u as prefix argument, if interactive), the texinfoinsert-node-lines command not only inserts @node lines but also inserts the chapter or section titles as the names of the corresponding nodes. In addition, it inserts the titles as node names in pre-existing @node lines that lack names. Since node names should be more concise than section or chapter titles, you must manually edit node names so inserted.

For example, the following marks a whole buffer as a region and inserts @node lines and titles throughout:

     C-x h C-u M-x texinfo-insert-node-lines

This command inserts titles as node names in @node lines; the texinfo-start-menu-description command (see Section 2.2 [Inserting], page 15) inserts titles as descriptions in menu entries, a different action. However, in both cases, you need to edit the inserted text.

M-x texinfo-multiple-files-update

Update nodes and menus in a document built from several separate files. With C-u as a prefix argument, create and insert a master menu in the outer file. With a numeric prefix argument, such as C-u 2, first update all the menus and all the ‘Next’, ‘Previous’, and ‘Up’ pointers of all the included files before creating and inserting a master menu in the outer file. The texinfo-multiple-files-update command is described in the appendix on @include files. See

Section D.2 [texinfo-multiple-files-update], page 196.
M-x texinfo-indent-menu-description

Indent every description in the menu following point to the specified column. You can use this command to give yourself more space for descriptions. With an argument (C-u as prefix argument, if interactive), the texinfo-indent-menu-description command indents every description in every menu in the region. However, this command does not indent the second and subsequent lines of a multi-line description.

M-x texinfo-sequential-node-update

Insert the names of the nodes immediately following and preceding the current node as the ‘Next’ or ‘Previous’ pointers regardless of those nodes’ hierarchical level. This means that the ‘Next’ node of a subsection may well be the next chapter. Sequentially ordered nodes are useful for novels and other documents that you read through sequentially. (However, in Info, the g * command lets you look through the file sequentially, so sequentially ordered nodes are not strictly necessary.) With an argument (prefix argument, if interactive), the texinfosequential-node-update command sequentially updates all the nodes in the region.

2.5 Formatting for Info

Texinfo mode provides several commands for formatting part or all of a Texinfo file for Info. Often, when you are writing a document, you want to format only part of a file—that is, a region.

    You can use either the texinfo-format-region or the makeinfo-region command

to format a region:

C-c C-e C-r
M-x texinfo-format-region
C-c C-m C-r
M-x makeinfo-region

Format the current region for Info.

    You can use either the texinfo-format-buffer or the makeinfo-buffer command

to format a whole buffer:

C-c C-e C-b
M-x texinfo-format-buffer
C-c C-m C-b
M-x makeinfo-buffer

Format the current buffer for Info.

For example, after writing a Texinfo file, you can type the following:

     C-u C-c C-u m

or

     C-u M-x texinfo-master-menu

This updates all the nodes and menus. Then type the following to create an Info file:

     C-c C-m C-b

or

     M-x makeinfo-buffer

For TEX or the Info formatting commands to work, the file must include a line that has @setfilename in its header.

See Section 20.1 [Creating an Info File], page 153, for details about Info formatting.

2.6 Formatting and Printing

Typesetting and printing a Texinfo file is a multi-step process in which you first create a file for printing (called a DVI file), and then print the file. Optionally, you may also create indices. To do this, you must run the texindex command after first running the tex typesetting command; and then you must run the tex command again. Or else run the texi2dvi command which automatically creates indices as needed (see Section 19.3 [Format with texi2dvi], page 144).

Often, when you are writing a document, you want to typeset and print only part of a file to see what it will look like. You can use the texinfo-tex-region and related commands for this purpose. Use the texinfo-tex-buffer command to format all of a buffer.

C-c C-t C-b
M-x texinfo-tex-buffer

Run texi2dvi on the buffer. In addition to running TEX on the buffer, this command automatically creates or updates indices as needed.

C-c C-t C-r
M-x texinfo-tex-region

Run TEX on the region.

C-c C-t C-i
M-x texinfo-texindex

Run texindex to sort the indices of a Texinfo file formatted with texinfo-

          tex-region. The texinfo-tex-region command does not run texindex au-

tomatically; it only runs the tex typesetting command. You must run the texinfo-tex-region command a second time after sorting the raw index files with the texindex command. (Usually, you do not format an index when you format a region, only when you format a buffer. Now that the texi2dvi command exists, there is little or no need for this command.)

C-c C-t C-p
M-x texinfo-tex-print

Print the file (or the part of the file) previously formatted with texinfo-tex-

          buffer or texinfo-tex-region.
    For texinfo-tex-region or texinfo-tex-buffer to work, the file must start with a

‘\input texinfo’ line and must include an @settitle line. The file must end with @bye on a line by itself. (When you use texinfo-tex-region, you must surround the @settitle line with start-of-header and end-of-header lines.)

See Chapter 19 [Hardcopy], page 142, for a description of the other TEX related com-

mands, such as tex-show-print-queue.

2.7 Texinfo Mode Summary

In Texinfo mode, each set of commands has default keybindings that begin with the same keys. All the commands that are custom-created for Texinfo mode begin with C-c. The keys are somewhat mnemonic.

Insert Commands

The insert commands are invoked by typing C-c twice and then the first letter of the @-command to be inserted. (It might make more sense mnemonically to use C-c C-i, for ‘custom insert’, but C-c C-c is quick to type.)

     C-c C-c c  Insert `@code'.
     C-c C-c d  Insert `@dfn'.
     C-c C-c e  Insert `@end'.
     C-c C-c i  Insert `@item'.
     C-c C-c n  Insert `@node'.
     C-c C-c s  Insert `@samp'.
     C-c C-c v  Insert `@var'.

C-c C-c { Insert braces.

C-c C-c ]

C-c C-c } Move out of enclosing braces.

C-c C-c C-d Insert a node’s section title in the space for the description in a menu entry line.

Show Structure

The texinfo-show-structure command is often used within a narrowed region. C-c C-s List all the headings.

The Master Update Command

The texinfo-master-menu command creates a master menu; and can be used to update every node and menu in a file as well.

C-c C-u m
M-x texinfo-master-menu

Create or update a master menu.

C-u C-c C-u m With C-u as a prefix argument, first create or update all nodes and regular menus, and then create a master menu.

Update Pointers

The update pointer commands are invoked by typing C-c C-u and then either C-n for

texinfo-update-node or C-e for texinfo-every-node-update.

C-c C-u C-n Update a node. C-c C-u C-e Update every node in the buffer.

Update Menus

Invoke the update menu commands by typing C-c C-u and then either C-m for texinfo-

make-menu or C-a for texinfo-all-menus-update. To update both nodes and menus at

the same time, precede C-c C-u C-a with C-u.

C-c C-u C-m Make or update a menu.

C-c C-u C-a Make or update all menus in a buffer.

C-u C-c C-u C-a With C-u as a prefix argument, first create or update all nodes and then create or update all menus.

Format for Info

The Info formatting commands that are written in Emacs Lisp are invoked by typing C-c C-e and then either C-r for a region or C-b for the whole buffer.

The Info formatting commands that are written in C and based on the makeinfo program are invoked by typing C-c C-m and then either C-r for a region or C-b for the whole buffer.

Use the texinfo-format... commands:

C-c C-e C-r Format the region. C-c C-e C-b Format the buffer.

Use makeinfo:

C-c C-m C-r Format the region. C-c C-m C-b Format the buffer. C-c C-m C-l Recenter the makeinfo output buffer. C-c C-m C-k Kill the makeinfo formatting job.

Typeset and Print

The TEX typesetting and printing commands are invoked by typing C-c C-t and then

another control command: C-r for texinfo-tex-region, C-b for texinfo-tex-buffer,

and so on.

C-c C-t C-r Run TEX on the region.

C-c C-t C-b  Run texi2dvi on the buffer.
C-c C-t C-i  Run texindex.

C-c C-t C-p Print the DVI file. C-c C-t C-q Show the print queue. C-c C-t C-d Delete a job from the print queue. C-c C-t C-k Kill the current TEX formatting job. C-c C-t C-x Quit a currently stopped TEX formatting job. C-c C-t C-l Recenter the output buffer.

Other Updating Commands

The remaining updating commands do not have standard keybindings because they are rarely used.

M-x texinfo-insert-node-lines

Insert missing @node lines in region. With C-u as a prefix argument, use section titles as node names.

M-x texinfo-multiple-files-update

Update a multi-file document. With C-u 2 as a prefix argument, create or update all nodes and menus in all included files first.

M-x texinfo-indent-menu-description

Indent descriptions.

M-x texinfo-sequential-node-update

Insert node pointers in strict sequence.

3 Beginning a Texinfo File

Certain pieces of information must be provided at the beginning of a Texinfo file, such as the name for the output file(s), the title of the document, and the Top node.

This chapter expands on the minimal complete Texinfo source file previously given (see Section 1.9 [Six Parts], page 10).

3.1 Sample Texinfo File Beginning

The following sample shows what is needed. The elements given here are explained in more detail in the following sections. Other commands are often included at the beginning of Texinfo files, but the ones here are the most critical.

See Section C.2 [GNU Sample Texts], page 192, for the full texts to be used in GNU manuals.

\input texinfo   @c -*-texinfo-*-
@c %**start of header
@setfilename infoname.info

@settitle name-of-manual version

@c %**end of header
@copying
This manual is for program, version version.
Copyright @copyright{} years copyright-owner.
@quotation
Permission is granted to ...
@end quotation
@end copying
@titlepage

@title name-of-manual-when-printed @subtitle subtitle-if-any @subtitle second-subtitle

@author author
@c  The following two commands
@c  start the copyright page.
@page
@vskip 0pt plus 1filll
@insertcopying
Published by ...
@end titlepage
@c So the toc is printed in the right place.
@contents
@ifnottex
@node Top

@top title

@insertcopying
@end ifnottex
@menu
* First Chapter::    Getting started ...
* Second Chapter::          ...
  ...
* Copying::          Your rights and freedoms.
@end menu
@node First Chapter
@chapter First Chapter
@cindex first chapter
@cindex chapter, first
...

3.2 Texinfo File Header

Texinfo files start with at least three lines that provide Info and TEX with necessary information. These are the \input texinfo line, the @settitle line, and the @setfilename line.

Also, if you want to format just part of the Texinfo file, you must write the @settitle and @setfilename lines between start-of-header and end-of-header lines. The start- and end-of-header lines are optional, but they do no harm, so you might as well always include them.

Any command that affects document formatting as a whole makes sense to include in the header. @synindex (see Section 12.4.2 [synindex], page 99), for instance, is another command often included in the header. See Section C.2 [GNU Sample Texts], page 192, for complete sample texts.

Thus, the beginning of a Texinfo file generally looks like this:

\input texinfo   @c -*-texinfo-*-
@c %**start of header
@setfilename sample.info
@settitle Sample Manual 1.0
@c %**end of header

3.2.1 The First Line of a Texinfo File

Every Texinfo file that is to be the top-level input to TEX must begin with a line that looks like this:

\input texinfo   @c -*-texinfo-*-

This line serves two functions:

1. When the file is processed by TEX, the ‘\input texinfo’ command tells TEX to load the macros needed for processing a Texinfo file. These are in a file called ‘texinfo.tex’, which should have been installed on your system along with either the TEX or Texinfo software. TEX uses the backslash, ‘\’, to mark the beginning of a command, exactly as Texinfo uses ‘@’. The ‘texinfo.tex’ file causes the switch from ‘\’ to ‘@’; before the switch occurs, TEX requires ‘\’, which is why it appears at the beginning of the file.

2. When the file is edited in GNU Emacs, the ‘-*-texinfo-*-’ mode specification tells Emacs to use Texinfo mode.

3.2.2 Start of Header

A start-of-header line is a Texinfo comment that looks like this:

     @c %**start of header

Write the start-of-header line on the second line of a Texinfo file. Follow the start-of-header line with @setfilename and @settitle lines and, optionally, with other commands that globally affect the document formatting, such as @synindex or @footnotestyle; and then by an end-of-header line (see Section 3.2.5 [End of Header], page 29).

The start- and end-of-header lines allow you to format only part of a Texinfo file for Info or printing. See Section 20.1.6 [texinfo-format commands], page 158.

The odd string of characters, ‘%**’, is to ensure that no other comment is accidentally taken for a start-of-header line. You can change it if you wish by setting the tex-start-of-header and/or tex-end-of-header Emacs variables. See Section 19.6 [Texinfo Mode Printing], page 145.

3.2.3 @setfilename: Set the output file name

In order to serve as the primary input file for either makeinfo or TEX, a Texinfo file must contain a line that looks like this:

@setfilename info-file-name Write the @setfilename command at the beginning of a line and follow it on the same

line by the Info file name. Do not write anything else on the line; anything on the line after the command is considered part of the file name, including what would otherwise be a comment.

The Info formatting commands ignore everything written before the @setfilename line, which is why the very first line of the file (the \input line) does not show up in the output.

The @setfilename line specifies the name of the output file to be generated. This name must be different from the name of the Texinfo file. There are two conventions for choosing the name: you can either remove the extension (such as ‘.texi’) entirely from the input file name, or, preferably, replace it with the ‘.info’ extension.

Although an explicit ‘.info’ extension is preferable, some operating systems cannot handle long file names. You can run into a problem even when the file name you specify is itself short enough. This occurs because the Info formatters split a long Info file into short indirect subfiles, and name them by appending ‘-1’, ‘-2’, …, ‘-10’, ‘-11’, and so on, to the original file name. (See Section 20.1.8 [Tag and Split Files], page 159.) The subfile name ‘texinfo.info-10’, for example, is too long for old systems with a 14-character limit on filenames; so the Info file name for this document is ‘texinfo’ rather than ‘texinfo.info’.

When makeinfo is running on operating systems such as MS-DOS which impose severe limits on file names, it may remove some characters from the original file name to leave enough space for the subfile suffix, thus producing files named ‘texin-10’, ‘gcc.i12’, etc.

When producing HTML output, makeinfo will replace any extension with ‘html’, or add ‘.html’ if the given name has no extension.

The @setfilename line produces no output when you typeset a manual with TEX, but it is nevertheless essential: it opens the index, cross-reference, and other auxiliary files used by Texinfo, and also reads ‘texinfo.cnf’ if that file is present on your system (see Section 19.9 [Preparing for TEX], page 148).

3.2.4 @settitle: Set the document title

In order to be made into a printed manual, a Texinfo file must contain a line that looks like this:

     @settitle title

Write the @settitle command at the beginning of a line and follow it on the same line by the title. This tells TEX the title to use in a header or footer. Do not write anything else on the line; anything on the line after the command is considered part of the title, including what would otherwise be a comment.

The @settitle command should precede everything that generates actual output in TEX.

In the HTML file produced by makeinfo, title also serves as the document ‘<title>’ and the default document description in the ‘<head>’ part; see Section 3.6.1 [documentdescription], page 38, for how to change that.

The title in the @settitle command does not affect the title as it appears on the title page. Thus, the two do not need not match exactly. A practice we recommend is to include the version or edition number of the manual in the @settitle title; on the title page, the version number generally appears as a @subtitle so it would be omitted from the @title. (See Section 3.4.1 [titlepage], page 31.)

Conventionally, when TEX formats a Texinfo file for double-sided output, the title is printed in the left-hand (even-numbered) page headings and the current chapter title is printed in the right-hand (odd-numbered) page headings. (TEX learns the title of each chapter from each @chapter command.) By default, no page footer is printed.

Even if you are printing in a single-sided style, TEX looks for an @settitle command line, in case you include the manual title in the heading.

TEX prints page headings only for that text that comes after the @end titlepage command in the Texinfo file, or that comes after an @headings command that turns on headings. (See Section 3.4.6 [The @headings Command], page 35, for more information.)

You may, if you wish, create your own, customized headings and footings. See Appendix E [Headings], page 200, for a detailed discussion of this.

3.2.5 End of Header

Follow the header lines with an end-of-header line, which is a Texinfo comment that looks like this:

     @c %**end of header

See Section 3.2.2 [Start of Header], page 28.

3.3 Document Permissions

The copyright notice and copying permissions for a document need to appear in several places in the various Texinfo output formats. Therefore, Texinfo provides a command (@copying) to declare this text once, and another command (@insertcopying) to insert the text at appropriate points.

3.3.1 @copying: Declare copying permissions

The @copying command should be given very early in the document; right after the header material (see Section 3.2 [Texinfo File Header], page 27) is the recommended location. It conventionally consists of a sentence or two about what the program is, the legal copyright line, and the copying permissions. Here is a skeletal example:

     @copying
     This manual is for program (version version),
     which ...
     Copyright @copyright{} years copyright-owner.
     @quotation
     Permission is granted to ...
     @end quotation
     @end copying

The @quotation has no legal significance; it’s there to improve readability in some contexts.

See Section C.2 [GNU Sample Texts], page 192, for the full text to be used in GNU manuals. See Section J.1 [GNU Free Documentation License], page 215, for the license itself under which GNU and other free manuals are distributed.

The text of @copying is output as a comment at the beginning of Info, HTML, and XML output files. It is not output implicitly in plain text or TEX; it’s up to you to use @insertcopying to emit the copying information. See the next section for details.

In output formats that support it (print and HTML), the @copyright{} command generates a ‘c’ inside a circle. In Info and plain text, it generates ‘(C)’. The copyright notice itself has the following legally defined sequence:

Copyright \begin{matrix}c \\ ◯ \end{matrix} years copyright-owner.

The word ‘Copyright’ must always be written in English, even if the manual is otherwise in another language. This is due to international law.

The list of years should include all years in which a version was completed (even if it was released in a subsequent year). Ranges are not allowed, each year must be written out individually, separated by commas.

The copyright owner (or owners) is whoever holds legal copyright on the work. In the case of works assigned to the FSF, the owner is ‘Free Software Foundation, Inc.’.

See section “Copyright Notices” in GNU Maintenance Instructions, for additional information.

3.3.2 @insertcopying: Include permissions text

The @insertcopying command is simply written on a line by itself, like this:

     @insertcopying

It inserts the text previously defined by @copying. Legally, it must be used on the copyright page in the printed manual (see Section 3.4.4 [Copyright], page 34).

Although it’s not a legal requirement, we also strongly recommend using @insertcopying in the Top node of your manual (see Section 3.5 [The Top Node], page 36). Here’s why:

The @copying command itself causes the permissions text to appear in an Info file before the first node. The text is also copied into the beginning of each split Info output file, as is legally necessary. This location implies a human reading the manual using Info does not see this text (except when using the advanced Info command g *). Therefore, an explicit @insertcopying in the Top node makes it apparent to readers that the manual is free.

Similarly, the @copying text is automatically included at the beginning of each HTML output file, as an HTML comment. Again, this text is not visible (unless the reader views the HTML source). And therefore again, the @insertcopying in the Top node is valuable because it makes the copying permissions visible and thus promotes freedom.

The permissions text defined by @copying also appears automatically at the beginning of the XML output file.

3.4 Title and Copyright Pages

In hard copy output, the manual’s name and author are usually printed on a title page. Copyright information is usually printed on the back of the title page.

The title and copyright pages appear in the printed manual, but not in the Info file. Because of this, it is possible to use several slightly obscure TEX typesetting commands that cannot be used in an Info file. In addition, this part of the beginning of a Texinfo file contains the text of the copying permissions that appears in the printed manual.

You may wish to include titlepage-like information for plain text output. Simply place any such leading material between @ifplaintext and @end ifplaintext; makeinfo includes this when writing plain text (‘--no-headers’), along with an @insertcopying.

3.4.1 @titlepage

Start the material for the title page and following copyright page with @titlepage on a line by itself and end it with @end titlepage on a line by itself.

The @end titlepage command starts a new page and turns on page numbering. (See Appendix E [Page Headings], page 200, for details about how to generate page headings.) All the material that you want to appear on unnumbered pages should be put between the @titlepage and @end titlepage commands. You can force the table of contents to appear there with the @setcontentsaftertitlepage command (see Section 4.2 [Contents], page 42).

By using the @page command you can force a page break within the region delineated by the @titlepage and @end titlepage commands and thereby create more than one unnumbered page. This is how the copyright page is produced. (The @titlepage command might perhaps have been better named the @titleandadditionalpages command, but that would have been rather long!)

When you write a manual about a computer program, you should write the version of the program to which the manual applies on the title page. If the manual changes more frequently than the program or is independent of it, you should also include an edition number^{1} for the manual. This helps readers keep track of which manual is for which version of the program. (The ‘Top’ node should also contain this information; see Section 3.5 [The Top Node], page 36.)

Texinfo provides two main methods for creating a title page. One method uses the @titlefont, @sp, and @center commands to generate a title page in which the words on the page are centered.

The second method uses the @title, @subtitle, and @author commands to create a title page with black rules under the title and author lines and the subtitle text set flush to the right hand side of the page. With this method, you do not specify any of the actual formatting of the title page. You specify the text you want, and Texinfo does the formatting.

You may use either method, or you may combine them; see the examples in the sections below.

For extremely simple applications, and for the bastard title page in traditional book front matter, Texinfo also provides a command @shorttitlepage which takes the rest of the line as the title. The argument is typeset on a page by itself and followed by a blank page.

3.4.2 @titlefont, @center, and @sp

You can use the @titlefont, @sp, and @center commands to create a title page for a printed document. (This is the first of the two methods for creating a title page in Texinfo.)

Use the @titlefont command to select a large font suitable for the title itself. You can use @titlefont more than once if you have an especially long title.

For example:

     @titlefont{Texinfo}

Use the @center command at the beginning of a line to center the remaining text on that line. Thus,

     @center @titlefont{Texinfo}

centers the title, which in this example is “Texinfo” printed in the title font.

Use the @sp command to insert vertical space. For example:

1 We have found that it is helpful to refer to versions of independent manuals as ‘editions’ and versions

of programs as ‘versions’; otherwise, we find we are liable to confuse each other in conversation by

referring to both the documentation and the software with the same words.

     @sp 2

This inserts two blank lines on the printed page. (See Section 14.4 [@sp], page 114, for more information about the @sp command.)

A template for this method looks like this:

     @titlepage
     @sp 10

@center @titlefont{name-of-manual-when-printed}

     @sp 2

@center subtitle-if-any

     @sp 2
     @center author
     ...
     @end titlepage

The spacing of the example fits an 8.5 by 11 inch manual.

3.4.3 @title, @subtitle, and @author

You can use the @title, @subtitle, and @author commands to create a title page in which the vertical and horizontal spacing is done for you automatically. This contrasts with the method described in the previous section, in which the @sp command is needed to adjust vertical spacing.

Write the @title, @subtitle, or @author commands at the beginning of a line followed by the title, subtitle, or author.

The @title command produces a line in which the title is set flush to the left-hand side of the page in a larger than normal font. The title is underlined with a black rule. Only a single line is allowed; the @* command may not be used to break the title into two lines. To handle very long titles, you may find it profitable to use both @title and @titlefont; see the final example in this section.

The @subtitle command sets subtitles in a normal-sized font flush to the right-hand side of the page.

The @author command sets the names of the author or authors in a middle-sized font flush to the left-hand side of the page on a line near the bottom of the title page. The names are underlined with a black rule that is thinner than the rule that underlines the title. (The black rule only occurs if the @author command line is followed by an @page command line.)

There are two ways to use the @author command: you can write the name or names on the remaining part of the line that starts with an @author command:

     @author by Jane Smith and John Doe

or you can write the names one above each other by using two (or more) @author commands:

     @author Jane Smith
     @author John Doe

(Only the bottom name is underlined with a black rule.)

A template for this method looks like this:

@titlepage

@title name-of-manual-when-printed @subtitle subtitle-if-any @subtitle second-subtitle

@author author
@page
...
@end titlepage

You may also combine the @titlefont method described in the previous section and @title method described in this one. This may be useful if you have a very long title. Here is a real-life example:

@titlepage
@titlefont{GNU Software}
@sp 1
@title for MS-Windows and MS-DOS
@subtitle Edition @value{e} for Release @value{cde}
@author by Daniel Hagerty, Melissa Weisshaus
@author and Eli Zaretskii

(The use of @value here is explained in Section 16.4.3 [value Example], page 133.

3.4.4 Copyright Page

By international treaty, the copyright notice for a book must be either on the title page or on the back of the title page. When the copyright notice is on the back of the title page, that page is customarily not numbered. Therefore, in Texinfo, the information on the copyright page should be within @titlepage and @end titlepage commands.

Use the @page command to cause a page break. To push the copyright notice and the other text on the copyright page towards the bottom of the page, use the following incantantion after @page:

@vskip 0pt plus 1filll

This is a TEX command that is not supported by the Info formatting commands. The @vskip command inserts whitespace. The ‘0pt plus 1filll’ means to put in zero points of mandatory whitespace, and as much optional whitespace as needed to push the following text to the bottom of the page. Note the use of three ‘l’s in the word ‘filll’; this is correct.

To insert the copyright text itself, write @insertcopying next (see Section 3.3 [Document Permissions], page 30):

@insertcopying

Follow the copying text by the publisher, ISBN numbers, cover art credits, and other such information.

Here is an example putting all this together:

@titlepage
...
@page
@vskip 0pt plus 1filll

     @insertcopying
     Published by ...
     Cover art by ...
     @end titlepage

3.4.5 Heading Generation

The @end titlepage command must be written on a line by itself. It not only marks the end of the title and copyright pages, but also causes TEX to start generating page headings and page numbers.

To repeat what is said elsewhere, Texinfo has two standard page heading formats, one for documents which are printed on one side of each sheet of paper (single-sided printing), and the other for documents which are printed on both sides of each sheet (double-sided printing). You can specify these formats in different ways:

• The conventional way is to write an @setchapternewpage command before the title page commands, and then have the @end titlepage command start generating page headings in the manner desired. (See Section 3.6.2 [setchapternewpage], page 38.)

• Alternatively, you can use the @headings command to prevent page headings from being generated or to start them for either single or double-sided printing. (Write an @headings command immediately after the @end titlepage command. See Section 3.4.6 [The @headings Command], page 35, for more information.)

• Or, you may specify your own page heading and footing format. See Appendix E [Page Headings], page 200, for detailed information about page headings and footings.

Most documents are formatted with the standard single-sided or double-sided format,

using @setchapternewpage odd for double-sided printing and no @setchapternewpage

command for single-sided printing.

3.4.6 The @headings Command

The @headings command is rarely used. It specifies what kind of page headings and footings to print on each page. Usually, this is controlled by the @setchapternewpage command. You need the @headings command only if the @setchapternewpage command does not do what you want, or if you want to turn off pre-defined page headings prior to defining your own. Write an @headings command immediately after the @end titlepage command.

You can use @headings as follows:

@headings off

Turn off printing of page headings.

@headings single

Turn on page headings appropriate for single-sided printing.

@headings double
@headings on

Turn on page headings appropriate for double-sided printing. The two com-

          mands, @headings on and @headings double, are synonymous.

@headings singleafter
@headings doubleafter

Turn on single or double headings, respectively, after the current page is output.

@headings on

Turn on page headings: single if ‘@setchapternewpage on’, double otherwise.

For example, suppose you write @setchapternewpage off before the @titlepage command to tell TEX to start a new chapter on the same page as the end of the last chapter. This command also causes TEX to typeset page headers for single-sided printing. To cause TEX to typeset for double sided printing, write @headings double after the @end titlepage command.

You can stop TEX from generating any page headings at all by writing @headings off on a line of its own immediately after the line containing the @end titlepage command, like this:

     @end titlepage
     @headings off

The @headings off command overrides the @end titlepage command, which would otherwise cause TEX to print page headings.

You can also specify your own style of page heading and footing. See Appendix E [Page Headings], page 200, for more information.

3.5 The ‘Top’ Node and Master Menu

The ‘Top’ node is the node in which a reader enters an Info manual. As such, it should begin with the @insertcopying command (see Section 3.3 [Document Permissions], page 30) to provide a brief description of the manual (including the version number) and copying permissions, and end with a master menu for the whole manual. Of course you should include any other general information you feel a reader would find helpful.

It is also conventional to write an @top sectioning command line containing the title of the document immediately after the @node Top line (see Section 6.3.6 [The @top Sectioning Command], page 55).

The contents of the ‘Top’ node should appear only in the online output; none of it should appear in printed output, so enclose it between @ifnottex and @end ifnottex commands. (TEX does not print either an @node line or a menu; they appear only in Info; strictly speaking, you are not required to enclose these parts between @ifnottex and @end ifnottext, but it is simplest to do so. See Chapter 16 [Conditionally Visible Text], page 129.)

3.5.1 Top Node Example

Here is an example of a Top node.

     @ifnottex
     @node Top
     @top Sample Title
     @insertcopying

Additional general information.
@menu
* First Chapter::
* Second Chapter::
...
* Index::
@end menu

3.5.2 Parts of a Master Menu

A master menu is a detailed main menu listing all the nodes in a file.

A master menu is enclosed in @menu and @end menu commands and does not appear in the printed document.

Generally, a master menu is divided into parts.

• The first part contains the major nodes in the Texinfo file: the nodes for the chapters, chapter-like sections, and the appendices.

• The second part contains nodes for the indices.

• The third and subsequent parts contain a listing of the other, lower level nodes, often ordered by chapter. This way, rather than go through an intermediary menu, an inquirer can go directly to a particular node when searching for specific information. These menu items are not required; add them if you think they are a convenience. If you do use them, put @detailmenu before the first one, and @end detailmenu after the last; otherwise, makeinfo will get confused.

Each section in the menu can be introduced by a descriptive line. So long as the line does not begin with an asterisk, it will not be treated as a menu entry. (See Section 7.1 [Writing a Menu], page 57, for more information.)

For example, the master menu for this manual looks like the following (but has many more entries):

@menu
* Copying Conditions::  Your rights.
* Overview::            Texinfo in brief.
...
* Command and Variable Index::
* Concept Index::
@detailmenu
 --- The Detailed Node Listing ---
Overview of Texinfo
* Reporting Bugs:: ...
...

     Beginning a Texinfo File
     * Sample Beginning:: ...
     ...
     @end detailmenu
     @end menu

3.6 Global Document Commands

Besides the basic commands mentioned in the previous sections, here are additional commands which affect the document as a whole. They are generally all given before the Top node, if they are given at all.

3.6.1 @documentdescription: Summary text

When producing HTML output for a document, makeinfo writes a ‘<meta>’ element in the ‘<head>’ to give some idea of the content of the document. By default, this description is the title of the document, taken from the @settitle command (see Section 3.2.4 [settitle], page 29). To change this, use the @documentdescription environment, as in:

     @documentdescription
     descriptive text.
     @end documentdescription

This will produce the following output in the ‘<head>’ of the HTML:

     <meta name=description content="descriptive text.">

@documentdescription must be specified before the first node of the document.

3.6.2 @setchapternewpage:

In an officially bound book, text is usually printed on both sides of the paper, chapters start on right-hand pages, and right-hand pages have odd numbers. But in short reports, text often is printed only on one side of the paper. Also in short reports, chapters sometimes do not start on new pages, but are printed on the same page as the end of the preceding chapter, after a small amount of vertical whitespace.

You can use the @setchapternewpage command with various arguments to specify how TEX should start chapters and whether it should format headers for printing on one or both sides of the paper (single-sided or double-sided printing).

Write the @setchapternewpage command at the beginning of a line followed by its argument.

For example, you would write the following to cause each chapter to start on a fresh odd-numbered page:

     @setchapternewpage odd

You can specify one of three alternatives with the @setchapternewpage command:

@setchapternewpage off

Cause TEX to typeset a new chapter on the same page as the last chapter, after skipping some vertical whitespace. Also, cause TEX to format page headers for single-sided printing.

@setchapternewpage on

Cause TEX to start new chapters on new pages and to format page headers for single-sided printing. This is the form most often used for short reports or personal printing. This is the default.

@setchapternewpage odd

Cause TEX to start new chapters on new, odd-numbered pages (right-handed pages) and to typeset for double-sided printing. This is the form most often used for books and manuals.

Texinfo does not have an @setchapternewpage even command, because there is no printing tradition of starting chapters or books on an even-numbered page.

If you don’t like the default headers that @setchapternewpage sets, you can explicit control them with the @headings command. See Section 3.4.6 [The @headings Command], page 35.

At the beginning of a manual or book, pages are not numbered—for example, the title and copyright pages of a book are not numbered. By convention, table of contents and frontmatter pages are numbered with roman numerals and not in sequence with the rest of the document.

Since an Info file does not have pages, the @setchapternewpage command has no effect on it.

We recommend not including any @setchapternewpage command in your manual sources at all, since the desired output is not intrinsic to the document. For a particular hard copy run, if you don’t want the default option (no blank pages, same headers on all pages) use the ‘--texinfo’ option to texi2dvi to specify the output you want.

3.6.3 Paragraph Indenting

The Texinfo processors may insert whitespace at the beginning of the first line of each paragraph, thereby indenting that paragraph. You can use the @paragraphindent command to specify this indentation. Write an @paragraphindent command at the beginning of a line followed by either ‘asis’ or a number:

     @paragraphindent indent

The indentation is according to the value of indent:

asis Do not change the existing indentation (not implemented in TEX).

none

0 Omit all indentation.

n Indent by n space characters in Info output, by n ems in TEX.

The default value of indent is 3. @paragraphindent is ignored for HTML output. It is best to write the @paragraphindent command before the end-of-header line at

the beginning of a Texinfo file, so the region formatting commands indent paragraphs as specified. See Section 3.2.2 [Start of Header], page 28.

    A peculiarity of the texinfo-format-buffer and texinfo-format-region commands

is that they do not indent (nor fill) paragraphs that contain @w or @* commands. See Appendix G [Refilling Paragraphs], page 212, for further information.

3.6.4 @exampleindent: Environment Indenting

The Texinfo processors indent each line of @example and similar environments. You can use the @exampleindent command to specify this indentation. Write an @exampleindent command at the beginning of a line followed by either ‘asis’ or a number:

     @exampleindent indent

The indentation is according to the value of indent:

asis Do not change the existing indentation (not implemented in TEX).

0 Omit all indentation.

n Indent environments by n space characters in Info output, by n ems in TEX.

The default value of indent is 5. @exampleindent is ignored for HTML output.

It is best to write the @exampleindent command before the end-of-header line at the beginning of a Texinfo file, so the region formatting commands indent paragraphs as specified. See Section 3.2.2 [Start of Header], page 28.

3.7 Software Copying Permissions

If the Texinfo file has a section containing the “General Public License” and the distribution information and a warranty disclaimer for the software that is documented, we recommend placing this right after the ‘Top’ node. The General Public License is very important to Project GNU software. It ensures that you and others will continue to have a right to use and share the software.

The copying and distribution information and the disclaimer are followed by an introduction or else by the first chapter of the manual.

Although an introduction is not a required part of a Texinfo file, it is very helpful. Ideally, it should state clearly and concisely what the file is about and who would be interested in reading it. In general, an introduction would follow the licensing and distribution information, although sometimes people put it earlier in the document.

4 Ending a Texinfo File

The end of a Texinfo file should include commands to create indices and (perhaps) to generate both the full and summary tables of contents. Finally, it must include the @bye command that marks the last line to be processed.

For example:

     @node Index
     @unnumbered Index
     @printindex cp
     @shortcontents
     @contents
     @bye

4.1 Printing Indices and Menus

To print an index means to include it as part of a manual or Info file. This does not happen automatically just because you use @cindex or other index-entry generating commands in the Texinfo file; those just cause the raw data for the index to be accumulated. To generate an index, you must include the @printindex command at the place in the document where you want the index to appear. Also, as part of the process of creating a printed manual, you must run a program called texindex (see Chapter 19 [Hardcopy], page 142) to sort the raw data to produce a sorted index file. The sorted index file is what is actually used to print the index.

Texinfo offers six separate types of predefined index, each with a two-letter abbreviation, as illustrated in the following table. However, you may merge indices (see Section 12.4 [Combining Indices], page 98) or define your own indices (see Section 12.5 [New Indices], page 99).

Here are the predefined indices, their abbreviations, and the corresponding index entry commands:

‘cp’ concept index (@cindex)

‘fn’ function index (@findex)

‘vr’ variable index (@index)

‘ky’ key index (@kindex)

‘pg’ program index (@pindex)

‘tp’ data type index (@tindex)

The @printindex command takes a two-letter index abbreviation, reads the corresponding sorted index file and formats it appropriately into an index.

The @printindex command does not generate a chapter heading for the index. Consequently, you should precede the @printindex command with a suitable section or chapter command (usually @appendix or @unnumbered) to supply the chapter heading and put the index into the table of contents. Precede the @unnumbered command with an @node line.

For example:

     @node Variable Index
     @unnumbered Variable Index
     @printindex vr
     @node Concept Index
     @unnumbered Concept Index
     @printindex cp

We recommend placing the concept index last, since that makes it easiest to find. We also recommend having a single index whenever possible, since then readers have only one place to look (see Section 12.4 [Combining Indices], page 98).

4.2 Generating a Table of Contents

The @chapter, @section, and other structuring commands supply the information to make up a table of contents, but they do not cause an actual table to appear in the manual. To do this, you must use the @contents and/or @summarycontents command(s).

@contents

Generate a table of contents in a printed manual, including all chapters, sections, subsections, etc., as well as appendices and unnumbered chapters. Headings generated by the @heading series of commands do not appear in the table of contents.

@shortcontents
@summarycontents
          (@summarycontents is a synonym for @shortcontents.)

Generate a short or summary table of contents that lists only the chapters, appendices, and unnumbered chapters. Sections, subsections and subsubsections are omitted. Only a long manual needs a short table of contents in addition to the full table of contents.

Both contents commands should be written on a line by themselves. The contents commands automatically generate a chapter-like heading at the top of the first table of contents page, so don’t include any sectioning command such as @unnumbered before them.

Since an Info file uses menus instead of tables of contents, the Info formatting commands ignore the contents commands. But the contents are included in plain text output (generated by makeinfo --no-headers), unless makeinfo is writing its output to standard output.

When makeinfo writes a short table of contents while producing html output, the links in the short table of contents point to corresponding entries in the full table of contents rather than the text of the document. The links in the full table of contents point to the main text of the document.

The contents commands can be placed either at the very end of the file, after any indices (see the previous section) and just before the @bye (see the next section), or near the beginning of the file, after the @end titlepage (see Section 3.4.1 [titlepage], page 31). The advantage to the former is that then the contents output is always up to date, because it reflects the processing just done. The advantage to the latter is that the contents are printed in the proper place, thus you do not need to rearrange the DVI file with dviselect or shuffle paper.

As an author, you can put the contents commands wherever you prefer. But if you are a user simply printing a manual, you may wish to print the contents after the title page even if the author put the contents commands at the end of the document (as is the case in most existing Texinfo documents, at this writing). You can do this by spec-

ifying @setcontentsaftertitlepage and/or @setshortcontentsaftertitlepage. The

first prints only the main contents after the @end titlepage; the second prints both the short contents and the main contents. In either case, any subsequent @contents or @shortcontents is ignored (unless no @end titlepage is ever encountered).

You need to include the @set...contentsaftertitlepage commands early in the document (just after @setfilename, for example). We recommend using texi2dvi (see Section 19.3 [Format with texi2dvi], page 144) to specify this without altering the source file at all. For example:

     texi2dvi --texinfo=@setcontentsaftertitlepage foo.texi

4.3 @bye File Ending

An @bye command terminates TEX or Info formatting. None of the formatting commands reading anything following @bye. The @bye command should be on a line by itself.

If you wish, you may follow the @bye line with notes. These notes will not be formatted and will not appear in either Info or a printed manual; it is as if text after @bye were within @ignore … @end ignore. Also, you may follow the @bye line with a local variables list for Emacs. See Section 19.7 [Using Local Variables and the Compile Command], page 147, for more information.

5 Chapter Structuring

The chapter structuring commands divide a document into a hierarchy of chapters, sections, subsections, and subsubsections. These commands generate large headings; they also provide information for the table of contents of a printed manual (see Section 4.2 [Generating a Table of Contents], page 42).

The chapter structuring commands do not create an Info node structure, so normally you should put an @node command immediately before each chapter structuring command (see Chapter 6 [Nodes], page 50). The only time you are likely to use the chapter structuring commands without using the node structuring commands is if you are writing a document that contains no cross references and will never be transformed into Info format.

It is unlikely that you will ever write a Texinfo file that is intended only as an Info file and not as a printable document. If you do, you might still use chapter structuring commands to create a heading at the top of each node—but you don’t need to.

5.1 Tree Structure of Sections

A Texinfo file is usually structured like a book with chapters, sections, subsections, and the like. This structure can be visualized as a tree (or rather as an upside-down tree) with the root at the top and the levels corresponding to chapters, sections, subsection, and subsubsections.

Here is a diagram that shows a Texinfo file with three chapters, each of which has two sections.

                               Top
                                |
              -------------------------------------
             |                  |                  |
          Chapter 1          Chapter 2          Chapter 3
             |                  |                  |
          --------           --------           --------
         |        |         |        |         |        |
      Section  Section   Section  Section   Section  Section
        1.1      1.2       2.1      2.2       3.1      3.2

In a Texinfo file that has this structure, the beginning of Chapter 2 looks like this:

     @node    Chapter 2,  Chapter 3, Chapter 1, top
     @chapter Chapter 2

The chapter structuring commands are described in the sections that follow; the @node and @menu commands are described in following chapters. (See Chapter 6 [Nodes], page 50, and see Chapter 7 [Menus], page 57.)

5.2 Structuring Command Types

The chapter structuring commands fall into four groups or series, each of which contains structuring commands corresponding to the hierarchical levels of chapters, sections, subsections, and subsubsections.

The four groups are the @chapter series, the @unnumbered series, the @appendix series, and the @heading series.

Each command produces titles that have a different appearance on the printed page or Info file; only some of the commands produce titles that are listed in the table of contents of a printed book or manual.

• The @chapter and @appendix series of commands produce numbered or lettered entries both in the body of a printed work and in its table of contents.

• The @unnumbered series of commands produce unnumbered entries both in the body of a printed work and in its table of contents. The @top command, which has a special use, is a member of this series (see Section 5.3 [@top], page 45).

• The @heading series of commands produce unnumbered headings that do not appear in a table of contents. The heading commands never start a new page.

• The @majorheading command produces results similar to using the @chapheading command but generates a larger vertical whitespace before the heading.

  • When an @setchapternewpage command says to do so, the @chapter, @unnumbered,

and @appendix commands start new pages in the printed manual; the @heading commands do not.

Here are the four groups of chapter structuring commands:

No new page

Numbered  Unnumbered  Lettered/numbered  Unnumbered

In contents In contents In contents Omitted from contents

                    @top                                                  @majorheading
@chapter            @unnumbered                 @appendix                  @chapheading
@section            @unnumberedsec              @appendixsec               @heading
@subsection         @unnumberedsubsec           @appendixsubsec            @subheading
@subsubsection      @unnumberedsubsubsec        @appendixsubsubsec         @subsubheading

5.3 @top

The @top command is a special sectioning command that you use only after an ‘@node Top’ line at the beginning of a Texinfo file. The @top command tells the makeinfo formatter which node is the ‘Top’ node, so it can use it as the root of the node tree if your manual uses implicit pointers. It has the same typesetting effect as @unnumbered (see Section 5.5 [@unnumbered and @appendix], page 46). For detailed information, see Section 6.3.6 [The @top Command], page 55.

The @top node and its menu (if any) is conventionally wrapped in an @ifnottex conditional so that it will appear only in Info and HTML output, not TEX.

5.4 @chapter

@chapter identifies a chapter in the document. Write the command at the beginning of a line and follow it on the same line by the title of the chapter.

For example, this chapter in this manual is entitled “Chapter Structuring”; the @chapter line looks like this:

     @chapter Chapter Structuring

In TEX, the @chapter command creates a chapter in the document, specifying the chapter title. The chapter is numbered automatically.

In Info, the @chapter command causes the title to appear on a line by itself, with a line of asterisks inserted underneath. Thus, in Info, the above example produces the following output:

     Chapter Structuring
     *******************

Texinfo also provides a command @centerchap, which is analogous to @unnumbered, but centers its argument in the printed output. This kind of stylistic choice is not usually offered by Texinfo.

5.5 @unnumbered and @appendix

Use the @unnumbered command to create a chapter that appears in a printed manual without chapter numbers of any kind. Use the @appendix command to create an appendix in a printed manual that is labelled by letter instead of by number.

For Info file output, the @unnumbered and @appendix commands are equivalent to @chapter: the title is printed on a line by itself with a line of asterisks underneath. (See Section 5.4 [@chapter], page 45.)

To create an appendix or an unnumbered chapter, write an @appendix or @unnumbered command at the beginning of a line and follow it on the same line by the title, as you would if you were creating a chapter.

5.6 @majorheading, @chapheading

The @majorheading and @chapheading commands put chapter-like headings in the body of a document.

However, neither command causes TEX to produce a numbered heading or an entry in the table of contents; and neither command causes TEX to start a new page in a printed manual.

In TEX, an @majorheading command generates a larger vertical whitespace before the heading than an @chapheading command but is otherwise the same.

In Info, the @majorheading and @chapheading commands are equivalent to @chapter: the title is printed on a line by itself with a line of asterisks underneath. (See Section 5.4 [@chapter], page 45.)

5.7 @section

In a printed manual, an @section command identifies a numbered section within a chapter. The section title appears in the table of contents. In Info, an @section command provides a title for a segment of text, underlined with ‘=’.

This section is headed with an @section command and looks like this in the Texinfo file:

     @section @code{@@section}

To create a section, write the @section command at the beginning of a line and follow it on the same line by the section title.

Thus,

     @section This is a section

produces

     This is a section
     =================

in Info.

5.8 @unnumberedsec, @appendixsec, @heading

The @unnumberedsec, @appendixsec, and @heading commands are, respectively, the unnumbered, appendix-like, and heading-like equivalents of the @section command. (See Section 5.7 [@section], page 46.)

@unnumberedsec

The @unnumberedsec command may be used within an unnumbered chapter or within a regular chapter or appendix to provide an unnumbered section.

@appendixsec
@appendixsection

@appendixsection is a longer spelling of the @appendixsec command; the two are synonymous.

Conventionally, the @appendixsec or @appendixsection command is used only within appendices.

@heading You may use the @heading command anywhere you wish for a section-style heading that will not appear in the table of contents.

5.9 The @subsection Command

Subsections are to sections as sections are to chapters. (See Section 5.7 [@section], page 46.) In Info, subsection titles are underlined with ‘-’. For example,

     @subsection This is a subsection

produces

     This is a subsection
     --------------------

In a printed manual, subsections are listed in the table of contents and are numbered three levels deep.

5.10 The @subsection-like Commands

    The @unnumberedsubsec, @appendixsubsec, and @subheading commands are, respec-

tively, the unnumbered, appendix-like, and heading-like equivalents of the @subsection command. (See Section 5.9 [@subsection], page 47.)

In Info, the @subsection-like commands generate a title underlined with hyphens. In a printed manual, an @subheading command produces a heading like that of a subsection except that it is not numbered and does not appear in the table of contents. Similarly, an @unnumberedsubsec command produces an unnumbered heading like that of a subsection and an @appendixsubsec command produces a subsection-like heading labelled with a letter and numbers; both of these commands produce headings that appear in the table of contents.

5.11 The ‘subsub’ Commands

The fourth and lowest level sectioning commands in Texinfo are the ‘subsub’ commands. They are:

@subsubsection

Subsubsections are to subsections as subsections are to sections. (See Section 5.9 [@subsection], page 47.) In a printed manual, subsubsection titles appear in the table of contents and are numbered four levels deep.

@unnumberedsubsubsec

Unnumbered subsubsection titles appear in the table of contents of a printed manual, but lack numbers. Otherwise, unnumbered subsubsections are the same as subsubsections. In Info, unnumbered subsubsections look exactly like ordinary subsubsections.

@appendixsubsubsec

Conventionally, appendix commands are used only for appendices and are lettered and numbered appropriately in a printed manual. They also appear in the table of contents. In Info, appendix subsubsections look exactly like ordinary subsubsections.

@subsubheading

The @subsubheading command may be used anywhere that you need a small heading that will not appear in the table of contents. In Info, subsubheadings look exactly like ordinary subsubsection headings.

In Info, ‘subsub’ titles are underlined with periods. For example,

     @subsubsection This is a subsubsection

produces

     This is a subsubsection
     .......................

5.12 @raisesections and @lowersections

The @raisesections and @lowersections commands raise and lower the hierarchical level of chapters, sections, subsections and the like. The @raisesections command changes sections to chapters, subsections to sections, and so on. The @lowersections command changes chapters to sections, sections to subsections, and so on.

An @lowersections command is useful if you wish to include text that is written as an outer or standalone Texinfo file in another Texinfo file as an inner, included file. If you write the command at the beginning of the file, all your @chapter commands are formatted as if they were @section commands, all your @section command are formatted as if they were @subsection commands, and so on.

@raisesections raises a command one level in the chapter structuring hierarchy: Change To

     @subsection     @section,
     @section        @chapter,
     @heading        @chapheading,

etc.

@lowersections lowers a command one level in the chapter structuring hierarchy: Change To

     @chapter        @section,
     @subsection     @subsubsection,
     @heading        @subheading,

etc.

An @raisesections or @lowersections command changes only those structuring commands that follow the command in the Texinfo file. Write an @raisesections or @lowersections command on a line of its own.

An @lowersections command cancels an @raisesections command, and vice versa. Typically, the commands are used like this:

     @lowersections
     @include somefile.texi
     @raisesections

Without the @raisesections, all the subsequent sections in your document will be lowered.

Repeated use of the commands continue to raise or lower the hierarchical level a step at a time.

An attempt to raise above ‘chapters’ reproduces chapter commands; an attempt to lower below ‘subsubsections’ reproduces subsubsection commands.

6 Nodes

Nodes are the primary segments of a Texinfo file. They do not themselves impose a hierarchical or any other kind of structure on a file. Nodes contain node pointers that name other nodes, and can contain menus which are lists of nodes. In Info, the movement commands can carry you to a pointed-to node or to a node listed in a menu. Node pointers and menus provide structure for Info files just as chapters, sections, subsections, and the like, provide structure for printed books.

6.1 Two Paths

The node and menu commands and the chapter structuring commands are technically independent of each other:

• In Info, node and menu commands provide structure. The chapter structuring commands generate headings with different kinds of underlining—asterisks for chapters, hyphens for sections, and so on; they do nothing else.

• In TEX, the chapter structuring commands generate chapter and section numbers and tables of contents. The node and menu commands provide information for cross references; they do nothing else.

You can use node pointers and menus to structure an Info file any way you want; and you can write a Texinfo file so that its Info output has a different structure than its printed output. However, virtually all Texinfo files are written such that the structure for the Info output corresponds to the structure for the printed output. It is neither convenient nor understandable to the reader to do otherwise.

Generally, printed output is structured in a tree-like hierarchy in which the chapters are the major limbs from which the sections branch out. Similarly, node pointers and menus are organized to create a matching structure in the Info output.

6.2 Node and Menu Illustration

Here is a copy of the diagram shown earlier that illustrates a Texinfo file with three chapters, each of which contains two sections.

The “root” is at the top of the diagram and the “leaves” are at the bottom. This is how such a diagram is drawn conventionally; it illustrates an upside-down tree. For this reason, the root node is called the ‘Top’ node, and ‘Up’ node pointers carry you closer to the root.

                               Top
                                |
              -------------------------------------
             |                  |                  |
          Chapter 1          Chapter 2          Chapter 3
             |                  |                  |
          --------           --------           --------
         |        |         |        |         |        |
      Section  Section   Section  Section   Section  Section
        1.1      1.2       2.1      2.2       3.1      3.2

The fully-written command to start Chapter 2 would be this:

     @node     Chapter 2,  Chapter 3, Chapter 1, Top
     @comment  node-name,  next,      previous,  up

This @node line says that the name of this node is “Chapter 2”, the name of the ‘Next’ node is “Chapter 3”, the name of the ‘Previous’ node is “Chapter 1”, and the name of the ‘Up’ node is “Top”. You can omit writing out these node names if your document is hierarchically organized (see Section 6.4 [makeinfo Pointer Creation], page 55), but the pointer relationships still obtain.

Please Note: ‘Next’ refers to the next node at the same hierarchical level in the manual, not necessarily to the next node within the Texinfo file. In the Texinfo file, the subsequent node may be at a lower level—a section-level node most often follows a chapter-level node, for example. ‘Next’ and ‘Previous’ refer to nodes at the same hierarchical level. (The ‘Top’ node contains the exception to this rule. Since the ‘Top’ node is the only node at that level, ‘Next’ refers to the first following node, which is almost always a chapter or chapter-level node.)

To go to Sections 2.1 and 2.2 using Info, you need a menu inside Chapter 2. (See Chapter 7 [Menus], page 57.) You would write the menu just before the beginning of Section 2.1, like this:

         @menu
         * Sect. 2.1::    Description of this section.
         * Sect. 2.2::
         @end menu

Write the node for Sect. 2.1 like this:

         @node     Sect. 2.1, Sect. 2.2, Chapter 2, Chapter 2
         @comment  node-name, next,      previous,  up

In Info format, the ‘Next’ and ‘Previous’ pointers of a node usually lead to other nodes at the same level—from chapter to chapter or from section to section (sometimes, as shown, the ‘Previous’ pointer points up); an ‘Up’ pointer usually leads to a node at the level above (closer to the ‘Top’ node); and a ‘Menu’ leads to nodes at a level below (closer to ‘leaves’). (A cross reference can point to a node at any level; see Chapter 8 [Cross References], page 61.)

Usually, an @node command and a chapter structuring command are used in sequence, along with indexing commands. (You may follow the @node line with a comment line that reminds you which pointer is which.)

Here is the beginning of the chapter in this manual called “Ending a Texinfo File”. This shows an @node line followed by a comment line, an @chapter line, and then by indexing lines.

     @node    Ending a File, Structuring, Beginning a File, Top
     @comment node-name,     next,        previous,         up
     @chapter Ending a Texinfo File
     @cindex Ending a Texinfo file
     @cindex Texinfo file ending
     @cindex File ending

6.3 The @node Command

A node is a segment of text that begins at an @node command and continues until the next @node command. The definition of node is different from that for chapter or section. A chapter may contain sections and a section may contain subsections; but a node cannot contain subnodes; the text of a node continues only until the next @node command in the file. A node usually contains only one chapter structuring command, the one that follows the @node line. On the other hand, in printed output nodes are used only for cross references, so a chapter or section may contain any number of nodes. Indeed, a chapter usually contains several nodes, one for each section, subsection, and subsubsection.

To create a node, write an @node command at the beginning of a line, and follow it with up to four arguments, separated by commas, on the rest of the same line. The first argument is required; it is the name of this node. The subsequent arguments are the names of the ‘Next’, ‘Previous’, and ‘Up’ pointers, in that order, and may be omitted if your Texinfo document is hierarchically organized (see Section 6.4 [makeinfo Pointer Creation], page 55).

You may insert spaces before each name if you wish; the spaces are ignored. You must write the name of the node and the names of the ‘Next’, ‘Previous’, and ‘Up’ pointers all on the same line. Otherwise, the formatters fail. (See Info file ‘info’, node ‘Top’, for more information about nodes in Info.)

Usually, you write one of the chapter-structuring command lines immediately after an @node line—for example, an @section or @subsection line. (See Section 5.2 [Structuring Command Types], page 44.)

Please note: The GNU Emacs Texinfo mode updating commands work only with Texinfo files in which @node lines are followed by chapter structuring lines. See Section 2.4.1 [Updating Requirements], page 20. TEX uses @node lines to identify the names to use for cross references. For this reason,

you must write @node lines in a Texinfo file that you intend to format for printing, even if you do not intend to format it for Info. (Cross references, such as the one at the end of this sentence, are made with @xref and related commands; see Chapter 8 [Cross References], page 61.)

6.3.1 Choosing Node and Pointer Names

The name of a node identifies the node. The pointers enable you to reach other nodes and consist of the names of those nodes.

Normally, a node’s ‘Up’ pointer contains the name of the node whose menu mentions that node. The node’s ‘Next’ pointer contains the name of the node that follows that node in that menu and its ‘Previous’ pointer contains the name of the node that precedes it in that menu. When a node’s ‘Previous’ node is the same as its ‘Up’ node, both node pointers name the same node.

Usually, the first node of a Texinfo file is the ‘Top’ node, and its ‘Up’ and ‘Previous’ pointers point to the ‘dir’ file, which contains the main menu for all of Info.

The ‘Top’ node itself contains the main or master menu for the manual. Also, it is helpful to include a brief description of the manual in the ‘Top’ node. See Section 6.3.5 [First Node], page 54, for information on how to write the first node of a Texinfo file.

Even when you explicitly specify all pointers, that does not mean you can write the nodes in the Texinfo source file in an arbitrary order! Because TEX processes the file sequentially, irrespective of node pointers, you must write the nodes in the order you wish them to appear in the printed output.

6.3.2 How to Write an @node Line

The easiest way to write an @node line is to write @node at the beginning of a line and then the name of the node, like this:

@node node-name

If you are using GNU Emacs, you can use the update node commands provided by Texinfo mode to insert the names of the pointers; or you can leave the pointers out of the Texinfo file and let makeinfo insert node pointers into the Info file it creates. (See Chapter 2 [Texinfo Mode], page 14, and Section 6.4 [makeinfo Pointer Creation], page 55.)

Alternatively, you can insert the ‘Next’, ‘Previous’, and ‘Up’ pointers yourself. If you do this, you may find it helpful to use the Texinfo mode keyboard command C-c C-c n. This command inserts ‘@node’ and a comment line listing the names of the pointers in their proper order. The comment line helps you keep track of which arguments are for which pointers. This comment line is especially useful if you are not familiar with Texinfo.

The template for a fully-written-out node line with ‘Next’, ‘Previous’, and ‘Up’ pointers looks like this:

@node node-name, next, previous, up If you wish, you can ignore @node lines altogether in your first draft and then use the

texinfo-insert-node-lines command to create @node lines for you. However, we do not recommend this practice. It is better to name the node itself at the same time that you write a segment so you can easily make cross references. A large number of cross references are an especially important feature of a good Info file.

After you have inserted an @node line, you should immediately write an @-command for the chapter or section and insert its name. Next (and this is important!), put in several index entries. Usually, you will find at least two and often as many as four or five ways of referring to the node in the index. Use them all. This will make it much easier for people to find the node.

6.3.3 @node Line Tips

Here are three suggestions:

• Try to pick node names that are informative but short.

In the Info file, the file name, node name, and pointer names are all inserted on one line, which may run into the right edge of the window. (This does not cause a problem with Info, but is ugly.)

• Try to pick node names that differ from each other near the beginnings of their names. This way, it is easy to use automatic name completion in Info.

• By convention, node names are capitalized just as they would be for section or chapter titles—initial and significant words are capitalized; others are not.

6.3.4 @node Line Requirements

Here are several requirements for @node lines:

• All the node names for a single Info file must be unique.

Duplicates confuse the Info movement commands. This means, for example, that if you end every chapter with a summary, you must name each summary node differently. You cannot just call each one “Summary”. You may, however, duplicate the titles of chapters, sections, and the like. Thus you can end each chapter in a book with a section called “Summary”, so long as the node names for those sections are all different.

• A pointer name must be the name of a node.

The node to which a pointer points may come before or after the node containing the pointer.

• @-commands used in node names generally confuse Info, so you should avoid them. This includes punctuation characters that are escaped with a ‘@’, such as @ and {. For a few rare cases when this is useful, Texinfo has limited support for using @-commands in node names; see Section 20.1.4 [Pointer Validation], page 156. Thus, the beginning of the section called @chapter looks like this:

     @node  chapter, unnumbered & appendix, makeinfo top, Structuring
     @comment  node-name,  next,  previous,  up
     @section @code{@@chapter}
     @findex chapter

• You cannot use parentheses in node names, because a node name such as ‘(foo)bar’ is interpreted by the Info readers as a node ‘bar’ in an Info file ‘foo’.

• Unfortunately, you cannot use periods, commas, colons or apostrophes within a node name; these confuse TEX or the Info formatters. For example, the following is a section title:

     @code{@@unnumberedsec}, @code{@@appendixsec}, @code{@@heading}

The corresponding node name is:

     unnumberedsec appendixsec heading

• Case is significant.

6.3.5 The First Node

The first node of a Texinfo file is the Top node, except in an included file (see Appendix D [Include Files], page 196). The Top node should contain a short summary, copying permissions, and a master menu. See Section 3.5 [The Top Node], page 36, for more information on the Top node contents and examples.

Here is a description of the node pointers to be used in the Top node:

• The Top node (which must be named ‘top’ or ‘Top’) should have as its ‘Up’ node the name of a node in another file, where there is a menu that leads to this file. Specify the file name in parentheses. Usually, all Info files are installed in the same Info directory tree; in this case, use ‘(dir)’ as the parent of the Top node; this is short for ‘(dir)top’, and specifies the Top node in the ‘dir’ file, which contains the main menu for the Info system as a whole.

• On the other hand, do not define the ‘Previous’ node of the Top node to be ‘(dir)’, as it causes confusing behavior for users: if you are in the Top node and hits ^{⟨}\underline{\overline{DEL}}^{⟩} to go backwards, you wind up in the middle of the some other entry in the ‘dir’ file, which has nothing to do with what you were reading.

• The ‘Next’ node of the Top node should be the first chapter in your document.

See Section 20.2 [Installing an Info File], page 161, for more information about installing an Info file in the ‘info’ directory.

For concreteness, here is an example with explicit pointers (which you can maintain automatically with the texinfo mode commands):

Or you can leave the pointers off entirely and let the tools implicitly define them. This is recommended. Thus:

     @node Top

6.3.6 The @top Sectioning Command

A special sectioning command, @top should be used with the @node Top line. The @top sectioning command tells makeinfo that it marks the ‘Top’ node in the file. It provides the information that makeinfo needs to insert node pointers automatically. Write the @top command at the beginning of the line immediately following the @node Top line. Write the title on the remaining part of the same line as the @top command.

In Info, the @top sectioning command causes the title to appear on a line by itself, with a line of asterisks inserted underneath, as other sectioning commands do.

In TEX and texinfo-format-buffer, the @top sectioning command is merely a synonym for @unnumbered. Neither of these formatters require an @top command, and do nothing special with it. You can use @chapter or @unnumbered after the @node Top line when you use these formatters. Also, you can use @chapter or @unnumbered when you use the Texinfo updating commands to create or update pointers and menus.

Thus, in practice, a Top node starts like this:

     @node Top
     @top Your Manual Title

6.4 Creating Pointers with makeinfo

The makeinfo program has a feature for automatically defining node pointers for a hierarchically organized file.

When you take advantage of this feature, you do not need to write the ‘Next’, ‘Previous’, and ‘Up’ pointers after the name of a node. However, you must write a sectioning command, such as @chapter or @section, on the line immediately following each truncated @node line (except that comment lines may intervene).

In addition, you must follow the ‘Top’ @node line with a line beginning with @top to mark the ‘Top’ node in the file. See Section 5.3 [@top], page 45.

Finally, you must write the name of each node (except for the ‘Top’ node) in a menu that is one or more hierarchical levels above the node’s hierarchical level.

This node pointer insertion feature in makeinfo relieves you from the need to update menus and pointers manually or with Texinfo mode commands. (See Section 2.4 [Updating Nodes and Menus], page 17.)

In most cases, you will want to take advantage of this feature and not redundantly specify node pointers. However, Texinfo documents are not required to be organized hierarchically or in fact contain sectioning commands at all. For example, if you never intend the document to be printed. In those cases, you will need to explicitly specify the pointers.

6.5 @anchor: Defining Arbitrary Cross-reference Targets

An anchor is a position in your document, labeled so that cross-references can refer to it, just as they can to nodes. You create an anchor with the @anchor command, and give the label as a normal brace-delimited argument. For example:

     This marks the @anchor{x-spot}spot.
     ...
     @xref{x-spot,,the spot}.

produces:

     This marks the spot.
     ...
     See [the spot], page 1.

As you can see, the @anchor command itself produces no output. This example defines an anchor ‘x-spot’ just before the word ‘spot’. You can refer to it later with an @xref or other cross-reference command, as shown. See Chapter 8 [Cross References], page 61, for details on the cross-reference commands.

It is best to put @anchor commands just before the position you wish to refer to; that way, the reader’s eye is led on to the correct text when they jump to the anchor. You can put the @anchor command on a line by itself if that helps readability of the source. Spaces are always ignored after @anchor.

Anchor names and node names may not conflict. Anchors and nodes are given similar treatment in some ways; for example, the goto-node command in standalone Info takes either an anchor name or a node name as an argument. (See section “goto-node” in GNU Info.)

7 Menus

Menus contain pointers to subordinate nodes.^{1} In Info, you use menus to go to such nodes. Menus have no effect in printed manuals and do not appear in them.

By convention, a menu is put at the end of a node since a reader who uses the menu may not see text that follows it. Furthermore, a node that has a menu should not contain much text. If you have a lot of text and a menu, move most of the text into a new subnode—all but a few lines. Otherwise, a reader with a terminal that displays only a few lines may miss the menu and its associated text. As a practical matter, you should locate a menu within 20 lines of the beginning of the node.

The short text before a menu may look awkward in a printed manual. To avoid this, you can write a menu near the beginning of its node and follow the menu by an @node line, and then an @heading line located within @ifinfo and @end ifinfo. This way, the menu, @node line, and title appear only in the Info file, not the printed document.

For example, the preceding two paragraphs follow an Info-only menu, @node line, and heading, and look like this:

     @menu
     * Menu Location::             Put a menu in a short node.
     * Writing a Menu::            What is a menu?
     * Menu Parts::                A menu entry has three parts.
     * Less Cluttered Menu Entry:: Two part menu entry.
     * Menu Example::              Two and three part entries.
     * Other Info Files::          How to refer to a different
                                     Info file.
     @end menu
     @node Menu Location, Writing a Menu, , Menus
     @ifinfo
     @heading Menus Need Short Nodes
     @end ifinfo

The Texinfo file for this document contains a number of examples of this procedure; one is at the beginning of this chapter.

7.1 Writing a Menu

A menu consists of an @menu command on a line by itself followed by menu entry lines or menu comment lines and then by an @end menu command on a line by itself.

A menu looks like this:

1 Menus can carry you to any node, regardless of the hierarchical structure; even to nodes in a different

Info file. However, the GNU Emacs Texinfo mode updating commands work only to create menus of

subordinate nodes. Conventionally, cross references are used to refer to other nodes.

     @menu
     Larger Units of Text
     * Files::                       All about handling files.
     * Multiples: Buffers.           Multiple buffers; editing
                                       several files at once.
     @end menu

In a menu, every line that begins with an ‘* ’ is a menu entry. (Note the space after the asterisk.) A line that does not start with an ‘* ’ may also appear in a menu. Such a line is not a menu entry but is a menu comment line that appears in the Info file. In the example above, the line ‘Larger Units of Text’ is a menu comment line; the two lines starting with ‘* ’ are menu entries. Space characters in a menu are preserved as-is; this allows you to format the menu as you wish.

7.2 The Parts of a Menu

A menu entry has three parts, only the second of which is required:

1. The menu entry name (optional).

2. The name of the node (required).

3. A description of the item (optional).

The template for a menu entry looks like this:

* menu-entry-name: node-name. description

Follow the menu entry name with a single colon and follow the node name with tab, comma, period, or newline.

In Info, a user selects a node with the m (Info-menu) command. The menu entry name is what the user types after the m command.

The third part of a menu entry is a descriptive phrase or sentence. Menu entry names and node names are often short; the description explains to the reader what the node is about. A useful description complements the node name rather than repeats it. The description, which is optional, can spread over two or more lines; if it does, some authors prefer to indent the second line while others prefer to align it with the first (and all others). It’s up to you.

7.3 Less Cluttered Menu Entry

When the menu entry name and node name are the same, you can write the name immediately after the asterisk and space at the beginning of the line and follow the name with two colons.

For example, write

* Name:: description instead of

     * Name: Name.  description

You should use the node name for the menu entry name whenever possible, since it reduces visual clutter in the menu.

7.4 A Menu Example

A menu looks like this in Texinfo:

@menu
* menu entry name: Node name.   A short description.
* Node name::                   This form is preferred.
@end menu

This produces:

* menu:
* menu entry name: Node name.   A short description.
* Node name::                   This form is preferred.

Here is an example as you might see it in a Texinfo file:

@menu
Larger Units of Text
* Files::                       All about handling files.
* Multiples: Buffers.           Multiple buffers; editing
                                  several files at once.
@end menu

This produces:

* menu:
Larger Units of Text
* Files::                       All about handling files.
* Multiples: Buffers.           Multiple buffers; editing
                                  several files at once.

In this example, the menu has two entries. ‘Files’ is both a menu entry name and the name of the node referred to by that name. ‘Multiples’ is the menu entry name; it refers to the node named ‘Buffers’. The line ‘Larger Units of Text’ is a comment; it appears in the menu, but is not an entry.

Since no file name is specified with either ‘Files’ or ‘Buffers’, they must be the names of nodes in the same Info file (see Section 7.5 [Referring to Other Info Files], page 59).

7.5 Referring to Other Info Files

You can create a menu entry that enables a reader in Info to go to a node in another Info file by writing the file name in parentheses just before the node name. In this case, you should use the three-part menu entry format, which saves the reader from having to type the file name.

The format looks like this:

@menu

* first-entry-name:(filename)nodename. description

* second-entry-name:(filename)second-node. description

@end menu

For example, to refer directly to the ‘Outlining’ and ‘Rebinding’ nodes in the Emacs Manual, you would write a menu like this:

     @menu
     * Outlining: (emacs)Outline Mode. The major mode for
                                       editing outlines.
     * Rebinding: (emacs)Rebinding.    How to redefine the
                                       meaning of a key.
     @end menu

If you do not list the node name, but only name the file, then Info presumes that you are referring to the ‘Top’ node.

The ‘dir’ file that contains the main menu for Info has menu entries that list only file names. These take you directly to the ‘Top’ nodes of each Info document. (See Section 20.2 [Installing an Info File], page 161.)

For example:

     * Info: (info).         Documentation browsing system.
     * Emacs: (emacs).       The extensible, self-documenting
                             text editor.

(The ‘dir’ top level directory for the Info system is an Info file, not a Texinfo file, but a menu entry looks the same in both types of file.)

The GNU Emacs Texinfo mode menu updating commands only work with nodes within the current buffer, so you cannot use them to create menus that refer to other files. You must write such menus by hand.

8 Cross References

Cross references are used to refer the reader to other parts of the same or different Texinfo files. In Texinfo, nodes and anchors are the places to which cross references can refer.

Often, but not always, a printed document should be designed so that it can be read sequentially. People tire of flipping back and forth to find information that should be presented to them as they need it.

However, in any document, some information will be too detailed for the current context, or incidental to it; use cross references to provide access to such information. Also, an online help system or a reference manual is not like a novel; few read such documents in sequence from beginning to end. Instead, people look up what they need. For this reason, such creations should contain many cross references to help readers find other information that they may not have read.

In a printed manual, a cross reference results in a page reference, unless it is to another manual altogether, in which case the cross reference names that manual.

In Info, a cross reference results in an entry that you can follow using the Info ‘f’ command. (See Info file ‘info’, node ‘Help-Adv’.)

The various cross reference commands use nodes (or anchors, see Section 6.5 [@anchor], page 56) to define cross reference locations. This is evident in Info, in which a cross reference takes you to the specified location. TEX also uses nodes to define cross reference locations, but the action is less obvious. When TEX generates a DVI file, it records each node’s page number and uses the page numbers in making references. Thus, if you are writing a manual that will only be printed, and will not be used online, you must nonetheless write @node lines to name the places to which you make cross references.

8.1 Different Cross Reference Commands

There are four different cross reference commands:

@xref Used to start a sentence in the printed manual saying ‘See …’ or an Info cross-reference saying ‘*Note name: node.’.

@ref Used within or, more often, at the end of a sentence; same as @xref for Info; produces just the reference in the printed manual without a preceding ‘See’.

@pxref Used within parentheses to make a reference that suits both an Info file and a printed book. Starts with a lower case ‘see’ within the printed manual. (‘p’ is for ‘parenthesis’.)

@inforef Used to make a reference to an Info file for which there is no printed manual.

(The @cite command is used to make references to books and manuals for which there is no corresponding Info file and, therefore, no node to which to point. See Section 9.1.12 [@cite], page 77.)

8.2 Parts of a Cross Reference

A cross reference command requires only one argument, which is the name of the node to which it refers. But a cross reference command may contain up to four additional arguments. By using these arguments, you can provide a cross reference name for Info, a topic description or section title for the printed output, the name of a different Info file, and the name of a different printed manual.

Here is a simple cross reference example:

     @xref{Node name}.

which produces

     *Note Node name::.

and See Section nnn [Node name], page ppp. Here is an example of a full five-part cross reference:

     @xref{Node name, Cross Reference Name, Particular Topic,
     info-file-name, A Printed Manual}, for details.

which produces

     *Note Cross Reference Name: (info-file-name)Node name,
     for details.

in Info and

See section “Particular Topic” in A Printed Manual, for details. in a printed book.

The five possible arguments for a cross reference are:

1. The node or anchor name (required). This is the location to which the cross reference takes you. In a printed document, the location of the node provides the page reference only for references within the same document.

2. The cross reference name for the Info reference, if it is to be different from the node name. If you include this argument, it becomes the first part of the cross reference. It is usually omitted.

3. A topic description or section name. Often, this is the title of the section. This is used as the name of the reference in the printed manual. If omitted, the node name is used.

4. The name of the Info file in which the reference is located, if it is different from the current file. You need not include any ‘.info’ suffix on the file name, since Info readers try appending it automatically.

5. The name of a printed manual from a different Texinfo file.

The template for a full five argument cross reference looks like this:

@xref{node-name, cross-reference-name, title-or-topic, info-file-name, printed-manual-title}. Cross references with one, two, three, four, and five arguments are described separately

following the description of @xref.

Write a node name in a cross reference in exactly the same way as in the @node line, including the same capitalization; otherwise, the formatters may not find the reference.

You can write cross reference commands within a paragraph, but note how Info and TEX format the output of each of the various commands: write @xref at the beginning of a sentence; write @pxref only within parentheses, and so on.

8.3 @xref

The @xref command generates a cross reference for the beginning of a sentence. The Info formatting commands convert it into an Info cross reference, which the Info ‘f’ command can use to bring you directly to another node. The TEX typesetting commands convert it into a page reference, or a reference to another book or manual.

Most often, an Info cross reference looks like this:

*Note node-name::. or like this

*Note cross-reference-name: node-name. In TEX, a cross reference looks like this:

See Section section-number [node-name], page page. or like this

See Section section-number [title-or-topic], page page.

The @xref command does not generate a period or comma to end the cross reference in either the Info file or the printed output. You must write that period or comma yourself; otherwise, Info will not recognize the end of the reference. (The @pxref command works differently. See Section 8.6 [@pxref], page 68.)

Please note: A period or comma must follow the closing brace of an @xref. It is required to terminate the cross reference. This period or comma will appear in the output, both in the Info file and in the printed manual. @xref must refer to an Info node by name. Use @node to define the node (see Sec-

tion 6.3.2 [Writing a Node], page 53).

@xref is followed by several arguments inside braces, separated by commas. Whitespace before and after these commas is ignored.

A cross reference requires only the name of a node; but it may contain up to four additional arguments. Each of these variations produces a cross reference that looks somewhat different.

Please note: Commas separate arguments in a cross reference; avoid including them in the title or other part lest the formatters mistake them for separators.

8.3.1 @xref with One Argument

The simplest form of @xref takes one argument, the name of another node in the same Info file. The Info formatters produce output that the Info readers can use to jump to the reference; TEX produces output that specifies the page and section number for you. For example,

     @xref{Tropical Storms}.

produces

     *Note Tropical Storms::.

and

See Section 3.1 [Tropical Storms], page 24. (Note that in the preceding example the closing brace is followed by a period.)

You can write a clause after the cross reference, like this:

     @xref{Tropical Storms}, for more info.

which produces

     *Note Tropical Storms::, for more info.

and

See Section 3.1 [Tropical Storms], page 24, for more info. (Note that in the preceding example the closing brace is followed by a comma, and then by the clause, which is followed by a period.)

8.3.2 @xref with Two Arguments

With two arguments, the second is used as the name of the Info cross reference, while the first is still the name of the node to which the cross reference points. The template is like this:

@xref{node-name, cross-reference-name}. For example,

     @xref{Electrical Effects, Lightning}.

produces:

     *Note Lightning: Electrical Effects.

and

See Section 5.2 [Electrical Effects], page 57. (Note that in the preceding example the closing brace is followed by a period; and that the node name is printed, not the cross reference name.)

You can write a clause after the cross reference, like this:

     @xref{Electrical Effects, Lightning}, for more info.

which produces

     *Note Lightning: Electrical Effects, for more info.

and

See Section 5.2 [Electrical Effects], page 57, for more info. (Note that in the preceding example the closing brace is followed by a comma, and then by the clause, which is followed by a period.)

8.3.3 @xref with Three Arguments

A third argument replaces the node name in the TEX output. The third argument should be the name of the section in the printed output, or else state the topic discussed by that section. Often, you will want to use initial upper case letters so it will be easier to read when the reference is printed. Use a third argument when the node name is unsuitable because of syntax or meaning.

Remember to avoid placing a comma within the title or topic section of a cross reference, or within any other section. The formatters divide cross references into arguments according to the commas; a comma within a title or other section will divide it into two arguments. In a reference, you need to write a title such as “Clouds, Mist, and Fog” without the commas.

Also, remember to write a comma or period after the closing brace of an @xref to terminate the cross reference. In the following examples, a clause follows a terminating comma.

The template is like this:

@xref{node-name, cross-reference-name, title-or-topic}.

For example,

     @xref{Electrical Effects, Lightning, Thunder and Lightning},
     for details.

produces

     *Note Lightning: Electrical Effects, for details.

and

See Section 5.2 [Thunder and Lightning], page 57, for details.

If a third argument is given and the second one is empty, then the third argument serves both. (Note how two commas, side by side, mark the empty second argument.)

     @xref{Electrical Effects, , Thunder and Lightning},
     for details.

produces

     *Note Thunder and Lightning: Electrical Effects, for details.

and

See Section 5.2 [Thunder and Lightning], page 57, for details.

As a practical matter, it is often best to write cross references with just the first argument if the node name and the section title are the same, and with the first and third arguments if the node name and title are different.

Here are several examples from The GNU Awk User’s Guide:

     @xref{Sample Program}.
     @xref{Glossary}.
     @xref{Case-sensitivity, ,Case-sensitivity in Matching}.
     @xref{Close Output, , Closing Output Files and Pipes},
        for more information.
     @xref{Regexp, , Regular Expressions as Patterns}.

8.3.4 @xref with Four and Five Arguments

In a cross reference, a fourth argument specifies the name of another Info file, different from the file in which the reference appears, and a fifth argument specifies its title as a printed manual.

Remember that a comma or period must follow the closing brace of an @xref command to terminate the cross reference. In the following examples, a clause follows a terminating comma. The template is:

@xref{node-name, cross-reference-name, title-or-topic,

info-file-name, printed-manual-title}. For example,

     @xref{Electrical Effects, Lightning, Thunder and Lightning,
     weather, An Introduction to Meteorology}, for details.

produces

     *Note Lightning: (weather)Electrical Effects, for details.

The name of the Info file is enclosed in parentheses and precedes the name of the node. In a printed manual, the reference looks like this:

See section “Thunder and Lightning” in An Introduction to Meteorology, for details.

The title of the printed manual is typeset in italics; and the reference lacks a page number since TEX cannot know to which page a reference refers when that reference is to another manual.

Often, you will leave out the second argument when you use the long version of @xref. In this case, the third argument, the topic description, will be used as the cross reference name in Info. The template looks like this:

@xref{node-name, , title-or-topic, info-file-name,

printed-manual-title}, for details. which produces

*Note title-or-topic: (info-file-name)node-name, for details. and

See section title-or-topic in printed-manual-title, for details. For example,

     @xref{Electrical Effects, , Thunder and Lightning,
     weather, An Introduction to Meteorology}, for details.

produces

     *Note Thunder and Lightning: (weather)Electrical Effects,
     for details.

and See section “Thunder and Lightning” in An Introduction to Meteorology, for details.

On rare occasions, you may want to refer to another Info file that is within a single printed manual—when multiple Texinfo files are incorporated into the same TEX run but make separate Info files. In this case, you need to specify only the fourth argument, and not the fifth.

8.4 Naming a ‘Top’ Node

In a cross reference, you must always name a node. This means that in order to refer to a whole manual, you must identify the ‘Top’ node by writing it as the first argument to the @xref command. (This is different from the way you write a menu entry; see Section 7.5 [Referring to Other Info Files], page 59.) At the same time, to provide a meaningful section topic or title in the printed cross reference (instead of the word ‘Top’), you must write an appropriate entry for the third argument to the @xref command.

Thus, to make a cross reference to The GNU Make Manual, write:

     @xref{Top, , Overview, make, The GNU Make Manual}.

which produces

     *Note Overview: (make)Top.

and See section “Overview” in The GNU Make Manual.

In this example, ‘Top’ is the name of the first node, and ‘Overview’ is the name of the first section of the manual.

8.5 @ref

@ref is nearly the same as @xref except that it does not generate a ‘See’ in the printed output, just the reference itself. This makes it useful as the last part of a sentence. For example,

     For more information, see @ref{Hurricanes}.

produces

     For more information, see *Note Hurricanes::.

and For more information, see Section 8.2 [Hurricanes], page 123.

The @ref command sometimes leads writers to express themselves in a manner that is suitable for a printed manual but looks awkward in the Info format. Bear in mind that your audience will be using both the printed and the Info format. For example,

     Sea surges are described in @ref{Hurricanes}.

produces

Sea surges are described in Section 6.7 [Hurricanes], page 72. in a printed document, and the following in Info:

     Sea surges are described in *Note Hurricanes::.

Caution: You must write a period, comma, or right parenthesis immediately after an @ref command with two or more arguments. Otherwise, Info will not find the end of the cross reference entry and its attempt to follow the cross reference will fail. As a general rule, you should write a period or comma after every @ref command. This looks best in both the printed and the Info output.

8.6 @pxref

The parenthetical reference command, @pxref, is nearly the same as @xref, but you use it only inside parentheses and you do not type a comma or period after the command’s closing brace. The command differs from @xref in two ways:

1. TEX typesets the reference for the printed manual with a lower case ‘see’ rather than an upper case ‘See’.

2. The Info formatting commands automatically end the reference with a closing colon or period.

Because one type of formatting automatically inserts closing punctuation and the other does not, you should use @pxref only inside parentheses as part of another sentence. Also, you yourself should not insert punctuation after the reference, as you do with @xref.

@pxref is designed so that the output looks right and works right between parentheses both in printed output and in an Info file. In a printed manual, a closing comma or period should not follow a cross reference within parentheses; such punctuation is wrong. But in an Info file, suitable closing punctuation must follow the cross reference so Info can recognize its end. @pxref spares you the need to use complicated methods to put a terminator into one form of the output and not the other. With one argument, a parenthetical cross reference looks like this:

     ... storms cause flooding (@pxref{Hurricanes}) ...

which produces

     ... storms cause flooding (*Note Hurricanes::) ...

and … storms cause flooding (see Section 6.7 [Hurricanes], page 72) … With two arguments, a parenthetical cross reference has this template: ... (@pxref{node-name, cross-reference-name}) ...

which produces

... (*Note cross-reference-name: node-name.) ... and

… (see Section nnn [node-name], page ppp) …

@pxref can be used with up to five arguments just like @xref (see Section 8.3 [@xref], page 63).

Please note: Use @pxref only as a parenthetical reference. Do not try to use @pxref as a clause in a sentence. It will look bad in either the Info file, the printed output, or both.

Also, parenthetical cross references look best at the ends of sentences. Although you may write them in the middle of a sentence, that location breaks up the flow of text.

8.7 @inforef

@inforef is used for cross references to Info files for which there are no printed manuals. Even in a printed manual, @inforef generates a reference directing the user to look in an Info file.

The command takes either two or three arguments, in the following order:

1. The node name.

2. The cross reference name (optional).

3. The Info file name. Separate the arguments with commas, as with @xref. Also, you must terminate the reference with a comma or period after the ‘}’, as you do with @xref. The template is:

@inforef{node-name, cross-reference-name, info-file-name}, Thus,

     @inforef{Expert, Advanced Info commands, info},
     for more information.

produces

     *Note Advanced Info commands: (info)Expert,
     for more information.

and

See Info file ‘info’, node ‘Expert’, for more information. Similarly,

     @inforef{Expert, , info}, for more information.

produces

     *Note (info)Expert::, for more information.

and See Info file ‘info’, node ‘Expert’, for more information. The converse of @inforef is @cite, which is used to refer to printed works for which

no Info form exists. See Section 9.1.12 [@cite], page 77.

8.8 @uref{url[, text][, replacement]}

@uref produces a reference to a uniform resource locator (url). It takes one mandatory argument, the url, and two optional arguments which control the text that is displayed. In HTML output, @uref produces a link you can follow.

The second argument, if specified, is the text to display (the default is the url itself); in Info and DVI output, but not in HTML output, the url is also output.

The third argument, on the other hand, if specified is also the text to display, but the url is not output in any format. This is useful when the text is already sufficiently referential, as in a man page. If the third argument is given, the second argument is ignored.

The simple one argument form, where the url is both the target and the text of the link:

     The official GNU ftp site is @uref{ftp://ftp.gnu.org/gnu}.

produces:

The official GNU ftp site is ftp://ftp.gnu.org/gnu.

An example of the two-argument form:

     The official @uref{ftp://ftp.gnu.org/gnu, GNU ftp site}
     holds programs and texts.

produces:

     The official GNU ftp site (ftp://ftp.gnu.org/gnu)

holds programs and texts. that is, the Info output is this:

     The official GNU ftp site (ftp://ftp.gnu.org/gnu)
     holds programs and texts.

and the HTML output is this:

     The official <a href="ftp://ftp.gnu.org/gnu">GNU ftp site</a>
     holds programs and texts.

An example of the three-argument form:

     The @uref{/man.cgi/1/ls,,ls(1)} program ...

produces:

The ls(1) program … but with HTML:

     The <a href="/man.cgi/1/ls">ls(1)</a> program ...

To merely indicate a url without creating a link people can follow, use @url (see Section 9.1.14 [url], page 78).

Some people prefer to display url’s in the unambiguous format:

<URL:http://host/path> You can use this form in the input file if you wish. We feel it’s not necessary to clutter up the output with the extra ‘<URL:’ and ‘>’, since any software that tries to detect url’s in text already has to detect them without the ‘<URL:’ to be useful.

9 Marking Words and Phrases

In Texinfo, you can mark words and phrases in a variety of ways. The Texinfo formatters use this information to determine how to highlight the text. You can specify, for example, whether a word or phrase is a defining occurrence, a metasyntactic variable, or a symbol used in a program. Also, you can emphasize text, in several different ways.

9.1 Indicating Definitions, Commands, etc.

Texinfo has commands for indicating just what kind of object a piece of text refers to. For example, metasyntactic variables are marked by @var, and code by @code. Since the pieces of text are labelled by commands that tell what kind of object they are, it is easy to change the way the Texinfo formatters prepare such text. (Texinfo is an intentional formatting language rather than a typesetting formatting language.)

For example, in a printed manual, code is usually illustrated in a typewriter font; @code tells TEX to typeset this text in this font. But it would be easy to change the way TEX highlights code to use another font, and this change would not affect how keystroke examples are highlighted. If straight typesetting commands were used in the body of the file and you wanted to make a change, you would need to check every single occurrence to make sure that you were changing code and not something else that should not be changed.

The highlighting commands can be used to extract useful information from the file, such as lists of functions or file names. It is possible, for example, to write a program in Emacs Lisp (or a keyboard macro) to insert an index entry after every paragraph that contains words or phrases marked by a specified command. You could do this to construct an index of functions if you had not already made the entries.

The commands serve a variety of purposes:

@code{sample-code} Indicate text that is a literal example of a piece of a program.

@kbd{keyboard-characters} Indicate keyboard input.

@key{key-name} Indicate the conventional name for a key on a keyboard.

@samp{text}

Indicate text that is a literal example of a sequence of characters.

@var{metasyntactic-variable} Indicate a metasyntactic variable.

@env{environment-variable} Indicate an environment variable.

@file{file-name} Indicate the name of a file.

@command{command-name} Indicate the name of a command.

@option{option}

Indicate a command-line option.

@dfn{term}

Indicate the introductory or defining use of a term.

@cite{reference} Indicate the name of a book.

@acronym{acronym}

Indicate an acronym.

@url{uniform-resource-locator} Indicate a uniform resource locator for the World Wide Web.

@email{email-address[, displayed-text]} Indicate an electronic mail address.

9.1.1 @code{sample-code}

Use the @code command to indicate text that is a piece of a program and which consists of entire syntactic tokens. Enclose the text in braces.

Thus, you should use @code for an expression in a program, for the name of a variable or function used in a program, or for a keyword in a programming language.

Use @code for command names in languages that resemble programming languages, such as Texinfo. For example, @code and @samp are produced by writing ‘@code{@@code}’ and ‘@code{@@samp}’ in the Texinfo source, respectively.

It is incorrect to alter the case of a word inside an @code command when it appears at the beginning of a sentence. Most computer languages are case sensitive. In C, for example, Printf is different from the identifier printf, and most likely is a misspelling of it. Even in languages which are not case sensitive, it is confusing to a human reader to see identifiers spelled in different ways. Pick one spelling and always use that. If you do not want to start a sentence with a command name written all in lower case, you should rearrange the sentence.

In the printed manual, @code causes TEX to typeset the argument in a typewriter face. In the Info file, it causes the Info formatting commands to use single quotation marks around the text.

For example,

     The function returns @code{nil}.

produces this in the printed manual:

The function returns nil. and this in the Info file:

     The function returns `nil'.

Here are some cases for which it is preferable not to use @code:

• For shell command names such as ls (use @command).

• For shell options such as ‘-c’ when such options stand alone (use @option).

• Also, an entire shell command often looks better if written using @samp rather than @code. In this case, the rule is to choose the more pleasing format.

• For environment variable such as TEXINPUTS (use @env).

• For a string of characters shorter than a syntactic token. For example, if you are writing about ‘goto-ch’, which is just a part of the name for the goto-char Emacs Lisp function, you should use @samp.

• In general, when writing about the characters used in a token; for example, do not use @code when you are explaining what letters or printable symbols can be used in the names of functions. (Use @samp.) Also, you should not use @code to mark text that is considered input to programs unless the input is written in a language that is like a programming language. For example, you should not use @code for the keystroke commands of GNU Emacs (use @kbd instead) although you may use @code for the names of the Emacs Lisp functions that the keystroke commands invoke.

Since @command, @option, and @env were introduced relatively recently, it is acceptable to use @code or @samp for command names, options, and environment variables. The new commands allow you to express the markup more precisely, but there is no real harm in using the older commands, and of course the long-standing manuals do so.

9.1.2 @kbd{keyboard-characters}

Use the @kbd command for characters of input to be typed by users. For example, to refer to the characters M-a, write

     @kbd{M-a}

and to refer to the characters M-x shell, write

     @kbd{M-x shell}

The @kbd command has the same effect as @code in Info, but by default produces a different font (slanted typewriter instead of normal typewriter) in the printed manual, so users can distinguish the characters they are supposed to type from those the computer outputs.

Since the usage of @kbd varies from manual to manual, you can control the font switching with the @kbdinputstyle command. This command has no effect on Info output. Write this command at the beginning of a line with a single word as an argument, one of the following:

‘code’ Always use the same font for @kbd as @code.

‘example’ Use the distinguishing font for @kbd only in @example and similar environments.

‘distinct’

(the default) Always use the distinguishing font for @kbd.

You can embed another @-command inside the braces of an @kbd command. Here, for example, is the way to describe a command that would be described more verbosely as “press an ‘r’ and then press the ^{⟨}\underline{\overline{RET}}^{⟩} key”:

     @kbd{r @key{RET}}

This produces: r ^{⟨}\underline{\overline{RET}}^{⟩} You also use the @kbd command if you are spelling out the letters you type; for example:

To give the @code{logout} command,
type the characters @kbd{l o g o u t @key{RET}}.

This produces:

To give the logout command, type the characters l o g o u t ^{⟨}\underline{\overline{RET}}^{⟩}.

(Also, this example shows that you can add spaces for clarity. If you really want to mention a space character as one of the characters of input, write @key{SPC} for it.)

9.1.3 @key{key-name}

Use the @key command for the conventional name for a key on a keyboard, as in:

@key{RET}

You can use the @key command within the argument of an @kbd command when the sequence of characters to be typed includes one or more keys that are described by name.

For example, to produce C-x ^{⟨}\underline{\overline{ESC}}^{⟩} you would type:

@kbd{C-x @key{ESC}}

Here is a list of the recommended names for keys:

SPC Space

RET Return

LFD Linefeed (however, since most keyboards nowadays do not have a Linefeed key, it might be better to call this character C-j.

TAB  Tab

BS Backspace

ESC Escape

DEL Delete

SHIFT  Shift

CTRL Control

META  Meta

There are subtleties to handling words like ‘meta’ or ‘ctrl’ that are names of modifier keys. When mentioning a character in which the modifier key is used, such as Meta-a, use the @kbd command alone; do not use the @key command; but when you are referring to the modifier key in isolation, use the @key command. For example, write ‘@kbd{Meta-a}’ to produce Meta-a and ‘@key\{META\}’ to produce ^{⟨}\underline{\overline{META}}^{⟩}.

9.1.4 @samp{text}

Use the @samp command to indicate text that is a literal example or ‘sample’ of a sequence of characters in a file, string, pattern, etc. Enclose the text in braces. The argument appears within single quotation marks in both the Info file and the printed manual; in addition, it is printed in a fixed-width font.

     To match @samp{foo} at the end of the line,
     use the regexp @samp{foo$}.

produces To match ‘foo’ at the end of the line, use the regexp ‘foo$’. Any time you are referring to single characters, you should use @samp unless @kbd or

@key is more appropriate. Also, you may use @samp for entire statements in C and for entire shell commands—in this case, @samp often looks better than @code. Basically, @samp is a catchall for whatever is not covered by @code, @kbd, or @key.

Only include punctuation marks within braces if they are part of the string you are specifying. Write punctuation marks outside the braces if those punctuation marks are part of the English text that surrounds the string. In the following sentence, for example, the commas and period are outside of the braces:

     In English, the vowels are @samp{a}, @samp{e},
     @samp{i}, @samp{o}, @samp{u}, and sometimes
     @samp{y}.

This produces:

In English, the vowels are ‘a’, ‘e’, ‘i’, ‘o’, ‘u’, and sometimes ‘y’.

9.1.5 @verb{<char>text<char>}

Use the @verb command to print a verbatim sequence of characters.

Like LaTEX’s \verb command, the verbatim text can be quoted using any unique delimiter character. Enclose the verbatim text, including the delimiters, in braces. Text is printed in a fixed-width font:

     How many @verb{|@|}-escapes does one need to print this
     @verb{.@a @b @c.} string or @verb{+@'e?`!`{}\+} this?

produces

     How many @-escapes does one need to print this
     @a @b @c string or these @'e?`{}!`\ this?

This is in contrast to @samp (see the previous section), whose argument is normal Texinfo text, where the characters @{} are special; with @verb, nothing is special except the delimiter character you choose.

9.1.6 @var{metasyntactic-variable}

Use the @var command to indicate metasyntactic variables. A metasyntactic variable is something that stands for another piece of text. For example, you should use a metasyntactic variable in the documentation of a function to describe the arguments that are passed to that function.

Do not use @var for the names of particular variables in programming languages. These are specific names from a program, so @code is correct for them (see Section 9.1.1 [code], page 72). For example, the Emacs Lisp variable texinfo-tex-command is not a metasyntactic variable; it is properly formatted using @code.

Do not use @var for environment variables either; @env is correct for them (see the next section).

The effect of @var in the Info file is to change the case of the argument to all upper case. In the printed manual and HTML output, the argument is printed in slanted type.

For example,

     To delete file @var{filename},
     type @samp{rm @var{filename}}.

produces

To delete file filename, type ‘rm filename’. (Note that @var may appear inside @code, @samp, @file, etc.)

Write a metasyntactic variable all in lower case without spaces, and use hyphens to make it more readable. Thus, the Texinfo source for the illustration of how to begin a Texinfo manual looks like this:

     \input texinfo
     @@setfilename @var{info-file-name}
     @@settitle @var{name-of-manual}

This produces:

     \input texinfo

@setfilename info-file-name @settitle name-of-manual In some documentation styles, metasyntactic variables are shown with angle brackets,

for example:

     ..., type rm <filename>

However, that is not the style that Texinfo uses. (You can, of course, modify the sources to ‘texinfo.tex’ and the Info formatting commands to output the <...> format if you wish.)

9.1.7 @env{environment-variable}

Use the @env command to indicate environment variables, as used by many operating systems, including GNU. Do not use it for metasyntactic variables; use @var instead (see the previous section).

@env is equivalent to @code in its effects. For example:

     The @env{PATH} environment variable ...

produces The PATH environment variable …

9.1.8 @file{file-name}

Use the @file command to indicate text that is the name of a file, buffer, or directory, or is the name of a node in Info. You can also use the command for file name suffixes. Do not use @file for symbols in a programming language; use @code.

Currently, @file is equivalent to @samp in its effects. For example,

     The @file{.el} files are in
     the @file{/usr/local/emacs/lisp} directory.

produces The ‘.el’ files are in the ‘/usr/local/emacs/lisp’ directory.

9.1.9 @command{command-name}

Use the @command command to indicate command names, such as ls or cc.

@command is equivalent to @code in its effects. For example:

     The command @command{ls} lists directory contents.

produces The command ls lists directory contents.

You should write the name of a program in the ordinary text font, rather than using @command, if you regard it as a new English word, such as ‘Emacs’ or ‘Bison’.

When writing an entire shell command invocation, as in ‘ls -l’, you should use either @samp or @code at your discretion.

9.1.10 @option{option-name}

Use the @option command to indicate a command-line option; for example, ‘-l’ or

‘--version’ or ‘--output=filename’.

@option is equivalent to @samp in its effects. For example:

     The option @option{-l} produces a long listing.

produces The option ‘-l’ produces a long listing.

In tables, putting options inside @code produces a more pleasing effect.

9.1.11 @dfn{term}

Use the @dfn command to identify the introductory or defining use of a technical term. Use the command only in passages whose purpose is to introduce a term which will be used again or which the reader ought to know. Mere passing mention of a term for the first time does not deserve @dfn. The command generates italics in the printed manual, and double quotation marks in the Info file. For example:

     Getting rid of a file is called @dfn{deleting} it.

produces Getting rid of a file is called deleting it.

As a general rule, a sentence containing the defining occurrence of a term should be a definition of the term. The sentence does not need to say explicitly that it is a definition, but it should contain the information of a definition—it should make the meaning clear.

9.1.12 @cite{reference}

Use the @cite command for the name of a book that lacks a companion Info file. The command produces italics in the printed manual, and quotation marks in the Info file.

If a book is written in Texinfo, it is better to use a cross reference command since a reader can easily follow such a reference in Info. See Section 8.3 [@xref], page 63.

9.1.13 @acronym{acronym}

Use the @acronym command for abbreviations written in all capital letters, such as ‘nasa’. The abbreviation is given as the single argument in braces, as in ‘@acronym{NASA}’. As a matter of style, or for particular abbreviations, you may prefer to use periods, as in

‘@acronym{F.B.I.}’.

In TEX and HTML, the argument is printed in a slightly smaller font size. In Info or plain text output, this command changes nothing.

9.1.14 @url{uniform-resource-locator}

Use the @url command to indicate a uniform resource locator on the World Wide Web. This is analogous to @file, @var, etc., and is purely for markup purposes. It does not produce a link you can follow in HTML output (use the @uref command for that, see Section 8.8 [@uref], page 69). It is useful for url’s which do not actually exist. For example:

     For example, the url might be @url{http://example.org/path}.

which produces:

For example, the url might be http://example.org/path.

9.1.15 @email{email-address[, displayed-text]}

Use the @email command to indicate an electronic mail address. It takes one mandatory argument, the address, and one optional argument, the text to display (the default is the address itself).

In Info and TEX, the address is shown in angle brackets, preceded by the text to display if any. In HTML output, @email produces a ‘mailto’ link that usually brings up a mail composition window. For example:

     Send bug reports to @email{bug-texinfo@@gnu.org},
     suggestions to the @email{bug-texinfo@@gnu.org, same place}.

produces

     Send bug reports to bug-texinfo@gnu.org,

suggestions to the same place (bug-texinfo@gnu.org).

9.2 Emphasizing Text

Usually, Texinfo changes the font to mark words in the text according to what category the words belong to; an example is the @code command. Most often, this is the best way to mark words. However, sometimes you will want to emphasize text without indicating a category. Texinfo has two commands to do this. Also, Texinfo has several commands that specify the font in which TEX will typeset text. These commands have no effect on Info and only one of them, the @r command, has any regular use.

9.2.1 @emph{text} and @strong{text}

The @emph and @strong commands are for emphasis; @strong is stronger. In printed output, @emph produces italics and @strong produces bold.

For example,

     @quotation
     @strong{Caution:} @samp{rm * .[^.]*} removes @emph{all}
     files in the directory.
     @end quotation

produces the following in printed output:

Caution: ‘rm * .[^.]*’ removes all files in the directory.

and the following in Info:

          *Caution*: `rm * .[^.]*' removes _all_
          files in the directory.

The @strong command is seldom used except to mark what is, in effect, a typographical element, such as the word ‘Caution’ in the preceding example.

In the Info output, @emph surrounds the text with underscores (‘_’), and @strong puts asterisks around the text.

Caution: Do not use @strong with the word ‘Note’; Info will mistake the combination for a cross reference. Use a phrase such as Please note or Caution instead.

9.2.2 @sc{text}: The Small Caps Font

Use the ‘@sc’ command to set text in the printed and the HTML output in a small caps font and set text in the Info file in upper case letters. Write the text you want to be in small caps (where possible) between braces in lower case, like this:

     The @sc{acm} and @sc{ieee} are technical societies.

This produces:

The acm and ieee are technical societies.

TEX typesets the small caps font in a manner that prevents the letters from ‘jumping out at you on the page’. This makes small caps text easier to read than text in all upper case— but it’s usually better to use regular mixed case anyway. The Info formatting commands set all small caps text in upper case. In HTML, the text is upper-cased and a smaller font is used to render it.

If the text between the braces of an @sc command is uppercase, TEX typesets in FULL-SIZE CAPITALS. Use full-size capitals sparingly, if ever, and since it’s redundant to mark all-uppercase text with @sc, makeinfo warns about such usage.

You may also use the small caps font for a jargon word such as ato (a nasa word meaning ‘abort to orbit’).

There are subtleties to using the small caps font with a jargon word such as cdr, a word used in Lisp programming. In this case, you should use the small caps font when the word refers to the second and subsequent elements of a list (the cdr of the list), but you should use ‘@code’ when the word refers to the Lisp function of the same spelling.

9.2.3 Fonts for Printing, Not Info

Texinfo provides four font commands that specify font changes in the printed manual but have no effect in the Info file. @i requests italic font (in some versions of TEX, a slanted font is used), @b requests bold face, @t requests the fixed-width, typewriter-style font used by @code, and @r requests a roman font, which is the usual font in which text is printed. All four commands apply to an argument that follows, surrounded by braces.

Only the @r command has much use: in example programs, you can use the @r command to convert code comments from the fixed-width font to a roman font. This looks better in printed output.

For example,

     @lisp
     (+ 2 2)    ; @r{Add two plus two.}
     @end lisp

produces (+ 2 2) ; Add two plus two. If possible, you should avoid using the other three font commands. If you need to use

one, it probably indicates a gap in the Texinfo language.

10 Quotations and Examples

Quotations and examples are blocks of text consisting of one or more whole paragraphs that are set off from the bulk of the text and treated differently. They are usually indented.

In Texinfo, you always begin a quotation or example by writing an @-command at the beginning of a line by itself, and end it by writing an @end command that is also at the beginning of a line by itself. For instance, you begin an example by writing @example by itself at the beginning of a line and end the example by writing @end example on a line by itself, at the beginning of that line.

10.1 Block Enclosing Commands

Here are commands for quotations and examples, explained further in the following sections:

@quotation

Indicate text that is quoted. The text is filled, indented, and printed in a roman font by default.

@example Illustrate code, commands, and the like. The text is printed in a fixed-width font, and indented but not filled.

@verbatim

Mark a piece of text that is to be printed verbatim; no character substitutions are made and all commands are ignored, until the next @end verbatim. The text is printed in a fixed-width font, and not indented or filled. Extra spaces and blank lines are significant, and tabs are expanded.

@smallexample

Same as @example, except that in TEX this command typesets text in a smaller font.

@lisp Like @example, but specifically for illustrating Lisp code. The text is printed in a fixed-width font, and indented but not filled.

@smalllisp
          Is to @lisp as @smallexample is to @example.

@display Display illustrative text. The text is indented but not filled, and no font is selected (so, by default, the font is roman).

@smalldisplay
          Is to @display as @smallexample is to @example.

@format Like @display (the text is not filled and no font is selected), but the text is not indented.

@smallformat
          Is to @format as @smallexample is to @example.

The @exdent command is used within the above constructs to undo the indentation of a line.

The @flushleft and @flushright commands are used to line up the left or right margins of unfilled text.

The @noindent command may be used after one of the above constructs to prevent the following text from being indented as a new paragraph.

You can use the @cartouche command within one of the above constructs to highlight the example or quotation by drawing a box with rounded corners around it. See Section 10.13 [Drawing Cartouches Around Examples], page 87.

10.2 @quotation

The text of a quotation is processed normally except that:

• the margins are closer to the center of the page, so the whole of the quotation is indented;

• the first lines of paragraphs are indented no more than other lines;

• in the printed output, interparagraph spacing is reduced.

This is an example of text written between an @quotation command and an @end quotation command. An @quotation command is most often used to indicate text that is excerpted from another (real or hypothetical) printed work.

Write an @quotation command as text on a line by itself. This line will disappear from the output. Mark the end of the quotation with a line beginning with and containing only @end quotation. The @end quotation line will likewise disappear from the output. Thus, the following,

@quotation
This is
a foo.
@end quotation

produces This is a foo.

10.3 @example: Example Text

The @example command is used to indicate an example that is not part of the running text, such as computer input or output.

This is an example of text written between an
@example command
and an @end example command.
The text is indented but not filled.
In the printed manual, the text is typeset in a
fixed-width font, and extra spaces and blank lines are
significant.  In the Info file, an analogous result is
obtained by indenting each line with five spaces.

Write an @example command at the beginning of a line by itself. Mark the end of the example with an @end example command, also written at the beginning of a line by itself.

For example,

     @example
     mv foo bar
     @end example

produces

     mv foo bar

The lines containing @example and @end example will disappear from the output. To make the output look good, you should put a blank line before the @example and another blank line after the @end example. Note that blank lines inside the beginning @example and the ending @end example will appear in the output.

Caution: Do not use tabs in the lines of an example or anywhere else in Texinfo (except in verbatim environments)! The TEX implementation of Texinfo treats tabs as single spaces, and that is not what they look like. (If necessary, in Emacs, you can use M-x untabify to convert tabs in a region to multiple spaces.) Examples are often, logically speaking, “in the middle” of a paragraph, and the text

that continues after an example should not be indented. The @noindent command prevents a piece of text from being indented as if it were a new paragraph.

(The @code command is used for examples of code that are embedded within sentences, not set off from preceding and following text. See Section 9.1.1 [@code], page 72.)

10.4 @verbatim: Literal Text

Use the @verbatim environment for printing of text that may contain special characters or commands that should not be interpreted, such as computer input or output (@example interprets its text as regular Texinfo commands). This is especially useful for including automatically generated output in a Texinfo manual. Here is an example; the output you see is just the same as the input, with a line @verbatim before and a line @end verbatim after.

This is an example of text written in a @verbatim
block.  No character substitutions are made.  All commands
are ignored, until `<at>end verbatim'.
In the printed manual, the text is typeset in a
fixed-width font, and not indented or filled.  All
spaces and blank lines are significant, including tabs.

Write a @verbatim command at the beginning of a line by itself. This line will disappear from the output. Mark the end of the verbatim block with a @end verbatim command, also written at the beginning of a line by itself. The @end verbatim will also disappear from the output.

For example: @verbatim

{

<tab>@command with strange characters: @’e expand<tab>me

}

@end verbatim produces

{
        @command with strange characters: @'e

expand  me

}

Since the lines containing @verbatim and @end verbatim produce no output, tyically you should put a blank line before the @verbatim and another blank line after the @end verbatim. Blank lines between the beginning @verbatim and the ending @end verbatim will appear in the output.

10.5 @verbatiminclude file: Include a File Verbatim

You can include the exact contents of a file in the document with the @verbatiminclude command:

     @verbatiminclude filename

The contents of filename is printed in a verbatim environment (see Section 10.4 [@verbatim], page 83). Generally, the file is printed exactly as it is, with all special characters and white space retained.

10.6 @lisp: Marking a Lisp Example

The @lisp command is used for Lisp code. It is synonymous with the @example command.

     This is an example of text written between an
     @lisp command and an @end lisp command.

Use @lisp instead of @example to preserve information regarding the nature of the example. This is useful, for example, if you write a function that evaluates only and all the Lisp code in a Texinfo file. Then you can use the Texinfo file as a Lisp library.^{1}

Mark the end of @lisp with @end lisp on a line by itself.

10.7 @small... Block Commands

In addition to the regular @example and @lisp commands, Texinfo has “small”

example-style commands. These are @smalldisplay, @smallexample, @smallformat, and
@smalllisp.

In TEX, the @small... commands typeset text in a smaller font than the non-small example commands. Consequently, many examples containing long lines fit on a page without needing to be shortened.

1 It would be straightforward to extend Texinfo to work in a similar fashion for C, Fortran, or other

languages.

In Info, the @small... commands are equivalent to their non-small companion commands.

Mark the end of an @small... block with a corresponding @end small.... For exam-

ple, pair @smallexample with @end smallexample.

Here is an example written in the small font used by the @smallexample and

@smalllisp commands:
      . . . to make sure that you have the freedom to
      distribute copies of free software (and charge for
      this service if you wish), that you receive source
      code or can get it if you want it, that you can
      change the software or use pieces of it in new free
      programs; and that you know you can do these things.

The @small... commands make it easier to prepare manuals without forcing you to edit examples by hand to fit them onto narrower pages.

As a general rule, a printed document looks better if you use only one of (for example) @example or in @smallexample consistently within a chapter. Only occasionally should you mix the two formats.

See Section 19.11 [Printing “Small” Books], page 150, for more information about the

@smallbook command.

10.8 @display and @smalldisplay

The @display command begins a kind of example. It is like the @example command except that, in a printed manual, @display does not select the fixed-width font. In fact, it does not specify the font at all, so that the text appears in the same font it would have appeared in without the @display command.

This is an example of text written between an @display command and an @end display command. The @display command indents the text, but does not fill it.

Texinfo also provides a command @smalldisplay, which is like @display but uses a smaller font in @smallbook format. See Section 10.7 [small], page 84.

10.9 @format and @smallformat

The @format command is similar to @example except that, in the printed manual, @format does not select the fixed-width font and does not narrow the margins.

This is an example of text written between an @format command and an @end format command. As you can see from this example, the @format command does not fill the text.

Texinfo also provides a command @smallformat, which is like @format but uses a smaller font in @smallbook format. See Section 10.7 [small], page 84.

10.10 @exdent: Undoing a Line’s Indentation

The @exdent command removes any indentation a line might have. The command is written at the beginning of a line and applies only to the text that follows the command that is on the same line. Do not use braces around the text. In a printed manual, the text on an @exdent line is printed in the roman font.

@exdent is usually used within examples. Thus,

@example
This line follows an @@example command.
@exdent This line is exdented.
This line follows the exdented line.
The @@end example comes on the next line.
@end group

produces

This line follows an @example command.

This line is exdented.

This line follows the exdented line.
The @end example comes on the next line.

In practice, the @exdent command is rarely used. Usually, you un-indent text by ending the example and returning the page to its normal width.

10.11 @flushleft and @flushright

The @flushleft and @flushright commands line up the ends of lines on the left and right margins of a page, but do not fill the text. The commands are written on lines of their own, without braces. The @flushleft and @flushright commands are ended by @end flushleft and @end flushright commands on lines of their own.

For example,

@flushleft
This text is
written flushleft.
@end flushleft

produces This text is written flushleft. @flushright produces the type of indentation often used in the return address of

letters. For example,

@flushright
Here is an example of text written
flushright.  The @code{@flushright} command
right justifies every line but leaves the
left end ragged.
@end flushright

produces

Here is an example of text written flushright. The @flushright command right justifies every line but leaves the

left end ragged.

10.12 @noindent: Omitting Indentation

An example or other inclusion can break a paragraph into segments. Ordinarily, the formatters indent text that follows an example as a new paragraph. However, you can prevent this by writing @noindent at the beginning of a line by itself preceding the continuation text.

For example:

     @example
     This is an example
     @end example
     @noindent
     This line is not indented.  As you can see, the
     beginning of the line is fully flush left with the line
     that follows after it.  (This whole example is between
     @code{@@display} and @code{@@end display}.)

produces

          This is an example

This line is not indented. As you can see, the beginning of the line is fully flush left with the line that follows after it. (This whole example is between

     @display and @end display.)

To adjust the number of blank lines properly in the Info file output, remember that the line containing @noindent does not generate a blank line, and neither does the @end

example line.

In the Texinfo source file for this manual, each line that says ‘produces’ is preceded by a line containing @noindent.

Do not put braces after an @noindent command; they are not necessary, since @noindent is a command used outside of paragraphs (see Appendix H [Command Syntax], page 213).

10.13 @cartouche: Rounded Rectangles Around Examples

In a printed manual, the @cartouche command draws a box with rounded corners around its contents. You can use this command to further highlight an example or quotation. For instance, you could write a manual in which one type of example is surrounded by a cartouche for emphasis.

@cartouche affects only the printed manual; it has no effect in other output files.

For example,

@example
@cartouche
% pwd
/usr/local/share/emacs
@end cartouche
@end example

surrounds the two-line example with a box with rounded corners, in the printed manual. In a printed manual, the example looks like this:

 % pwd
 /usr/local/lib/emacs/info

11 Lists and Tables

Texinfo has several ways of making lists and tables. Lists can be bulleted or numbered; two-column tables can highlight the items in the first column; multi-column tables are also supported.

Texinfo automatically indents the text in lists or tables, and numbers an enumerated list. This last feature is useful if you modify the list, since you do not need to renumber it yourself.

Numbered lists and tables begin with the appropriate @-command at the beginning of a line, and end with the corresponding @end command on a line by itself. The table and itemized-list commands also require that you write formatting information on the same line as the beginning @-command.

Begin an enumerated list, for example, with an @enumerate command and end the list with an @end enumerate command. Begin an itemized list with an @itemize command, followed on the same line by a formatting command such as @bullet, and end the list with

an @end itemize command.

Precede each element of a list with an @item or @itemx command.

Here is an itemized list of the different kinds of table and lists:

• Itemized lists with and without bullets.

• Enumerated lists, using numbers or letters.

• Two-column tables with highlighting.

Here is an enumerated list with the same items:

1. Itemized lists with and without bullets.

2. Enumerated lists, using numbers or letters.

3. Two-column tables with highlighting.

And here is a two-column table with the same items and their @-commands:

@itemize Itemized lists with and without bullets.

@enumerate

Enumerated lists, using numbers or letters.

@table
@ftable

@vtable Two-column tables, optionally with indexing.

11.1 @itemize: Making an Itemized List

The @itemize command produces sequences of indented paragraphs, with a bullet or other mark inside the left margin at the beginning of each paragraph for which such a mark is desired.

Begin an itemized list by writing @itemize at the beginning of a line. Follow the command, on the same line, with a character or a Texinfo command that generates a mark. Usually, you will write @bullet after @itemize, but you can use @minus, or any command or character that results in a single character in the Info file. If you don’t want any mark at all, use @w. (When you write the mark command such as @bullet after an @itemize command, you may omit the ‘{}’.) If you don’t specify a mark command, the default is

@bullet.

Write the text of the indented paragraphs themselves after the @itemize, up to another line that says @end itemize.

Before each paragraph for which a mark in the margin is desired, write a line that says just @item. It is ok to have text following the @item.

Usually, you should put a blank line before an @item. This puts a blank line in the Info file. (TEX inserts the proper interline whitespace in either case.) Except when the entries are very brief, these blank lines make the list look better.

Here is an example of the use of @itemize, followed by the output it produces. @bullet produces an ‘*’ in Info and a round dot in TEX.

@itemize @bullet
@item
Some text for foo.
@item
Some text
for bar.
@end itemize

This produces:

• Some text for foo.

• Some text for bar.

Itemized lists may be embedded within other itemized lists. Here is a list marked with dashes embedded in a list marked with bullets:

@itemize @bullet
@item
First item.
@itemize @minus
@item
Inner item.
@item
Second inner item.
@end itemize
@item
Second outer item.
@end itemize

This produces:

• First item.

− Inner item.

− Second inner item.

• Second outer item.

11.2 @enumerate: Making a Numbered or Lettered List

@enumerate is like @itemize (see Section 11.1 [@itemize], page 89), except that the labels on the items are successive integers or letters instead of bullets.

Write the @enumerate command at the beginning of a line. The command does not require an argument, but accepts either a number or a letter as an option. Without an argument, @enumerate starts the list with the number ‘1’. With a numeric argument, such as ‘3’, the command starts the list with that number. With an upper or lower case letter, such as ‘a’ or ‘A’, the command starts the list with that letter.

Write the text of the enumerated list in the same way you write an itemized list: put @item on a line of its own before the start of each paragraph that you want enumerated. Do not write any other text on the line beginning with @item.

You should put a blank line between entries in the list. This generally makes it easier to read the Info file.

Here is an example of @enumerate without an argument:

@enumerate
@item
Underlying causes.
@item
Proximate causes.
@end enumerate

This produces:

1. Underlying causes.

2. Proximate causes.

Here is an example with an argument of 3:

@enumerate 3
@item
Predisposing causes.
@item
Precipitating causes.
@item
Perpetuating causes.
@end enumerate

This produces:

3. Predisposing causes.

4. Precipitating causes.

5. Perpetuating causes.

Here is a brief summary of the alternatives. The summary is constructed using @enumerate with an argument of a.

 a. @enumerate

Without an argument, produce a numbered list, starting with the number 1. b. @enumerate positive-integer

With a (positive) numeric argument, start a numbered list with that number. You can use this to continue a list that you interrupted with other text.

c. @enumerate upper-case-letter With an upper case letter as argument, start a list in which each item is marked by a letter, beginning with that upper case letter.

d. @enumerate lower-case-letter With a lower case letter as argument, start a list in which each item is marked by a letter, beginning with that lower case letter.

You can also nest enumerated lists, as in an outline.

11.3 Making a Two-column Table

@table is similar to @itemize (see Section 11.1 [@itemize], page 89), but allows you to specify a name or heading line for each item. The @table command is used to produce twocolumn tables, and is especially useful for glossaries, explanatory exhibits, and commandline option summaries.

Write the @table command at the beginning of a line and follow it on the same line with an argument that is a Texinfo “indicating” command such as @code, @samp, @var, or @kbd (see Section 9.1 [Indicating], page 71). Although these commands are usually followed by arguments in braces, in this case you use the command name without an argument because @item will supply the argument. This command will be applied to the text that goes into the first column of each item and determines how it will be highlighted. For example, @code will cause the text in the first column to be highlighted with an @code command. (We recommend @code for @table’s of command-line options.)

You may also choose to use the @asis command as an argument to @table. @asis is a command that does nothing; if you use this command after @table, TEX and the Info formatting commands output the first column entries without added highlighting (“as is”).

(The @table command may work with other commands besides those listed here. However, you can only use commands that normally take arguments in braces.)

Begin each table entry with an @item command at the beginning of a line. Write the first column text on the same line as the @item command. Write the second column text on the line following the @item line and on subsequent lines. (You do not need to type anything for an empty second column entry.) You may write as many lines of supporting text as you wish, even several paragraphs. But only text on the same line as the @item will be placed in the first column, including any footnote.

Normally, you should put a blank line before an @item line. This puts a blank like in the Info file. Except when the entries are very brief, a blank line looks better.

The following table, for example, highlights the text in the first column with an @samp command:

     @table @samp
     @item foo
     This is the text for
     @samp{foo}.
     @item bar
     Text for @samp{bar}.
     @end table

This produces:

‘foo’ This is the text for ‘foo’.

‘bar’ Text for ‘bar’.

If you want to list two or more named items with a single block of text, use the @itemx command. (See Section 11.3.2 [@itemx], page 93.)

11.3.1 @ftable and @vtable

The @ftable and @vtable commands are the same as the @table command except that @ftable automatically enters each of the items in the first column of the table into the index of functions and @vtable automatically enters each of the items in the first column of the table into the index of variables. This simplifies the task of creating indices. Only the items on the same line as the @item commands are indexed, and they are indexed in exactly the form that they appear on that line. See Chapter 12 [Indices], page 96, for more information about indices.

Begin a two-column table using @ftable or @vtable by writing the @-command at the beginning of a line, followed on the same line by an argument that is a Texinfo command such as @code, exactly as you would for an @table command; and end the table with an @end ftable or @end vtable command on a line by itself.

See the example for @table in the previous section.

11.3.2 @itemx

Use the @itemx command inside a table when you have two or more first column entries for the same item, each of which should appear on a line of its own. Use @itemx for all but the first entry; @itemx should always follow an @item command. The @itemx command works exactly like @item except that it does not generate extra vertical space above the first column text.

For example,

     @table @code
     @item upcase
     @itemx downcase
     These two functions accept a character or a string as
     argument, and return the corresponding upper case (lower
     case) character or string.
     @end table

This produces:

upcase

downcase These two functions accept a character or a string as argument, and return the corresponding upper case (lower case) character or string.

(Note also that this example illustrates multi-line supporting text in a two-column table.)

11.4 Multi-column Tables

@multitable allows you to construct tables with any number of columns, with each column having any width you like.

You define the column widths on the @multitable line itself, and write each row of the actual table following an @item command, with columns separated by an @tab command. Finally, @end multitable completes the table. Details in the sections below.

11.4.1 Multitable Column Widths

You can define the column widths for a multitable in two ways: as fractions of the line length; or with a prototype row. Mixing the two methods is not supported. In either case, the widths are defined entirely on the same line as the @multitable command.

1. To specify column widths as fractions of the line length, write @columnfractions and

the decimal numbers (presumably less than 1) after the @multitable command, as in:

         @multitable @columnfractions .33 .33 .33

The fractions need not add up exactly to 1.0, as these do not. This allows you to produce tables that do not need the full line length. You can use a leading zero if you wish.

2. To specify a prototype row, write the longest entry for each column enclosed in braces

after the @multitable command. For example:

         @multitable {some text for column one} {for column two}

The first column will then have the width of the typeset ‘some text for column one’, and the second column the width of ‘for column two’. The prototype entries need not appear in the table itself. Although we used simple text in this example, the prototype entries can contain Texinfo commands; markup commands such as @code are particularly likely to be useful.

11.4.2 Multitable Rows

After the @multitable command defining the column widths (see the previous section), you begin each row in the body of a multitable with @item, and separate the column entries with @tab. Line breaks are not special within the table body, and you may break input lines in your source file as necessary.

Here is a complete example of a multi-column table (the text is from The GNU Emacs Manual, see section “Splitting Windows” in The GNU Emacs Manual):

@multitable @columnfractions .15 .45 .4
@item Key @tab Command @tab Description
@item C-x 2
@tab @code{split-window-vertically}
@tab Split the selected window into two windows,
with one above the other.
@item C-x 3
@tab @code{split-window-horizontally}
@tab Split the selected window into two windows
positioned side by side.
@item C-Mouse-2
@tab
@tab In the mode line or scroll bar of a window,
split that window.
@end multitable

produces:

Key Command Description

C-x 2

        split-window-vertically

Split the selected window into two windows, with one above the other.

C-x 3

        split-window-horizontally

Split the selected window into two windows positioned side by side.

C-Mouse-2 In the mode line or scroll bar of a window, split that window.

12 Indices

Using Texinfo, you can generate indices without having to sort and collate entries manually. In an index, the entries are listed in alphabetical order, together with information on how to find the discussion of each entry. In a printed manual, this information consists of page numbers. In an Info file, this information is a menu entry leading to the first node referenced.

Texinfo provides several predefined kinds of index: an index for functions, an index for variables, an index for concepts, and so on. You can combine indices or use them for other than their canonical purpose. If you wish, you can define your own indices.

12.1 Making Index Entries

When you are making index entries, it is good practice to think of the different ways people may look for something. Different people do not think of the same words when they look something up. A helpful index will have items indexed under all the different words that people may use. For example, one reader may think it obvious that the twoletter names for indices should be listed under “Indices, two-letter names”, since the word “Index” is the general concept. But another reader may remember the specific concept of two-letter names and search for the entry listed as “Two letter names for indices”. A good index will have both entries and will help both readers.

Like typesetting, the construction of an index is a highly skilled, professional art, the subtleties of which are not appreciated until you need to do it yourself.

See Section 4.1 [Printing Indices & Menus], page 41, for information about printing an index at the end of a book or creating an index menu in an Info file.

12.2 Predefined Indices

Texinfo provides six predefined indices:

• A concept index listing concepts that are discussed.

• A function index listing functions (such as entry points of libraries).

• A variables index listing variables (such as global variables of libraries).

• A keystroke index listing keyboard commands.

• A program index listing names of programs.

• A data type index listing data types (such as structures defined in header files).

Not every manual needs all of these, and most manuals use two or three of them. This manual has two indices: a concept index and an @-command index (that is actually the function index but is called a command index in the chapter heading). Two or more indices can be combined into one using the @synindex or @syncodeindex commands. See Section 12.4 [Combining Indices], page 98.

12.3 Defining the Entries of an Index

The data to make an index come from many individual indexing commands scattered throughout the Texinfo source file. Each command says to add one entry to a particular index; after formatting, the index will give the current page number or node name as the reference.

An index entry consists of an indexing command at the beginning of a line followed, on the rest of the line, by the entry.

For example, this section begins with the following five entries for the concept index:

     @cindex Defining indexing entries
     @cindex Index entries
     @cindex Entries for an index
     @cindex Specifying index entries
     @cindex Creating index entries

Each predefined index has its own indexing command—@cindex for the concept index, @findex for the function index, and so on.

Concept index entries consist of text. The best way to write an index is to choose entries that are terse yet clear. If you can do this, the index often looks better if the entries are not capitalized, but written just as they would appear in the middle of a sentence. (Capitalize proper names and acronyms that always call for upper case letters.) This is the case convention we use in most GNU manuals’ indices.

If you don’t see how to make an entry terse yet clear, make it longer and clear—not terse and confusing. If many of the entries are several words long, the index may look better if you use a different convention: to capitalize the first word of each entry. But do not capitalize a case-sensitive name such as a C or Lisp function name or a shell command; that would be a spelling error.

Whichever case convention you use, please use it consistently!

Entries in indices other than the concept index are symbol names in programming languages, or program names; these names are usually case-sensitive, so use upper and lower case as required for them.

By default, entries for a concept index are printed in a small roman font and entries for the other indices are printed in a small @code font. You may change the way part of an entry is printed with the usual Texinfo commands, such as @file for file names and @emph for emphasis (see Chapter 9 [Marking Text], page 71).

The six indexing commands for predefined indices are:

@cindex concept

Make an entry in the concept index for concept.

@findex function Make an entry in the function index for function.

@vindex variable Make an entry in the variable index for variable.

@kindex keystroke Make an entry in the key index for keystroke.

@pindex program

Make an entry in the program index for program.

@tindex data type Make an entry in the data type index for data type.

Caution: Do not use a colon in an index entry. In Info, a colon separates the menu entry name from the node name, so a colon in the entry itself confuses Info. See Section 7.2 [The Parts of a Menu], page 58, for more information about the structure of a menu entry. You are not actually required to use the predefined indices for their canonical purposes.

For example, suppose you wish to index some C preprocessor macros. You could put them in the function index along with actual functions, just by writing @findex commands for them; then, when you print the “Function Index” as an unnumbered chapter, you could give it the title ‘Function and Macro Index’ and all will be consistent for the reader. Or you could put the macros in with the data types by writing @tindex commands for them, and give that index a suitable title so the reader will understand. (See Section 4.1 [Printing Indices & Menus], page 41.)

12.4 Combining Indices

Sometimes you will want to combine two disparate indices such as functions and concepts, perhaps because you have few enough of one of them that a separate index for them would look silly.

You could put functions into the concept index by writing @cindex commands for them instead of @findex commands, and produce a consistent manual by printing the concept index with the title ‘Function and Concept Index’ and not printing the ‘Function Index’ at all; but this is not a robust procedure. It works only if your document is never included as part of another document that is designed to have a separate function index; if your document were to be included with such a document, the functions from your document and those from the other would not end up together. Also, to make your function names appear in the right font in the concept index, you would need to enclose every one of them between the braces of @code.

12.4.1 @syncodeindex

When you want to combine functions and concepts into one index, you should index the functions with @findex and index the concepts with @cindex, and use the @syncodeindex command to redirect the function index entries into the concept index.

The @syncodeindex command takes two arguments; they are the name of the index to redirect, and the name of the index to redirect it to. The template looks like this:

     @syncodeindex from to

For this purpose, the indices are given two-letter names:

‘cp’ concept index

‘fn’ function index

‘vr’ variable index ‘ky’ key index

‘pg’ program index

‘tp’ data type index

Write an @syncodeindex command before or shortly after the end-of-header line at the beginning of a Texinfo file. For example, to merge a function index with a concept index, write the following:

     @syncodeindex fn cp

This will cause all entries designated for the function index to merge in with the concept index instead.

To merge both a variables index and a function index into a concept index, write the following:

     @syncodeindex vr cp
     @syncodeindex fn cp

The @syncodeindex command puts all the entries from the ‘from’ index (the redirected index) into the @code font, overriding whatever default font is used by the index to which the entries are now directed. This way, if you direct function names from a function index into a concept index, all the function names are printed in the @code font as you would expect.

12.4.2 @synindex

The @synindex command is nearly the same as the @syncodeindex command, except that it does not put the ‘from’ index entries into the @code font; rather it puts them in the roman font. Thus, you use @synindex when you merge a concept index into a function index.

See Section 4.1 [Printing Indices & Menus], page 41, for information about printing an index at the end of a book or creating an index menu in an Info file.

12.5 Defining New Indices

In addition to the predefined indices, you may use the @defindex and @defcodeindex commands to define new indices. These commands create new indexing @-commands with which you mark index entries. The @defindex command is used like this:

     @defindex name

The name of an index should be a two letter word, such as ‘au’. For example:

     @defindex au

This defines a new index, called the ‘au’ index. At the same time, it creates a new indexing command, @auindex, that you can use to make index entries. Use the new indexing command just as you would use a predefined indexing command.

For example, here is a section heading followed by a concept index entry and two ‘au’ index entries.

     @section Cognitive Semantics
     @cindex kinesthetic image schemas
     @auindex Johnson, Mark
     @auindex Lakoff, George

(Evidently, ‘au’ serves here as an abbreviation for “author”.) Texinfo constructs the new indexing command by concatenating the name of the index with ‘index’; thus, defining an ‘au’ index leads to the automatic creation of an @auindex command.

Use the @printindex command to print the index, as you do with the predefined indices. For example:

     @node Author Index, Subject Index, , Top
     @unnumbered Author Index
     @printindex au

The @defcodeindex is like the @defindex command, except that, in the printed output, it prints entries in an @code font instead of a roman font. Thus, it parallels the @findex command rather than the @cindex command.

You should define new indices within or right after the end-of-header line of a Texinfo file, before any @synindex or @syncodeindex commands (see Section 3.2 [Texinfo File Header], page 27).

13 Special Insertions

Texinfo provides several commands for inserting characters that have special meaning in Texinfo, such as braces, and for other graphic elements that do not correspond to simple characters you can type.

These are:

• Braces and ‘@’.

• Whitespace within and around a sentence.

• Accents.

• Dots and bullets.

• The TEX logo and the copyright symbol.

• The pounds currency symbol.

• The minus sign.

• Mathematical expressions.

• Glyphs for evaluation, macros, errors, etc.

• Footnotes.

• Images.

13.1 Inserting @ and Braces

‘@’ and curly braces are special characters in Texinfo. To insert these characters so they appear in text, you must put an ‘@’ in front of these characters to prevent Texinfo from misinterpreting them.

Do not put braces after any of these commands; they are not necessary.

13.1.1 Inserting ‘@’ with @@

@@ stands for a single ‘@’ in either printed or Info output.

Do not put braces after an @@ command.

13.1.2 Inserting ‘{’ and ‘}’with @{ and @}

@{ stands for a single ‘{’ in either printed or Info output.

@} stands for a single ‘}’ in either printed or Info output.

Do not put braces after either an @{ or an @} command.

13.2 Inserting Space

The following sections describe commands that control spacing of various kinds within and after sentences.

13.2.1 Not Ending a Sentence

Depending on whether a period or exclamation point or question mark is inside or at the end of a sentence, less or more space is inserted after a period in a typeset manual. Since it is not always possible to determine when a period ends a sentence and when it is used in an abbreviation, special commands are needed in some circumstances. Usually, Texinfo can guess how to handle periods, so you do not need to use the special commands; you just enter a period as you would if you were using a typewriter, which means you put two spaces after the period, question mark, or exclamation mark that ends a sentence.

Use the @: command after a period, question mark, exclamation mark, or colon that should not be followed by extra space. For example, use @: after periods that end abbreviations which are not at the ends of sentences.

For example,

     The s.o.p.@: has three parts ...
     The s.o.p. has three parts ...

produces the following. If you look carefully at this printed output, you will see a little more whitespace after ‘s.o.p.’ in the second line.

The s.o.p. has three parts … The s.o.p. has three parts …

(Incidentally, ‘s.o.p.’ is an abbreviation for “Standard Operating Procedure”.) @: has no effect on the Info output. Do not put braces after @:.

13.2.2 Ending a Sentence

Use @. instead of a period, @! instead of an exclamation point, and @? instead of a question mark at the end of a sentence that ends with a single capital letter. Otherwise, TEX will think the letter is an abbreviation and will not insert the correct end-of-sentence spacing. Here is an example:

     Give it to M.I.B. and to M.E.W@.  Also, give it to R.J.C@.
     Give it to M.I.B. and to M.E.W.  Also, give it to R.J.C.

produces the following. If you look carefully at this printed output, you will see a little more whitespace after the ‘W’ in the first line.

Give it to M.I.B. and to M.E.W. Also, give it to R.J.C. Give it to M.I.B. and to M.E.W. Also, give it to R.J.C. In the Info file output, @. is equivalent to a simple ‘.’; likewise for @! and @?. The meanings of @: and @. in Texinfo are designed to work well with the Emacs

sentence motion commands (see section “Sentences” in The GNU Emacs Manual). Do not put braces after any of these commands.

13.2.3 Multiple Spaces

Ordinarily, TEX collapses multiple whitespace characters (space, tab, and newline) into a single space. Info output, on the other hand, preserves whitespace as you type it, except for changing a newline into a space; this is why it is important to put two spaces at the end of sentences in Texinfo documents.

Occasionally, you may want to actually insert several consecutive spaces, either for purposes of example (what your program does with multiple spaces as input), or merely for purposes of appearance in headings or lists. Texinfo supports three commands: @SPACE, @TAB, and @NL, all of which insert a single space into the output. (Here, @SPACE represents an ‘@’ character followed by a space, i.e., ‘@ ’, and TAB and NL represent the tab character and end-of-line, i.e., when ‘@’ is the last character on a line.)

For example,

     Spacey@ @ @ @
     example.

produces

     Spacey    example.

Other possible uses of @SPACE have been subsumed by @multitable (see Section 11.4 [Multi-column Tables], page 94).

Do not follow any of these commands with braces.

To produce a non-breakable space, see Section 14.3 [w], page 114.

13.2.4 @dmn{dimension}: Format a Dimension

At times, you may want to write ‘12pt’ or ‘8.5in’ with little or no space between the number and the abbreviation for the dimension. You can use the @dmn command to do this. On seeing the command, TEX inserts just enough space for proper typesetting; the Info formatting commands insert no space at all, since the Info file does not require it.

To use the @dmn command, write the number and then follow it immediately, with no intervening space, by @dmn, and then by the dimension within braces. For example,

     A4 paper is 8.27@dmn{in} wide.

produces A4 paper is 8.27 in wide. Not everyone uses this style. Some people prefer ‘8.27 in.@:’ or ‘8.27 inches’ to

‘8.27@dmn{in}’ in the Texinfo file. In these cases, however, the formatters may insert a line break between the number and the dimension, so use @w (see Section 14.3 [w], page 114). Also, if you write a period after an abbreviation within a sentence, you should write ‘@:’ after the period to prevent TEX from inserting extra whitespace, as shown here. See Section 13.2.1 [Not Ending a Sentence], page 102.

13.3 Inserting Accents

Here is a table with the commands Texinfo provides for inserting floating accents. The commands with non-alphabetic names do not take braces around their argument (which is taken to be the next character). (Exception: @, does take braces around its argument.) This is so as to make the source as convenient to type and read as possible, since accented characters are very common in some languages.

Command  Output What
@"o  ö  umlaut accent
@'o  ó  acute accent

@,{c}  ç  cedilla accent
@=o  ō  macron/overbar accent
@^o  ô  circumflex accent
@`o  ò  grave accent
@~o  õ  tilde accent
@dotaccent{o} ȯ  overdot accent
@H{o}  ő  long Hungarian umlaut

@ringaccent{o} o̊  ring accent

@tieaccent{oo} oo  tie-after accent
@u{o} ŏ breve accent
@ubaraccent{o}  ō  underbar accent
@udotaccent{o}  o.  underdot accent
@v{o} ǒ hacek or check accent

This table lists the Texinfo commands for inserting other characters commonly used in languages other than English.

@exclamdown{}  ¡  upside-down !
@questiondown{} ¿  upside-down ?

@aa{},@AA{}  å,Å  a,A with circle
@ae{},@AE{}  æ,Æ ae,AE ligatures

@dotless{i}  ı  dotless i
@dotless{j}  ȷ  dotless j

@l{},@L{}  ł,Ł  suppressed-L,l
@o{},@O{}  ø,Ø  O,o with slash
@oe{},@OE{}  œ,Œ oe,OE ligatures
@ss{}  ß  es-zet or sharp S

13.4 Inserting Ellipsis and Bullets

An ellipsis (a line of dots) is not typeset as a string of periods, so a special command is used for ellipsis in Texinfo. The @bullet command is special, too. Each of these commands is followed by a pair of braces, ‘{}’, without any whitespace between the name of the command and the braces. (You need to use braces with these commands because you can use them next to other text; without the braces, the formatters would be confused. See Appendix H [@-Command Syntax], page 213, for further information.)

13.4.1 @dots{} (…) and @enddots{} (…. )

Use the @dots{} command to generate an ellipsis, which is three dots in a row, appropriately spaced, like this: ‘…’. Do not simply write three periods in the input file; that would work for the Info file output, but would produce the wrong amount of space between the periods in the printed manual.

Similarly, the @enddots{} command generates an end-of-sentence ellipsis (four dots) ….

Here is an ellipsis: … Here are three periods in a row: ...

In printed output, the three periods in a row are closer together than the dots in the ellipsis.

13.4.2 @bullet{} (•)

Use the @bullet{} command to generate a large round dot, or the closest possible thing to one. In Info, an asterisk is used.

Here is a bullet: •

When you use @bullet in @itemize, you do not need to type the braces, because @itemize supplies them. (See Section 11.1 [@itemize], page 89.)

13.5 Inserting TEX and the Copyright Symbol

The logo ‘TEX’ is typeset in a special fashion and it needs an @-command. The copyright symbol, ‘\begin{matrix}c \\ ◯ \end{matrix}’, is also special. Each of these commands is followed by a pair of braces, ‘\{\}’, without any whitespace between the name of the command and the braces.

13.5.1 @TeX{} (TEX)

Use the @TeX{} command to generate ‘TEX’. In a printed manual, this is a special logo that is different from three ordinary letters. In Info, it just looks like ‘TeX’. The @TeX{} command is unique among Texinfo commands in that the ‘T’ and the ‘X’ are in upper case.

13.5.2 @copyright\{\} (\begin{matrix}c \\ ◯ \end{matrix})

Use the @copyright\{\} command to generate ‘\begin{matrix}c \\ ◯ \end{matrix}’. In a printed manual, this is a ‘c’ inside a circle, and in Info, this is ‘(C)’.

13.6 @pounds{} (£): Pounds Sterling

Use the @pounds{} command to generate ‘£’. In a printed manual, this is the symbol for the currency pounds sterling. In Info, it is a ‘#’. Other currency symbols are unfortunately not available.

13.7 @minus{} (−): Inserting a Minus Sign

Use the @minus{} command to generate a minus sign. In a fixed-width font, this is a single hyphen, but in a proportional font, the symbol is the customary length for a minus sign—a little longer than a hyphen, shorter than an em-dash:

‘−’ is a minus sign generated with ‘@minus{}’,

‘-’ is a hyphen generated with the character ‘-’,

‘—’ is an em-dash for text. In the fixed-width font used by Info, @minus{} is the same as a hyphen.

You should not use @minus{} inside @code or @example because the width distinction is not made in the fixed-width font they use.

When you use @minus to specify the mark beginning each entry in an itemized list, you do not need to type the braces (see Section 11.1 [@itemize], page 89.)

13.8 @math: Inserting Mathematical Expressions

You can write a short mathematical expression with the @math command. Write the mathematical expression between braces, like this:

     @math{(a + b)(a + b) = a^2 + 2ab + b^2}

This produces the following in TEX:

(a + b)(a + b) = a^{2} + 2ab + b^{2} and the following in Info:

     (a + b)(a + b) = a^2 + 2ab + b^2

Thus, the @math command has no effect on the Info output.

@math implies @tex. This not only makes it possible to write superscripts and subscripts (as in the above example), but also allows you to use any of the plain TEX math control sequences. It’s conventional to use ‘\’ instead of ‘@’ for these commands. As in:

     @math{\sin 2\pi \equiv \cos 3\pi}

which looks like this in TEX:

sin 2π ≡ cos 3π and which looks like the input in Info and HTML:

     \sin 2\pi \equiv \cos 3\pi

Since ‘\’ is an escape character inside @math, you can use @\ to get a literal backslash (\\ will work in TEX, but you’ll get the literal ‘\\’ in Info). @\ is not defined outside of @math, since a ‘\’ ordinarily produces a literal ‘\’.

For displayed equations, you must at present use TEX directly (see Section 16.3 [Raw Formatter Commands], page 130).

13.9 Glyphs for Examples

In Texinfo, code is often illustrated in examples that are delimited by @example and @end example, or by @lisp and @end lisp. In such examples, you can indicate the results of evaluation or an expansion using ‘⇒’ or ‘↦’. Likewise, there are commands to insert glyphs to indicate printed output, error messages, equivalence of expressions, and the location of point.

The glyph-insertion commands do not need to be used within an example, but most often they are. Every glyph-insertion command is followed by a pair of left- and right-hand braces.

13.9.1 Glyphs Summary

Here are the different glyph commands:

⇒ @result{} points to the result of an expression.

↦ @expansion{} shows the results of a macro expansion.

⊣ @print{} indicates printed output.

error @error{} indicates that the following text is an error message.

≡ @equiv{} indicates the exact equivalence of two forms.

⋆ @point{} shows the location of point.

13.9.2 @result{} (⇒): Indicating Evaluation

Use the @result{} command to indicate the result of evaluating an expression.

The @result{} command is displayed as ‘=>’ in Info and as ‘⇒’ in the printed output.

Thus, the following,

     (cdr '(1 2 3))
          ⇒ (2 3)

may be read as “(cdr '(1 2 3)) evaluates to (2 3)”.

13.9.3 @expansion{} (↦): Indicating an Expansion

When an expression is a macro call, it expands into a new expression. You can indicate the result of the expansion with the @expansion{} command.

The @expansion{} command is displayed as ‘==>’ in Info and as ‘↦’ in the printed output.

For example, the following

     @lisp
     (third '(a b c))
          @expansion{} (car (cdr (cdr '(a b c))))
          @result{} c
     @end lisp

produces

     (third '(a b c))
          ↦ (car (cdr (cdr '(a b c))))
          ⇒ c

which may be read as:

     (third '(a b c)) expands to (car (cdr (cdr '(a b c)))); the result of eval-

uating the expression is c.

Often, as in this case, an example looks better if the @expansion{} and @result{} commands are indented five spaces.

13.9.4 @print{} ( ⊣ ): Indicating Printed Output

Sometimes an expression will print output during its execution. You can indicate the printed output with the @print{} command.

The @print{} command is displayed as ‘-|’ in Info and as ‘ ⊣ ’ in the printed output.

In the following example, the printed text is indicated with ‘ ⊣ ’, and the value of the expression follows on the last line.

     (progn (print 'foo) (print 'bar))
          ⊣ foo

          ⊣ bar
          ⇒ bar

In a Texinfo source file, this example is written as follows:

@lisp
(progn (print 'foo) (print 'bar))
     @print{} foo
     @print{} bar
     @result{} bar
@end lisp

13.9.5 @error{} ( error ): Indicating an Error Message

A piece of code may cause an error when you evaluate it. You can designate the error message with the @error{} command.

The @error{} command is displayed as ‘error-->’ in Info and as ‘ error ’ in the printed output.

Thus,

@lisp
(+ 23 'x)
@error{} Wrong type argument: integer-or-marker-p, x
@end lisp

produces

(+ 23 'x)
 error Wrong type argument: integer-or-marker-p, x

This indicates that the following error message is printed when you evaluate the expression:

Wrong type argument: integer-or-marker-p, x

‘ error ’ itself is not part of the error message.

13.9.6 @equiv{} ( ≡): Indicating Equivalence

Sometimes two expressions produce identical results. You can indicate the exact equivalence of two forms with the @equiv{} command.

The @equiv{} command is displayed as ‘==’ in Info and as ‘ ≡ ’ in the printed output. Thus,

@lisp
(make-sparse-keymap) @equiv{} (list 'keymap)
@end lisp

produces

(make-sparse-keymap) ≡ (list 'keymap)

This indicates that evaluating (make-sparse-keymap) produces identical results to evalu-

ating (list 'keymap).

13.9.7 @point{} (⋆): Indicating Point in a Buffer

Sometimes you need to show an example of text in an Emacs buffer. In such examples, the convention is to include the entire contents of the buffer in question between two lines of dashes containing the buffer name.

You can use the ‘@point{}’ command to show the location of point in the text in the buffer. (The symbol for point, of course, is not part of the text in the buffer; it indicates the place between two characters where point is located.)

The @point{} command is displayed as ‘-!-’ in Info and as ‘⋆’ in the printed output. The following example shows the contents of buffer ‘foo’ before and after evaluating a

Lisp command to insert the word changed.

---------- Buffer: foo ----------
This is the ⋆contents of foo.
---------- Buffer: foo ----------
(insert "changed ")
     ⇒ nil
---------- Buffer: foo ----------
This is the changed ⋆contents of foo.
---------- Buffer: foo ----------

In a Texinfo source file, the example is written like this:

@example
---------- Buffer: foo ----------
This is the @point{}contents of foo.
---------- Buffer: foo ----------
(insert "changed ")
     @result{} nil
---------- Buffer: foo ----------
This is the changed @point{}contents of foo.
---------- Buffer: foo ----------
@end example

13.10 Footnotes

A footnote is for a reference that documents or elucidates the primary text.^{1}

13.10.1 Footnote Commands

In Texinfo, footnotes are created with the @footnote command. This command is followed immediately by a left brace, then by the text of the footnote, and then by a terminating right brace. Footnotes may be of any length (they will be broken across pages if necessary), but are usually short. The template is:

ordinary text@footnote{text of footnote}

As shown here, the @footnote command should come right after the text being footnoted, with no intervening space; otherwise, the footnote marker might end up starting a line.

1 A footnote should complement or expand upon the primary text, but a reader should not need to read

a footnote to understand the primary text. For a thorough discussion of footnotes, see The Chicago Manual of Style, which is published by the University of Chicago Press.

For example, this clause is followed by a sample footnote^{2}; in the Texinfo source, it looks like this:

 ...a sample footnote@footnote{Here is the sample
 footnote.}; in the Texinfo source...

As you can see, the source includes two punctuation marks next to each other; in this case, ‘.};’ is the sequence. This is normal (the first ends the footnote and the second belongs to the sentence being footnoted), so don’t worry that it looks odd.

In a printed manual or book, the reference mark for a footnote is a small, superscripted number; the text of the footnote appears at the bottom of the page, below a horizontal line.

In Info, the reference mark for a footnote is a pair of parentheses with the footnote number between them, like this: ‘(1)’. The reference mark is followed by a cross-reference link to the footnote’s text.

In the HTML output, footnote references are marked with a small, superscripted number which is rendered as a hypertext link to the footnote text.

By the way, footnotes in the argument of an @item command for a @table must be on the same line as the @item (as usual). See Section 11.3 [Two-column Tables], page 92.

13.10.2 Footnote Styles

Info has two footnote styles, which determine where the text of the footnote is located:

• In the ‘End’ node style, all the footnotes for a single node are placed at the end of that node. The footnotes are separated from the rest of the node by a line of dashes with the word ‘Footnotes’ within it. Each footnote begins with an ‘(n)’ reference mark. Here is an example of a single footnote in the end of node style:

      --------- Footnotes ---------
     (1)  Here is a sample footnote.

• In the ‘Separate’ node style, all the footnotes for a single node are placed in an automatically constructed node of their own. In this style, a “footnote reference” follows each ‘(n)’ reference mark in the body of the node. The footnote reference is actually a cross reference which you use to reach the footnote node. The name of the node with the footnotes is constructed by appending ‘-Footnotes’ to the name of the node that contains the footnotes. (Consequently, the footnotes’ node for the ‘Footnotes’ node is ‘Footnotes-Footnotes’!) The footnotes’ node has an ‘Up’ node pointer that leads back to its parent node. Here is how the first footnote in this manual looks after being formatted for Info in the separate node style:

     File: texinfo.info  Node: Overview-Footnotes, Up: Overview
     (1) The first syllable of "Texinfo" is pronounced like "speck", not
     "hex". ...

2 Here is the sample footnote.

A Texinfo file may be formatted into an Info file with either footnote style.

Use the @footnotestyle command to specify an Info file’s footnote style. Write this command at the beginning of a line followed by an argument, either ‘end’ for the end node style or ‘separate’ for the separate node style.

For example,

     @footnotestyle end

or

     @footnotestyle separate

Write an @footnotestyle command before or shortly after the end-of-header line at the beginning of a Texinfo file. (If you include the @footnotestyle command between the start-of-header and end-of-header lines, the region formatting commands will format footnotes as specified.)

If you do not specify a footnote style, the formatting commands use their default style.

Currently, texinfo-format-buffer and texinfo-format-region use the ‘separate’ style

and makeinfo uses the ‘end’ style.

13.11 Inserting Images

You can insert an image given in an external file with the @image command:

@image{filename, [width], [height], [alttext], [extension]} The filename argument is mandatory, and must not have an extension, because the

different processors support different formats:

• TEX reads the file ‘filename.eps’ (Encapsulated PostScript format).

• PDFTEX reads ‘filename.pdf’ (Adobe’s Portable Document Format).

• makeinfo uses ‘filename.txt’ verbatim for Info output (more or less as if it was an

    @example).

• makeinfo uses the optional fifth argument to @image for the extension if you supply

it. For example:

         @image{foo,,,,xpm}
    will cause ‘makeinfo --html’ to try ‘foo.xpm’.

If you do not supply the optional fifth argument, ‘makeinfo ---html’ first tries ‘filename.png’; if that does not exist, it tries ‘filename.jpg’. If that does not exist either, it complains. (We cannot support GIF format directly due to software patents.)

The optional width and height arguments specify the size to scale the image to (they are ignored for Info output). If neither is specified, the image is presented in its natural size (given in the file); if only one is specified, the other is scaled proportionately; and if both are specified, both are respected, thus possibly distorting the original image by changing its aspect ratio.

The width and height may be specified using any valid TEX dimension, namely:

pt point (72.27pt = 1in)

pc pica (1pc = 12pt) bp big point (72bp = 1in)

in inch

cm centimeter (2.54cm = 1in)

mm millimeter (10mm = 1cm)

dd didôt point (1157dd = 1238pt)

cc cicero (1cc = 12dd)

sp scaled point (65536sp = 1pt)

For example, the following will scale a file ‘ridt.eps’ to one inch vertically, with the width scaled proportionately:

     @image{ridt,,1in}

For @image to work with TEX, the file ‘epsf.tex’ must be installed somewhere that TEX can find it. (The standard location is ‘texmf /tex/generic/dvips/epsf.tex’, where texmf is a root of your TEX directory tree.) This file is included in the Texinfo distribution and is also available from ftp://tug.org/tex/epsf.tex, among other places.

@image can be used within a line as well as for displayed figures. Therefore, if you intend it to be displayed, be sure to leave a blank line before the command, or the output will run into the preceding text.

When producing html, makeinfo sets the alt attribute for inline images to the optional fourth argument to @image, if supplied. If not supplied, makeinfo uses the full file name of the image being displayed.

14 Making and Preventing Breaks

Usually, a Texinfo file is processed both by TEX and by one of the Info formatting commands. Line, paragraph, or page breaks sometimes occur in the ‘wrong’ place in one or other form of output. You must ensure that text looks right both in the printed manual and in the Info file.

For example, in a printed manual, page breaks may occur awkwardly in the middle of an example; to prevent this, you can hold text together using a grouping command that keeps the text from being split across two pages. Conversely, you may want to force a page break where none would occur normally. Fortunately, problems like these do not often arise. When they do, use the break, break prevention, or pagination commands.

The break commands create or allow line and paragraph breaks:

@* Force a line break.

@sp n Skip n blank lines.

@- Insert a discretionary hyphen.

@hyphenation{hy-phen-a-ted words} Define hyphen points in hy-phen-a-ted words.

The line-break-prevention command holds text together all on one line:

@w{text} Prevent text from being split and hyphenated across two lines.

The pagination commands apply only to printed output, since Info files do not have pages.

@page Start a new page in the printed manual.

@group Hold text together that must appear on one printed page.

@need mils

Start a new printed page if not enough space on this one.

14.1 @*: Generate Line Breaks

The @* command forces a line break in both the printed manual and in Info.

For example,

     This line @* is broken @*in two places.

produces

     This line
      is broken
     in two places.

(Note that the space after the first @* command is faithfully carried down to the next line.) The @* command is often used in a file’s copyright page:

     This is edition 2.0 of the Texinfo documentation,@*
     and is for ...

In this case, the @* command keeps TEX from stretching the line across the whole page in an ugly manner.

Please note: Do not write braces after an @* command; they are not needed. Do not write an @refill command at the end of a paragraph containing an @* command; it will cause the paragraph to be refilled after the line break occurs, negating the effect of the line break.

14.2 @- and @hyphenation: Helping TEX hyphenate

Although TEX’s hyphenation algorithm is generally pretty good, it does miss useful hyphenation points from time to time. (Or, far more rarely, insert an incorrect hyphenation.) So, for documents with an unusual vocabulary or when fine-tuning for a printed edition, you may wish to help TEX out. Texinfo supports two commands for this:

@- Insert a discretionary hyphen, i.e., a place where TEX can (but does not have to) hyphenate. This is especially useful when you notice an overfull hbox is due to TEX missing a hyphenation (see Section 19.10 [Overfull hboxes], page 149). TEX will not insert any hyphenation points itself into a word containing @-.

@hyphenation{hy-phen-a-ted words} Tell TEX how to hyphenate hy-phen-a-ted words. As shown, you put a ‘-’ at each hyphenation point. For example:

               @hyphenation{man-u-script man-u-scripts}

TEX only uses the specified hyphenation points when the words match exactly, so give all necessary variants.

Info output is not hyphenated, so these commands have no effect there.

14.3 @w{text}: Prevent Line Breaks

@w{text} outputs text and prohibits line breaks within text.

You can use the @w command to prevent TEX from automatically hyphenating a long name or phrase that happens to fall near the end of a line. For example:

     You can copy GNU software from @w{@samp{ftp.gnu.org}}.

produces You can copy GNU software from ‘ftp.gnu.org’. You can also use @w to produce a non-breakable space:

     None of the formatters will break at this@w{ }space.

14.4 @sp n: Insert Blank Lines

A line beginning with and containing only @sp n generates n blank lines of space in both the printed manual and the Info file. @sp also forces a paragraph break. For example,

     @sp 2

generates two blank lines.

The @sp command is most often used in the title page.

14.5 @page: Start a New Page

A line containing only @page starts a new page in a printed manual. The command has no effect on Info files since they are not paginated. An @page command is often used in the @titlepage section of a Texinfo file to start the copyright page.

14.6 @group: Prevent Page Breaks

The @group command (on a line by itself) is used inside an @example or similar construct to begin an unsplittable vertical group, which will appear entirely on one page in the printed output. The group is terminated by a line containing only @end group. These two lines produce no output of their own, and in the Info file output they have no effect at all.

Although @group would make sense conceptually in a wide variety of contexts, its current implementation works reliably only within @example and variants, and within

@display, @format, @flushleft and @flushright. See Chapter 10 [Quotations and Ex-

amples], page 81. (What all these commands have in common is that each line of input produces a line of output.) In other contexts, @group can cause anomalous vertical spacing.

This formatting requirement means that you should write:

     @example
     @group
     ...
     @end group
     @end example
with the @group and @end group commands inside the @example and @end example com-

mands.

The @group command is most often used to hold an example together on one page. In this Texinfo manual, more than 100 examples contain text that is enclosed between @group

and @end group.

If you forget to end a group, you may get strange and unfathomable error messages when you run TEX. This is because TEX keeps trying to put the rest of the Texinfo file onto the one page and does not start to generate error messages until it has processed considerable text. It is a good rule of thumb to look for a missing @end group if you get incomprehensible error messages in TEX.

14.7 @need mils: Prevent Page Breaks

A line containing only @need n starts a new page in a printed manual if fewer than n mils (thousandths of an inch) remain on the current page. Do not use braces around the argument n. The @need command has no effect on Info files since they are not paginated.

This paragraph is preceded by an @need command that tells TEX to start a new page if fewer than 800 mils (eight-tenths inch) remain on the page. It looks like this:

     @need 800
     This paragraph is preceded by ...

The @need command is useful for preventing orphans (single lines at the bottoms of printed pages).

15 Definition Commands

The @deffn command and the other definition commands enable you to describe functions, variables, macros, commands, user options, special forms and other such artifacts in a uniform format.

In the Info file, a definition causes the entity category—‘Function’, ‘Variable’, or whatever—to appear at the beginning of the first line of the definition, followed by the entity’s name and arguments. In the printed manual, the command causes TEX to print the entity’s name and its arguments on the left margin and print the category next to the right margin. In both output formats, the body of the definition is indented. Also, the name of the entity is entered into the appropriate index: @deffn enters the name into the index of functions, @defvr enters it into the index of variables, and so on.

A manual need not and should not contain more than one definition for a given name. An appendix containing a summary should use @table rather than the definition commands.

15.1 The Template for a Definition

The @deffn command is used for definitions of entities that resemble functions. To write a definition using the @deffn command, write the @deffn command at the beginning of a line and follow it on the same line by the category of the entity, the name of the entity itself, and its arguments (if any). Then write the body of the definition on succeeding lines. (You may embed examples in the body.) Finally, end the definition with an @end deffn command written on a line of its own. (The other definition commands follow the same format.)

The template for a definition looks like this: @deffn category name arguments... body-of-definition

     @end deffn

For example,

     @deffn Command forward-word count
     This command moves point forward @var{count} words
     (or backward if @var{count} is negative). ...
     @end deffn

produces

Command forward-word count

This function moves point forward count words (or backward if count is negative)….

Capitalize the category name like a title. If the name of the category contains spaces, as in the phrase ‘Interactive Command’, write braces around it. For example:

     @deffn {Interactive Command} isearch-forward
     ...
     @end deffn

Otherwise, the second word will be mistaken for the name of the entity.

Some of the definition commands are more general than others. The @deffn command, for example, is the general definition command for functions and the like—for entities that may take arguments. When you use this command, you specify the category to which the entity belongs. The @deffn command possesses three predefined, specialized variations, @defun, @defmac, and @defspec, that specify the category for you: “Function”, “Macro”, and “Special Form” respectively. (In Lisp, a special form is an entity much like a function.) The @defvr command also is accompanied by several predefined, specialized variations for describing particular kinds of variables.

The template for a specialized definition, such as @defun, is similar to the template for a generalized definition, except that you do not need to specify the category:

@defun name arguments...

body-of-definition

     @end defun

Thus,

     @defun buffer-end flag
     This function returns @code{(point-min)} if @var{flag}
     is less than 1, @code{(point-max)} otherwise.
     ...
     @end defun

produces

Function buffer-end flag

This function returns (point-min) if flag is less than 1, (point-max) otherwise. …

See Section 15.6 [A Sample Function Definition], page 126, for a more detailed example of a function definition, including the use of @example inside the definition.

The other specialized commands work like @defun.

Note that, due to implementation difficulties, macros are not expanded in @deffn and all the other definition commands.

15.2 Optional and Repeated Arguments

Some entities take optional or repeated arguments, which may be specified by a distinctive glyph that uses square brackets and ellipses. For example, a special form often breaks its argument list into separate arguments in more complicated ways than a straightforward function.

An argument enclosed within square brackets is optional. Thus, the phrase ‘[optionalarg]’ means that optional-arg is optional. An argument followed by an ellipsis is optional and may be repeated more than once. Thus, ‘repeated-args...’ stands for zero or more arguments. Parentheses are used when several arguments are grouped into additional levels of list structure in Lisp.

Here is the @defspec line of an example of an imaginary special form:

Special Form foobar (var [from to [inc]]) body …

In this example, the arguments from and to are optional, but must both be present or both absent. If they are present, inc may optionally be specified as well. These arguments are grouped with the argument var into a list, to distinguish them from body, which includes all remaining elements of the form.

In a Texinfo source file, this @defspec line is written like this (except it would not be split over two lines, as it is in this example).

     @defspec foobar (@var{var} [@var{from} @var{to}
          [@var{inc}]]) @var{body}@dots{}

The function is listed in the Command and Variable Index under ‘foobar’.

15.3 Two or More ‘First’ Lines

To create two or more ‘first’ or header lines for a definition, follow the first @deffn line by a line beginning with @deffnx. The @deffnx command works exactly like @deffn except that it does not generate extra vertical white space between it and the preceding line.

For example,

     @deffn {Interactive Command} isearch-forward
     @deffnx {Interactive Command} isearch-backward
     These two search commands are similar except ...
     @end deffn

produces

Interactive Command isearch-forward Interactive Command isearch-backward

These two search commands are similar except …

Each definition command has an ‘x’ form: @defunx, @defvrx, @deftypefunx, etc.

The ‘x’ forms work just like @itemx; see Section 11.3.2 [@itemx], page 93.

15.4 The Definition Commands

Texinfo provides more than a dozen definition commands, all of which are described in this section.

The definition commands automatically enter the name of the entity in the appropriate index: for example, @deffn, @defun, and @defmac enter function names in the index of functions; @defvr and @defvar enter variable names in the index of variables.

Although the examples that follow mostly illustrate Lisp, the commands can be used for other programming languages.

15.4.1 Functions and Similar Entities

This section describes the commands for describing functions and similar entities:

@deffn category name arguments...

The @deffn command is the general definition command for functions, interactive commands, and similar entities that may take arguments. You must choose a term to describe the category of entity being defined; for example, “Function” could be used if the entity is a function. The @deffn command is written at the beginning of a line and is followed on the same line by the category of entity being described, the name of this particular entity, and its arguments, if any. Terminate the definition with @end deffn on a line of its own. For example, here is a definition:

     @deffn Command forward-char nchars
     Move point forward @var{nchars} characters.
     @end deffn

This shows a rather terse definition for a “command” named forward-char with one argument, nchars. @deffn prints argument names such as nchars in italics or upper case, as if @var had been used, because we think of these names as metasyntactic variables— they stand for the actual argument values. Within the text of the description, write an argument name explicitly with @var to refer to the value of the argument. In the example above, we used ‘@var{nchars}’ in this way. The template for @deffn is:

@deffn category name arguments...

body-of-definition

     @end deffn

@defun name arguments...

The @defun command is the definition command for functions. @defun is equiv-

alent to ‘@deffn Function ...’.

For example,

     @defun set symbol new-value
     Change the value of the symbol @var{symbol}
     to @var{new-value}.
     @end defun

shows a rather terse definition for a function set whose arguments are symbol and new-value. The argument names on the @defun line automatically appear in italics or upper case as if they were enclosed in @var. Terminate the definition with @end defun on a line of its own. The template is:

@defun function-name arguments...

body-of-definition

     @end defun

@defun creates an entry in the index of functions.

@defmac name arguments...

The @defmac command is the definition command for macros. @defmac is equivalent to ‘@deffn Macro ...’ and works like @defun.

@defspec name arguments...

The @defspec command is the definition command for special forms. (In Lisp, a special form is an entity much like a function, see section “Special Forms” in GNU Emacs Lisp Reference Manual.) @defspec is equivalent to ‘@deffn

{Special Form} ...’ and works like @defun.

15.4.2 Variables and Similar Entities

Here are the commands for defining variables and similar entities:

@defvr category name The @defvr command is a general definition command for something like a variable—an entity that records a value. You must choose a term to describe the category of entity being defined; for example, “Variable” could be used if the entity is a variable. Write the @defvr command at the beginning of a line and follow it on the same line by the category of the entity and the name of the entity. Capitalize the category name like a title. If the name of the category contains spaces, as in the name “User Option”, enclose it in braces. Otherwise, the second word will be mistaken for the name of the entity. For example,

     @defvr {User Option} fill-column
     This buffer-local variable specifies
     the maximum width of filled lines.
     ...
     @end defvr

Terminate the definition with @end defvr on a line of its own. The template is:

@defvr category name

body-of-definition

     @end defvr

@defvr creates an entry in the index of variables for name.

@defvar name

The @defvar command is the definition command for variables. @defvar is

equivalent to ‘@defvr Variable ...’.

For example:

     @defvar kill-ring
     ...
     @end defvar

The template is:

     @defvar name

body-of-definition

     @end defvar

@defvar creates an entry in the index of variables for name.

@defopt name

The @defopt command is the definition command for user options, i.e., variables intended for users to change according to taste; Emacs has many such (see section “Variables” in The GNU Emacs Manual). @defopt is equivalent to

‘@defvr {User Option} ...’ and works like @defvar.

15.4.3 Functions in Typed Languages

The @deftypefn command and its variations are for describing functions in languages in which you must declare types of variables and functions, such as C and C++.

@deftypefn category data-type name arguments...

The @deftypefn command is the general definition command for functions and similar entities that may take arguments and that are typed. The @deftypefn command is written at the beginning of a line and is followed on the same line by the category of entity being described, the type of the returned value, the name of this particular entity, and its arguments, if any. For example,

     @deftypefn {Library Function} int foobar
        (int @var{foo}, float @var{bar})
     ...
     @end deftypefn

(where the text before the “…”, shown above as two lines, would actually be a single line in a real Texinfo file) produces the following in Info:

     -- Library Function: int foobar (int FOO, float BAR)
     ...

In a printed manual, it produces:

Library Function int foobar (int foo, float bar)

…

This means that foobar is a “library function” that returns an int, and its arguments are foo (an int) and bar (a float). The argument names that you write in @deftypefn are not subject to an implicit @var—since the actual names of the arguments in @deftypefn are typically scattered among data type names and keywords, Texinfo cannot find them without help. Instead, you must write @var explicitly around the argument names. In the example above, the argument names are ‘foo’ and ‘bar’. The template for @deftypefn is:

@deftypefn category data-type name arguments ...

body-of-description

     @end deftypefn

Note that if the category or data type is more than one word then it must be enclosed in braces to make it a single argument. If you are describing a procedure in a language that has packages, such as Ada, you might consider using @deftypefn in a manner somewhat contrary to the convention described in the preceding paragraphs.

For example:

     @deftypefn stacks private push
             (@var{s}:in out stack;
             @var{n}:in integer)
     ...
     @end deftypefn

(The @deftypefn arguments are shown split into three lines, but would be a single line in a real Texinfo file.) In this instance, the procedure is classified as belonging to the package stacks rather than classified as a ‘procedure’ and its data type is described as private. (The name of the procedure is push, and its arguments are s and n.) @deftypefn creates an entry in the index of functions for name.

@deftypefun data-type name arguments...

The @deftypefun command is the specialized definition command for functions in typed languages. The command is equivalent to ‘@deftypefn Function ...’. Thus,

     @deftypefun int foobar (int @var{foo}, float @var{bar})
     ...
     @end deftypefun

produces the following in Info:

     -- Function: int foobar (int FOO, float BAR)
     ...

and the following in a printed manual:

Function int foobar (int foo, float bar)

…

The template is:

@deftypefun type name arguments...

body-of-description

     @end deftypefun

@deftypefun creates an entry in the index of functions for name.

15.4.4 Variables in Typed Languages

Variables in typed languages are handled in a manner similar to functions in typed languages. See Section 15.4.3 [Typed Functions], page 121. The general definition command @deftypevr corresponds to @deftypefn and the specialized definition command

@deftypevar corresponds to @deftypefun.

@deftypevr category data-type name The @deftypevr command is the general definition command for something like a variable in a typed language—an entity that records a value. You must choose a term to describe the category of the entity being defined; for example, “Variable” could be used if the entity is a variable.

The @deftypevr command is written at the beginning of a line and is followed on the same line by the category of the entity being described, the data type, and the name of this particular entity. For example:

@deftypevr {Global Flag} int enable
...
@end deftypevr

produces the following in Info:

-- Global Flag: int enable
...

and the following in a printed manual:

Global Flag int enable

…

The template is:

@deftypevr category data-type name body-of-description

@end deftypevr

@deftypevr creates an entry in the index of variables for name.

@deftypevar data-type name The @deftypevar command is the specialized definition command for variables in typed languages. @deftypevar is equivalent to ‘@deftypevr Variable ...’. For example:

@deftypevar int fubar
...
@end deftypevar

produces the following in Info:

-- Variable: int fubar
...

and the following in a printed manual:

Variable int fubar

…

The template is:

@deftypevar data-type name body-of-description

@end deftypevar

@deftypevar creates an entry in the index of variables for name.

15.4.5 Object-Oriented Programming

Here are the commands for formatting descriptions about abstract objects, such as are used in object-oriented programming. A class is a defined type of abstract object. An instance of a class is a particular object that has the type of the class. An instance variable is a variable that belongs to the class but for which each instance has its own value.

In a definition, if the name of a class is truly a name defined in the programming system for a class, then you should write an @code around it. Otherwise, it is printed in the usual text font.

@defcv category class name The @defcv command is the general definition command for variables associated with classes in object-oriented programming. The @defcv command is followed by three arguments: the category of thing being defined, the class to which it belongs, and its name. Thus,

     @defcv {Class Option} Window border-pattern
     ...
     @end defcv

illustrates how you would write the first line of a definition of the borderpattern class option of the class Window. The template is:

@defcv category class name

     ...
     @end defcv

@defcv creates an entry in the index of variables.

@defivar class name The @defivar command is the definition command for instance variables in object-oriented programming. @defivar is equivalent to ‘@defcv {Instance

Variable} ...’

The template is:

@defivar class instance-variable-name

body-of-definition

     @end defivar

@defivar creates an entry in the index of variables.

@deftypeivar class data-type name The @deftypeivar command is the definition command for typed instance variables in object-oriented programming. It is similar to @defivar with the addition of the data-type parameter to specify the type of the instance variable. @deftypeivar creates an entry in the index of variables.

@defop category class name arguments...

The @defop command is the general definition command for entities that may resemble methods in object-oriented programming. These entities take arguments, as functions do, but are associated with particular classes of objects. For example, some systems have constructs called wrappers that are associated with classes as methods are, but that act more like macros than like functions. You could use @defop Wrapper to describe one of these. Sometimes it is useful to distinguish methods and operations. You can think of an operation as the specification for a method. Thus, a window system might specify that all window classes have a method named expose; we would say that this window system defines an expose operation on windows in general. Typically, the operation has a name and also specifies the pattern of arguments; all methods that implement the operation must accept the same arguments, since applications that use the operation do so without knowing which method will implement it. Often it makes more sense to document operations than methods. For example, window application developers need to know about the expose operation, but need not be concerned with whether a given class of windows has its own method to implement this operation. To describe this operation, you would write:

     @defop Operation windows expose

The @defop command is written at the beginning of a line and is followed on the same line by the overall name of the category of operation, the name of the class of the operation, the name of the operation, and its arguments, if any. The template is:

@defop category class name arguments... body-of-definition

     @end defop

@defop creates an entry, such as ‘expose on windows’, in the index of functions.

@deftypeop category class data-type name arguments...

The @deftypeop command is the definition command for typed operations in object-oriented programming. It is similar to @defop with the addition of the data-type parameter to specify the return type of the method. @deftypeop creates an entry in the index of functions.

@defmethod class name arguments...

The @defmethod command is the definition command for methods in objectoriented programming. A method is a kind of function that implements an operation for a particular class of objects and its subclasses. @defmethod is equivalent to ‘@defop Method ...’. The command is written at the beginning of a line and is followed by the name of the class of the method, the name of the method, and its arguments, if any. For example:

     @defmethod bar-class bar-method argument
     ...
     @end defmethod

illustrates the definition for a method called bar-method of the class bar-class. The method takes an argument. The template is:

@defmethod class method-name arguments... body-of-definition

     @end defmethod

@defmethod creates an entry, such as ‘bar-method on bar-class’, in the index of functions.

@deftypemethod class data-type name arguments...

The @deftypemethod command is the definition command for methods in object-oriented typed languages, such as C++ and Java. It is similar to the @defmethod command with the addition of the data-type parameter to specify the return type of the method.

15.4.6 Data Types

Here is the command for data types:

@deftp category name attributes...

The @deftp command is the generic definition command for data types. The command is written at the beginning of a line and is followed on the same line by the category, by the name of the type (which is a word like int or float), and then by names of attributes of objects of that type. Thus, you could use this command for describing int or float, in which case you could use data type as the category. (A data type is a category of certain objects for purposes of deciding which operations can be performed on them.) In Lisp, for example, pair names a particular data type, and an object of that type has two slots called the car and the cdr. Here is how you would write the first line of a definition of pair.

     @deftp {Data type} pair car cdr
     ...
     @end deftp

The template is:

@deftp category name-of-type attributes... body-of-definition

     @end deftp

@deftp creates an entry in the index of data types.

15.5 Conventions for Writing Definitions

When you write a definition using @deffn, @defun, or one of the other definition commands, please take care to use arguments that indicate the meaning, as with the count argument to the forward-word function. Also, if the name of an argument contains the name of a type, such as integer, take care that the argument actually is of that type.

15.6 A Sample Function Definition

A function definition uses the @defun and @end defun commands. The name of the function follows immediately after the @defun command and it is followed, on the same line, by the parameter list.

Here is a definition from section “Calling Functions” in The GNU Emacs Lisp Reference Manual.

Function apply function &rest arguments apply calls function with arguments, just like funcall but with one difference: the last of arguments is a list of arguments to give to function, rather than a single argument. We also say that this list is appended to the other arguments. apply returns the result of calling function. As with funcall, function must either be a Lisp function or a primitive function; special forms and macros do not make sense in apply.

          (setq f 'list)
               ⇒ list
          (apply f 'x 'y 'z)
           error Wrong type argument: listp, z
          (apply '+ 1 2 '(3 4))
               ⇒ 10
          (apply '+ '(1 2 3 4))
               ⇒ 10
          (apply 'append '((a b c) nil (x y z) nil))
               ⇒ (a b c x y z)

An interesting example of using apply is found in the description of

     mapcar.

In the Texinfo source file, this example looks like this:

@defun apply function &rest arguments
@code{apply} calls @var{function} with
@var{arguments}, just like @code{funcall} but with one
difference: the last of @var{arguments} is a list of
arguments to give to @var{function}, rather than a single
argument.  We also say that this list is @dfn{appended}
to the other arguments.
@code{apply} returns the result of calling
@var{function}.  As with @code{funcall},
@var{function} must either be a Lisp function or a
primitive function; special forms and macros do not make
sense in @code{apply}.
@example
(setq f 'list)
     @result{} list
(apply f 'x 'y 'z)
@error{} Wrong type argument: listp, z
(apply '+ 1 2 '(3 4))
     @result{} 10
(apply '+ '(1 2 3 4))
     @result{} 10
(apply 'append '((a b c) nil (x y z) nil))
     @result{} (a b c x y z)
@end example

     An interesting example of using @code{apply} is found
     in the description of @code{mapcar}.
     @end defun

In this manual, this function is listed in the Command and Variable Index under apply.

Ordinary variables and user options are described using a format like that for functions except that variables do not take arguments.

16 Conditionally Visible Text

Sometimes it is good to use different text for different output formats. For example, you can use the conditional commands to specify different text for the printed manual and the Info output.

Conditional commands may not be nested.

The conditional commands comprise the following categories.

• Commands for HTML, Info, or TEX.

• Commands for not HTML, Info, or TEX.

• Raw TEX or HTML commands.

• Substituting text for all formats, and testing if a flag is set or clear.

16.1 Conditional Commands

Texinfo has a pair of commands for each output format, to allow conditional inclusion of text for a particular output format.

@ifinfo begins segments of text that should be ignored by TEX when it typesets the printed manual. The segment of text appears only in the Info file and (for historical compatibility) the plain text output. The @ifinfo command should appear on a line by itself; end the Info-only text with a line containing @end ifinfo by itself.

The @iftex and @end iftex commands are analogous to the @ifinfo and @end ifinfo commands; they specify text that will appear in the printed manual but not in the Info file. Likewise for @ifhtml and @end ifhtml, which specify text to appear only in HTML output. And for @ifplaintext and @end ifplaintext, which specify text to appear only in plain text output.

For example,

     @iftex
     This text will appear only in the printed manual.
     @end iftex
     @ifinfo
     However, this text will appear only in Info (or plain text).
     @end ifinfo
     @ifhtml
     And this text will only appear in HTML.
     @end ifhtml
     @ifplaintext
     Whereas this text will only appear in plain text.
     @end ifplaintext

The preceding example produces the following line: This text will appear only in the printed manual.

Notice that you only see one of the input lines, depending on which version of the manual you are reading.

16.2 Conditional Not Commands

You can specify text to be included in any output format other than some given one with the @ifnot... commands:

     @ifnothtml ... @end ifnothtml
     @ifnotinfo ... @end ifnotinfo
     @ifnotplaintext ... @end ifnotplaintext
     @ifnottex ... @end ifnottex

(The @ifnot... command and the @end command must appear on lines by themselves in your actual source file.)

If the output file is not being made for the given format, the region is included. Otherwise, it is ignored.

With one exception (for historical compatibility): @ifnotinfo text is omitted for both Info and plain text output, not just Info. To specify text which appears only in Info and not in plain text, use @ifnotplaintext, like this:

The regions delimited by these commands are ordinary Texinfo source as with @iftex, not raw formatter source as with @tex (see Section 16.3 [Raw Formatter Commands], page 130).

16.3 Raw Formatter Commands

Inside a region delineated by @iftex and @end iftex, you can embed some raw TEX commands. Info will ignore these commands since they are only in that part of the file which is seen by TEX. You can write the TEX commands as you would write them in a normal TEX file, except that you must replace the ‘\’ used by TEX with an ‘@’. For example, in the @titlepage section of a Texinfo file, you can use the TEX command @vskip to format the copyright page. (The @titlepage command causes Info to ignore the region automatically, as it does with the @iftex command.)

However, many features of plain TEX will not work, as they are overridden by Texinfo features.

You can enter plain TEX completely, and use ‘\’ in the TEX commands, by delineating a region with the @tex and @end tex commands. (The @tex command also causes Info to ignore the region, like the @iftex command.) The sole exception is that the @ character still introduces a command, so that @end tex can be recognized properly.

For example, here is a mathematical expression written in plain TEX:

     @tex
     $$ \chi^2 = \sum_{i=1}^N
               \left (y_i - (a + b x_i)
               \over \sigma_i\right)^2 $$
     @end tex

The output of this example will appear only in a printed manual. If you are reading this in Info, you will not see the equation that appears in the printed manual. In a printed manual, the above expression looks like this:

χ^{2} = ∑_{i=1}^{N} (\frac{y_{i} − (a + bx_{i})}{σ_{i}})^{2}

Analogously, you can use @ifhtml ... @end ifhtml to delimit a region to be included in HTML output only, and @html ... @end html for a region of raw HTML (again, except that @ is still the escape character, so the @end command can be recognized.)

16.4 @set, @clear, and @value

You can direct the Texinfo formatting commands to format or ignore parts of a Texinfo file with the @set, @clear, @ifset, and @ifclear commands.

Brief descriptions:

@set flag [value] Set the variable flag, to the optional value if specifed.

@clear flag

Undefine the variable flag, whether or not it was previously defined.

@ifset flag

If flag is set, text through the next @end ifset command is formatted. If flag is clear, text through the following @end ifset command is ignored.

@ifclear flag

If flag is set, text through the next @end ifclear command is ignored. If flag is clear, text through the following @end ifclear command is formatted.

16.4.1 @set and @value

You use the @set command to specify a value for a flag, which is later expanded by the @value command.

A flag is an identifier. In general, it is best to use only letters and numerals in a flag name, not ‘-’ or ‘_’—they will work in some contexts, but not all, due to limitations in TEX.

The value is the remainder of the input line, and can contain anything.

Write the @set command like this:

     @set foo This is a string.

This sets the value of the flag foo to “This is a string.”.

The Texinfo formatters then replace an @value{flag} command with the string to which flag is set. Thus, when foo is set as shown above, the Texinfo formatters convert this:

     @value{foo}

to this:

     This is a string.

You can write an @value command within a paragraph; but you must write an @set command on a line of its own.

If you write the @set command like this:

@set foo

without specifying a string, the value of foo is the empty string.

If you clear a previously set flag with @clear flag, a subsequent @value{flag} command will report an error.

For example, if you set foo as follows:

@set how-much very, very, very

then the formatters transform

It is a @value{how-much} wet day.

into

It is a very, very, very wet day.

If you write

@clear how-much

then the formatters transform

It is a @value{how-much} wet day.

into

It is a {No value for "how-much"} wet day.
16.4.2 @ifset and @ifclear

When a flag is set, the Texinfo formatting commands format text between subsequent pairs of @ifset flag and @end ifset commands. When the flag is cleared, the Texinfo formatting commands do not format the text. @ifclear operates analogously.

Write the conditionally formatted text between @ifset flag and @end ifset commands, like this:

@ifset flag

conditional-text

@end ifset

For example, you can create one document that has two variants, such as a manual for a ‘large’ and ‘small’ model:

You can use this machine to dig up shrubs
without hurting them.
@set large
@ifset large
It can also dig up fully grown trees.
@end ifset
Remember to replant promptly ...

In the example, the formatting commands will format the text between @ifset large and @end ifset because the large flag is set.

When flag is cleared, the Texinfo formatting commands do not format the text between @ifset flag and @end ifset; that text is ignored and does not appear in either printed or Info output.

For example, if you clear the flag of the preceding example by writing an @clear large command after the @set large command (but before the conditional text), then the Texinfo formatting commands ignore the text between the @ifset large and @end ifset commands. In the formatted output, that text does not appear; in both printed and Info output, you see only the lines that say, “You can use this machine to dig up shrubs without hurting them. Remember to replant promptly …”.

If a flag is cleared with an @clear flag command, then the formatting commands format text between subsequent pairs of @ifclear and @end ifclear commands. But if the flag is set with @set flag, then the formatting commands do not format text between an @ifclear and an @end ifclear command; rather, they ignore that text. An @ifclear command looks like this:

@ifclear flag

16.4.3 @value Example

You can use the @value command to minimize the number of places you need to change when you record an update to a manual. See Section C.2 [GNU Sample Texts], page 192, for an example of this same principle can work with Automake distributions, and full texts.

Here is an example adapted from section “Overview” in The GNU Make Manual):

1. Set the flags:

@set EDITION 0.35 Beta
@set VERSION 3.63 Beta
@set UPDATED 14 August 1992
@set UPDATE-MONTH August 1992

2. Write text for the @copying section (see Section 3.3.1 [copying], page 30):

@copying
This is Edition @value{EDITION},
last updated @value{UPDATED},
of @cite{The GNU Make Manual},
for @code{make}, version @value{VERSION}.
Copyright ...
Permission is granted ...
@end copying

3. Write text for the title page, for people reading the printed manual:

@titlepage
@title GNU Make
@subtitle A Program for Directing Recompilation
@subtitle Edition @value{EDITION}, ...
@subtitle @value{UPDATE-MONTH}
@page
@insertcopying
...
@end titlepage

(On a printed cover, a date listing the month and the year looks less fussy than a date listing the day as well as the month and year.)

4. Write text for the Top node, for people reading the Info file:

@ifnottex
@node Top
@top Make
@insertcopying
...
@end ifnottex

After you format the manual, the @value constructs have been expanded, so the output contains text like this:

This is Edition 0.35 Beta, last updated 14 August 1992,
of `The GNU Make Manual', for `make', Version 3.63 Beta.

When you update the manual, you change only the values of the flags; you do not need to edit the three sections.

17 Internationalization

Texinfo has some support for writing in languages other than English, although this area still needs considerable work.

For a list of the various accented and special characters Texinfo supports, see Section 13.3 [Inserting Accents], page 103.

17.1 @documentlanguage cc: Set the Document Language

The @documentlanguage command declares the current document language. Write it on a line by itself, with a two-letter ISO-639 language code following (list is included below). If you have a multilingual document, the intent is to be able to use this command multiple times, to declare each language change. If the command is not used at all, the default is en for English.

At present, this command is ignored in Info and HTML output. For TEX, it causes the file ‘txi-cc.tex’ to be read (if it exists). Such a file appropriately redefines the various English words used in TEX output, such as ‘Chapter’, ‘See’, and so on.

It would be good if this command also changed TEX’s ideas of the current hyphenation patterns (via the TEX primitive \language), but this is unfortunately not currently implemented.

Hereare the valid language codes, from ISO-639.

aa  Afar  ab Abkhazian  af Afrikaans
am  Amharic  ar Arabic  as Assamese
ay  Aymara  az Azerbaijani  ba Bashkir
be  Byelorussian  bg Bulgarian  bh Bihari
bi  Bislama  bn Bengali; Bangla  bo Tibetan
br  Breton  ca Catalan  co Corsican
cs  Czech  cy Welsh  da Danish
de  German  dz Bhutani  el Greek
en  English  eo Esperanto  es Spanish
et  Estonian  eu Basque  fa Persian
fi  Finnish  fj Fiji  fo Faroese
fr  French  fy Frisian  ga Irish
gd  Scots Gaelic  gl Galician  gn Guarani
gu  Gujarati  ha Hausa  he Hebrew
hi  Hindi  hr Croatian  hu Hungarian
hy  Armenian  ia Interlingua  id Indonesian
ie  Interlingue  ik Inupiak  is Icelandic
it  Italian  iu Inuktitut  ja Japanese
jw  Javanese  ka Georgian  kk Kazakh
kl  Greenlandic  km Cambodian  kn Kannada
ks  Kashmiri  ko Korean  ku Kurdish
ky  Kirghiz  la Latin  ln Lingala
lt  Lithuanian  lo Laothian  lv Latvian, Lettish
mg  Malagasy  mi Maori  mk Macedonian

ml  Malayalam  mn Mongolian  mo Moldavian
mr  Marathi  ms Malay  mt Maltese
my  Burmese  na Nauru  ne Nepali
nl  Dutch  no Norwegian  oc Occitan
om  (Afan) Oromo  or Oriya  pa Punjabi
pl  Polish  ps Pashto, Pushto  pt Portuguese
qu  Quechua  rm Rhaeto-Romance  rn Kirundi
ro  Romanian  ru Russian  rw Kinyarwanda
sa  Sanskrit  sd Sindhi  sg Sangro
sh  Serbo-Croatian  si Sinhalese  sk Slovak
sl  Slovenian  sm Samoan  sn Shona
so  Somali  sq Albanian  sr Serbian
ss  Siswati  st Sesotho  su Sundanese
sv  Swedish  sw Swahili  ta Tamil
te  Telugu  tg Tajik  th Thai
ti  Tigrinya  tk Turkmen  tl Tagalog
tn  Setswana  to Tonga  tr Turkish
ts  Tsonga  tt Tatar  tw Twi
ug  Uighur  uk Ukrainian  ur Urdu
uz  Uzbek  vi Vietnamese  vo Volapuk
wo  Wolof  xh Xhosa  yi Yiddish
yo  Yoruba  za Zhuang  zh Chinese
zu  Zulu

17.2 @documentencoding enc: Set Input Encoding

The @documentencoding command declares the input document encoding. Write it on a line by itself, with a valid encoding specification following, such as ‘ISO-8859-1’.

At present, this is used only in HTML output from makeinfo. If a document encoding enc is specified, it is used in a ‘<meta>’ tag included in the ‘<head>’ of the output:

     <meta http-equiv="Content-Type" content="text/html;
           charset=enc">

18 Defining New Texinfo Commands

Texinfo provides several ways to define new commands:

• A Texinfo macro allows you to define a new Texinfo command as any sequence of text and/or existing commands (including other macros). The macro can have any number of parameters—text you supply each time you use the macro. Incidentally, these macros have nothing to do with the @defmac command, which is for documenting macros in the subject of the manual (see Section 15.1 [Def Cmd Template], page 116).

• ‘@alias’ is a convenient way to define a new name for an existing command.

• ‘@definfoenclose’ allows you to define new commands with customized output in the Info file.

18.1 Defining Macros

You use the Texinfo @macro command to define a macro, like this:

@macro macroname{param1, param2, ...} text ... \param1\ ...

     @end macro

The parameters param1, param2, … correspond to arguments supplied when the macro is subsequently used in the document (described in the next section).

For a macro to work with TEX, macroname must consist entirely of letters: no digits, hyphens, underscores, or other special characters.

If a macro needs no parameters, you can define it either with an empty list (‘@macro foo {}’) or with no braces at all (‘@macro foo’).

The definition or body of the macro can contain most Texinfo commands, including previously-defined macros. Not-yet-defined macro invocations are not allowed; thus, it is not possible to have mutually recursive Texinfo macros. Also, a macro definition that defines another macro does not work in TEX due to limitations in the design of @macro.

In the macro body, instances of a parameter name surrounded by backslashes, as in ‘\param1\’ in the example above, are replaced by the corresponding argument from the macro invocation. You can use parameter names any number of times in the body, including zero.

To get a single ‘\’ in the macro expansion, use ‘\\’. Any other use of ‘\’ in the body yields a warning.

The newlines after the @macro line and before the @end macro line are ignored, that is, not included in the macro body. All other whitespace is treated according to the usual Texinfo rules.

To allow a macro to be used recursively, that is, in an argument to a call to itself, you must define it with ‘@rmacro’, like this:

     @rmacro rmac {arg}
     a\arg\b
     @end rmacro
     ...

@rmac{1@rmac{text}2}

This produces the output ‘a1atextb2b’. With ‘@macro’ instead of ‘@rmacro’, an error message is given.

You can undefine a macro foo with @unmacro foo. It is not an error to undefine a macro that is already undefined. For example:

@unmacro foo

18.2 Invoking Macros

After a macro is defined (see the previous section), you can use (invoke) it in your document like this:

@macroname {arg1, arg2, ...} and the result will be just as if you typed the body of macroname at that spot. For example:

@macro foo {p, q}
Together: \p\ & \q\.
@end macro
@foo{a, b}

produces:

Together: a & b. Thus, the arguments and parameters are separated by commas and delimited by braces;

any whitespace after (but not before) a comma is ignored. The braces are required in the invocation (but not the definition), even when the macro takes no arguments, consistent with all other Texinfo commands. For example:

@macro argless {}
No arguments here.
@end macro
@argless{}

produces:

No arguments here. To insert a comma, brace, or backslash in an argument, prepend a backslash, as in

@macname {\\\{\}\,}

which will pass the (almost certainly error-producing) argument ‘\{},’ to macname. However, commas in parameters, even if escaped by a backslash, might cause trouble in TEX.

If the macro is defined to take a single argument, and is invoked without any braces, the entire rest of the line after the macro name is supplied as the argument. For example:

@macro bar {p}
Twice: \p\ & \p\.
@end macro
@bar aah

produces:

Twice: aah & aah. If the macro is defined to take a single argument, and is invoked with braces, the braced

text is passed as the argument, regardless of commas. For example:

@macro bar {p}
Twice: \p\ & \p\.
@end macro
@bar{a,b}

produces:

Twice: a,b & a,b.

18.3 Macro Details

Due to unavoidable disparities in the TEX and makeinfo implementations, Texinfo macros have the following limitations.

• All macros are expanded inside at least one TEX group. This means that @set and other such commands will have no effect inside a macro.

• Macros containing a command which must be on a line by itself, such as a conditional, cannot be invoked in the middle of a line.

• Commas in macro arguments, even if escaped by a backslash, don’t always work.

• The TEX implementation cannot construct macros that define macros in the natural

way. To do this, you must use conditionals and raw TEX. For example:

@ifnottex
@macro ctor {name, arg}
@macro \name\
something involving \arg\ somehow
@end macro
@end macro
@end ifnottex
@tex
\gdef\ctor#1{\ctorx#1,}
\gdef\ctorx#1,#2,{\def#1{something involving #2 somehow}}
@end tex

• It is best to avoid comments inside macro definitions.

If some macro feature causes errors when producing the printed version of a manual, try expanding the macros with makeinfo by invoking texi2dvi with the ‘-e’ option; see Section 19.3 [Format with texi2dvi], page 144.

18.4 ‘@alias new=existing’

The ‘@alias’ command defines a new command to be just like an existing one. This is useful for defining additional markup names, thus preserving semantic information in the input even though the output result may be the same.

Write the ‘@alias’ command on a line by itself, followed by the new command name, an equals sign, and the existing command name. Whitespace around the equals sign is ignored. Thus:

@alias new = existing For example, if your document contains citations for both books and some other media

(movies, for example), you might like to define a macro @moviecite{} that does the same thing as an ordinary @cite{} but conveys the extra semantic information as well. You’d do this as follows:

     @alias moviecite = cite

Macros do not always have the same effect due to vagaries of argument parsing. Also, aliases are much simpler to define than macros. So the command is not redundant. (It was also heavily used in the Jargon File!)

Aliases must not be recursive, directly or indirectly.

18.5 ‘definfoenclose’: Customized Highlighting

A @definfoenclose command may be used to define a highlighting command for Info, but not for TEX. A command defined using @definfoenclose marks text by enclosing it in strings that precede and follow the text. You can use this to get closer control of your Info output.

Presumably, if you define a command with @definfoenclose for Info, you will create a corresponding command for TEX, either in ‘texinfo.tex’, ‘texinfo.cnf’, or within an ‘@iftex’ in your document.

Write a @definfoenclose command on a line and follow it with three arguments separated by commas. The first argument to @definfoenclose is the @-command name (without the @); the second argument is the Info start delimiter string; and the third argument is the Info end delimiter string. The latter two arguments enclose the highlighted text in the Info file. A delimiter string may contain spaces. Neither the start nor end delimiter is required. If you do not want a start delimiter but do want an end delimiter, you must follow the command name with two commas in a row; otherwise, the Info formatting commands will naturally misinterpret the end delimiter string you intended as the start delimiter string.

If you do a @definfoenclose on the name of a pre-defined macro (such as @emph, @strong, @t, or @i), the enclosure definition will override the built-in definition.

An enclosure command defined this way takes one argument in braces; this is intended for new markup commands (see Chapter 9 [Marking Text], page 71).

For example, you can write:

     @definfoenclose phoo,//,\\

near the beginning of a Texinfo file to define @phoo as an Info formatting command that inserts ‘//’ before and ‘\\’ after the argument to @phoo. You can then write @phoo{bar} wherever you want ‘//bar\\’ highlighted in Info.

Also, for TEX formatting, you could write

     @iftex
     @global@let@phoo=@i
     @end iftex

to define @phoo as a command that causes TEX to typeset the argument to @phoo in italics.

Each definition applies to its own formatter: one for TEX, the other for texinfo-

format-buffer or texinfo-format-region. The @definfoenclose command need not

be within ‘@ifinfo’, but the raw TEX commands do need to be in ‘@iftex’.

Here is another example: write

     @definfoenclose headword, , :

near the beginning of the file, to define @headword as an Info formatting command that inserts nothing before and a colon after the argument to @headword.

‘@definfoenclose’ definitions must not be recursive, directly or indirectly.

19 Formatting and Printing Hardcopy

There are three major shell commands for making a printed manual from a Texinfo file: one for converting the Texinfo file into a file that will be printed, a second for sorting indices, and a third for printing the formatted document. When you use the shell commands, you can either work directly in the operating system shell or work within a shell inside GNU Emacs.

If you are using GNU Emacs, you can use commands provided by Texinfo mode instead of shell commands. In addition to the three commands to format a file, sort the indices, and print the result, Texinfo mode offers key bindings for commands to recenter the output buffer, show the print queue, and delete a job from the print queue.

19.1 Use TEX

The typesetting program called TEX is used for formatting a Texinfo file. TEX is a very powerful typesetting program and, if used correctly, does an exceptionally good job. (See Appendix I [How to Obtain TEX], page 214, for information on how to obtain TEX.)

The standalone makeinfo program and Emacs functions texinfo-format-region and texinfo-format-buffer commands read the very same @-commands in the Texinfo file as does TEX, but process them differently to make an Info file (see Section 20.1 [Creating an Info File], page 153).

19.2 Format with tex and texindex

Format the Texinfo file with the shell command tex followed by the name of the Texinfo file. For example:

     tex foo.texi

TEX will produce a DVI file as well as several auxiliary files containing information for indices, cross references, etc. The DVI file (for DeVice Independent file) can be printed on virtually any device (see the following sections).

The tex formatting command itself does not sort the indices; it writes an output file of unsorted index data. (The texi2dvi command automatically generates indices; see Section 19.3 [Format with texi2dvi], page 144.) To generate a printed index after running the tex command, you first need a sorted index to work from. The texindex command sorts indices. (The source file ‘texindex.c’ comes as part of the standard Texinfo distribution, among other places.)

The tex formatting command outputs unsorted index files under names that obey a standard convention: the name of your main input file with any ‘.tex’ (or similar, see section “tex invocation” in Web2c) extension removed, followed by the two letter names of indices. For example, the raw index output files for the input file ‘foo.texinfo’ would be ‘foo.cp’, ‘foo.vr’, ‘foo.fn’, ‘foo.tp’, ‘foo.pg’ and ‘foo.ky’. Those are exactly the arguments to give to texindex.

Instead of specifying all the unsorted index file names explicitly, you can use ‘??’ as shell wildcards and give the command in this form:

     texindex foo.??

This command will run texindex on all the unsorted index files, including any that you have defined yourself using @defindex or @defcodeindex. (You may execute ‘texindex foo.??’ even if there are similarly named files with two letter extensions that are not index files, such as ‘foo.el’. The texindex command reports but otherwise ignores such files.)

For each file specified, texindex generates a sorted index file whose name is made by appending ‘s’ to the input file name. The @printindex command looks for a file with that name (see Section 4.1 [Printing Indices & Menus], page 41). texindex does not alter the raw index output file.

After you have sorted the indices, you need to rerun the tex formatting command on the Texinfo file. This regenerates the DVI file, this time with up-to-date index entries.

Finally, you may need to run tex one more time, to get the page numbers in the cross-references correct.

To summarize, this is a five step process:

1. Run tex on your Texinfo file. This generates a DVI file (with undefined cross-references and no indices), and the raw index files (with two letter extensions).

2. Run texindex on the raw index files. This creates the corresponding sorted index files (with three letter extensions).

3. Run tex again on your Texinfo file. This regenerates the DVI file, this time with indices and defined cross-references, but with page numbers for the cross-references from last time, generally incorrect.

4. Sort the indices again, with texindex.

5. Run tex one last time. This time the correct page numbers are written for the crossreferences.

Alternatively, it’s a one-step process: run texi2dvi (see Section 19.3 [Format with texi2dvi], page 144).

You need not run texindex each time after you run tex. If you do not, on the next run, the tex formatting command will use whatever sorted index files happen to exist from the previous use of texindex. This is usually ok while you are debugging.

Sometimes you may wish to print a document while you know it is incomplete, or to print just one chapter of a document. In that case, the usual auxiliary files that TEX creates and warnings TEX gives when cross-references are not satisfied are just nuisances. You can avoid them with the @novalidate command, which you must give before the @setfilename command (see Section 3.2.3 [@setfilename], page 28). Thus, the beginning of your file would look approximately like this:

     \input texinfo
     @novalidate
     @setfilename myfile.info
     ...

@novalidate also turns off validation in makeinfo, just like its --no-validate option (see Section 20.1.4 [Pointer Validation], page 156).

19.3 Format with texi2dvi

The texi2dvi command automatically runs both tex and texindex as many times as necessary to produce a DVI file with sorted indices and all cross-references resolved. It simplifies the tex—texindex—tex—tex sequence described in the previous section.

To run texi2dvi on an input file ‘foo.texi’, do this (where ‘prompt$ ’ is your shell prompt):

     prompt$ texi2dvi foo.texi

As shown in this example, the input filenames to texi2dvi must include any extension (‘.texi’, ‘.texinfo’, etc.). Under MS-DOS and perhaps in other circumstances, you may need to run ‘sh texi2dvi foo.texi’ instead of relying on the operating system to invoke the shell on the ‘texi2dvi’ script.

Perhaps the most useful option to texi2dvi is ‘--texinfo=cmd’. This inserts cmd on a line by itself after the @setfilename in a temporary copy of the input file before running TEX. With this, you can specify different printing formats, such as @smallbook (see Section 19.11 [smallbook], page 150), @afourpaper (see Section 19.12 [A4 Paper], page 150), or @pagesizes (see Section 19.13 [pagesizes], page 150), without actually changing the document source. (You can also do this on a site-wide basis with ‘texinfo.cnf’; see Section 19.9 [Preparing for TEX], page 148).

For a list of other options, run ‘texi2dvi --help’.

19.4 Shell Print Using lpr -d

The precise command to print a DVI file depends on your system installation. Two

common ones are ‘dvips foo.dvi -o’ and ‘lpr -d foo.dvi’.

For example, the following commands will (perhaps) suffice to sort the indices, format, and print the Bison Manual:

     tex bison.texinfo
     texindex bison.??
     tex bison.texinfo
     lpr -d bison.dvi

(Remember that the shell commands may be different at your site; but these are commonly used versions.)

Using the texi2dvi shell script (see the previous section):

     texi2dvi bison.texinfo
     lpr -d bison.dvi
     # or perhaps dvips bison.dvi -o

lpr is a standard program on Unix systems, but it is usually absent on MS-DOS/MS-Windows. Some network packages come with a program named lpr, but these are usually limited to sending files to a print server over the network, and generally don’t support the ‘-d’ option. If you are unfortunate enough to work on one of these systems, you have several alternative ways of printing DVI files:

• Find and install a Unix-like lpr program, or its clone. If you can do that, you will be able to print DVI files just like described above.

• Send the DVI files to a network printer queue for DVI files. Some network printers have special queues for printing DVI files. You should be able to set up your network software to send files to that queue. In some cases, the version of lpr which comes with your network software will have a special option to send a file to specific queues, like this:

         lpr -Qdvi -hprint.server.domain bison.dvi

• Convert the DVI file to a Postscript or PCL file and send it to your local printer. See section “dvips invocation” in Dvips, and the man pages for dvilj, for detailed description of these tools. Once the DVI file is converted to the format your local printer understands directly, just send it to the appropriate port, usually ‘PRN’.

19.5 From an Emacs Shell

You can give formatting and printing commands from a shell within GNU Emacs. To create a shell within Emacs, type M-x shell. In this shell, you can format and print the document. See Chapter 19 [Format and Print Hardcopy], page 142, for details.

You can switch to and from the shell buffer while tex is running and do other editing. If you are formatting a long document on a slow machine, this can be very convenient.

You can also use texi2dvi from an Emacs shell. For example, here is how to use texi2dvi to format and print Using and Porting GNU CC from a shell within Emacs:

     texi2dvi gcc.texinfo
     lpr -d gcc.dvi

19.6 Formatting and Printing in Texinfo Mode

Texinfo mode provides several predefined key commands for TEX formatting and printing. These include commands for sorting indices, looking at the printer queue, killing the formatting job, and recentering the display of the buffer in which the operations occur.

C-c C-t C-b
M-x texinfo-tex-buffer

Run texi2dvi on the current buffer.

C-c C-t C-r
M-x texinfo-tex-region

Run TEX on the current region.

C-c C-t C-i
M-x texinfo-texindex

Sort the indices of a Texinfo file formatted with texinfo-tex-region.

C-c C-t C-p
M-x texinfo-tex-print
          Print a DVI file that was made with texinfo-tex-region or texinfo-tex-
          buffer.
C-c C-t C-q
M-x tex-show-print-queue

Show the print queue.

C-c C-t C-d
M-x texinfo-delete-from-print-queue

Delete a job from the print queue; you will be prompted for the job num-

          ber shown by a preceding C-c C-t C-q command (texinfo-show-tex-print-
          queue).
C-c C-t C-k
M-x tex-kill-job

Kill the currently running TEX job started by either texinfo-tex-region or texinfo-tex-buffer, or any other process running in the Texinfo shell buffer.

C-c C-t C-x
M-x texinfo-quit-job

Quit a TEX formatting job that has stopped because of an error by sending an ⟨\underline{\overline{x}}^{⟩} to it. When you do this, TEX preserves a record of what it did in a ‘.log’ file.

C-c C-t C-l
M-x tex-recenter-output-buffer

Redisplay the shell buffer in which the TEX printing and formatting commands are run to show its most recent output.

Thus, the usual sequence of commands for formatting a buffer is as follows (with comments to the right):

     C-c C-t C-b  Run texi2dvi on the buffer.

C-c C-t C-p Print the DVI file.

C-c C-t C-q Display the printer queue.

The Texinfo mode TEX formatting commands start a subshell in Emacs called the

‘*tex-shell*’. The texinfo-tex-command, texinfo-texindex-command, and tex-dvi-

print-command commands are all run in this shell.

You can watch the commands operate in the ‘*tex-shell*’ buffer, and you can switch to and from and use the ‘*tex-shell*’ buffer as you would any other shell buffer.

The formatting and print commands depend on the values of several variables. The default values are:

Variable Default value

     texinfo-texi2dvi-command                  "texi2dvi"
     texinfo-tex-command                       "tex"
     texinfo-texindex-command                  "texindex"
     texinfo-delete-from-print-queue-command   "lprm"
     texinfo-tex-trailer                       "@bye"
     tex-start-of-header                       "%**start"
     tex-end-of-header                         "%**end"
     tex-dvi-print-command                     "lpr -d"
     tex-show-queue-command                    "lpq"

You can change the values of these variables with the M-x edit-options command (see section “Editing Variable Values” in The GNU Emacs Manual), with the M-x setvariable command (see section “Examining and Setting Variables” in The GNU Emacs Manual), or with your ‘.emacs’ initialization file (see section “Init File” in The GNU Emacs Manual).

Beginning with version 20, GNU Emacs offers a user-friendly interface, called Customize, for changing values of user-definable variables. See section “Easy Customization Interface” in The GNU Emacs Manual, for more details about this. The Texinfo variables can be found in the ‘Development/Docs/Texinfo’ group, once you invoke the M-x

customize command.

19.7 Using the Local Variables List

Yet another way to apply the TEX formatting command to a Texinfo file is to put that command in a local variables list at the end of the Texinfo file. You can then specify the tex or texi2dvi commands as a compile-command and have Emacs run it by typing M-x compile. This creates a special shell called the ‘*compilation*’ buffer in which Emacs runs the compile command. For example, at the end of the ‘gdb.texinfo’ file, after the @bye, you could put the following:

     Local Variables:
     compile-command: "texi2dvi gdb.texinfo"
     End:

This technique is most often used by programmers who also compile programs this way; see section “Compilation” in The GNU Emacs Manual.

19.8 TEX Formatting Requirements Summary

Every Texinfo file that is to be input to TEX must begin with a \input command and must contain an @setfilename command:

     \input texinfo

@setfilename arg-not-used-by-TEX The first command instructs TEX to load the macros it needs to process a Texinfo file and the second command opens auxiliary files.

Every Texinfo file must end with a line that terminates TEX’s processing and forces out unfinished pages:

     @bye

Strictly speaking, these lines are all a Texinfo file needs to be processed successfully by TEX.

Usually, however, the beginning includes an @settitle command to define the title of the printed manual, an @setchapternewpage command, a title page, a copyright page, and permissions. Besides an @bye, the end of a file usually includes indices and a table of contents. (And of course most manuals contain a body of text as well.)

For more information, see:

• Section 3.2.4 [@settitle], page 29

• Section 3.6.2 [@setchapternewpage], page 38

• Appendix E [Page Headings], page 200

• Section 3.4 [Titlepage & Copyright Page], page 31 • Section 4.1 [Printing Indices & Menus], page 41

• Section 4.2 [Contents], page 42

19.9 Preparing for TEX

TEX needs to know where to find the ‘texinfo.tex’ file that the ‘\input texinfo’ command on the first line reads. The ‘texinfo.tex’ file tells TEX how to handle @-commands; it is included in all standard GNU distributions.

Usually, the installer has put the ‘texinfo.tex’ file in the default directory that contains TEX macros when GNU Texinfo, Emacs or other GNU software is installed. In this case, TEX will find the file and you do not need to do anything special. If this has not been done, you can put ‘texinfo.tex’ in the current directory when you run TEX, and TEX will find it there.

Also, you should install ‘epsf.tex’, if it is not already installed from another distribution. More details are at the end of the description of the @image command (see Section 13.11 [Images], page 111).

Likewise for ‘pdfcolor.tex’, if it is not already installed and you use pdftex.

Optionally, you may create an additional ‘texinfo.cnf’, and install it as well. This file is read by TEX when the @setfilename command is executed (see Section 3.2.3 [@setfilename], page 28). You can put any commands you like there, according to local site-wide conventions. They will be read by TEX when processing any Texinfo document. For example, if ‘texinfo.cnf’ contains the line ‘@afourpaper’ (see Section 19.12 [A4 Paper], page 150), then all Texinfo documents will be processed with that page size in effect. If you have nothing to put in ‘texinfo.cnf’, you do not need to create it.

If neither of the above locations for these system files suffice for you, you can specify the directories explicitly. For ‘texinfo.tex’, you can do this by writing the complete path for the file after the \input command. Another way, that works for both ‘texinfo.tex’ and ‘texinfo.cnf’ (and any other file TEX might read), is to set the TEXINPUTS environment variable in your ‘.cshrc’ or ‘.profile’ file.

Which you use of ‘.cshrc’ or ‘.profile’ depends on whether you use a Bourne shell-compatible (sh, bash, ksh, …) or C shell-compatible (csh, tcsh) command interpreter. The latter read the ‘.cshrc’ file for initialization information, and the former read

‘.profile’.

In a ‘.cshrc’ file, you could use the following csh command sequence:

     setenv TEXINPUTS .:/home/me/mylib:/usr/lib/tex/macros

In a ‘.profile’ file, you could use the following sh command sequence:

     TEXINPUTS=.:/home/me/mylib:/usr/lib/tex/macros
     export TEXINPUTS

On MS-DOS/MS-Windows, you would say it like this^{1}:

     set TEXINPUTS=.;d:/home/me/mylib;c:/usr/lib/tex/macros

It is customary for DOS/Windows users to put such commands in the ‘autoexec.bat’ file, or in the Windows Registry.

1 Note the use of the ‘;’ character, instead of ‘:’, as directory separator on these systems.

These settings would cause TEX to look for ‘\input’ file first in the current directory, indicated by the ‘.’, then in a hypothetical user’s ‘me/mylib’ directory, and finally in a

system directory ‘/usr/lib/tex/macros’.

Finally, you may wish to dump a ‘.fmt’ file (see section “Memory dumps” in Web2c) so that TEX can load Texinfo faster. (The disadvantage is that then updating ‘texinfo.tex’ requires redumping.) You can do this by running this command, assuming ‘epsf.tex’ is findable by TEX:

     initex texinfo @dump

(dump is a TEX primitive.) Then, move ‘texinfo.fmt’ to wherever your .fmt files are found; typically, this will be in the subdirectory ‘web2c’ of your TEX installation.

19.10 Overfull “hboxes”

TEX is sometimes unable to typeset a line without extending it into the right margin. This can occur when TEX comes upon what it interprets as a long word that it cannot hyphenate, such as an electronic mail network address or a very long title. When this happens, TEX prints an error message like this:

     Overfull @hbox (20.76302pt too wide)

(In TEX, lines are in “horizontal boxes”, hence the term, “hbox”. ‘@hbox’ is a TEX primitive not needed in the Texinfo language.)

TEX also provides the line number in the Texinfo source file and the text of the offending line, which is marked at all the places that TEX considered hyphenation. See Section F.2 [Catching Errors with TEX Formatting], page 205, for more information about typesetting errors.

If the Texinfo file has an overfull hbox, you can rewrite the sentence so the overfull hbox does not occur, or you can decide to leave it. A small excursion into the right margin often does not matter and may not even be noticeable.

If you have many overfull boxes and/or an antipathy to rewriting, you can coerce TEX into greatly increasing the allowable interword spacing, thus (if you’re lucky) avoiding many of the bad line breaks, like this:

     @tex
     \global\emergencystretch = .9\hsize
     @end tex

(You should adjust the fraction as needed.) This huge value for \emergencystretch cannot be the default, since then the typeset output would generally be of noticeably lower quality; the default is ‘.15\hsize’. \hsize is the TEX dimension containing the current line width.

For what overfull boxes you have, however, TEX will print a large, ugly, black rectangle beside the line that contains the overfull hbox unless told otherwise. This is so you will notice the location of the problem if you are correcting a draft.

To prevent such a monstrosity from marring your final printout, write the following in the beginning of the Texinfo file on a line of its own, before the @titlepage command:

     @finalout

19.11 Printing “Small” Books

By default, TEX typesets pages for printing in an 8.5 by 11 inch format. However, you can direct TEX to typeset a document in a 7 by 9.25 inch format that is suitable for bound books by inserting the following command on a line by itself at the beginning of the Texinfo file, before the title page:

     @smallbook

(Since many books are about 7 by 9.25 inches, this command might better have been called the @regularbooksize command, but it came to be called the @smallbook command by comparison to the 8.5 by 11 inch format.)

If you write the @smallbook command between the start-of-header and end-of-header lines, the Texinfo mode TEX region formatting command, texinfo-tex-region, will format the region in “small” book size (see Section 3.2.2 [Start of Header], page 28).

See Section 10.7 [small], page 84, for information about commands that make it easier to produce examples for a smaller manual.

See Section 19.3 [Format with texi2dvi], page 144, and Section 19.9 [Preparing for TEX], page 148, for other ways to format with @smallbook that do not require changing the source file.

19.12 Printing on A4 Paper

You can tell TEX to format a document for printing on European size A4 paper (or A5) with the @afourpaper (or @afivepaper) command. Write the command on a line by itself near the beginning of the Texinfo file, before the title page. For example, this is how you would write the header for this manual:

     \input texinfo    @c -*-texinfo-*-
     @c %**start of header
     @setfilename texinfo
     @settitle Texinfo
     @afourpaper
     @c %**end of header

See Section 19.3 [Format with texi2dvi], page 144, and Section 19.9 [Preparing for TEX], page 148, for other ways to format for different paper sizes that do not require changing the source file.

You may or may not prefer the formatting that results from the command @afourlatex. There’s also @afourwide for A4 paper in wide format.

19.13 @pagesizes [width][, height]: Custom page sizes

You can explicitly specify the height and (optionally) width of the main text area on the page with the @pagesizes command. Write this on a line by itself near the beginning of the Texinfo file, before the title page. The height comes first, then the width if desired, separated by a comma. Examples:

     @pagesizes 200mm,150mm

and

     @pagesizes 11.5in

This would be reasonable for printing on B5-size paper. To emphasize, this command specifies the size of the text area, not the size of the paper (which is 250 mm by 177 mm for B5, 14 in by 8.5 in for legal).

To make more elaborate changes, such as changing any of the page margins, you must define a new command in ‘texinfo.tex’ (or ‘texinfo.cnf’, see Section 19.9 [Preparing for TEX], page 148).

See Section 19.3 [Format with texi2dvi], page 144, and Section 19.9 [Preparing for TEX], page 148, for other ways to specify @pagesizes that do not require changing the source file.

    @pagesizes is ignored by makeinfo.

19.14 Cropmarks and Magnification

You can (attempt to) direct TEX to print cropmarks at the corners of pages with the @cropmarks command. Write the @cropmarks command on a line by itself between @iftex and @end iftex lines near the beginning of the Texinfo file, before the title page, like this:

     @iftex @cropmarks @end iftex

This command is mainly for printers that typeset several pages on one sheet of film; but you can attempt to use it to mark the corners of a book set to 7 by 9.25 inches with the @smallbook command. (Printers will not produce cropmarks for regular sized output that is printed on regular sized paper.) Since different printing machines work in different ways, you should explore the use of this command with a spirit of adventure. You may have to redefine the command in ‘texinfo.tex’.

You can attempt to direct TEX to typeset pages larger or smaller than usual with the \mag TEX command. Everything that is typeset is scaled proportionally larger or smaller. (\mag stands for “magnification”.) This is not a Texinfo @-command, but is a plain TEX command that is prefixed with a backslash. You have to write this command between @tex and @end tex (see Section 16.3 [Raw Formatter Commands], page 130).

Follow the \mag command with an ‘=’ and then a number that is 1000 times the magnification you desire. For example, to print pages at 1.2 normal size, write the following near the beginning of the Texinfo file, before the title page:

     @tex \mag=1200 @end tex

With some printing technologies, you can print normal-sized copies that look better than usual by giving a larger-than-normal master to your print shop. They do the reduction, thus effectively increasing the resolution.

Depending on your system, DVI files prepared with a nonstandard-\mag may not print or may print only with certain magnifications. Be prepared to experiment.

19.15 PDF Output

You can generate a PDF output file from Texinfo source by using the pdftex program to process your file instead of plain tex. Just run ‘pdftex foo.texi’ instead of ‘tex foo.texi’, or give the ‘--pdf’ option to texi2dvi.

PDF stands for ‘Portable Document Format’. It was invented by Adobe Systems some years ago for document interchange, based on their PostScript language. A PDF reader (http://www.foolabs.com/xpdf/) for the X window system is freely available, as is the

definition of the file format (http://partners.adobe.com/asn/developer/technotes/).

Since PDF is a binary format, there are no ‘@ifpdf’ or ‘@pdf’ commands as with the other output formats.

Despite the ‘portable’ in the name, PDF files are nowhere near as portable in practice as the plain ASCII formats (Info, HTML) that Texinfo supports (DVI portability is arguable). They also tend to be much larger and do not support the bitmap fonts used by TEX (by default) very well. Nevertheless, a PDF file does preserve an actual printed document on a screen as faithfully as possible, so it has its place.

PDF support in Texinfo is fairly rudimentary.

20 Creating and Installing Info Files

This chapter describes how to create and install Info files. See Section 1.3 [Info Files], page 5, for general information about the file format itself.

20.1 Creating an Info File

makeinfo is a program that converts a Texinfo file into an Info file, HTML file, or plain

text. texinfo-format-region and texinfo-format-buffer are GNU Emacs functions

that convert Texinfo to Info.

For information on installing the Info file in the Info system, see Section 20.2 [Installing an Info File], page 161.

20.1.1 makeinfo Preferred

The makeinfo utility creates an Info file from a Texinfo source file more quickly than either of the Emacs formatting commands and provides better error messages. We recommend it. makeinfo is a C program that is independent of Emacs. You do not need to run Emacs to use makeinfo, which means you can use makeinfo on machines that are too small to run Emacs. You can run makeinfo in any one of three ways: from an operating system shell, from a shell inside Emacs, or by typing the C-c C-m C-r or the C-c C-m C-b command in Texinfo mode in Emacs.

    The texinfo-format-region and the texinfo-format-buffer commands are useful

if you cannot run makeinfo. Also, in some circumstances, they format short regions or buffers more quickly than makeinfo.

20.1.2 Running makeinfo from a Shell

To create an Info file from a Texinfo file, type makeinfo followed by the name of the Texinfo file. Thus, to create the Info file for Bison, type the following to the shell:

     makeinfo bison.texinfo

(You can run a shell inside Emacs by typing M-x shell.)

20.1.3 Options for makeinfo

The makeinfo command takes a number of options. Most often, options are used to set the value of the fill column and specify the footnote style. Each command line option is a word preceded by ‘--’ or a letter preceded by ‘-’. You can use abbreviations for the long option names as long as they are unique.

For example, you could use the following shell command to create an Info file for ‘bison.texinfo’ in which each line is filled to only 68 columns:

     makeinfo --fill-column=68 bison.texinfo

You can write two or more options in sequence, like this:

     makeinfo --no-split --fill-column=70 ...

This would keep the Info file together as one possibly very long file and would also set the fill column to 70.

The options are:

-D var Cause the variable var to be defined. This is equivalent to @set var in the Texinfo file (see Section 16.4 [set clear value], page 131).

--commands-in-node-names

Allow @-commands in node names. This is not recommended, as it can probably never be implemented in TEX. It also makes makeinfo much slower. Also, this option is ignored when ‘--no-validate’ is used. See Section 20.1.4 [Pointer Validation], page 156, for more details.

--docbook

Generate DocBook output rather than Info.

--error-limit=limit

-e limit Set the maximum number of errors that makeinfo will report before exiting (on the assumption that continuing would be useless); default 100.

--fill-column=width

-f width Specify the maximum number of columns in a line; this is the right-hand edge of a line. Paragraphs that are filled will be filled to this width. (Filling is the process of breaking up and connecting lines so that lines are the same length as or shorter than the number specified as the fill column. Lines are broken between words.) The default value is 72. Ignored with ‘--html’.

--footnote-style=style

-s style Set the footnote style to style, either ‘end’ for the end node style (the default) or ‘separate’ for the separate node style. The value set by this option overrides the value set in a Texinfo file by an @footnotestyle command (see Section 13.10 [Footnotes], page 109). When the footnote style is ‘separate’, makeinfo makes a new node containing the footnotes found in the current node. When the footnote style is ‘end’, makeinfo places the footnote references at the end of the current node. Ignored with ‘--html’.

--force

-F Ordinarily, if the input file has errors, the output files are not created. With this option, they are preserved.

--help

-h Print a usage message listing all available options, then exit successfully.

--html Generate HTML output rather than Info. See Section 20.1.9 [makeinfo html], page 160. By default, the HTML output is split into one output file per source node, and the split output is written into a subdirectory with the name of the top-level info file.

-I dir Append dir to the directory search list for finding files that are included using the @include command. By default, makeinfo searches only the current directory. If dir is not given, the current directory ‘.’ is appended. Note that dir can actually be a list of several directories separated by the usual path separator character (‘:’ on Unix, ‘;’ on MS-DOS/MS-Windows).

--macro-expand=file

-E file Output the Texinfo source with all the macros expanded to the named file. Normally, the results of macro expansion are used internally by makeinfo and then discarded. This option is used by texi2dvi if you are using an old version of ‘texinfo.tex’ that does not support @macro.

--no-headers

For Info output, do not include menus or node separator lines in the output. This results in a simple plain text file that you can (for example) send in email without complications, or include in a distribution (as in an ‘INSTALL’ file). For HTML output, likewise omit menus. And if ‘--no-split’ is also specified, do not include a navigation links at the top of each node (these are never included in the default case of split output). See Section 20.1.9 [makeinfo html], page 160. In both cases, write to standard output by default (can still be overridden by ‘-o’).

--no-split

Suppress the splitting stage of makeinfo. By default, large output files (where the size is greater than 70k bytes) are split into smaller subfiles. For Info output, each one is approximately 50k bytes. For HTML output, each file contains one node (see Section 20.1.9 [makeinfo html], page 160).

--no-pointer-validate
--no-validate

Suppress the pointer-validation phase of makeinfo. This can also be done with the @novalidate command (see Section 19.1 [Use TEX], page 142). Normally, after a Texinfo file is processed, some consistency checks are made to ensure that cross references can be resolved, etc. See Section 20.1.4 [Pointer Validation], page 156.

--no-warn

Suppress warning messages (but not error messages). You might want this if the file you are creating has examples of Texinfo cross references within it, and the nodes that are referenced do not actually exist.

--number-sections

Output chapter, section, and appendix numbers as in printed manuals.

--no-number-footnotes

Suppress automatic footnote numbering. By default, makeinfo numbers each footnote sequentially in a single node, resetting the current footnote number to 1 at the start of each node.

--output=file

-o file Specify that the output should be directed to file and not to the file name specified in the @setfilename command found in the Texinfo source (see Section 3.2.3 [setfilename], page 28). If file is ‘-’, output goes to standard output and ‘--no-split’ is implied. For split HTML output, file is the name for the directory into which all HTML nodes are written (see Section 20.1.9 [makeinfo html], page 160).

-P dir Prepend dir to the directory search list for @include. If dir is not given, the current directory ‘.’ is prepended. See ‘-I’ for more details.

--paragraph-indent=indent

-p indent Set the paragraph indentation style to indent. The value set by this option overrides the value set in a Texinfo file by an @paragraphindent command (see Section 3.6.3 [paragraphindent], page 39). The value of indent is interpreted as follows:

‘asis’ Preserve any existing indentation at the starts of paragraphs.

‘0’ or ‘none’ Delete any existing indentation.

num Indent each paragraph by num spaces.

--reference-limit=limit

-r limit Set the value of the number of references to a node that makeinfo will make without reporting a warning. If a node has more than this number of references in it, makeinfo will make the references but also report a warning. The default is 1000.

-U var Cause var to be undefined. This is equivalent to @clear var in the Texinfo file (see Section 16.4 [set clear value], page 131).

--verbose

Cause makeinfo to display messages saying what it is doing. Normally, makeinfo only outputs messages if there are errors or warnings.

--version

-V Print the version number, then exit successfully.

--xml Generate XML output rather than Info.

20.1.4 Pointer Validation

If you do not suppress pointer validation with the ‘--no-validate’ option or the @novalidate command in the source file (see Section 19.1 [Use TEX], page 142), makeinfo will check the validity of the final Info file. Mostly, this means ensuring that nodes you have referenced really exist. Here is a complete list of what is checked:

1. If a ‘Next’, ‘Previous’, or ‘Up’ node reference is a reference to a node in the current file and is not an external reference such as to ‘(dir)’, then the referenced node must exist.

2. In every node, if the ‘Previous’ node is different from the ‘Up’ node, then the node pointed to by the ‘Previous’ field must have a ‘Next’ field which points back to this node.

3. Every node except the ‘Top’ node must have an ‘Up’ pointer.

4. The node referenced by an ‘Up’ pointer must itself reference the current node through a menu item, unless the node referenced by ‘Up’ has the form ‘(file)’.

5. If the ‘Next’ reference of a node is not the same as the ‘Next’ reference of the ‘Up’ reference, then the node referenced by the ‘Next’ pointer must have a ‘Previous’ pointer that points back to the current node. This rule allows the last node in a section to point to the first node of the next chapter.

6. Every node except ‘Top’ should be referenced by at least one other node, either via the ‘Previous’ or ‘Next’ links, or via a menu or a cross-reference.

Some Texinfo documents might fail during the validation phase because they use commands like @value and @definfoenclose in node definitions and cross-references inconsistently. Consider the following example:

     @set nodename Node 1
     @node @value{nodename}, Node 2, Top, Top
     This is node 1.
     @node Node 2, , Node 1, Top
     This is node 2.

Here, the node “Node 1” was referenced both verbatim and through @value.

By default, makeinfo fails such cases, because node names are not fully expanded until they are written to the output file. You should always try to reference nodes consistently; e.g., in the above example, the second @node line should have also used @value. However, if, for some reason, you must reference node names inconsistently, and makeinfo fails to validate the file, you can use the ‘--commands-in-node-names’ option to force makeinfo to perform the expensive expansion of all node names it finds in the document. This might considerably slow down the program, though; twofold increase in conversion time was measured for large documents such as the Jargon file.

The support for @-commands in @node directives is not general enough to be freely used. For example, if the example above redefined nodename somewhere in the document, makeinfo will fail to convert it, even if invoked with the ‘--commands-in-node-names’ option.

    ‘--commands-in-node-names’ has no effect if the ‘--no-validate’ option is given.

20.1.5 Running makeinfo inside Emacs

You can run makeinfo in GNU Emacs Texinfo mode by using either the makeinforegion or the makeinfo-buffer commands. In Texinfo mode, the commands are bound

to C-c C-m C-r and C-c C-m C-b by default.
C-c C-m C-r
M-x makeinfo-region

Format the current region for Info.

C-c C-m C-b
M-x makeinfo-buffer

Format the current buffer for Info.

When you invoke either makeinfo-region or makeinfo-buffer, Emacs prompts for a file name, offering the name of the visited file as the default. You can edit the default file name in the minibuffer if you wish, before pressing ^{⟨}\underline{\overline{RET}}^{⟩} to start the makeinfo process.

    The Emacs makeinfo-region and makeinfo-buffer commands run the makeinfo

program in a temporary shell buffer. If makeinfo finds any errors, Emacs displays the error messages in the temporary buffer.

You can parse the error messages by typing C-x ` (next-error). This causes Emacs to go to and position the cursor on the line in the Texinfo source that makeinfo thinks caused the error. See section “Running make or Compilers Generally” in The GNU Emacs Manual, for more information about using the next-error command.

In addition, you can kill the shell in which the makeinfo command is running or make the shell buffer display its most recent output.

C-c C-m C-k
M-x makeinfo-kill-job
          Kill the current running makeinfo job (from makeinfo-region or makeinfo-
          buffer).
C-c C-m C-l
M-x makeinfo-recenter-output-buffer

Redisplay the makeinfo shell buffer to display its most recent output.

(Note that the parallel commands for killing and recentering a TEX job are C-c C-t C-k and C-c C-t C-l. See Section 19.6 [Texinfo Mode Printing], page 145.)

You can specify options for makeinfo by setting the makeinfo-options variable with either the M-x edit-options or the M-x set-variable command, or by setting the variable in your ‘.emacs’ initialization file.

For example, you could write the following in your ‘.emacs’ file:

     (setq makeinfo-options
           "--paragraph-indent=0 --no-split
            --fill-column=70 --verbose")

For more information, see Section 20.1.3 [Options for makeinfo], page 153, as well as “Editing Variable Values,” “Examining and Setting Variables,” and “Init File” in The GNU Emacs Manual.

20.1.6 The texinfo-format... Commands

In GNU Emacs in Texinfo mode, you can format part or all of a Texinfo file with the texinfo-format-region command. This formats the current region and displays the formatted text in a temporary buffer called ‘*Info Region*’.

Similarly, you can format a buffer with the texinfo-format-buffer command. This command creates a new buffer and generates the Info file in it. Typing C-x C-s will save the Info file under the name specified by the @setfilename line which must be near the beginning of the Texinfo file.

C-c C-e C-r
texinfo-format-region

Format the current region for Info.

C-c C-e C-b
texinfo-format-buffer

Format the current buffer for Info.

    The texinfo-format-region and texinfo-format-buffer commands provide you

with some error checking, and other functions can provide you with further help in finding formatting errors. These procedures are described in an appendix; see Appendix F [Catching Mistakes], page 204. However, the makeinfo program is often faster and provides better error checking (see Section 20.1.5 [makeinfo in Emacs], page 157).

20.1.7 Batch Formatting

You can format Texinfo files for Info using batch-texinfo-format and Emacs Batch mode. You can run Emacs in Batch mode from any shell, including a shell inside of Emacs. (See section “Command Line Switches and Arguments” in The GNU Emacs Manual.)

Here is a shell command to format all the files that end in ‘.texinfo’ in the current directory:

     emacs -batch -funcall batch-texinfo-format *.texinfo

Emacs processes all the files listed on the command line, even if an error occurs while attempting to format some of them.

Run batch-texinfo-format only with Emacs in Batch mode as shown; it is not interactive. It kills the Batch mode Emacs on completion.

batch-texinfo-format is convenient if you lack makeinfo and want to format several Texinfo files at once. When you use Batch mode, you create a new Emacs process. This frees your current Emacs, so you can continue working in it. (When you run texinfoformat-region or texinfo-format-buffer, you cannot use that Emacs for anything else until the command finishes.)

20.1.8 Tag Files and Split Files

If a Texinfo file has more than 30,000 bytes, texinfo-format-buffer automatically creates a tag table for its Info file; makeinfo always creates a tag table. With a tag table, Info can jump to new nodes more quickly than it can otherwise.

In addition, if the Texinfo file contains more than about 70,000 bytes, texinfo-format-buffer and makeinfo split the large Info file into shorter indirect subfiles of about 50,000 bytes each. Big files are split into smaller files so that Emacs does not need to make a large buffer to hold the whole of a large Info file; instead, Emacs allocates just enough memory for the small, split-off file that is needed at the time. This way, Emacs avoids wasting memory when you run Info. (Before splitting was implemented, Info files were always kept short and include files were designed as a way to create a single, large printed manual out of the smaller Info files. See Appendix D [Include Files], page 196, for more information. Include files are still used for very large documents, such as The Emacs Lisp Reference Manual, in which each chapter is a separate file.)

When a file is split, Info itself makes use of a shortened version of the original file that contains just the tag table and references to the files that were split off. The split-off files are called indirect files.

The split-off files have names that are created by appending ‘-1’, ‘-2’, ‘-3’ and so on to the file name specified by the @setfilename command. The shortened version of the original file continues to have the name specified by @setfilename.

At one stage in writing this document, for example, the Info file was saved as the file ‘test-texinfo’ and that file looked like this:

     Info file: test-texinfo,    -*-Text-*-
     produced by texinfo-format-buffer
     from file: new-texinfo-manual.texinfo
     ^_
     Indirect:
     test-texinfo-1: 102
     test-texinfo-2: 50422
     test-texinfo-3: 101300
     ^_^L
     Tag table:
     (Indirect)
     Node: overview^?104
     Node: info file^?1271
     Node: printed manual^?4853
     Node: conventions^?6855
     ...

(But ‘test-texinfo’ had far more nodes than are shown here.) Each of the split-off,

indirect files, ‘test-texinfo-1’, ‘test-texinfo-2’, and ‘test-texinfo-3’, is listed in this

file after the line that says ‘Indirect:’. The tag table is listed after the line that says ‘Tag

table:’.

In the list of indirect files, the number following the file name records the cumulative number of bytes in the preceding indirect files, not counting the file list itself, the tag table, or the permissions text in each file. In the tag table, the number following the node name records the location of the beginning of the node, in bytes from the beginning of the (unsplit) output.

If you are using texinfo-format-buffer to create Info files, you may want to run the Info-validate command. (The makeinfo command does such a good job on its own, you do not need Info-validate.) However, you cannot run the M-x Info-validate nodechecking command on indirect files. For information on how to prevent files from being split and how to validate the structure of the nodes, see Section F.5.1 [Using Info-validate], page 209.

20.1.9 Generating HTML

Besides generating output in the Info format, you can use the ‘--html’ option to generate output in HTML format, for installation on a web site (for example). By default, the HTML output is split at node level.

When splitting, the HTML output files are written into a subdirectory. The subdirectory is named according to the name from @setfilename with any extension removed; for example, HTML output for @setfilename emacs.info would be written into a subdirectory named ‘emacs’. If that directory cannot be created for any reason, then ‘.html’ is appended to the directory name, as in ‘emacs.html’ (this is necessary because sometimes the info file is named without an extension, e.g., ‘texinfo’). If the ‘name.html’ directory can’t be created either, makeinfo gives up. In any case, the top-level output file within the directory is always named ‘index.html’.

Monolithic output (--no-split) is named according to @setfilename or --outfile. Cross-document node references are not supported in monolithic HTML.

Texinfo input marked up with the @ifhtml command will produce output only with the ‘--html’ option supplied. Input marked up with the @html is passed literally to the output (suppressing the normal escaping of input ‘<’, ‘>’ and ‘&’ characters which have special significance in HTML).

The ‘--footnote-style’ option is currently ignored for HTML output; footnotes are linked to the end of the output file.

The HTML generated is mostly standard (i.e., HTML 2.0, RFC-1866). The exception is that HTML 3.2 tables are generated from the @multitable command, but tagged to degrade as well as possible in browsers without table support. The HTML 4 ‘lang’ attribute on the ‘<html>’ attribute is also used. Please report output from an error-free run of makeinfo which has browser portability problems as a bug.

Navigation bars are inserted at the start of nodes, similarly to Info output. The ‘--no-headers’ option will suppress this if used with ‘--no-split’. Header <link> elements in split output can support info-like navigation with browsers like Lynx and Emacs W3 which implement this HTML 1.0 feature. ‘@xref’ commands to other documents are generated assuming the other document is available in split HTML form, and installed in the same HTML documentation tree, at ‘../<info-document>/’.

20.2 Installing an Info File

Info files are usually kept in the ‘info’ directory. You can read Info files using the standalone Info program or the Info reader built into Emacs. (See Info file ‘info’, node ‘Top’, for an introduction to Info.)

20.2.1 The Directory File ‘dir’

For Info to work, the ‘info’ directory must contain a file that serves as a top level directory for the Info system. By convention, this file is called ‘dir’. (You can find the location of this file within Emacs by typing C-h i to enter Info and then typing C-x C-f to see the pathname to the ‘info’ directory.)

The ‘dir’ file is itself an Info file. It contains the top level menu for all the Info files in the system. The menu looks like this:

     * Menu:
     * Info:    (info).     Documentation browsing system.
     * Emacs:   (emacs).    The extensible, self-documenting
                            text editor.
     * Texinfo: (texinfo).  With one source file, make
                            either a printed manual using
                            @TeX{} or an Info file.
     ...

Each of these menu entries points to the ‘Top’ node of the Info file that is named in parentheses. (The menu entry does not need to specify the ‘Top’ node, since Info goes to the ‘Top’ node if no node name is mentioned. See Section 7.5 [Nodes in Other Info Files], page 59.)

Thus, the ‘Info’ entry points to the ‘Top’ node of the ‘info’ file and the ‘Emacs’ entry points to the ‘Top’ node of the ‘emacs’ file.

In each of the Info files, the ‘Up’ pointer of the ‘Top’ node refers back to the dir file. For example, the line for the ‘Top’ node of the Emacs manual looks like this in Info:

     File: emacs  Node: Top, Up: (DIR), Next: Distrib

In this case, the ‘dir’ file name is written in upper case letters—it can be written in either upper or lower case. This is not true in general, it is a special case for ‘dir’.

20.2.2 Listing a New Info File

To add a new Info file to your system, you must write a menu entry to add to the menu in the ‘dir’ file in the ‘info’ directory. For example, if you were adding documentation for GDB, you would write the following new entry:

     * GDB: (gdb).           The source-level C debugger.

The first part of the menu entry is the menu entry name, followed by a colon. The second part is the name of the Info file, in parentheses, followed by a period. The third part is the description.

The name of an Info file often has a ‘.info’ extension. Thus, the Info file for GDB might be called either ‘gdb’ or ‘gdb.info’. The Info reader programs automatically try the file name both with and without ‘.info’^{1}; so it is better to avoid clutter and not to write ‘.info’ explicitly in the menu entry. For example, the GDB menu entry should use just ‘gdb’ for the file name, not ‘gdb.info’.

20.2.3 Info Files in Other Directories

If an Info file is not in the ‘info’ directory, there are three ways to specify its location:

1. Write the pathname in the ‘dir’ file as the second part of the menu.

2. If you are using Emacs, list the name of the file in a second ‘dir’ file, in its directory; and then add the name of that directory to the Info-directory-list variable in your personal or site initialization file. This variable tells Emacs where to look for ‘dir’ files (the files must be named ‘dir’). Emacs merges the files named ‘dir’ from each of the listed directories. (In Emacs version 18, you can set the Info-directory variable to the name of only one directory.)

3. Specify the Info directory name in the INFOPATH environment variable in your ‘.profile’ or ‘.cshrc’ initialization file. (Only you and others who set this environment variable will be able to find Info files whose location is specified this way.)

For example, to reach a test file in the ‘/home/bob/info’ directory, you could add an entry like this to the menu in the standard ‘dir’ file:

     * Test: (/home/bob/info/info-test).  Bob's own test file.

In this case, the absolute file name of the ‘info-test’ file is written as the second part of the menu entry.

1 On MS-DOS/MS-Windows systems, Info will try the ‘.inf’ extension as well.

Alternatively, you could write the following in your ‘.emacs’ file:

     (require 'info)
     (setq Info-directory-list
       (cons (expand-file-name "/home/bob/info")
             Info-directory-list))

This tells Emacs to merge the system ‘dir’ file with the ‘dir’ file in ‘/home/bob/info’. Thus, Info will list the ‘/home/bob/info/info-test’ file as a menu entry in the ‘/home/bob/info/dir’ file. Emacs does the merging only when M-x info is first run, so if you want to set Info-directory-list in an Emacs session where you’ve already run info, you must (setq Info-dir-contents nil) to force Emacs to recompose the ‘dir’ file.

Finally, you can tell Info where to look by setting the INFOPATH environment variable in your shell startup file, such as ‘.cshrc’, ‘.profile’ or ‘autoexec.bat’. If you use a Bourne-compatible shell such as sh or bash for your shell command interpreter, you set the INFOPATH environment variable in the ‘.profile’ initialization file; but if you use csh or tcsh, you set the variable in the ‘.cshrc’ initialization file. On MS-DOS/MS-Windows systems, you must set INFOPATH in your ‘autoexec.bat’ file or in the Registry. Each type of shell uses a different syntax.

• In a ‘.cshrc’ file, you could set the INFOPATH variable as follows:

         setenv INFOPATH .:~/info:/usr/local/emacs/info

• In a ‘.profile’ file, you would achieve the same effect by writing:

         INFOPATH=.:$HOME/info:/usr/local/emacs/info
         export INFOPATH

• In a ‘autoexec.bat’ file, you write this command^{2}:

         set INFOPATH=.;%HOME%/info;c:/usr/local/emacs/info

The ‘.’ indicates the current directory as usual. Emacs uses the INFOPATH environment variable to initialize the value of Emacs’s own Info-directory-list variable. The standalone Info reader merges any files named ‘dir’ in any directory listed in the INFOPATH variable into a single menu presented to you in the node called ‘(dir)Top’.

However you set INFOPATH, if its last character is a colon^{3}, this is replaced by the default (compiled-in) path. This gives you a way to augment the default path with new directories without having to list all the standard places. For example (using sh syntax):

     INFOPATH=/local/info:
     export INFOPATH

will search ‘/local/info’ first, then the standard directories. Leading or doubled colons are not treated specially.

When you create your own ‘dir’ file for use with Info-directory-list or INFOPATH, it’s easiest to start by copying an existing ‘dir’ file and replace all the text after the ‘* Menu:’ with your desired entries. That way, the punctuation and special CTRL- characters that Info needs will be present.

2 Note the use of ‘;’ as the directory separator, and a different syntax for using values of other environ-

ment variables.

3 On MS-DOS/MS-Windows systems, use semi-colon instead.

20.2.4 Installing Info Directory Files

When you install an Info file onto your system, you can use the program installinfo to update the Info directory file ‘dir’. Normally the makefile for the package runs install-info, just after copying the Info file into its proper installed location.

In order for the Info file to work with install-info, you include the commands

@dircategory and @direntry. . .@end direntry in the Texinfo source file. Use @direntry

to specify the menu entries to add to the Info directory file, and use @dircategory to specify which part of the Info directory to put it in. Here is how these commands are used in this manual:

@dircategory Texinfo documentation system
@direntry
* Texinfo: (texinfo).           The GNU documentation format.
* install-info: (texinfo)Invoking install-info. ...
...
@end direntry

Here’s what this produces in the Info file:

INFO-DIR-SECTION Texinfo documentation system
START-INFO-DIR-ENTRY
* Texinfo: (texinfo).           The GNU documentation format.
* install-info: (texinfo)Invoking install-info. ...
...
END-INFO-DIR-ENTRY

The install-info program sees these lines in the Info file, and that is how it knows what to do.

Always use the @direntry and @dircategory commands near the beginning of the Texinfo input, before the first @node command. If you use them later on in the input, install-info will not notice them.

If you use @dircategory more than once in the Texinfo source, each usage specifies the ‘current’ category; any subsequent @direntry commands will add to that category.

Here are some recommended @dircategory categories: GNU packages GNU programming tools GNU programming documentation GNU Emacs Lisp GNU libraries TeX Individual utilities The idea is to include the ‘Invoking’ node for every program installed by a package

under ‘Individual utilities’, and an entry for the manual as a whole in the appropriate other category.

20.2.5 Invoking install-info

install-info inserts menu entries from an Info file into the top-level ‘dir’ file in the Info system (see the previous sections for an explanation of how the ‘dir’ file works). It’s most often run as part of software installation, or when constructing a ‘dir’ file for all manuals on a system. Synopsis:

     install-info [option]... [info-file [dir-file]]

If info-file or dir-file are not specified, the options (described below) that define them must be. There are no compile-time defaults, and standard input is never used. installinfo can read only one Info file and write only one ‘dir’ file per invocation.

If dir-file (however specified) does not exist, install-info creates it if possible (with no entries).

If any input file is compressed with gzip (see section “Invoking gzip” in Gzip), install-info automatically uncompresses it for reading. And if dir-file is compressed, install-info also automatically leaves it compressed after writing any changes. If dir-file itself does not exist, install-info tries to open ‘dir-file.gz’.

Options:

--delete Delete the entries in info-file from dir-file. The file name in the entry in dir-file must be info-file (except for an optional ‘.info’ in either one). Don’t insert any new entries.

--dir-file=name

-d name Specify file name of the Info directory file. This is equivalent to using the dir-file argument.

--entry=text

-e text Insert text as an Info directory entry; text should have the form of an Info menu item line plus zero or more extra lines starting with whitespace. If you specify more than one entry, they are all added. If you don’t specify any entries, they are determined from information in the Info file itself.

--help

-h Display a usage message listing basic usage and all available options, then exit successfully.

--info-file=file

-i file Specify Info file to install in the directory. Equivalent to using the info-file argument.

--info-dir=dir

-D dir Specify the directory where ‘dir’ resides. Equivalent to ‘--dir-file=dir/dir’.

--item=text

Same as ‘--entry=text’. An Info directory entry is actually a menu item.

--quiet Suppress warnings.

--remove
-r  Same as ‘--delete’.
--section=sec

-s sec Put this file’s entries in section sec of the directory. If you specify more than one section, all the entries are added in each of the sections. If you don’t specify any sections, they are determined from information in the Info file itself.

Chapter 20: Creating and Installing Info Files 166

--version

-V Display version information and exit successfully.

Appendix A @-Command List

Here is an alphabetical list of the @-commands in Texinfo. Square brackets, [ ], indicate optional arguments; an ellipsis, ‘...’, indicates repeated text.

@whitespace An @ followed by a space, tab, or newline produces a normal, stretchable, interword space. See Section 13.2.3 [Multiple Spaces], page 102.

@! Generate an exclamation point that really does end a sentence (usually after an end-of-sentence capital letter). See Section 13.2.2 [Ending a Sentence], page 102.

@"

@' Generate an umlaut or acute accent, respectively, over the next character, as in ö and ó. See Section 13.3 [Inserting Accents], page 103.

@* Force a line break. Do not end a paragraph that uses @* with an @refill command. See Section 14.1 [Line Breaks], page 113.

@,{c} Generate a cedilla accent under c, as in ç. See Section 13.3 [Inserting Accents], page 103.

@- Insert a discretionary hyphenation point. See Section 14.2 [- and hyphenation], page 114.

@. Produce a period that really does end a sentence (usually after an end-of-sentence capital letter). See Section 13.2.2 [Ending a Sentence], page 102.

@: Indicate to TEX that an immediately preceding period, question mark, exclamation mark, or colon does not end a sentence. Prevent TEX from inserting extra whitespace as it does at the end of a sentence. The command has no effect on the Info file output. See Section 13.2.1 [Not Ending a Sentence], page 102.

@= Generate a macron (bar) accent over the next character, as in ō. See Section 13.3 [Inserting Accents], page 103.

@? Generate a question mark that really does end a sentence (usually after an endof-sentence capital letter). See Section 13.2.2 [Ending a Sentence], page 102.

@@ Stands for an at sign, ‘@’. See Section 13.1 [Inserting @ and braces], page 101.

@\ Stands for a backslash (‘\’) inside @math. See Section 13.8 [math], page 106.

@^

@` Generate a circumflex (hat) or grave accent, respectively, over the next character, as in ô and è. See Section 13.3 [Inserting Accents], page 103.

@{ Stands for a left brace, ‘{’. See Section 13.1 [Inserting @ and braces], page 101.

@} Stands for a right-hand brace, ‘}’.

See Section 13.1 [Inserting @ and braces], page 101.

@~ Generate a tilde accent over the next character, as in Ñ. See Section 13.3 [Inserting Accents], page 103.

@AA{}

@aa{} Generate the uppercase and lowercase Scandinavian A-ring letters, respectively: Å, å. See Section 13.3 [Inserting Accents], page 103.

@acronym{abbrev}

Tag abbrev as an acronym, that is, an abbreviation written in all capital letters, such as ‘NASA’. See Section 9.1.13 [acronym], page 78.

@AE{}

@ae{} Generate the uppercase and lowercase AE ligatures, respectively: Æ, æ. See

Section 13.3 [Inserting Accents], page 103.

@afivepaper

Change page dimensions for the A5 paper size. See Section 19.12 [A4 Paper], page 150.

@afourlatex
@afourpaper
@afourwide

Change page dimensions for the A4 paper size. See Section 19.12 [A4 Paper], page 150.

@alias new=existing Make the command ‘@new’ an alias for the existing command ‘@existing’. See Section 18.4 [alias], page 139.

@anchor{name}

Define name as the current location for use as a cross-reference target. See Section 6.5 [@anchor], page 56.

@appendix title

Begin an appendix. The title appears in the table of contents of a printed manual. In Info, the title is underlined with asterisks. See Section 5.5 [The

@unnumbered and @appendix Commands], page 46.
@appendixsec title
@appendixsection title

Begin an appendix section within an appendix. The section title appears in the table of contents of a printed manual. In Info, the title is underlined with equal signs. @appendixsection is a longer spelling of the @appendixsec command. See Section 5.8 [Section Commands], page 47.

@appendixsubsec title

Begin an appendix subsection within an appendix. The title appears in the table of contents of a printed manual. In Info, the title is underlined with hyphens. See Section 5.10 [Subsection Commands], page 48.

@appendixsubsubsec title

Begin an appendix subsubsection within an appendix subsection. The title appears in the table of contents of a printed manual. In Info, the title is underlined with periods. See Section 5.11 [The ‘subsub’ Commands], page 48.

@asis Used following @table, @ftable, and @vtable to print the table’s first column without highlighting (“as is”). See Section 11.3 [Making a Two-column Table], page 92.

@author author

Typeset author flushleft and underline it. See Section 3.4.3 [The @title and @author Commands], page 33.

@b{text} Print text in bold font. No effect in Info. See Section 9.2.3 [Fonts], page 80.

@bullet{}

Generate a large round dot, or the closest possible thing to one. See Section 13.4.2 [@bullet], page 105.

@bye Stop formatting a file. The formatters do not see the contents of a file following an @bye command. See Chapter 4 [Ending a File], page 41.

@c comment Begin a comment in Texinfo. The rest of the line does not appear in either the Info file or the printed manual. A synonym for @comment. See Section 1.7 [Comments], page 9.

@cartouche

Highlight an example or quotation by drawing a box with rounded corners around it. Pair with @end cartouche. No effect in Info. See Section 10.13 [Drawing Cartouches Around Examples], page 87.)

@center line-of-text Center the line of text following the command. See Section 3.4.2 [@center], page 32.

@centerchap line-of-text Like @chapter, but centers the chapter title. See Section 5.4 [@chapter], page 45.

@chapheading title

Print a chapter-like heading in the text, but not in the table of contents of a printed manual. In Info, the title is underlined with asterisks. See Section 5.6

[@majorheading and @chapheading], page 46.
@chapter title

Begin a chapter. The chapter title appears in the table of contents of a printed manual. In Info, the title is underlined with asterisks. See Section 5.4 [@chapter], page 45.

@cindex entry

Add entry to the index of concepts. See Section 12.1 [Defining the Entries of an Index], page 96.

@cite{reference} Highlight the name of a book or other reference that lacks a companion Info file. See Section 9.1.12 [@cite], page 77.

@clear flag

Unset flag, preventing the Texinfo formatting commands from formatting text between subsequent pairs of @ifset flag and @end ifset commands, and preventing @value{flag} from expanding to the value to which flag is set. See Section 16.4 [@set @clear @value], page 131.

@code{sample-code} Highlight text that is an expression, a syntactically complete token of a program, or a program name. See Section 9.1.1 [@code], page 72.

@command{command-name} Indicate a command name, such as ls. See Section 9.1.9 [@command], page 77.

@comment comment

Begin a comment in Texinfo. The rest of the line does not appear in either the Info file or the printed manual. A synonym for @c. See Section 1.7 [Comments], page 9.

@contents

Print a complete table of contents. Has no effect in Info, which uses menus instead. See Section 4.2 [Generating a Table of Contents], page 42.

@copyright{}

Generate a copyright symbol. See Section 13.5.2 [@copyright], page 105.

@defcodeindex index-name

Define a new index and its indexing command. Print entries in an @code font. See Section 12.5 [Defining New Indices], page 99.

@defcv category class name @defcvx category class name

Format a description for a variable associated with a class in object-oriented programming. Takes three arguments: the category of thing being defined, the class to which it belongs, and its name. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@deffn category name arguments... @deffnx category name arguments...

Format a description for a function, interactive command, or similar entity that may take arguments. @deffn takes as arguments the category of entity being described, the name of this particular entity, and its arguments, if any. See Chapter 15 [Definition Commands], page 116.

@defindex index-name Define a new index and its indexing command. Print entries in a roman font. See Section 12.5 [Defining New Indices], page 99.

@definfoenclose newcmd, before, after,

Create new @-command newcmd for Info that marks text by enclosing it in strings that precede and follow the text. See Section 18.5 [definfoenclose], page 140.

@defivar class instance-variable-name @defivarx class instance-variable-name

This command formats a description for an instance variable in object-oriented programming. The command is equivalent to ‘@defcv {Instance Variable} ...’. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@defmac macroname arguments... @defmacx macroname arguments...

Format a description for a macro. The command is equivalent to ‘@deffn Macro ...’. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@defmethod class method-name arguments... @defmethodx class method-name arguments...

Format a description for a method in object-oriented programming. The command is equivalent to ‘@defop Method ...’. Takes as arguments the name of the class of the method, the name of the method, and its arguments, if any. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@defop category class name arguments... @defopx category class name arguments...

Format a description for an operation in object-oriented programming. @defop takes as arguments the overall name of the category of operation, the name of the class of the operation, the name of the operation, and its arguments, if any. See Chapter 15 [Definition Commands], page 116, and Section 15.4.5 [Abstract Objects], page 123.

@defopt option-name @defoptx option-name

Format a description for a user option. The command is equivalent to ‘@defvr {User Option} ...’. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@defspec special-form-name arguments... @defspecx special-form-name arguments...

Format a description for a special form. The command is equivalent to ‘@deffn {Special Form} ...’. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@deftp category name-of-type attributes... @deftpx category name-of-type attributes...

Format a description for a data type. @deftp takes as arguments the category, the name of the type (which is a word like ‘int’ or ‘float’), and then the names of attributes of objects of that type. See Chapter 15 [Definition Commands], page 116, and Section 15.4.6 [Data Types], page 126.

@deftypefn classification data-type name arguments... @deftypefnx classification data-type name arguments...

Format a description for a function or similar entity that may take arguments and that is typed. @deftypefn takes as arguments the classification of entity being described, the type, the name of the entity, and its arguments, if any. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@deftypefun data-type function-name arguments... @deftypefunx data-type function-name arguments...

Format a description for a function in a typed language. The command is equivalent to ‘@deftypefn Function ...’. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@deftypeivar class data-type variable-name @deftypeivarx class data-type variable-name

Format a description for a typed instance variable in object-oriented programming. See Chapter 15 [Definition Commands], page 116, and Section 15.4.5 [Abstract Objects], page 123.

@deftypemethod class data-type method-name arguments... @deftypemethodx class data-type method-name arguments...

Format a description for a typed method in object-oriented programming. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@deftypeop category class data-type name arguments... @deftypeopx category class data-type name arguments...

Format a description for a typed operation in object-oriented programming. See Chapter 15 [Definition Commands], page 116, and Section 15.4.5 [Abstract Objects], page 123.

@deftypevar data-type variable-name @deftypevarx data-type variable-name

Format a description for a variable in a typed language. The command is equivalent to ‘@deftypevr Variable ...’. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@deftypevr classification data-type name @deftypevrx classification data-type name

Format a description for something like a variable in a typed language—an entity that records a value. Takes as arguments the classification of entity being described, the type, and the name of the entity. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@defun function-name arguments... @defunx function-name arguments...

Format a description for functions. The command is equivalent to ‘@deffn Function ...’. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@defvar variable-name @defvarx variable-name

Format a description for variables. The command is equivalent to ‘@defvr Variable ...’. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@defvr category name @defvrx category name

Format a description for any kind of variable. @defvr takes as arguments the category of the entity and the name of the entity. See Chapter 15 [Definition Commands], page 116, and Section 15.3 [Def Cmds in Detail], page 118.

@detailmenu

Avoid makeinfo confusion stemming from the detailed node listing in a master menu. See Section 3.5.2 [Master Menu Parts], page 37.

@dfn{term}

Highlight the introductory or defining use of a term. See Section 9.1.11 [@dfn], page 77.

@dircategory dirpart

Specify a part of the Info directory menu where this file’s entry should go. See Section 20.2.4 [Installing Dir Entries], page 164.

@direntry

Begin the Info directory menu entry for this file. Pair with @end direntry. See Section 20.2.4 [Installing Dir Entries], page 164.

@display Begin a kind of example. Like @example (indent text, do not fill), but do not select a new font. Pair with @end display. See Section 10.8 [@display], page 85.

@dmn{dimension} Format a unit of measure, as in 12 pt. Causes TEX to insert a thin space before dimension. No effect in Info. See Section 13.2.4 [@dmn], page 103.

@documentdescription

Set the document description text, included in the HTML output. Pair with

@end documentdescription. See Section 3.6.1 [@documentdescription],

page 38.

@documentencoding enc

Declare the input encoding to be enc. See Section 17.2 [@documentencoding], page 136.

@documentlanguage CC

Declare the document language as the two-character ISO-639 abbreviation CC. See Section 17.1 [@documentlanguage], page 135.

@dotaccent{c}

Generate a dot accent over the character c, as in ȯ. See Section 13.3 [Inserting Accents], page 103.

@dots{} Insert an ellipsis: ‘...’. See Section 13.4.1 [@dots], page 104.

@email{address[, displayed-text]} Indicate an electronic mail address. See Section 9.1.15 [@email], page 78.

@emph{text}

Highlight text; text is displayed in italics in printed output, and surrounded by asterisks in Info. See Section 9.2 [Emphasizing Text], page 78.

@end environment Ends environment, as in ‘@end example’. See Section 1.5 [@-commands], page 7.

@env{environment-variable} Indicate an environment variable name, such as PATH. See Section 9.1.7 [@env], page 76.

@enddots{}

Generate an end-of-sentence of ellipsis, like this …. See Section 13.4.1 [@dots{}], page 104.

@enumerate [number-or-letter] Begin a numbered list, using @item for each entry. Optionally, start list with number-or-letter. Pair with @end enumerate. See Section 11.2 [@enumerate], page 91.

@equiv{} Indicate to the reader the exact equivalence of two forms with a glyph: ‘ ≡ ’. See Section 13.9.6 [Equivalence], page 108.

@error{} Indicate to the reader with a glyph that the following text is an error message: ‘ error ’. See Section 13.9.5 [Error Glyph], page 108.

@evenfooting [left] @| [center] @| [right]
@evenheading [left] @| [center] @| [right]

Specify page footings resp. headings for even-numbered (left-hand) pages. See Section E.3 [How to Make Your Own Headings], page 202.

@everyfooting [left] @| [center] @| [right]
@everyheading [left] @| [center] @| [right]

Specify page footings resp. headings for every page. Not relevant to Info. See Section E.3 [How to Make Your Own Headings], page 202.

@example Begin an example. Indent text, do not fill, and select fixed-width font. Pair with @end example. See Section 10.3 [@example], page 82.

@exampleindent indent

Indent example-like environments by indent number of spaces (perhaps 0). See Section 3.6.4 [Paragraph Indenting], page 40.

@exclamdown{}

Produce an upside-down exclamation point. See Section 13.3 [Inserting Accents], page 103.

@exdent line-of-text Remove any indentation a line might have. See Section 10.10 [Undoing the Indentation of a Line], page 86.

@expansion{}

Indicate the result of a macro expansion to the reader with a special glyph: ‘↦’. See Section 13.9.3 [↦ Indicating an Expansion], page 107.

@file{filename}

Highlight the name of a file, buffer, node, or directory. See Section 9.1.8 [@file], page 76.

@finalout

Prevent TEX from printing large black warning rectangles beside over-wide lines. See Section 19.10 [Overfull hboxes], page 149.

@findex entry

Add entry to the index of functions. See Section 12.1 [Defining the Entries of an Index], page 96.

@flushleft
@flushright

Left justify every line but leave the right end ragged. Leave font as is. Pair with

@end flushleft. @flushright analogous. See Section 10.11 [@flushleft and
@flushright], page 86.

@footnote{text-of-footnote} Enter a footnote. Footnote text is printed at the bottom of the page by TEX; Info may format in either ‘End’ node or ‘Separate’ node style. See Section 13.10 [Footnotes], page 109.

@footnotestyle style

Specify an Info file’s footnote style, either ‘end’ for the end node style or ‘separate’ for the separate node style. See Section 13.10 [Footnotes], page 109.

@format Begin a kind of example. Like @display, but do not narrow the margins. Pair with @end format. See Section 10.3 [@example], page 82.

@ftable formatting-command Begin a two-column table, using @item for each entry. Automatically enter each of the items in the first column into the index of functions. Pair with @end ftable. The same as @table, except for indexing. See Section 11.3.1 [@ftable and @vtable], page 93.

@group Hold text together that must appear on one printed page. Pair with @end group. Not relevant to Info. See Section 14.6 [@group], page 115.

@H{c} Generate the long Hungarian umlaut accent over c, as in ő.

@heading title

Print an unnumbered section-like heading in the text, but not in the table of contents of a printed manual. In Info, the title is underlined with equal signs. See Section 5.8 [Section Commands], page 47.

@headings on-off-single-double Turn page headings on or off, and/or specify single-sided or double-sided page headings for printing. See Section 3.4.6 [The @headings Command], page 35.

@html Enter HTML completely. Pair with @end html. See Section 16.3 [Raw Formatter Commands], page 130.

@hyphenation{hy-phen-a-ted words} Explicitly define hyphenation points. See Section 14.2 [@- and @hyphenation], page 114.

@i{text} Print text in italic font. No effect in Info. See Section 9.2.3 [Fonts], page 80.

@ifclear flag

If flag is cleared, the Texinfo formatting commands format text between @ifclear flag and the following @end ifclear command. See Section 16.4

[@set @clear @value], page 131.
@ifhtml

@ifinfo Begin a stretch of text that will be ignored by TEX when it typesets the printed manual. @ifhtml text appears only in the HTML output. @ifinfo output appears in both Info and (for historical compatibility) plain text output . Pair with @end ifhtml resp. @end ifinfo. See Chapter 16 [Conditionals], page 129.

@ifnothtml
@ifnotinfo
@ifnotplaintext
@ifnottex

Begin a stretch of text that will be ignored in one output format but not the others. The text appears in the formats not specified: @ifnothtml text is omitted from html output, etc. The exception is @ifnotinfo text, which is omitted from plain text output as well as Info output. Pair with @end ifnothtml resp.

@end ifnotinfo resp. @end ifnotplaintext resp. @end ifnottex. See Chap-

ter 16 [Conditionals], page 129.

@ifplaintext

Begin a stretch of text that appears only in the plain text output. Pair with @end ifplaintext. See Chapter 16 [Conditionals], page 129.

@ifset flag

If flag is set, the Texinfo formatting commands format text between @ifset flag and the following @end ifset command. See Section 16.4 [@set @clear @value], page 131.

@iftex Begin a stretch of text that will not appear in the Info file, but will be processed only by TEX. Pair with @end iftex. See Chapter 16 [Conditionally Visible Text], page 129.

@ignore Begin a stretch of text that will not appear in either the Info file or the printed output. Pair with @end ignore. See Section 1.7 [Comments and Ignored Text], page 9.

@image{filename, [width], [height], [alt], [ext]} Include graphics image in external filename scaled to the given width and/or height, using alt text and looking for ‘filename.ext’ in HTML. See Section 13.11 [Images], page 111.

@include filename

Incorporate the contents of the file filename into the Info file or printed document. See Appendix D [Include Files], page 196.

@inforef{node-name, [entry-name], info-file-name} Make a cross reference to an Info file for which there is no printed manual. See Section 8.7 [Cross references using @inforef], page 69.

\input macro-definitions-file Use the specified macro definitions file. This command is used only in the first line of a Texinfo file to cause TEX to make use of the ‘texinfo’ macro definitions file. The backslash in \input is used instead of an @ because TEX does not recognize @ until after it has read the definitions file. See Section 3.2 [Texinfo File Header], page 27.

@item Indicate the beginning of a marked paragraph for @itemize and @enumerate; indicate the beginning of the text of a first column entry for @table, @ftable, and @vtable. See Chapter 11 [Lists and Tables], page 89.

@itemize mark-generating-character-or-command Produce a sequence of indented paragraphs, with a mark inside the left margin at the beginning of each paragraph. Pair with @end itemize. See Section 11.1 [@itemize], page 89.

@itemx Like @item but do not generate extra vertical space above the item text. See Section 11.3.2 [@itemx], page 93.

@kbd{keyboard-characters} Indicate text that is characters of input to be typed by users. See Section 9.1.2 [@kbd], page 73.

@kbdinputstyle style

Specify when @kbd should use a font distinct from @code. See Section 9.1.2 [@kbd], page 73.

@key{key-name} Indicate a name for a key on a keyboard. See Section 9.1.3 [@key], page 74.

@kindex entry

Add entry to the index of keys. See Section 12.1 [Defining the Entries of an Index], page 96.

@L{}

@l{} Generate the uppercase and lowercase Polish suppressed-L letters, respectively: Ł, ł.

@lisp Begin an example of Lisp code. Indent text, do not fill, and select fixed-width font. Pair with @end lisp. See Section 10.6 [@lisp], page 84.

@lowersections

Change subsequent chapters to sections, sections to subsections, and so on. See

Section 5.12 [@raisesections and @lowersections], page 49.

@macro macroname {params} Define a new Texinfo command @macroname{params}. Only supported by makeinfo and texi2dvi. See Section 18.1 [Defining Macros], page 137.

@majorheading title

Print a chapter-like heading in the text, but not in the table of contents of a printed manual. Generate more vertical whitespace before the heading than the @chapheading command. In Info, the chapter heading line is underlined with asterisks. See Section 5.6 [@majorheading and @chapheading], page 46.

@math{mathematical-expression} Format a mathematical expression. See Section 13.8 [@math: Inserting Mathematical Expressions], page 106.

@menu Mark the beginning of a menu of nodes in Info. No effect in a printed manual. Pair with @end menu. See Chapter 7 [Menus], page 57.

@minus{} Generate a minus sign, ‘−’. See Section 13.7 [@minus], page 105.

@multitable column-width-spec Begin a multi-column table. Pair with @end multitable. See Section 11.4.1 [Multitable Column Widths], page 94.

@need n Start a new page in a printed manual if fewer than n mils (thousandths of an inch) remain on the current page. See Section 14.7 [@need], page 115.

@node name, next, previous, up Define the beginning of a new node in Info, and serve as a locator for references for TEX. See Section 6.3 [@node], page 52.

@noindent

Prevent text from being indented as if it were a new paragraph. See Section 10.12 [@noindent], page 87.

@novalidate

Suppress validation of node references, omit creation of auxiliary files with TEX. Use before @setfilename. See Section 20.1.4 [Pointer Validation], page 156.

@O{}

@o{} Generate the uppercase and lowercase O-with-slash letters, respectively: Ø, ø.

@oddfooting [left] @| [center] @| [right]
@oddheading [left] @| [center] @| [right]

Specify page footings resp. headings for odd-numbered (right-hand) pages. See Section E.3 [How to Make Your Own Headings], page 202.

@OE{}

@oe{} Generate the uppercase and lowercase OE ligatures, respectively: Œ, œ. See Section 13.3 [Inserting Accents], page 103.

@option{option-name} Indicate a command-line option, such as ‘-l’ or ‘--help’. See Section 9.1.10 [@option], page 77.

@page Start a new page in a printed manual. No effect in Info. See Section 14.5 [@page], page 115.

@pagesizes [width][, height]

Change page dimensions. See Section 19.13 [pagesizes], page 150.

@paragraphindent indent

Indent paragraphs by indent number of spaces (perhaps 0); preserve source file indentation if indent is asis. See Section 3.6.3 [Paragraph Indenting], page 39.

@pindex entry

Add entry to the index of programs. See Section 12.1 [Defining the Entries of an Index], page 96.

@point{} Indicate the position of point in a buffer to the reader with a glyph: ‘⋆’. See Section 13.9.7 [Indicating Point in a Buffer], page 108.

@pounds{}

Generate the pounds sterling currency sign. See Section 13.6 [@pounds{}], page 105.

@print{} Indicate printed output to the reader with a glyph: ‘ ⊣ ’. See Section 13.9.4 [Print Glyph], page 107.

@printindex index-name

Print an alphabetized two-column index in a printed manual or generate an alphabetized menu of index entries for Info. See Section 4.1 [Printing Indices & Menus], page 41.

@pxref{node-name, [entry], [topic-or-title], [info-file], [manual]} Make a reference that starts with a lower case ‘see’ in a printed manual. Use within parentheses only. Do not follow command with a punctuation mark— the Info formatting commands automatically insert terminating punctuation as needed. Only the first argument is mandatory. See Section 8.6 [@pxref], page 68.

@questiondown{}

Generate an upside-down question mark. See Section 13.3 [Inserting Accents], page 103.

@quotation

Narrow the margins to indicate text that is quoted from another real or imaginary work. Write command on a line of its own. Pair with @end quotation. See Section 10.2 [@quotation], page 82.

@r{text} Print text in roman font. No effect in Info. See Section 9.2.3 [Fonts], page 80.

@raisesections

Change subsequent sections to chapters, subsections to sections, and so on. See

Section 5.12 [@raisesections and @lowersections], page 49.

@ref{node-name, [entry], [topic-or-title], [info-file], [manual]} Make a reference. In a printed manual, the reference does not start with a ‘See’. Follow command with a punctuation mark. Only the first argument is mandatory. See Section 8.5 [@ref], page 67.

@refill In Info, refill and indent the paragraph after all the other processing has been done. No effect on TEX, which always refills. This command is no longer needed, since all formatters now automatically refill. See Appendix G [Refilling Paragraphs], page 212.

@result{}

Indicate the result of an expression to the reader with a special glyph: ‘⇒’. See Section 13.9.2 [@result], page 107.

@ringaccent{c}

Generate a ring accent over the next character, as in o̊. See Section 13.3 [Inserting Accents], page 103.

@samp{text}

Highlight text that is a literal example of a sequence of characters. Used for single characters, for statements, and often for entire shell commands. See Section 9.1.4 [@samp], page 74.

@sc{text} Set text in a printed output in the small caps font and set text in the Info file in uppercase letters. See Section 9.2.2 [Smallcaps], page 79.

@section title

Begin a section within a chapter. In a printed manual, the section title is numbered and appears in the table of contents. In Info, the title is underlined with equal signs. See Section 5.7 [@section], page 46.

@set flag [string] Make flag active, causing the Texinfo formatting commands to format text between subsequent pairs of @ifset flag and @end ifset commands. Optionally, set value of flag to string. See Section 16.4 [@set @clear @value], page 131.

@setchapternewpage on-off-odd

Specify whether chapters start on new pages, and if so, whether on oddnumbered (right-hand) new pages. See Section 3.6.2 [@setchapternewpage], page 38.

@setcontentsaftertitlepage

Put the table of contents after the ‘@end titlepage’ even if the @contents command is not there. See Section 4.2 [Contents], page 42.

@setfilename info-file-name Provide a name to be used by the Info file. This command is essential for TEX formatting as well, even though it produces no output. See Section 3.2.3

[@setfilename], page 28.
@setshortcontentsaftertitlepage

Place the short table of contents after the ‘@end titlepage’ command even if the @shortcontents command is not there. See Section 4.2 [Contents], page 42.

@settitle title

Provide a title for page headers in a printed manual, and the default document description for HTML ‘<head>’. See Section 3.2.4 [@settitle], page 29.

@shortcontents

Print a short table of contents. Not relevant to Info, which uses menus rather than tables of contents. A synonym for @summarycontents. See Section 4.2 [Generating a Table of Contents], page 42.

@shorttitlepage title

Generate a minimal title page. See Section 3.4.1 [@titlepage], page 31.

@smallbook

Cause TEX to produce a printed manual in a 7 by 9.25 inch format rather than the regular 8.5 by 11 inch format. See Section 19.11 [Printing Small Books], page 150. Also, see Section 10.7 [small], page 84.

@smalldisplay

Begin a kind of example. Like @smallexample (narrow margins, no filling), but do not select the fixed-width font. Pair with @end smalldisplay. See Section 10.7 [small], page 84.

@smallexample

Indent text to indicate an example. Do not fill, select fixed-width font, narrow the margins. In printed manuals, print text in a smaller font than with @example. Pair with @end smallexample. See Section 10.7 [small], page 84.

@smallformat

Begin a kind of example. Like @smalldisplay, but do not narrow the margins. Pair with @end smallformat. See Section 10.7 [small], page 84.

@smalllisp

Begin an example of Lisp code. Same as @smallexample. Pair with @end smalllisp. See Section 10.7 [small], page 84.

@sp n Skip n blank lines. See Section 14.4 [@sp], page 114.

@ss{} Generate the German sharp-S es-zet letter, ß. See Section 13.3 [Inserting Accents], page 103.

@strong {text}

Emphasize text by typesetting it in a bold font for the printed manual and by surrounding it with asterisks for Info. See Section 9.2.1 [Emphasizing Text], page 78.

@subheading title

Print an unnumbered subsection-like heading in the text, but not in the table of contents of a printed manual. In Info, the title is underlined with hyphens. See

Section 5.10 [@unnumberedsubsec @appendixsubsec @subheading], page 48.
@subsection title

Begin a subsection within a section. In a printed manual, the subsection title is numbered and appears in the table of contents. In Info, the title is underlined with hyphens. See Section 5.9 [@subsection], page 47.

@subsubheading title

Print an unnumbered subsubsection-like heading in the text, but not in the table of contents of a printed manual. In Info, the title is underlined with periods. See Section 5.11 [The ‘subsub’ Commands], page 48.

@subsubsection title

Begin a subsubsection within a subsection. In a printed manual, the subsubsection title is numbered and appears in the table of contents. In Info, the title is underlined with periods. See Section 5.11 [The ‘subsub’ Commands], page 48.

@subtitle title

In a printed manual, set a subtitle in a normal sized font flush to the righthand side of the page. Not relevant to Info, which does not have title pages. See Section 3.4.3 [@title @subtitle and @author Commands], page 33.

@summarycontents

Print a short table of contents. Not relevant to Info, which uses menus rather than tables of contents. A synonym for @shortcontents. See Section 4.2 [Generating a Table of Contents], page 42.

@syncodeindex from-index into-index Merge the index named in the first argument into the index named in the second argument, printing the entries from the first index in @code font. See Section 12.4 [Combining Indices], page 98.

@synindex from-index into-index Merge the index named in the first argument into the index named in the second argument. Do not change the font of from-index entries. See Section 12.4 [Combining Indices], page 98.

@t{text} Print text in a fixed-width, typewriter-like font. No effect in Info. See Section 9.2.3 [Fonts], page 80.

@tab Separate columns in a multitable. See Section 11.4.2 [Multitable Rows], page 94.

@table formatting-command Begin a two-column table, using @item for each entry. Write each first column entry on the same line as @item. First column entries are printed in the font resulting from formatting-command. Pair with @end table. See Section 11.3 [Making a Two-column Table], page 92. Also see Section 11.3.1 [@ftable and @vtable], page 93, and Section 11.3.2 [@itemx], page 93.

@TeX\{\} Insert the logo TEX. See Section 13.5 [Inserting TEX and \begin{matrix}c \\ ◯ \end{matrix}], page 105.

@tex Enter TEX completely. Pair with @end tex. See Section 16.3 [Raw Formatter Commands], page 130.

@thischapter
@thischaptername
@thisfile
@thispage
@thistitle

Only allowed in a heading or footing. Stands for the number and name of the current chapter (in the format ‘Chapter 1: Title’), the chapter name only, the filename, the current page number, and the title of the document, respectively. See Section E.3 [How to Make Your Own Headings], page 202.

@tieaccent{cc}

Generate a tie-after accent over the next two characters cc, as in ‘oo’. See Section 13.3 [Inserting Accents], page 103.

@tindex entry

Add entry to the index of data types. See Section 12.1 [Defining the Entries of an Index], page 96.

@title title

In a printed manual, set a title flush to the left-hand side of the page in a larger than normal font and underline it with a black rule. Not relevant to Info, which does not have title pages. See Section 3.4.3 [The @title @subtitle and @author Commands], page 33.

@titlefont{text}

In a printed manual, print text in a larger than normal font. Not relevant to Info, which does not have title pages. See Section 3.4.2 [The @titlefont @center and @sp Commands], page 32.

@titlepage

Indicate to Texinfo the beginning of the title page. Write command on a line of its own. Pair with @end titlepage. Nothing between @titlepage and @end titlepage appears in Info. See Section 3.4.1 [@titlepage], page 31.

@today{} Insert the current date, in ‘1 Jan 1900’ style. See Section E.3 [How to Make Your Own Headings], page 202.

@top title In a Texinfo file to be formatted with makeinfo, identify the topmost @node in the file, which must be written on the line immediately preceding the @top command. Used for makeinfo’s node pointer insertion feature. The title is underlined with asterisks. Both the @node line and the @top line normally should be enclosed by @ifnottex and @end ifnottex. In TEX and texinfoformat-buffer, the @top command is merely a synonym for @unnumbered. See Section 6.4 [Creating Pointers with makeinfo], page 55.

@u{c}
@ubaraccent{c}
@udotaccent{c}

Generate a breve, underbar, or underdot accent, respectively, over or under the character c, as in ŏ, ō, o.. See Section 13.3 [Inserting Accents], page 103.

@unnumbered title

In a printed manual, begin a chapter that appears without chapter numbers of any kind. The title appears in the table of contents of a printed manual. In Info, the title is underlined with asterisks. See Section 5.5 [@unnumbered and

@appendix], page 46.
@unnumberedsec title

In a printed manual, begin a section that appears without section numbers of any kind. The title appears in the table of contents of a printed manual. In Info, the title is underlined with equal signs. See Section 5.8 [Section Commands], page 47.

@unnumberedsubsec title

In a printed manual, begin an unnumbered subsection within a chapter. The title appears in the table of contents of a printed manual. In Info, the title is underlined with hyphens. See Section 5.10 [@unnumberedsubsec

@appendixsubsec @subheading], page 48.
@unnumberedsubsubsec title

In a printed manual, begin an unnumbered subsubsection within a chapter. The title appears in the table of contents of a printed manual. In Info, the title is underlined with periods. See Section 5.11 [The ‘subsub’ Commands], page 48.

@uref{url[, displayed-text][, replacement} Define a cross reference to an external uniform resource locator for the World Wide Web. See Section 8.8 [@uref], page 69.

@url{url} Indicate text that is a uniform resource locator for the World Wide Web. See Section 9.1.14 [@url], page 78.

@v{c} Generate check accent over the character c, as in ǒ. See Section 13.3 [Inserting Accents], page 103.

@value{flag}

Replace flag with the value to which it is set by @set flag. See Section 16.4

[@set @clear @value], page 131.

@var{metasyntactic-variable} Highlight a metasyntactic variable, which is something that stands for another piece of text. See Section 9.1.6 [Indicating Metasyntactic Variables], page 75.

@verb{delim literal delim} Output literal, delimited by the single character delim, exactly as is (in the fixed-width font), including any whitespace or Texinfo special characters. See Section 9.1.5 [verb], page 75.

@verbatim

Output the text of the environment exactly as is (in the fixed-width font). Pair with @end verbatim. See Section 10.4 [verbatim], page 83.

@verbatiminclude filename

Output the contents of filename exactly as is (in the fixed-width font). See Section 10.5 [verbatiminclude], page 84.

@vindex entry

Add entry to the index of variables. See Section 12.1 [Defining the Entries of an Index], page 96.

@vskip amount

In a printed manual, insert whitespace so as to push text on the remainder of the page towards the bottom of the page. Used in formatting the copyright page with the argument ‘0pt plus 1filll’. (Note spelling of ‘filll’.) @vskip may be used only in contexts ignored for Info. See Section 3.4.4 [Copyright], page 34.

@vtable formatting-command Begin a two-column table, using @item for each entry. Automatically enter each of the items in the first column into the index of variables. Pair with @end vtable. The same as @table, except for indexing. See Section 11.3.1 [@ftable and @vtable], page 93.

@w{text} Prevent text from being split across two lines. Do not end a paragraph that uses @w with an @refill command. See Section 14.3 [@w], page 114.

@xref{node-name, [entry], [topic-or-title], [info-file], [manual]} Make a reference that starts with ‘See’ in a printed manual. Follow command with a punctuation mark. Only the first argument is mandatory. See Section 8.3 [@xref], page 63.

Appendix B Tips and Hints

Here are some tips for writing Texinfo documentation:

• Write in the present tense, not in the past or the future.

• Write actively! For example, write “We recommend that …” rather than “It is recommended that …”.

• Use 70 or 72 as your fill column. Longer lines are hard to read.

• Include a copyright notice and copying permissions.

Index, Index, Index!

Write many index entries, in different ways. Readers like indices; they are helpful and convenient.

Although it is easiest to write index entries as you write the body of the text, some people prefer to write entries afterwards. In either case, write an entry before the paragraph to which it applies. This way, an index entry points to the first page of a paragraph that is split across pages.

Here are more hints we have found valuable:

• Write each index entry differently, so each entry refers to a different place in the document.

• Write index entries only where a topic is discussed significantly. For example, it is not useful to index “debugging information” in a chapter on reporting bugs. Someone who wants to know about debugging information will certainly not find it in that chapter.

• Consistently capitalize the first word of every concept index entry, or else consistently use lower case. Terse entries often call for lower case; longer entries for capitalization. Whichever case convention you use, please use one or the other consistently! Mixing the two styles looks bad.

• Always capitalize or use upper case for those words in an index for which this is proper, such as names of countries or acronyms. Always use the appropriate case for casesensitive names, such as those in C or Lisp.

• Write the indexing commands that refer to a whole section immediately after the section command, and write the indexing commands that refer to a paragraph before that paragraph. In the example that follows, a blank line comes after the index entry for “Leaping”:

     @section The Dog and the Fox
     @cindex Jumping, in general
     @cindex Leaping
     @cindex Dog, lazy, jumped over
     @cindex Lazy dog jumped over
     @cindex Fox, jumps over dog
     @cindex Quick fox jumps over dog
     The quick brown fox jumps over the lazy dog.

(Note that the example shows entries for the same concept that are written in different ways—‘Lazy dog’, and ‘Dog, lazy’—so readers can look up the concept in different ways.)

Blank Lines

• Insert a blank line between a sectioning command and the first following sentence or paragraph, or between the indexing commands associated with the sectioning command and the first following sentence or paragraph, as shown in the tip on indexing. Otherwise, a formatter may fold title and paragraph together.

• Always insert a blank line before an @table command and after an @end table command; but never insert a blank line after an @table command or before an @end table command. For example,

     Types of fox:
     @table @samp
     @item Quick
     Jump over lazy dogs.
     @item Brown
     Also jump over lazy dogs.
     @end table
     @noindent
     On the other hand, ...

Insert blank lines before and after @itemize … @end itemize and @enumerate …

@end enumerate in the same way.

Complete Phrases

Complete phrases are easier to read than …

• Write entries in an itemized list as complete sentences; or at least, as complete phrases. Incomplete expressions … awkward … like this.

• Write the prefatory sentence or phrase for a multi-item list or table as a complete expression. Do not write “You can set:”; instead, write “You can set these variables:”. The former expression sounds cut off.

Editions, Dates and Versions

Include edition numbers, version numbers, and dates in the @copying text (for people reading the Texinfo file, and for the legal copyright in the output files). Then use @insertcopying in the @titlepage section (for people reading the printed output) and the Top node (for people reading the online output).

It is easiest to do this using @set and @value. See Section 16.4.3 [@value Example], page 133, and Section C.2 [GNU Sample Texts], page 192.

Definition Commands

Definition commands are @deffn, @defun, @defmac, and the like, and enable you to write descriptions in a uniform format.

• Write just one definition command for each entity you define with a definition command. The automatic indexing feature creates an index entry that leads the reader to the definition.

• Use @table … @end table in an appendix that contains a summary of functions, not @deffn or other definition commands.

Capitalization

• Capitalize “Texinfo”; it is a name. Do not write the ‘x’ or ‘i’ in upper case.

• Capitalize “Info”; it is a name.

• Write TEX using the @TeX{} command. Note the uppercase ‘T’ and ‘X’. This command causes the formatters to typeset the name according to the wishes of Donald Knuth, who wrote TEX.

Spaces

Do not use spaces to format a Texinfo file, except inside of @example … @end example and similar commands.

For example, TEX fills the following:

         @kbd{C-x v}
         @kbd{M-x vc-next-action}
            Perform the next logical operation
            on the version-controlled file
            corresponding to the current buffer.

so it looks like this:

C-x v M-x vc-next-action Perform the next logical operation on the versioncontrolled file corresponding to the current buffer.

In this case, the text should be formatted with @table, @item, and @itemx, to create a table.

@code, @samp, @var, and ‘---’

• Use @code around Lisp symbols, including command names. For example,

         The main function is @code{vc-next-action}, ...

• Avoid putting letters such as ‘s’ immediately after an ‘@code’. Such letters look bad.

• Use @var around meta-variables. Do not write angle brackets around them.

• Use three hyphens in a row, ‘---’, to indicate a long dash. TEX typesets these as a long dash and the Info formatters reduce three hyphens to two.

Periods Outside of Quotes

Place periods and other punctuation marks outside of quotations, unless the punctuation is part of the quotation. This practice goes against publishing conventions in the United States, but enables the reader to distinguish between the contents of the quotation and the whole passage.

For example, you should write the following sentence with the period outside the end quotation marks:

     Evidently, `au' is an abbreviation for ``author''.

since ‘au’ does not serve as an abbreviation for ‘author.’ (with a period following the word).

Introducing New Terms

• Introduce new terms so that a reader who does not know them can understand them from context; or write a definition for the term. For example, in the following, the terms “check in”, “register” and “delta” are all appearing for the first time; the example sentence should be rewritten so they are understandable.

The major function assists you in checking in a file to your version control system and registering successive sets of changes to it as deltas.

• Use the @dfn command around a word being introduced, to indicate that the reader should not expect to know the meaning already, and should expect to learn the meaning from this passage.

@pxref

Absolutely never use @pxref except in the special context for which it is designed: inside parentheses, with the closing parenthesis following immediately after the closing brace. One formatter automatically inserts closing punctuation and the other does not. This means that the output looks right both in printed output and in an Info file, but only when the command is used inside parentheses.

Invoking from a Shell

You can invoke programs such as Emacs, GCC, and gawk from a shell. The documentation for each program should contain a section that describes this. Unfortunately, if the node names and titles for these sections are all different, they are difficult for users to find.

So, there is a convention to name such sections with a phrase beginning with the word ‘Invoking’, as in ‘Invoking Emacs’; this way, users can find the section easily.

ANSI C Syntax

When you use @example to describe a C function’s calling conventions, use the ANSI C syntax, like this:

     void dld_init (char *@var{path});

And in the subsequent discussion, refer to the argument values by writing the same argument names, again highlighted with @var.

Avoid the obsolete style that looks like this:

     #include <dld.h>
     dld_init (path)
     char *path;

Also, it is best to avoid writing #include above the declaration just to indicate that the function is declared in a header file. The practice may give the misimpression that the #include belongs near the declaration of the function. Either state explicitly which header file holds the declaration or, better yet, name the header file used for a group of functions at the beginning of the section that describes the functions.

Bad Examples

Here are several examples of bad writing to avoid: In this example, say, “ … you must @dfn{check in} the new version.” That flows

better.

When you are done editing the file, you must perform a @dfn{check in}. In the following example, say, “… makes a unified interface such as VC mode possible.” SCCS, RCS and other version-control systems all perform similar functions in broadly similar ways (it is this resemblance which makes a unified control mode like this possible). And in this example, you should specify what ‘it’ refers to: If you are working with other people, it assists in coordinating everyone’s changes so they do not step on each other.

And Finally …

• Pronounce TEX as if the ‘X’ were a Greek ‘chi’, as the last sound in the name ‘Bach’. But pronounce Texinfo as in ‘speck’: “teckinfo”.

• Write notes for yourself at the very end of a Texinfo file after the @bye. None of the formatters process text after the @bye; it is as if the text were within @ignore … @end

ignore.

Appendix C Sample Texinfo Files

The first example is from the first chapter (see Section 1.10 [Short Sample], page 11), given here in its entirety, without commentary. The second sample includes the full texts to be used in GNU manuals.

C.1 Short Sample

Here is a complete, short sample Texinfo file, without any commentary. You can see this file, with comments, in the first chapter. See Section 1.10 [Short Sample], page 11.

In a nutshell: The makeinfo program transforms a Texinfo source file such as this into an Info file or HTML; and TEX typesets it for a printed manual.

\input texinfo   @c -*-texinfo-*-
@c %**start of header
@setfilename sample.info
@settitle Sample Manual 1.0
@c %**end of header
@copying
This is a short example of a complete Texinfo file.
Copyright (C) 2002 Free Software Foundation, Inc.
@end copying
@titlepage
@title Sample Title
@page
@vskip 0pt plus 1filll
@insertcopying
@end titlepage
@c Output the table of the contents at the beginning.
@contents
@ifnottex
@node Top
@insertcopying
@end ifnottex
@menu
* First Chapter::    The first chapter is the
                       only chapter in this sample.
* Index::            Complete index.
@end menu

@node First Chapter
@chapter First Chapter
@cindex chapter, first
This is the first chapter.
@cindex index entry, another
Here is a numbered list.
@enumerate
@item
This is the first item.
@item
This is the second item.
@end enumerate

@node Index
@unnumbered Index
@printindex cp
@bye

C.2 GNU Sample Texts

Here is a sample Texinfo document with the full texts that should be used in GNU manuals.

As well as the legal texts, it also serves as a practical example of how many elements in a GNU system can affect the manual. If you’re not familiar with all these different elements, don’t worry. They’re not required and a perfectly good manual can be written without them. They’re included here nonetheless because many manuals do (or could) benefit from them.

See Section 1.10 [Short Sample], page 11, for a minimal example of a Texinfo file. See Chapter 3 [Beginning a File], page 26, for a full explanation of that minimal example.

Here are some notes on the example:

• The ‘$Id: texinfo.txi,v 1.219 2002/03/28 16:36:00 karl Exp $’ comment is for

CVS (see section “Overview” in Concurrent Versions System) or RCS (see rcsintro(1)) version control systems, which expand it into a string such as:

    $Id: texinfo.txi,v 1.219 2002/03/28 16:36:00 karl Exp $

(This is useful in all sources that use version control, not just manuals.)

• The ‘version.texi’ in the @include command is maintained automatically by Automake (see section “Introduction” in GNU Automake). It sets the ‘VERSION’ and ‘UPDATED’ values used elsewhere. If your distribution doesn’t use Automake, you can mimic these or equivalent settings.

• The @syncodeindex command reflects the recommendation to use only one index if at all possible, to make it easier for readers.

• The @dircategory is for constructing the Info directory. See Section 20.2.4 [Installing Dir Entries], page 164, which includes a variety of recommended category names.

• The ‘Invoking’ node is a GNU standard to help users find the basic information about command-line usage of a given program. See section “Manual Structure Details” in GNU Coding Standards.

• It is best to include the entire GNU Free Documentation License in a GNU manual, unless the manual is only a few pages long. Of course this sample is even shorter than that, but it includes the FDL anyway in order to show one conventional way of doing so. The ‘fdl.texi’ file is available on the GNU machines (and in the Texinfo and other GNU distributions). The FDL provides for omitting itself under certain conditions, but in that case the sample texts given here have to be modified. See Section J.1 [GNU Free Documentation License], page 215.

• If your manual has invariant sections (again, see the license itself for details), then don’t forget to include them.

Here is the sample document:

\input texinfo    @c -*-texinfo-*-
@comment $Id: texinfo.txi,v 1.219 2002/03/28 16:36:00 karl Exp $
@comment %**start of header
@setfilename sample.info
@include version.texi
@settitle GNU Sample @value{VERSION}
@syncodeindex pg cp
@comment %**end of header
@copying
This manual is for GNU Sample
(version @value{VERSION}, @value{UPDATED}),
which is an example in the Texinfo documentation.
Copyright @copyright{} 2002 Free Software Foundation, Inc.
@quotation
Permission is granted to copy, distribute and/or modify this document
under the terms of the GNU Free Documentation License, Version 1.1 or
any later version published by the Free Software Foundation; with no
Invariant Sections, with the Front-Cover texts being ``A GNU Manual,''
and with the Back-Cover Texts as in (a) below.  A copy of the
license is included in the section entitled ``GNU Free Documentation
License.''
(a) The FSF's Back-Cover Text is: ``You have freedom to copy and modify
this GNU Manual, like GNU software.  Copies published by the Free
Software Foundation raise funds for GNU development.''
@end quotation
@end copying

@dircategory Texinfo documentation system
@direntry
* sample: (sample)Invoking sample.
@end direntry
@titlepage
@title GNU Sample
@subtitle for version @value{VERSION}, @value{UPDATED}
@author A.U. Thor (@email{bug-texinfo@@gnu.org})
@page
@vskip 0pt plus 1filll
@insertcopying
@end titlepage
@contents
@ifnottex
@node Top
@top GNU Sample
@insertcopying
@end ifnottex
@menu
* Invoking sample::
* Copying This Manual::
* Index::
@end menu

@node Invoking sample
@chapter Invoking sample
@pindex sample
@cindex invoking @command{sample}
This is a sample manual.  There is no sample program to
invoke, but if there was, you could see its basic usage
and command line options here.

@node Copying This Manual
@appendix Copying This Manual
@menu
* GNU Free Documentation License::  License for copying this manual.
@end menu

@include fdl.texi
@node Index
@unnumbered Index
@printindex cp
@bye

Appendix D Include Files

When TEX or an Info formatting command sees an @include command in a Texinfo file, it processes the contents of the file named by the command and incorporates them into the DVI or Info file being created. Index entries from the included file are incorporated into the indices of the output file.

Include files let you keep a single large document as a collection of conveniently small parts.

D.1 How to Use Include Files

To include another file within a Texinfo file, write the @include command at the beginning of a line and follow it on the same line by the name of a file to be included. For example:

     @include buffers.texi

An included file should simply be a segment of text that you expect to be included as is into the overall or outer Texinfo file; it should not contain the standard beginning and end parts of a Texinfo file. In particular, you should not start an included file with a line saying ‘\input texinfo’; if you do, that phrase is inserted into the output file as is. Likewise, you should not end an included file with an @bye command; nothing after @bye is formatted.

In the past, you were required to write an @setfilename line at the beginning of an included file, but no longer. Now, it does not matter whether you write such a line. If an @setfilename line exists in an included file, it is ignored.

Conventionally, an included file begins with an @node line that is followed by an @chapter line. Each included file is one chapter. This makes it easy to use the regular node and menu creating and updating commands to create the node pointers and menus within the included file. However, the simple Emacs node and menu creating and updating commands do not work with multiple Texinfo files. Thus you cannot use these commands to fill in the ‘Next’, ‘Previous’, and ‘Up’ pointers of the @node line that begins the included file. Also, you cannot use the regular commands to create a master menu for the whole file. Either you must insert the menus and the ‘Next’, ‘Previous’, and ‘Up’ pointers by hand, or you must use the GNU Emacs Texinfo mode command, texinfo-multiple-files-update, that is designed for @include files.

D.2 texinfo-multiple-files-update

GNU Emacs Texinfo mode provides the texinfo-multiple-files-update command. This command creates or updates ‘Next’, ‘Previous’, and ‘Up’ pointers of included files as well as those in the outer or overall Texinfo file, and it creates or updates a main menu in the outer file. Depending whether you call it with optional arguments, the command updates only the pointers in the first @node line of the included files or all of them:

M-x texinfo-multiple-files-update

Called without any arguments:

− Create or update the ‘Next’, ‘Previous’, and ‘Up’ pointers of the first @node line in each file included in an outer or overall Texinfo file.

− Create or update the ‘Top’ level node pointers of the outer or overall file. − Create or update a main menu in the outer file.

C-u M-x texinfo-multiple-files-update

Called with C-u as a prefix argument:

− Create or update pointers in the first @node line in each included file. − Create or update the ‘Top’ level node pointers of the outer file. − Create and insert a master menu in the outer file. The master menu is

made from all the menus in all the included files.

C-u 8 M-x texinfo-multiple-files-update

Called with a numeric prefix argument, such as C-u 8:

− Create or update all the ‘Next’, ‘Previous’, and ‘Up’ pointers of all the

included files. − Create or update all the menus of all the included files. − Create or update the ‘Top’ level node pointers of the outer or overall file. − And then create a master menu in the outer file. This is similar to invoking texinfo-master-menu with an argument when you are working with just one file.

Note the use of the prefix argument in interactive use: with a regular prefix argument, just C-u, the texinfo-multiple-files-update command inserts a master menu; with a numeric prefix argument, such as C-u 8, the command updates every pointer and menu in all the files and then inserts a master menu.

D.3 Include File Requirements

If you plan to use the texinfo-multiple-files-update command, the outer Texinfo file that lists included files within it should contain nothing but the beginning and end parts of a Texinfo file, and a number of @include commands listing the included files. It should not even include indices, which should be listed in an included file of their own.

Moreover, each of the included files must contain exactly one highest level node (conventionally, @chapter or equivalent), and this node must be the first node in the included file. Furthermore, each of these highest level nodes in each included file must be at the same hierarchical level in the file structure. Usually, each is an @chapter, an @appendix, or an @unnumbered node. Thus, normally, each included file contains one, and only one, chapter or equivalent-level node.

The outer file should contain only one node, the ‘Top’ node. It should not contain any nodes besides the single ‘Top’ node. The texinfo-multiple-files-update command will not process them.

D.4 Sample File with @include

Here is an example of a complete outer Texinfo file with @include files within it before running texinfo-multiple-files-update, which would insert a main or master menu:

\input texinfo @c -*-texinfo-*-
@setfilename  include-example.info
@settitle Include Example
@setchapternewpage odd
@titlepage
@sp 12
@center @titlefont{Include Example}
@sp 2
@center by Whom Ever
@page
@vskip 0pt plus 1filll
Copyright @copyright{} 2002 Free Software Foundation, Inc.
@end titlepage
@ifinfo
@node Top, First, , (dir)
@top Master Menu
@end ifinfo
@include foo.texinfo
@include bar.texinfo
@include concept-index.texinfo
@summarycontents
@contents
@bye

An included file, such as ‘foo.texinfo’, might look like this:

@node First, Second, , Top
@chapter First Chapter
Contents of first chapter ...

The full contents of ‘concept-index.texinfo’ might be as simple as this:

@node Concept Index
@unnumbered Concept Index
@printindex cp

The outer Texinfo source file for The GNU Emacs Lisp Reference Manual is named ‘elisp.texi’. This outer file contains a master menu with 417 entries and a list of 41

@include files.

D.5 Evolution of Include Files

When Info was first created, it was customary to create many small Info files on one subject. Each Info file was formatted from its own Texinfo source file. This custom meant that Emacs did not need to make a large buffer to hold the whole of a large Info file when someone wanted information; instead, Emacs allocated just enough memory for the small Info file that contained the particular information sought. This way, Emacs could avoid wasting memory.

References from one file to another were made by referring to the file name as well as the node name. (See Section 7.5 [Referring to Other Info Files], page 59. Also, see Section 8.3.4 [@xref with Four and Five Arguments], page 66.)

Include files were designed primarily as a way to create a single, large printed manual out of several smaller Info files. In a printed manual, all the references were within the same document, so TEX could automatically determine the references’ page numbers. The Info formatting commands used include files only for creating joint indices; each of the individual Texinfo files had to be formatted for Info individually. (Each, therefore, required its own

@setfilename line.)

However, because large Info files are now split automatically, it is no longer necessary to keep them small.

Nowadays, multiple Texinfo files are used mostly for large documents, such as The GNU Emacs Lisp Reference Manual, and for projects in which several different people write different sections of a document simultaneously.

In addition, the Info formatting commands have been extended to work with the @include command so as to create a single large Info file that is split into smaller files if necessary. This means that you can write menus and cross references without naming the different Texinfo files.

Appendix E Page Headings

Most printed manuals contain headings along the top of every page except the title and copyright pages. Some manuals also contain footings. (Headings and footings have no meaning to Info, which is not paginated.)

Texinfo provides standard page heading formats for manuals that are printed on one side of each sheet of paper and for manuals that are printed on both sides of the paper. Typically, you will use these formats, but you can specify your own format if you wish.

In addition, you can specify whether chapters should begin on a new page, or merely continue the same page as the previous chapter; and if chapters begin on new pages, you can specify whether they must be odd-numbered pages.

By convention, a book is printed on both sides of each sheet of paper. When you open a book, the right-hand page is odd-numbered, and chapters begin on right-hand pages—a preceding left-hand page is left blank if necessary. Reports, however, are often printed on just one side of paper, and chapters begin on a fresh page immediately following the end of the preceding chapter. In short or informal reports, chapters often do not begin on a new page at all, but are separated from the preceding text by a small amount of whitespace.

The @setchapternewpage command controls whether chapters begin on new pages, and whether one of the standard heading formats is used. In addition, Texinfo has several heading and footing commands that you can use to generate your own heading and footing formats.

In Texinfo, headings and footings are single lines at the tops and bottoms of pages; you cannot create multiline headings or footings. Each header or footer line is divided into three parts: a left part, a middle part, and a right part. Any part, or a whole line, may be left blank. Text for the left part of a header or footer line is set flushleft; text for the middle part is centered; and, text for the right part is set flushright.

E.1 Standard Heading Formats

Texinfo provides two standard heading formats, one for manuals printed on one side of each sheet of paper, and the other for manuals printed on both sides of the paper.

By default, nothing is specified for the footing of a Texinfo file, so the footing remains blank.

The standard format for single-sided printing consists of a header line in which the lefthand part contains the name of the chapter, the central part is blank, and the right-hand part contains the page number.

A single-sided page looks like this:

        _______________________
       |                       |
       | chapter   page number |
       |                       |
       | Start of text ...     |
       | ...                   |
       |                       |

The standard format for two-sided printing depends on whether the page number is even or odd. By convention, even-numbered pages are on the left- and odd-numbered pages are on the right. (TEX will adjust the widths of the left- and right-hand margins. Usually, widths are correct, but during double-sided printing, it is wise to check that pages will bind properly—sometimes a printer will produce output in which the even-numbered pages have a larger right-hand margin than the odd-numbered pages.)

In the standard double-sided format, the left part of the left-hand (even-numbered) page contains the page number, the central part is blank, and the right part contains the title (specified by the @settitle command). The left part of the right-hand (odd-numbered) page contains the name of the chapter, the central part is blank, and the right part contains the page number.

Two pages, side by side as in an open book, look like this:

        ______________________________________________
       |                       |   |                       |
       | page number     title |   | chapter   page number |
       |                       |   |                       |
       | Start of text ...     |   | More  text ...        |
       | ...                   |   | ...                   |
       |                       |   |                       |

The chapter name is preceded by the word “Chapter”, the chapter number and a colon. This makes it easier to keep track of where you are in the manual.

E.2 Specifying the Type of Heading

TEX does not begin to generate page headings for a standard Texinfo file until it reaches the @end titlepage command. Thus, the title and copyright pages are not numbered. The @end titlepage command causes TEX to begin to generate page headings according to a standard format specified by the @setchapternewpage command that precedes the

@titlepage section.

There are four possibilities:

No @setchapternewpage command

Cause TEX to specify the single-sided heading format, with chapters on new pages. This is the same as @setchapternewpage on.

@setchapternewpage on

Specify the single-sided heading format, with chapters on new pages.

@setchapternewpage off

Cause TEX to start a new chapter on the same page as the last page of the preceding chapter, after skipping some vertical whitespace. Also cause TEX to typeset for single-sided printing. (You can override the headers format with the @headings double command; see Section 3.4.6 [The @headings Command], page 35.)

@setchapternewpage odd

Specify the double-sided heading format, with chapters on new pages.

Texinfo lacks an @setchapternewpage even command.

E.3 How to Make Your Own Headings

You can use the standard headings provided with Texinfo or specify your own. By default, Texinfo has no footers, so if you specify them, the available page size for the main text will be slightly reduced.

Texinfo provides six commands for specifying headings and footings:

• @everyheading @everyfooting generate page headers and footers that are the same for both even- and odd-numbered pages.

• @evenheading and @evenfooting command generate headers and footers for evennumbered (left-hand) pages.

• @oddheading and @oddfooting generate headers and footers for odd-numbered (righthand) pages.

Write custom heading specifications in the Texinfo file immediately after the @end titlepage command. You must cancel the predefined heading commands with the @headings off command before defining your own specifications.

Here is how to tell TEX to place the chapter name at the left, the page number in the center, and the date at the right of every header for both even- and odd-numbered pages:

     @headings off
     @everyheading @thischapter @| @thispage @| @today{}

You need to divide the left part from the central part and the central part from the right part by inserting ‘@|’ between parts. Otherwise, the specification command will not be able to tell where the text for one part ends and the next part begins.

Each part can contain text or @-commands. The text is printed as if the part were within an ordinary paragraph in the body of the page. The @-commands replace themselves with the page number, date, chapter name, or whatever.

Here are the six heading and footing commands:

@everyheading left @| center @| right
@everyfooting left @| center @| right

The ‘every’ commands specify the format for both even- and odd-numbered pages. These commands are for documents that are printed on one side of each sheet of paper, or for documents in which you want symmetrical headers or footers.

@evenheading left @| center @| right
@oddheading left @| center @| right
@evenfooting left @| center @| right
@oddfooting left @| center @| right

The ‘even’ and ‘odd’ commands specify the format for even-numbered pages and odd-numbered pages. These commands are for books and manuals that are printed on both sides of each sheet of paper.

Use the ‘@this...’ series of @-commands to provide the names of chapters and sections and the page number. You can use the ‘@this...’ commands in the left, center, or right portions of headers and footers, or anywhere else in a Texinfo file so long as they are between

@iftex and @end iftex commands.

Here are the ‘@this...’ commands:

@thispage

Expands to the current page number.

@thischaptername

Expands to the name of the current chapter.

@thischapter

Expands to the number and name of the current chapter, in the format ‘Chapter 1: Title’.

@thistitle

Expands to the name of the document, as specified by the @settitle command.

@thisfile

For @include files only: expands to the name of the current @include file. If the current Texinfo source file is not an @include file, this command has no effect. This command does not provide the name of the current Texinfo source file unless it is an @include file. (See Appendix D [Include Files], page 196, for more information about @include files.)

You can also use the @today{} command, which expands to the current date, in ‘1 Jan 1900’ format.

Other @-commands and text are printed in a header or footer just as if they were in the body of a page. It is useful to incorporate text, particularly when you are writing drafts:

     @headings off
     @everyheading @emph{Draft!} @| @thispage @| @thischapter
     @everyfooting @| @| Version: 0.27: @today{}

Beware of overlong titles: they may overlap another part of the header or footer and blot it out.

Appendix F Formatting Mistakes

Besides mistakes in the content of your documentation, there are two kinds of mistake you can make with Texinfo: you can make mistakes with @-commands, and you can make mistakes with the structure of the nodes and chapters.

Emacs has two tools for catching the @-command mistakes and two for catching structuring mistakes.

For finding problems with @-commands, you can run TEX or a region formatting command on the region that has a problem; indeed, you can run these commands on each region as you write it.

For finding problems with the structure of nodes and chapters, you can use C-c Cs (texinfo-show-structure) and the related occur command and you can use the M-x

Info-validate command.

The makeinfo program does an excellent job of catching errors and reporting them—far

better than texinfo-format-region or texinfo-format-buffer. In addition, the various

functions for automatically creating and updating node pointers and menus remove many opportunities for human error.

If you can, use the updating commands to create and insert pointers and menus. These prevent many errors. Then use makeinfo (or its Texinfo mode manifestations, makeinforegion and makeinfo-buffer) to format your file and check for other errors. This is the best way to work with Texinfo. But if you cannot use makeinfo, or your problem is very puzzling, then you may want to use the tools described in this appendix.

F.1 Catching Errors with Info Formatting

After you have written part of a Texinfo file, you can use the texinfo-format-region or the makeinfo-region command to see whether the region formats properly.

Most likely, however, you are reading this section because for some reason you cannot use the makeinfo-region command; therefore, the rest of this section presumes that you

are using texinfo-format-region.

If you have made a mistake with an @-command, texinfo-format-region will stop processing at or after the error and display an error message. To see where in the buffer the error occurred, switch to the ‘*Info Region*’ buffer; the cursor will be in a position that is after the location of the error. Also, the text will not be formatted after the place where the error occurred (or more precisely, where it was detected).

For example, if you accidentally end a menu with the command @end menus with an ‘s’ on the end, instead of with @end menu, you will see an error message that says:

     @end menus is not handled by texinfo

The cursor will stop at the point in the buffer where the error occurs, or not long after it. The buffer will look like this:

     ---------- Buffer: *Info Region* ----------
     * Menu:
     * Using texinfo-show-structure::  How to use
                                       `texinfo-show-structure'
                                       to catch mistakes.
     * Running Info-Validate::         How to check for
                                       unreferenced nodes.
     @end menus

⋆

     ---------- Buffer: *Info Region* ----------

The texinfo-format-region command sometimes provides slightly odd error messages. For example, the following cross reference fails to format:

     (@xref{Catching Mistakes, for more info.)

In this case, texinfo-format-region detects the missing closing brace but displays a

message that says ‘Unbalanced parentheses’ rather than ‘Unbalanced braces’. This is

because the formatting command looks for mismatches between braces as if they were parentheses.

Sometimes texinfo-format-region fails to detect mistakes. For example, in the following, the closing brace is swapped with the closing parenthesis:

     (@xref{Catching Mistakes), for more info.}

Formatting produces:

     (*Note for more info.: Catching Mistakes)

The only way for you to detect this error is to realize that the reference should have looked like this:

     (*Note Catching Mistakes::, for more info.)

Incidentally, if you are reading this node in Info and type f ^{⟨}\underline{\overline{RET}}^{⟩} (Info-follow-reference), you will generate an error message that says:

     No such node: "Catching Mistakes) The only way ...

This is because Info perceives the example of the error as the first cross reference in this node and if you type a^{⟨}\underline{\overline{RET}}^{⟩} immediately after typing the Info f command, Info will attempt to go to the referenced node. If you type f catch ^{⟨}\underline{\overline{TAB}}^{⟩} ^{⟨}\underline{\overline{RET}}^{⟩}, Info will complete the node name of the correctly written example and take you to the ‘Catching Mistakes’ node. (If you try this, you can return from the ‘Catching Mistakes’ node by typing l (Info-last).)

F.2 Catching Errors with TEX Formatting

You can also catch mistakes when you format a file with TEX.

Usually, you will want to do this after you have run texinfo-format-buffer (or, better, makeinfo-buffer) on the same file, because texinfo-format-buffer sometimes displays error messages that make more sense than TEX. (See Section F.1 [Debugging with Info], page 204, for more information.)

For example, TEX was run on a Texinfo file, part of which is shown here:

---------- Buffer: texinfo.texi ----------
name of the Texinfo file as an extension.  The
@samp{??} are `wildcards' that cause the shell to
substitute all the raw index files.  (@xref{sorting
indices, for more information about sorting
indices.)@refill
---------- Buffer: texinfo.texi ----------

(The cross reference lacks a closing brace.) TEX produced the following output, after which it stopped:

---------- Buffer: *tex-shell* ----------
Runaway argument?
{sorting indices, for more information about sorting
indices.) @refill @ETC.
! Paragraph ended before @xref was complete.
<to be read again>
                   @par
l.27
?
---------- Buffer: *tex-shell* ----------

In this case, TEX produced an accurate and understandable error message:

Paragraph ended before @xref was complete.

‘@par’ is an internal TEX command of no relevance to Texinfo. ‘l.27’ means that TEX detected the problem on line 27 of the Texinfo file. The ‘?’ is the prompt TEX uses in this circumstance.

Unfortunately, TEX is not always so helpful, and sometimes you must truly be a Sherlock Holmes to discover what went wrong.

In any case, if you run into a problem like this, you can do one of three things.

1. You can tell TEX to continue running and ignore just this error by typing ^{⟨}\underline{\overline{RET}}^{⟩} at the ‘?’ prompt.

2. You can tell TEX to continue running and to ignore all errors as best it can by typing r ^{⟨}\underline{\overline{RET}}^{⟩} at the ‘?’ prompt. This is often the best thing to do. However, beware: the one error may produce a cascade of additional error messages as its consequences are felt through the rest of the file. To stop TEX when it is producing such an avalanche of error messages, type C-c (or C-c C-c, if you are running a shell inside Emacs).

3. You can tell TEX to stop this run by typing x ^{⟨}\underline{\overline{RET}}^{⟩} at the ‘?’ prompt.

If you are running TEX inside Emacs, you need to switch to the shell buffer and line at which TEX offers the ‘?’ prompt.

Sometimes TEX will format a file without producing error messages even though there is a problem. This usually occurs if a command is not ended but TEX is able to continue processing anyhow. For example, if you fail to end an itemized list with the @end itemize command, TEX will write a DVI file that you can print out. The only error message that TEX will give you is the somewhat mysterious comment that

     (@end occurred inside a group at level 1)

However, if you print the DVI file, you will find that the text of the file that follows the itemized list is entirely indented as if it were part of the last item in the itemized list. The error message is the way TEX says that it expected to find an @end command somewhere in the file; but that it could not determine where it was needed.

Another source of notoriously hard-to-find errors is a missing @end group command. If you ever are stumped by incomprehensible errors, look for a missing @end group command first.

If the Texinfo file lacks header lines, TEX may stop in the beginning of its run and display output that looks like the following. The ‘*’ indicates that TEX is waiting for input.

     This is TeX, Version 3.14159 (Web2c 7.0)
     (test.texinfo [1])
     *

In this case, simply type \end ^{⟨}\underline{\overline{RET}}^{⟩} after the asterisk. Then write the header lines in the Texinfo file and run the TEX command again. (Note the use of the backslash, ‘\’. TEX uses ‘\’ instead of ‘@’; and in this circumstance, you are working directly with TEX, not with Texinfo.)

F.3 Using texinfo-show-structure

It is not always easy to keep track of the nodes, chapters, sections, and subsections of a Texinfo file. This is especially true if you are revising or adding to a Texinfo file that someone else has written.

In GNU Emacs, in Texinfo mode, the texinfo-show-structure command lists all the lines that begin with the @-commands that specify the structure: @chapter, @section, @appendix, and so on. With an argument (C-u as prefix argument, if interactive), the command also shows the @node lines. The texinfo-show-structure command is bound to C-c C-s in Texinfo mode, by default.

The lines are displayed in a buffer called the ‘*Occur*’ buffer, indented by hierarchical level. For example, here is a part of what was produced by running texinfo-show-structure on this manual:

      Lines matching "^@\\(chapter \\|sect\\|subs\\|subh\\|
      unnum\\|major\\|chapheading \\|heading \\|appendix\\)"
      in buffer texinfo.texi.
      ...
      4177:@chapter Nodes
      4198:    @heading Two Paths
      4231:    @section Node and Menu Illustration
      4337:    @section The @code{@@node} Command
      4393:        @subheading Choosing Node and Pointer Names
      4417:        @subsection How to Write an @code{@@node} Line
      4469:        @subsection @code{@@node} Line Tips
      ...

This says that lines 4337, 4393, and 4417 of ‘texinfo.texi’ begin with the @section, @subheading, and @subsection commands respectively. If you move your cursor into the ‘*Occur*’ window, you can position the cursor over one of the lines and use the C-c Cc command (occur-mode-goto-occurrence), to jump to the corresponding spot in the Texinfo file. See section “Using Occur” in The GNU Emacs Manual, for more information

about occur-mode-goto-occurrence.

The first line in the ‘*Occur*’ window describes the regular expression specified by texinfo-heading-pattern. This regular expression is the pattern that texinfo-show-structure looks for. See section “Using Regular Expressions” in The GNU Emacs Manual, for more information.

When you invoke the texinfo-show-structure command, Emacs will display the structure of the whole buffer. If you want to see the structure of just a part of the buffer, of one chapter, for example, use the C-x n n (narrow-to-region) command to mark the region. (See section “Narrowing” in The GNU Emacs Manual.) This is how the example used above was generated. (To see the whole buffer again, use C-x n w (widen).)

If you call texinfo-show-structure with a prefix argument by typing C-u C-c C-s, it will list lines beginning with @node as well as the lines beginning with the @-sign commands

for @chapter, @section, and the like.

You can remind yourself of the structure of a Texinfo file by looking at the list in the ‘*Occur*’ window; and if you have mis-named a node or left out a section, you can correct the mistake.

F.4 Using occur

Sometimes the texinfo-show-structure command produces too much information. Perhaps you want to remind yourself of the overall structure of a Texinfo file, and are overwhelmed by the detailed list produced by texinfo-show-structure. In this case, you can use the occur command directly. To do this, type

     M-x occur

and then, when prompted, type a regexp, a regular expression for the pattern you want to match. (See section “Regular Expressions” in The GNU Emacs Manual.) The occur command works from the current location of the cursor in the buffer to the end of the buffer. If you want to run occur on the whole buffer, place the cursor at the beginning of the buffer.

For example, to see all the lines that contain the word ‘@chapter’ in them, just type ‘@chapter’. This will produce a list of the chapters. It will also list all the sentences with ‘@chapter’ in the middle of the line.

If you want to see only those lines that start with the word ‘@chapter’, type ‘^@chapter’ when prompted by occur. If you want to see all the lines that end with a word or phrase, end the last word with a ‘$’; for example, ‘catching mistakes$’. This can be helpful when you want to see all the nodes that are part of the same chapter or section and therefore have the same ‘Up’ pointer.

See section “Using Occur” in The GNU Emacs Manual, for more information.

F.5 Finding Badly Referenced Nodes

You can use the Info-validate command to check whether any of the ‘Next’, ‘Previous’, ‘Up’ or other node pointers fail to point to a node. This command checks that every node pointer points to an existing node. The Info-validate command works only on Info files, not on Texinfo files.

The makeinfo program validates pointers automatically, so you do not need to use the Info-validate command if you are using makeinfo. You only may need to use Infovalidate if you are unable to run makeinfo and instead must create an Info file using

texinfo-format-region or texinfo-format-buffer, or if you write an Info file from

scratch.

F.5.1 Running Info-validate

To use Info-validate, visit the Info file you wish to check and type:

     M-x Info-validate

Note that the Info-validate command requires an upper case ‘I’. You may also need to create a tag table before running Info-validate. See Section F.5.3 [Tagifying], page 210.

If your file is valid, you will receive a message that says “File appears valid”. However, if you have a pointer that does not point to a node, error messages will be displayed in a

buffer called ‘*problems in info file*’.

For example, Info-validate was run on a test file that contained only the first node of this manual. One of the messages said:

     In node "Overview", invalid Next: Texinfo Mode

This meant that the node called ‘Overview’ had a ‘Next’ pointer that did not point to anything (which was true in this case, since the test file had only one node in it).

Now suppose we add a node named ‘Texinfo Mode’ to our test case but we do not specify a ‘Previous’ for this node. Then we will get the following error message:

     In node "Texinfo Mode", should have Previous: Overview

This is because every ‘Next’ pointer should be matched by a ‘Previous’ (in the node where the ‘Next’ points) which points back.

Info-validate also checks that all menu entries and cross references point to actual nodes.

Info-validate requires a tag table and does not work with files that have been split. (The texinfo-format-buffer command automatically splits large files.) In order to use Info-validate on a large file, you must run texinfo-format-buffer with an argument so that it does not split the Info file; and you must create a tag table for the unsplit file.

F.5.2 Creating an Unsplit File

You can run Info-validate only on a single Info file that has a tag table. The command will not work on the indirect subfiles that are generated when a master file is split. If you have a large file (longer than 70,000 bytes or so), you need to run the texinfoformat-buffer or makeinfo-buffer command in such a way that it does not create indirect subfiles. You will also need to create a tag table for the Info file. After you have done this, you can run Info-validate and look for badly referenced nodes.

The first step is to create an unsplit Info file. To prevent texinfo-format-buffer from splitting a Texinfo file into smaller Info files, give a prefix to the M-x texinfo-format-buffer command:

     C-u M-x texinfo-format-buffer

or else

     C-u C-c C-e C-b

When you do this, Texinfo will not split the file and will not create a tag table for it.

F.5.3 Tagifying a File

After creating an unsplit Info file, you must create a tag table for it. Visit the Info file you wish to tagify and type:

     M-x Info-tagify

(Note the upper case ‘I’ in Info-tagify.) This creates an Info file with a tag table that you can validate.

The third step is to validate the Info file:

     M-x Info-validate

(Note the upper case ‘I’ in Info-validate.) In brief, the steps are:

     C-u M-x texinfo-format-buffer
     M-x Info-tagify
     M-x Info-validate

After you have validated the node structure, you can rerun texinfo-format-buffer in the normal way so it will construct a tag table and split the file automatically, or you can make the tag table and split the file manually.

F.5.4 Splitting a File Manually

You should split a large file or else let the texinfo-format-buffer or makeinfobuffer command do it for you automatically. (Generally you will let one of the formatting commands do this job for you. See Section 20.1 [Creating an Info File], page 153.)

The split-off files are called the indirect subfiles.

Info files are split to save memory. With smaller files, Emacs does not have make such a large buffer to hold the information.

If an Info file has more than 30 nodes, you should also make a tag table for it. See Section F.5.1 [Using Info-validate], page 209, for information about creating a tag table. (Again, tag tables are usually created automatically by the formatting command; you only need to create a tag table yourself if you are doing the job manually. Most likely, you will do this for a large, unsplit file on which you have run Info-validate.)

Visit the Info file you wish to tagify and split and type the two commands:

     M-x Info-tagify
     M-x Info-split

(Note that the ‘I’ in ‘Info’ is upper case.)

When you use the Info-split command, the buffer is modified into a (small) Info file which lists the indirect subfiles. This file should be saved in place of the original visited file. The indirect subfiles are written in the same directory the original file is in, with names generated by appending ‘-’ and a number to the original file name.

The primary file still functions as an Info file, but it contains just the tag table and a directory of subfiles.

Appendix G Refilling Paragraphs

The @refill command refills and, optionally, indents the first line of a paragraph.^{1} The @refill command is no longer important, but we describe it here because you once needed it. You will see it in many old Texinfo files.

Without refilling, paragraphs containing long @-constructs may look bad after formatting because the formatter removes @-commands and shortens some lines more than others. In the past, neither the texinfo-format-region command nor the texinfo-format-buffer command refilled paragraphs automatically. The @refill command had to be written at the end of every paragraph to cause these formatters to fill them. (Both TEX and makeinfo have always refilled paragraphs automatically.) Now, all the Info formatters automatically fill and indent those paragraphs that need to be filled and indented.

    The @refill command causes texinfo-format-region and texinfo-format-buffer

to refill a paragraph in the Info file after all the other processing has been done. For this reason, you can not use @refill with a paragraph containing either @* or @w{ ... } since the refilling action will override those two commands.

    The texinfo-format-region and texinfo-format-buffer commands now automat-

ically append @refill to the end of each paragraph that should be filled. They do not append @refill to the ends of paragraphs that contain @* or @w{ ...} and therefore do not refill or indent them.

1 Perhaps the command should have been called the @refillandindent command, but @refill is shorter

and the name was chosen before indenting was possible.

Appendix H @-Command Syntax

The character ‘@’ is used to start special Texinfo commands. (It has the same meaning that ‘\’ has in plain TEX.) Texinfo has four types of @-command:

1. Non-alphabetic commands.

These commands consist of an @ followed by a punctuation mark or other character that is not part of the alphabet. Non-alphabetic commands are almost always part of the text within a paragraph, and never take any argument. The two characters (@ and the other one) are complete in themselves; none is followed by braces. The non-alphabetic commands are: @., @:, @*, @SPACE, @TAB,

@NL, @@, @{, and @}.

2. Alphabetic commands that do not require arguments.

These commands start with @ followed by a word followed by left- and righthand braces. These commands insert special symbols in the document; they do not require arguments. For example, @dots{} ⇒ ‘...’, @equiv{} ⇒ ‘ ≡ ’,

@TeX{} ⇒ ‘TEX’, and @bullet{} ⇒ ‘•’.

3. Alphabetic commands that require arguments within braces.

These commands start with @ followed by a letter or a word, followed by an argument within braces. For example, the command @dfn indicates the introductory or defining use of a term; it is used as follows: ‘In Texinfo, @@-commands are

@dfn{mark-up} commands.’

4. Alphabetic commands that occupy an entire line.

These commands occupy an entire line. The line starts with @, followed by the name of the command (a word); for example, @center or @cindex. If no argument is needed, the word is followed by the end of the line. If there is an argument, it is separated from the command name by a space. Braces are not used.

Thus, the alphabetic commands fall into classes that have different argument syntaxes. You cannot tell to which class a command belongs by the appearance of its name, but you can tell by the command’s meaning: if the command stands for a glyph, it is in class 2 and does not require an argument; if it makes sense to use the command together with other text as part of a paragraph, the command is in class 3 and must be followed by an argument in braces; otherwise, it is in class 4 and uses the rest of the line as its argument.

The purpose of having a different syntax for commands of classes 3 and 4 is to make Texinfo files easier to read, and also to help the GNU Emacs paragraph and filling commands work properly. There is only one exception to this rule: the command @refill, which is always used at the end of a paragraph immediately following the final period or other punctuation character. @refill takes no argument and does not require braces. @refill never confuses the Emacs paragraph commands because it cannot appear at the beginning of a line.

Appendix I How to Obtain TEX

TEX is freely redistributable. You can obtain TEX for Unix systems via anonymous ftp or on physical media. The core material consists of the Web2c TEX distribution

(http://tug.org/web2c).

Instructions for retrieval by anonymous ftp and information on other available distributions:

ftp://tug.org/tex/unixtex.ftp
http://tug.org/unixtex.ftp

The Free Software Foundation provides a core distribution on its Source Code CD-ROM suitable for printing Texinfo manuals. To order it, contact:

Free Software Foundation, Inc. 59 Temple Place Suite 330 Boston, MA 02111-1307 USA Telephone: +1-617-542-5942 Fax: (including Japan) +1-617-542-2652 Free Dial Fax (in Japan):

0031-13-2473 (KDD) 0066-3382-0158 (IDC)

Electronic mail: gnu@gnu.org Many other TEX distributions are available; see http://tug.org/.

Appendix J Copying This Manual

J.1 GNU Free Documentation License

Version 1.1, March 2000

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000 Free Software Foundation, Inc. 59 Temple Place, Suite 330, Boston, MA 02111-1307, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other written document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (For example, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, whose contents can be viewed and edited directly and straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup has been designed to thwart or discourage subsequent modification by readers is not Transparent. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, sgml or xml using a publicly available dtd, and standard-conforming simple html designed for human modification. Opaque formats include PostScript, pdf, proprietary formats that can be read and edited only by proprietary word processors, sgml or xml for which the dtd and/or processing tools are not generally available, and the machine-generated html produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies of the Document numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects.

If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a publicly-accessible computer-network location containing a complete Transparent copy of the Document, free of added material, which the general network-using public has access to download anonymously at no charge using public-standard network protocols. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has less than five).

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section entitled “History”, and its title, and add to it an item stating

at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. In any section entitled “Acknowledgments” or “Dedications”, preserve the section’s title, and preserve in the section all the substance and tone of each of the contributor acknowledgments and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section as “Endorsements” or to conflict in title with

any Invariant Section.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice.

The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections entitled “History” in the various original documents, forming one section entitled “History”; likewise combine any sections entitled “Acknowledgments”, and any sections entitled “Dedications”. You must delete all sections entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, does not as a whole count as a Modified Version of the Document, provided no compilation copyright is claimed for the compilation. Such a compilation is called an “aggregate”, and this License does not apply to the other self-contained works thus compiled with the Document, on account of their being thus compiled, if they are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one quarter of the entire aggregate, the Document’s Cover Texts may be placed on covers that surround only the Document within the aggregate. Otherwise they must appear on covers around the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License provided that you also include the original English version of this License. In case of a disagreement between the translation and the original English version of this License, the original English version will prevail.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

J.1.1 ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

Copyright (C) year your name.
Permission is granted to copy, distribute and/or modify this document
under the terms of the GNU Free Documentation License, Version 1.1
or any later version published by the Free Software Foundation;
with the Invariant Sections being list their titles, with the
Front-Cover Texts being list, and with the Back-Cover Texts being list.
A copy of the license is included in the section entitled ``GNU
Free Documentation License''.

If you have no Invariant Sections, write “with no Invariant Sections” instead of saying which ones are invariant. If you have no Front-Cover Texts, write “no Front-Cover Texts” instead of “Front-Cover Texts being list”; likewise for Back-Cover Texts.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Command and Variable Index 222

Command and Variable Index

This is an alphabetical list of all the @-commands, assorted Emacs Lisp functions, and several variables. To make the list easier to use, the commands are listed without their preceding ‘@’.

!

! (end of sentence). . . . . . . . . . . . . . . . . . . . . . . . . . 102

"

" (umlaut accent). . . . . . . . . . . . . . . . . . . . . . . . . . . 103

’

' (umlaut accent). . . . . . . . . . . . . . . . . . . . . . . . . . . 103

(

(newline) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 (space). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 (tab) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

*

* (force line break) . . . . . . . . . . . . . . . . . . . . . . . . . 113

,

, (cedilla accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

-

- (discretionary hyphen) . . . . . . . . . . . . . . . . . . . . 114

.

. (end of sentence). . . . . . . . . . . . . . . . . . . . . . . . . . 102

<

<colon> (suppress widening) . . . . . . . . . . . . . . . . 102

=

= (macron accent) . . . . . . . . . . . . . . . . . . . . . . . . . . 103

?

? (end of sentence). . . . . . . . . . . . . . . . . . . . . . . . . . 102

@

@ (literal ‘@’) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

^

^(circumflex accent) . . . . . . . . . . . . . . . . . . . . . . . . 103

‘

` (grave accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

{

{ (literal ‘{’) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

}

} (literal ‘}’) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

\

\ (literal \ in @math) . . . . . . . . . . . . . . . . . . . . . . . . 106 \emergencystretch . . . . . . . . . . . . . . . . . . . . . . . . . 149 \input (raw TEX startup) . . . . . . . . . . . . . . . . . . . . . 9 \mag (raw TEX magnification) . . . . . . . . . . . . . . . 151

~

~ (tilde accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

A

aa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 AA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 acronym. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 ae . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 AE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 afourlatex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 afourpaper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 afourwide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 anchor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 appendix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 appendixsec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 appendixsection . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 appendixsubsec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 appendixsubsubsec . . . . . . . . . . . . . . . . . . . . . . . . . . 48 apply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 asis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

B

b (bold font) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 buffer-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 bullet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 bye. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 43

C

c (comment) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 cartouche . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 center . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 centerchap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 chapheading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 chapter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 cindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 cite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 columnfractions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 comment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 contents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 copying. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 105 cropmarks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151

D

defcodeindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 defcv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 deffn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 deffnx. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 defindex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 definfoenclose . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 defivar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 defmac. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 defmethod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125, 126 defop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 defopt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 defspec. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 deftp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 deftypefn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 deftypefun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 deftypeivar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 deftypeop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 deftypevar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 deftypevr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 defun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 defvar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 defvr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 dfn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 dircategory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 direntry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 display. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 dmn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 documentdescription. . . . . . . . . . . . . . . . . . . . . . . . 38 documentencoding . . . . . . . . . . . . . . . . . . . . . . . . . . 136 documentlanguage . . . . . . . . . . . . . . . . . . . . . . . . . . 135 dotaccent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 dotless. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 dots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

E

email . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 emph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 end. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81, 89 end titlepage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 enddots. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 enumerate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 env . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 equiv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 evenfooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 evenheading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 everyfooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 everyheading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 exampleindent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 exclamdown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 exdent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

F

file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 filll TEX dimension . . . . . . . . . . . . . . . . . . . . . . . . 34 finalout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 findex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 flushleft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 flushright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 foobar . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118, 121, 122 footnote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 footnotestyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 forward-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 ftable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

G

group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

H

H (Hungarian umlaut accent) . . . . . . . . . . . . . . . . 103 hbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 heading. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 headings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 headword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 html . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 hyphenation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

I

i (italic font) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 ifclear. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 ifhtml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129, 131 ifinfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 ifnothtml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 ifnotinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 ifnotplaintext . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 ifnottex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 ifplaintext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 ifset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 iftex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 ignore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 image . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 include. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 Info-validate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 inforef. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 insertcopying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 isearch-backward . . . . . . . . . . . . . . . . . . . . . . . . . . 118 isearch-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90, 92, 94 itemize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 itemx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

K

kbd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 kbdinputstyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 kindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

L

l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 lowersections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

M

macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 majorheading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 makeinfo-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 makeinfo-kill-job . . . . . . . . . . . . . . . . . . . . . . . . . 158

makeinfo-recenter-output-buffer. . . . . . . . . . 158

makeinfo-region . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 math . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 minus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 multitable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

N

need . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 next-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 noindent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 novalidate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143

O

o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 occur . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208

                                                occur-mode-goto-occurrence . . . . . . . . . . . . . . . . 17

oddfooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 oddheading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 oe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 OE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

P

page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 page, within @titlepage . . . . . . . . . . . . . . . . . . . . . 32 pagesizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 paragraphindent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 phoo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 pindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 pounds. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 printindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 pxref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

Q

questiondown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 quotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

R

r (Roman font) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 raisesections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 refill. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 result. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 ringaccent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 rmacro. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

S

samp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 sc (small caps font) . . . . . . . . . . . . . . . . . . . . . . . . . . 79 section. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 setchapternewpage . . . . . . . . . . . . . . . . . . . . . . . . . . 38

                                                setcontentsaftertitlepage . . . . . . . . . . . . . . . . . 43

setfilename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

                                                setshortcontentsaftertitlepage . . . . . . . . . . . . 43

settitle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 shortcontents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 shorttitlepage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 smallbook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 smalldisplay . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84, 85 smallexample . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 smallformat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84, 85 smalllisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 sp (line spacing) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 sp (titlepage line spacing) . . . . . . . . . . . . . . . . . . . . 32 ss . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 strong . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 subheading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 subsection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 subsubheading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 subsubsection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 subtitle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 summarycontents . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 syncodeindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 synindex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

T

t (typewriter font) . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 tab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 tex (command) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

texinfo-all-menus-update . . . . . . . . . . . . . . . . . . 19
texinfo-every-node-update . . . . . . . . . . . . . . . . . 19

texinfo-format-buffer . . . . . . . . . . . . . . . . . 22, 158 texinfo-format-region . . . . . . . . . . . . . . . . . 22, 158

texinfo-indent-menu-description . . . . . . . . . . . 21

texinfo-insert-@code. . . . . . . . . . . . . . . . . . . . . . . 15 texinfo-insert-@dfn. . . . . . . . . . . . . . . . . . . . . . . . 15 texinfo-insert-@end. . . . . . . . . . . . . . . . . . . . . . . . 15

texinfo-insert-@example . . . . . . . . . . . . . . . . . . . 16

texinfo-insert-@item. . . . . . . . . . . . . . . . . . . . . . . 15 texinfo-insert-@kbd. . . . . . . . . . . . . . . . . . . . . . . . 15 texinfo-insert-@node. . . . . . . . . . . . . . . . . . . . . . . 15

texinfo-insert-@noindent . . . . . . . . . . . . . . . . . . 15

texinfo-insert-@samp. . . . . . . . . . . . . . . . . . . . . . . 15 texinfo-insert-@table . . . . . . . . . . . . . . . . . . . . . 15 texinfo-insert-@var. . . . . . . . . . . . . . . . . . . . . . . . 15 texinfo-insert-braces . . . . . . . . . . . . . . . . . . . . . 16

texinfo-insert-node-lines . . . . . . . . . . . . . . . . . 21

texinfo-make-menu . . . . . . . . . . . . . . . . . . . . . . . . . . 19 texinfo-master-menu. . . . . . . . . . . . . . . . . . . . . . . . 18

texinfo-multiple-files-update . . . . . . . . . . . . 196
texinfo-multiple-files-update (in brief) . . . . 21
texinfo-sequential-node-update . . . . . . . . . . . . 21
texinfo-show-structure . . . . . . . . . . . . . . . . 17, 207
texinfo-start-menu-description . . . . . . . . . . . . 16

texinfo-tex-buffer . . . . . . . . . . . . . . . . . . . . . . . . . 23 texinfo-tex-print . . . . . . . . . . . . . . . . . . . . . . . . . . 23 texinfo-tex-region . . . . . . . . . . . . . . . . . . . . . . . . . 23 texinfo-update-node. . . . . . . . . . . . . . . . . . . . . . . . 19 thischapter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 thischaptername . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 thisfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 thispage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 thistitle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 tieaccent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 tindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 title . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 titlefont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 titlepage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 today . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 top . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 top (@-command) . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

U

u (breve accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 ubaraccent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 udotaccent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 unmacro. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 unnumbered . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 unnumberedsec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 unnumberedsubsec . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 unnumberedsubsubsec. . . . . . . . . . . . . . . . . . . . . . . . 48 up-list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 uref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 url . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

V

v (check accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 var . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 verb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 verbatim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 verbatiminclude . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 vindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 vskip TEX vertical skip . . . . . . . . . . . . . . . . . . . . . . 34 vtable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

W

w (prevent line break) . . . . . . . . . . . . . . . . . . . . . . . 114

X

xref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

Concept Index 226

Concept Index

!

¡. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

$

$Id: texinfo.txi,v 1.219 2002/03/28 16:36:00 karl Exp $ comment . . . . . . . . . . . . . . . . . . . . . . . . 192

(

(dir) as Up node of Top node . . . . . . . . . . . . . . . . . 54

-

--commands-in-node-names . . . . . . . . . . . . . . . . . 154

--delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 --dir-file=name . . . . . . . . . . . . . . . . . . . . . . . . . . 165 --docbook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 --entry=text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 --error-limit=limit . . . . . . . . . . . . . . . . . . . . . . . . 154 --fill-column=width . . . . . . . . . . . . . . . . . . . . . . . 154 --footnote-style=style . . . . . . . . . . . . . . . . . . . . 154 --force. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 --help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154, 165 --html. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 --info-dir=dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 --info-file=file . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 --item=text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 --no-headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 --no-number-footnotes . . . . . . . . . . . . . . . . . . . . 155 --no-pointer-validate . . . . . . . . . . . . . . . . . . . . 155 --no-split . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 --no-validate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 --no-warn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 --number-sections . . . . . . . . . . . . . . . . . . . . . . . . . 155 --output=file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 --paragraph-indent=indent . . . . . . . . . . . . . . . . . 156 --quiet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 --reference-limit=limit . . . . . . . . . . . . . . . . . . . 156 --remove . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 --section=sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 --verbose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 --version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156, 166 --xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 -D dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 -d name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 -D var . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 -e limit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 -e text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 -F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 -f width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 -h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154, 165 -I dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 -i file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 -o file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 -P dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 -p indent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 -r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 -r limit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 -s sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 -s style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 -V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156, 166

.

.cshrc initialization file . . . . . . . . . . . . . . . . . . . . . . 148 .profile initialization file . . . . . . . . . . . . . . . . . . . . . 148

                              <

<meta> HTML tag, and document description. . 38 <title> HTML tag. . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 <URL convention, not used . . . . . . . . . . . . . . . . . . . 70

?

¿ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

@

@-command list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 @-command syntax . . . . . . . . . . . . . . . . . . . . . . . . . 213 @-commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 @-commands in @node, limited support . . . . . . 157 @-commands in nodename. . . . . . . . . . . . . . . . . . . . 54 @include file sample . . . . . . . . . . . . . . . . . . . . . . . . 197 @menu parts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 @node line writing . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 @value in @node lines. . . . . . . . . . . . . . . . . . . . . . . 157 @w, for blank items . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

                              \

‘\input’ source line ignored. . . . . . . . . . . . . . . . . . . 28

A

A4 paper, printing on . . . . . . . . . . . . . . . . . . . . . . . 150 A5 paper, printing on . . . . . . . . . . . . . . . . . . . . . . . 150 å . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Å. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Abbreviations for keys . . . . . . . . . . . . . . . . . . . . . . . 74 Abbreviations, tagging . . . . . . . . . . . . . . . . . . . . . . . 78 Abstract of document . . . . . . . . . . . . . . . . . . . . . . . . 38 Accents, inserting. . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Acronyms, tagging . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Acute accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Adding a new Info file. . . . . . . . . . . . . . . . . . . . . . . 162 æ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Æ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Aliases, command . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Alphabetical @-command list . . . . . . . . . . . . . . . . 167 alt attribute for images. . . . . . . . . . . . . . . . . . . . . . 112 alternate text for images . . . . . . . . . . . . . . . . . . . . 112 Anchors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Another Info directory . . . . . . . . . . . . . . . . . . . . . . 162 Apostrophe in nodename . . . . . . . . . . . . . . . . . . . . . 54 Arguments, repeated and optional . . . . . . . . . . . 117 ASCII text output . . . . . . . . . . . . . . . . . . . . . . . . . . 155 Aspect ratio of images . . . . . . . . . . . . . . . . . . . . . . 111 autoexec.bat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 automake, and version info . . . . . . . . . . . . . . . . . . 192 Automatic pointer creation with makeinfo . . . . . 55 Automatically insert nodes, menus . . . . . . . . . . . . 17 Auxiliary files, avoiding . . . . . . . . . . . . . . . . . . . . . 143

B

B5 paper, printing on . . . . . . . . . . . . . . . . . . . . . . . 151 Backslash in macros . . . . . . . . . . . . . . . . . . . . . . . . 137 Backslash, and macros . . . . . . . . . . . . . . . . . . . . . . 138 Badly referenced nodes . . . . . . . . . . . . . . . . . . . . . . 209 Bastard title page. . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Batch formatting for Info. . . . . . . . . . . . . . . . . . . . 159 Beginning a Texinfo file . . . . . . . . . . . . . . . . . . . . . . 26 Beginning line of a Texinfo file . . . . . . . . . . . . . . . . 27 Berry, Karl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Big points. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Black rectangle in hardcopy . . . . . . . . . . . . . . . . . 149 Blank lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 Body of a macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 Bolio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Book characteristics, printed. . . . . . . . . . . . . . . . . . . 6 Book, printing small . . . . . . . . . . . . . . . . . . . . . . . . 150 BoTEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Box with rounded corners . . . . . . . . . . . . . . . . . . . . 87 Box, ugly black in hardcopy . . . . . . . . . . . . . . . . . 149 Braces and argument syntax . . . . . . . . . . . . . . . . . 213 Braces, in macro arguments . . . . . . . . . . . . . . . . . 138 Braces, inserting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 Braces, when to use . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Breaks in a line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 Breve accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Buffer formatting and printing . . . . . . . . . . . . . . . . 22 Bugs, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Bullets, inserting . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

C

Case in node name. . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Case, not altering in @code . . . . . . . . . . . . . . . . . . . 72 Catching errors with Info formatting . . . . . . . . . 204 Catching errors with TEX formatting . . . . . . . . . 205 Catching mistakes . . . . . . . . . . . . . . . . . . . . . . . . . . 204 Cedilla accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Centimeters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Chapter structuring . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Chapters, formatting one at a time. . . . . . . . . . . 143 Characteristics, printed books or manuals. . . . . . . 6 Characters, invalid in node name. . . . . . . . . . . . . . 54 Chassell, Robert J.. . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Check accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Checking for badly referenced nodes. . . . . . . . . . 209 Checklist for bug reports . . . . . . . . . . . . . . . . . . . . . . 3 Ciceros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Circumflex accent. . . . . . . . . . . . . . . . . . . . . . . . . . . 103 code, arg to @kbdinputstyle . . . . . . . . . . . . . . . . . . 73 Colon in nodename . . . . . . . . . . . . . . . . . . . . . . . . . . 54 colon, last in INFOPATH . . . . . . . . . . . . . . . . . . . . . . 163 Column widths, defining for multitables . . . . . . . 94 Combining indices . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 Comma in nodename. . . . . . . . . . . . . . . . . . . . . . . . . 54 Comma, in macro arguments . . . . . . . . . . . . . . . . 138 Command aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Command definitions . . . . . . . . . . . . . . . . . . . . . . . 126 Command names, indicating . . . . . . . . . . . . . . . . . . 77 Command syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 Commands to insert special characters . . . . . . . 101 Commands using raw HTML . . . . . . . . . . . . . . . . 130 Commands using raw TEX. . . . . . . . . . . . . . . . . . . 130 Commands, inserting them . . . . . . . . . . . . . . . . . . . 15 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Compile command for formatting . . . . . . . . . . . . 147 Compressed files, reading . . . . . . . . . . . . . . . . . . . . 165 Conditionally visible text. . . . . . . . . . . . . . . . . . . . 129 Conditions for copying Texinfo. . . . . . . . . . . . . . . . . 2 Contents, after title page . . . . . . . . . . . . . . . . . . . . . 43 Contents, Table of . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Contents-like outline of file structure . . . . . . . . . . 16 Conventions for writing definitions . . . . . . . . . . . 126 Conventions, syntactic . . . . . . . . . . . . . . . . . . . . . . . . 8 Conversion, from Docbook to Texinfo . . . . . . . . . . 4 Copying conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Copying Permissions . . . . . . . . . . . . . . . . . . . . . . . . . 30 Copying software . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Copying text, including . . . . . . . . . . . . . . . . . . . . . . 31 Copyright owner for FSF works . . . . . . . . . . . . . . . 30 Copyright page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 Copyright page, for plain text. . . . . . . . . . . . . . . . . 31 Copyright word, always in English . . . . . . . . . . . . 30 Correcting mistakes . . . . . . . . . . . . . . . . . . . . . . . . . 204 Create nodes, menus automatically. . . . . . . . . . . . 17 Creating an Info file . . . . . . . . . . . . . . . . . . . . . . . . 153 Creating an unsplit file . . . . . . . . . . . . . . . . . . . . . . 209 Creating index entries . . . . . . . . . . . . . . . . . . . . . . . . 97 Creating pointers with makeinfo . . . . . . . . . . . . . . 55 Cropmarks for printing. . . . . . . . . . . . . . . . . . . . . . 151 Cross reference parts . . . . . . . . . . . . . . . . . . . . . . . . . 62 Cross references . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 Cross references using @inforef. . . . . . . . . . . . . . . 69 Cross references using @pxref. . . . . . . . . . . . . . . . . 68 Cross references using @ref . . . . . . . . . . . . . . . . . . . 67 Cross references using @xref . . . . . . . . . . . . . . . . . . 63 Cross-reference targets, arbitrary. . . . . . . . . . . . . . 56 Custom page sizes . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Customize Emacs package

(Development/Docs/Texinfo) . . . . . . . . . . . 147

Customized highlighting. . . . . . . . . . . . . . . . . . . . . 140 Customizing of TEX for Texinfo . . . . . . . . . . . . . . 148 CVS $Id: texinfo.txi,v 1.219 2002/03/28 16:36:00

karl Exp $, in Texinfo. . . . . . . . . . . . . . . . . . . 192

D

Debugging the Texinfo structure . . . . . . . . . . . . . 204 Debugging with Info formatting. . . . . . . . . . . . . . 204 Debugging with TEX formatting . . . . . . . . . . . . . 205 Defining indexing entries . . . . . . . . . . . . . . . . . . . . . 97 Defining macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 Defining new indices . . . . . . . . . . . . . . . . . . . . . . . . . 99 Defining new Texinfo commands . . . . . . . . . . . . . 137 Definition commands. . . . . . . . . . . . . . . . . . . . . . . . 116 Definition conventions. . . . . . . . . . . . . . . . . . . . . . . 126 Definition template . . . . . . . . . . . . . . . . . . . . . . . . . 116 Definitions grouped together. . . . . . . . . . . . . . . . . 118 Delimiter character, for verbatim. . . . . . . . . . . . . . 75 Depth of text area . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Description for menu, start . . . . . . . . . . . . . . . . . . . 16 Description of document . . . . . . . . . . . . . . . . . . . . . 38 Details of macro usage . . . . . . . . . . . . . . . . . . . . . . 139 Didôt points. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Different cross reference commands. . . . . . . . . . . . 61 Dimension formatting . . . . . . . . . . . . . . . . . . . . . . . 103 Dimensions and image sizes . . . . . . . . . . . . . . . . . 111 ‘dir’ directory for Info installation . . . . . . . . . . . 161 ‘dir’ file listing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 ‘dir’ file, creating your own . . . . . . . . . . . . . . . . . 163 ‘dir’ files and Info directories . . . . . . . . . . . . . . . . 162 Dir files, compressed . . . . . . . . . . . . . . . . . . . . . . . . 165 ‘dir’, created by install-info . . . . . . . . . . . . . . 165 Display formatting . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 Displayed equations . . . . . . . . . . . . . . . . . . . . . . . . . 106 distinct, arg to @kbdinputstyle. . . . . . . . . . . . . . 73 Distorting images . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Docbook, converting to Texinfo . . . . . . . . . . . . . . . . 4 Document description . . . . . . . . . . . . . . . . . . . . . . . . 38 Document input encoding . . . . . . . . . . . . . . . . . . . 136 Document language, declaring . . . . . . . . . . . . . . . 135 Document Permissions . . . . . . . . . . . . . . . . . . . . . . . 30 Dot accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Dotless i, j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Dots, inserting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Double-colon menu entries. . . . . . . . . . . . . . . . . . . . 58 Dumping a .fmt file . . . . . . . . . . . . . . . . . . . . . . . . . 149 DVI file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

E

Ellipsis, inserting . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 em-dash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 Emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Emacs shell, format, print from . . . . . . . . . . . . . . 145 Emphasizing text . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Emphasizing text, font for . . . . . . . . . . . . . . . . . . . . 78 enable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 Encoding, declaring . . . . . . . . . . . . . . . . . . . . . . . . . 136 ‘End’ node footnote style . . . . . . . . . . . . . . . . . . . . 110 End of header line . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 End titlepage starts headings . . . . . . . . . . . . . . . . . 35 Ending a Sentence . . . . . . . . . . . . . . . . . . . . . . . . . . 102 Ending a Texinfo file . . . . . . . . . . . . . . . . . . . . . . . . . 41 Entries for an index . . . . . . . . . . . . . . . . . . . . . . . . . . 97 Entries, making index . . . . . . . . . . . . . . . . . . . . . . . . 96 Enumeration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 Environment indentation . . . . . . . . . . . . . . . . . . . . . 40 epsf.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 epsf.tex, installing . . . . . . . . . . . . . . . . . . . . . . . . . 148 Equations, displayed . . . . . . . . . . . . . . . . . . . . . . . . 106 Equivalence, indicating it. . . . . . . . . . . . . . . . . . . . 108 Error message, indicating it . . . . . . . . . . . . . . . . . 108 Errors, parsing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 Es-zet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 European A4 paper . . . . . . . . . . . . . . . . . . . . . . . . . 150 Evaluation glyph . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 Example beginning of Texinfo file . . . . . . . . . . . . . 26 Example indentation . . . . . . . . . . . . . . . . . . . . . . . . . 40 Example menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 example, arg to @kbdinputstyle . . . . . . . . . . . . . . . 73 Examples in smaller fonts . . . . . . . . . . . . . . . . . . . . 84 Examples, formatting them . . . . . . . . . . . . . . . . . . . 82 Examples, glyphs for . . . . . . . . . . . . . . . . . . . . . . . . 106 Expanding macros . . . . . . . . . . . . . . . . . . . . . . . . . . 138 Expansion, indicating it . . . . . . . . . . . . . . . . . . . . . 107 Expressions in a program, indicating . . . . . . . . . . 72

F

F.B.I., as acronym . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 FDL, GNU Free Documentation License . . . . . . 215 File beginning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 File ending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 File name collision . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 File section structure, showing it . . . . . . . . . . . . . . 16 Filling paragraphs . . . . . . . . . . . . . . . . . . . . . . . . . . 212 Final output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 Finding badly referenced nodes . . . . . . . . . . . . . . 209 Fine-tuning, and hyphenation. . . . . . . . . . . . . . . . 114 First line of a Texinfo file. . . . . . . . . . . . . . . . . . . . . 27 First node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Floating accents, inserting . . . . . . . . . . . . . . . . . . . 103 Flooding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Fonts for indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 Fonts for printing, not for Info . . . . . . . . . . . . . . . . 80 Footings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 Footnotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 Format a dimension. . . . . . . . . . . . . . . . . . . . . . . . . 103 Format and print hardcopy . . . . . . . . . . . . . . . . . . 142 Format and print in Texinfo mode . . . . . . . . . . . 145 Format file, dumping. . . . . . . . . . . . . . . . . . . . . . . . 149 Format with the compile command. . . . . . . . . . . 147 Format, print from Emacs shell . . . . . . . . . . . . . . 145 Formats for images . . . . . . . . . . . . . . . . . . . . . . . . . 111 Formatting a file for Info . . . . . . . . . . . . . . . . . . . . 153 Formatting commands . . . . . . . . . . . . . . . . . . . . . . . . 7 Formatting examples . . . . . . . . . . . . . . . . . . . . . . . . . 82 Formatting for Info . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Formatting for printing. . . . . . . . . . . . . . . . . . . . . . . 22 Formatting headings and footings . . . . . . . . . . . . 200 Formatting requirements . . . . . . . . . . . . . . . . . . . . 147 Formatting with tex and texindex. . . . . . . . . . . 142 Formulas, mathematical . . . . . . . . . . . . . . . . . . . . . 106 Fox, Brian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Frequently used commands, inserting . . . . . . . . . . 15 fubar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 Full texts, GNU . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 Function definitions . . . . . . . . . . . . . . . . . . . . . . . . . 126

G

General syntactic conventions . . . . . . . . . . . . . . . . . . 8 Generating menus with indices. . . . . . . . . . . . . . . . 41 Generating plain text files . . . . . . . . . . . . . . . . . . . 155 German S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 GIF, unsupported due to patents . . . . . . . . . . . . 111 Global Document Commands . . . . . . . . . . . . . . . . . 38 Globbing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 Glyphs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 GNU Emacs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 GNU Emacs shell, format, print from . . . . . . . . 145 GNU sample texts . . . . . . . . . . . . . . . . . . . . . . . . . . 192 Going to other Info files’ nodes . . . . . . . . . . . . . . . 59 Grave accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Group (hold text together vertically) . . . . . . . . . 115 Grouping two definitions together . . . . . . . . . . . . 118

H

Hardcopy, printing it. . . . . . . . . . . . . . . . . . . . . . . . 142 ‘hboxes’, overfull . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 Header for Texinfo files. . . . . . . . . . . . . . . . . . . . . . . 27 Header of a Texinfo file. . . . . . . . . . . . . . . . . . . . . . . 27 Headings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 Headings, page, begin to appear . . . . . . . . . . . . . . 35 Height of images. . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 Height of text area. . . . . . . . . . . . . . . . . . . . . . . . . . 150 help2man. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Highlighting text . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 Highlighting, customized . . . . . . . . . . . . . . . . . . . . 140 Hints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 History of Texinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Holding text together vertically . . . . . . . . . . . . . . 115 href, producing HTML . . . . . . . . . . . . . . . . . . . . . . 69 HTML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 HTML commands, using ordinary. . . . . . . . . . . . 130 http-equiv, and charset. . . . . . . . . . . . . . . . . . . . . . 136 Hungarian umlaut accent. . . . . . . . . . . . . . . . . . . . 103 Hurricanes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 hyphen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 Hyphenation patterns, language-dependent … 135 Hyphenation, helping TEX do . . . . . . . . . . . . . . . . 114 Hyphenation, preventing . . . . . . . . . . . . . . . . . . . . 114

I

ı. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 If text conditionally visible . . . . . . . . . . . . . . . . . . 129 Ignored before @setfilename . . . . . . . . . . . . . . . . . 28 Ignored text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Image formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 Images, inserting . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 Inches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Include file requirements . . . . . . . . . . . . . . . . . . . . 197 Include file sample . . . . . . . . . . . . . . . . . . . . . . . . . . 197 Include files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 Include files, and section levels . . . . . . . . . . . . . . . . 49 Including a file verbatim . . . . . . . . . . . . . . . . . . . . . 84 Including permissions text . . . . . . . . . . . . . . . . . . . . 31 Indentation undoing . . . . . . . . . . . . . . . . . . . . . . . . . 86 Indenting environments . . . . . . . . . . . . . . . . . . . . . . 40 Indenting paragraphs, control of . . . . . . . . . . . . . . 39 Index entries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 Index entries, making . . . . . . . . . . . . . . . . . . . . . . . . 96 Index entry writing . . . . . . . . . . . . . . . . . . . . . . . . . . 97 Index file names . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Index font types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 Indexing commands, predefined . . . . . . . . . . . . . . . 97 Indexing table entries automatically . . . . . . . . . . . 93 Indicating commands, definitions, etc. . . . . . . . . . 71 Indicating evaluation. . . . . . . . . . . . . . . . . . . . . . . . 107 Indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 Indices, combining them. . . . . . . . . . . . . . . . . . . . . . 98 Indices, defining new . . . . . . . . . . . . . . . . . . . . . . . . . 99 Indices, printing and menus . . . . . . . . . . . . . . . . . . 41 Indices, sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Indices, two letter names . . . . . . . . . . . . . . . . . . . . . 98 Indirect subfiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 Info batch formatting . . . . . . . . . . . . . . . . . . . . . . . 159 Info file installation . . . . . . . . . . . . . . . . . . . . . . . . . 161 Info file name, choosing . . . . . . . . . . . . . . . . . . . . . . 28 Info file, listing a new . . . . . . . . . . . . . . . . . . . . . . . 162 Info file, splitting manually . . . . . . . . . . . . . . . . . . 210 Info files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Info formatting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Info installed in another directory . . . . . . . . . . . . 162 Info validating a large file . . . . . . . . . . . . . . . . . . . 209 Info, creating an online file . . . . . . . . . . . . . . . . . . 153 Info-directory-list. . . . . . . . . . . . . . . . . . . . . . . 163 Info; other files’ nodes. . . . . . . . . . . . . . . . . . . . . . . . 59 INFOPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 Initialization file for TEX input . . . . . . . . . . . . . . 148 Input encoding, declaring. . . . . . . . . . . . . . . . . . . . 136 Insert nodes, menus automatically . . . . . . . . . . . . 17 Inserting @, braces. . . . . . . . . . . . . . . . . . . . . . . . . . 101 Inserting accents. . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Inserting dots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Inserting ellipsis . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Inserting frequently used commands . . . . . . . . . . . 15 Inserting space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 Inserting special characters and symbols . . . . . . 101 ‘INSTALL’ file, generating . . . . . . . . . . . . . . . . . . . . 155 install-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 Installing an Info file . . . . . . . . . . . . . . . . . . . . . . . . 161 Installing Info in another directory . . . . . . . . . . . 162 Internationalization . . . . . . . . . . . . . . . . . . . . . . . . . 135 Introduction to Texinfo. . . . . . . . . . . . . . . . . . . . . . . . 3 Introduction, as part of file . . . . . . . . . . . . . . . . . . . 40 Invalid characters in node names . . . . . . . . . . . . . . 54 Invoking macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 ISO 639 codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 Itemization. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

J

ȷ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 JPG image format . . . . . . . . . . . . . . . . . . . . . . . . . . 111

K

Keyboard input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Keys, recommended names . . . . . . . . . . . . . . . . . . . 74 Keywords, indicating. . . . . . . . . . . . . . . . . . . . . . . . . 72 Knuth, Donald . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

L

ł. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Ł . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Language codes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 Language, declaring. . . . . . . . . . . . . . . . . . . . . . . . . 135 Larger or smaller pages . . . . . . . . . . . . . . . . . . . . . 151 Legal paper, printing on. . . . . . . . . . . . . . . . . . . . . 151 Length of file names . . . . . . . . . . . . . . . . . . . . . . . . . 28 Less cluttered menu entry . . . . . . . . . . . . . . . . . . . . 58 License agreement . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Line breaks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 Line breaks, preventing . . . . . . . . . . . . . . . . . . . . . 114 Line length, column widths as fraction of . . . . . . 94 Line spacing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 Lisp example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 Lisp examples in smaller fonts . . . . . . . . . . . . . . . . 84 List of @-commands . . . . . . . . . . . . . . . . . . . . . . . . 167 Listing a new Info file . . . . . . . . . . . . . . . . . . . . . . . 162 Lists and tables, making. . . . . . . . . . . . . . . . . . . . . . 89 Local variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 Location of menus . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Looking for badly referenced nodes. . . . . . . . . . . 209 lpr (DVI print command) . . . . . . . . . . . . . . . . . . . 144 lpr-d, replacements on MS-DOS/MS-Windows

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144

M

Macro definitions. . . . . . . . . . . . . . . . . . . . . . . 126, 137 Macro details. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Macro invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 Macron accent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 Macros in definition commands . . . . . . . . . . . . . . 117 Macros, undefining. . . . . . . . . . . . . . . . . . . . . . . . . . 138 Magnified printing . . . . . . . . . . . . . . . . . . . . . . . . . . 151 mailto link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 makeinfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 makeinfo inside Emacs . . . . . . . . . . . . . . . . . . . . . . 157 makeinfo options . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Making a printed manual. . . . . . . . . . . . . . . . . . . . 142 Making a tag table automatically . . . . . . . . . . . . 159 Making a tag table manually . . . . . . . . . . . . . . . . 210 Making cross references . . . . . . . . . . . . . . . . . . . . . . 61 Making line and page breaks . . . . . . . . . . . . . . . . 113 Making lists and tables. . . . . . . . . . . . . . . . . . . . . . . 89 Man page output, not supported . . . . . . . . . . . . . . . 4 Man page, reference to . . . . . . . . . . . . . . . . . . . . . . . 69 Manual characteristics, printed. . . . . . . . . . . . . . . . . 6 Margins on page, not controllable . . . . . . . . . . . . 151 Marking text within a paragraph. . . . . . . . . . . . . . 71 Marking words and phrases . . . . . . . . . . . . . . . . . . . 71 Master menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Mathematical expressions . . . . . . . . . . . . . . . 106, 130 Menu description, start. . . . . . . . . . . . . . . . . . . . . . . 16 Menu entries with two colons . . . . . . . . . . . . . . . . . 58 Menu example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Menu location . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Menu parts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Menu writing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Menu, master . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Menus generated with indices . . . . . . . . . . . . . . . . . 41 Menus, omitting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 meta HTML tag, and charset . . . . . . . . . . . . . . . . 136 META key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Meta-syntactic chars for arguments . . . . . . . . . . 117 Millimeters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Minimal requirements for formatting . . . . . . . . . 147 Minimal Texinfo file (requirements) . . . . . . . . . . . . 9 Mistakes, catching . . . . . . . . . . . . . . . . . . . . . . . . . . 204 Mode, using Texinfo . . . . . . . . . . . . . . . . . . . . . . . . . 14 Multiple spaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 Multitable column widths . . . . . . . . . . . . . . . . . . . . 94 Multitable rows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 Must have in Texinfo file . . . . . . . . . . . . . . . . . . . . . . 9 Mutually recursive macros. . . . . . . . . . . . . . . . . . . 137

N

Names for indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 Names of index files. . . . . . . . . . . . . . . . . . . . . . . . . 142 Names recommended for keys . . . . . . . . . . . . . . . . . 74 Naming a ‘Top’ Node in references . . . . . . . . . . . . 67 NASA, as acronym . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Navigation links, omitting . . . . . . . . . . . . . . . . . . . 155 Need space at page bottom . . . . . . . . . . . . . . . . . . 115 New index defining . . . . . . . . . . . . . . . . . . . . . . . . . . 99 New Info file, listing it in ‘dir’ file . . . . . . . . . . . 162 New Texinfo commands, defining . . . . . . . . . . . . 137 Next node of Top node . . . . . . . . . . . . . . . . . . . . . . . 55 Node line requirements . . . . . . . . . . . . . . . . . . . . . . . 54 Node line writing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Node name must be unique . . . . . . . . . . . . . . . . . . . 54 Node name, should not contain @-commands . . 54 Node names, choosing . . . . . . . . . . . . . . . . . . . . . . . . 52 Node separators, omitting . . . . . . . . . . . . . . . . . . . 155 Node, ‘Top’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Node, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Nodes for menus are short . . . . . . . . . . . . . . . . . . . . 57 Nodes in other Info files . . . . . . . . . . . . . . . . . . . . . . 59 Nodes, catching mistakes . . . . . . . . . . . . . . . . . . . . 204 Nodes, checking for badly referenced . . . . . . . . . 209 Non-breakable space . . . . . . . . . . . . . . . . . . . . . . . . 114 Not ending a sentence. . . . . . . . . . . . . . . . . . . . . . . 102

O

O’Dea, Brendan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ø . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Ø . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Obtaining TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 Occurrences, listing with @occur . . . . . . . . . . . . . 208 œ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Œ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Optional and repeated arguments . . . . . . . . . . . . 117 Options for makeinfo . . . . . . . . . . . . . . . . . . . . . . . 153 Ordinary HTML commands, using . . . . . . . . . . . 130 Ordinary TEX commands, using . . . . . . . . . . . . . 130 Other Info files’ nodes. . . . . . . . . . . . . . . . . . . . . . . . 59 Outline of file structure, showing it. . . . . . . . . . . . 16 Output file splitting. . . . . . . . . . . . . . . . . . . . . . . . . 155 Output formats, supporting more . . . . . . . . . . . . . . 4 Overfull ‘hboxes’ . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 Overview of Texinfo. . . . . . . . . . . . . . . . . . . . . . . . . . . 3

P

Page breaks. . . . . . . . . . . . . . . . . . . . . . . . . . . . 113, 115 Page delimiter in Texinfo mode . . . . . . . . . . . . . . . 17 Page headings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 Page numbering . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 Page sizes for books. . . . . . . . . . . . . . . . . . . . . . . . . 150 Page sizes, customized . . . . . . . . . . . . . . . . . . . . . . 150 page-delimiter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Pages, starting odd . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Paper size, A4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Paragraph indentation control . . . . . . . . . . . . . . . . 39 Paragraph, marking text within . . . . . . . . . . . . . . . 71 Paragraphs, filling . . . . . . . . . . . . . . . . . . . . . . . . . . 212 Parameters to macros . . . . . . . . . . . . . . . . . . . . . . . 137 Parentheses in nodename . . . . . . . . . . . . . . . . . . . . . 54 Parsing errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 Part of file formatting and printing. . . . . . . . . . . . 22 Parts of a cross reference . . . . . . . . . . . . . . . . . . . . . 62 Parts of a master menu . . . . . . . . . . . . . . . . . . . . . . 37 Parts of a menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Patches, contributing. . . . . . . . . . . . . . . . . . . . . . . . . . 3 PDF output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 pdfcolor.tex, installing . . . . . . . . . . . . . . . . . . . . 148 pdftex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 pdftex, and images . . . . . . . . . . . . . . . . . . . . . . . . . 111 Period in nodename. . . . . . . . . . . . . . . . . . . . . . . . . . 54 Periods, inserting . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 Permissions text, including . . . . . . . . . . . . . . . . . . . 31 Permissions, printed . . . . . . . . . . . . . . . . . . . . . . . . . 34 Picas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 Pictures, inserting . . . . . . . . . . . . . . . . . . . . . . . . . . 111 Pinard, François . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 plain TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 Plain text output . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 PNG image format . . . . . . . . . . . . . . . . . . . . . . . . . 111 Point, indicating in a buffer . . . . . . . . . . . . . . . . . 108 Pointer creation with makeinfo . . . . . . . . . . . . . . . 55 Pointer validation with makeinfo . . . . . . . . . . . . 156 Pointer validation, suppressing . . . . . . . . . . 143, 155 Points (dimension) . . . . . . . . . . . . . . . . . . . . . . . . . . 111 Predefined indexing commands . . . . . . . . . . . . . . . 97 Predefined names for indices . . . . . . . . . . . . . . . . . . 98 Preparing for TEX . . . . . . . . . . . . . . . . . . . . . . . . . . 148 Preventing line and page breaks . . . . . . . . . . . . . 113 Previous node of Top node . . . . . . . . . . . . . . . . . . . 55 Print and format in Texinfo mode. . . . . . . . . . . . 145 Print, format from Emacs shell . . . . . . . . . . . . . . 145 Printed book and manual characteristics . . . . . . . . 6 Printed output, indicating it . . . . . . . . . . . . . . . . . 107 Printed permissions . . . . . . . . . . . . . . . . . . . . . . . . . . 34 Printing a region or buffer . . . . . . . . . . . . . . . . . . . . 22 Printing an index . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Printing cropmarks . . . . . . . . . . . . . . . . . . . . . . . . . 151 Printing DVI files, on MS-DOS/MS-Windows

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 Printing hardcopy . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Problems, catching. . . . . . . . . . . . . . . . . . . . . . . . . . 204 Program names, indicating . . . . . . . . . . . . . . . . . . . 77 Prototype row, column widths defined by . . . . . . 94

Q

Quotations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

R

ragged left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 ragged right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 Raising and lowering sections . . . . . . . . . . . . . . . . . 49 Raw formatter commands . . . . . . . . . . . . . . . . . . . 130 RCS $Id: texinfo.txi,v 1.219 2002/03/28 16:36:00

karl Exp $, in Texinfo. . . . . . . . . . . . . . . . . . . 192 Recommended names for keys . . . . . . . . . . . . . . . . 74 Rectangle, black in hardcopy . . . . . . . . . . . . . . . . 149 Recursion, mutual . . . . . . . . . . . . . . . . . . . . . . . . . . 137 Recursive macro invocations . . . . . . . . . . . . . . . . . 137 Reference to @-commands . . . . . . . . . . . . . . . . . . . 167 References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 References using @inforef. . . . . . . . . . . . . . . . . . . . 69 References using @pxref . . . . . . . . . . . . . . . . . . . . . . 68 References using @ref . . . . . . . . . . . . . . . . . . . . . . . . 67 References using @xref . . . . . . . . . . . . . . . . . . . . . . . 63 Referring to other Info files . . . . . . . . . . . . . . . . . . . 59 Refilling paragraphs. . . . . . . . . . . . . . . . . . . . . . . . . 212 Region formatting and printing . . . . . . . . . . . . . . . 22 Region printing in Texinfo mode . . . . . . . . . . . . . 145 Reid, Brian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Repeated and optional arguments . . . . . . . . . . . . 117 Reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Required in Texinfo file . . . . . . . . . . . . . . . . . . . . . . . 9 Requirements for formatting . . . . . . . . . . . . . . . . . 147 Requirements for include files . . . . . . . . . . . . . . . . 197 Requirements for updating commands . . . . . . . . . 20 Reserved words, indicating . . . . . . . . . . . . . . . . . . . 72 Restrictions on node names. . . . . . . . . . . . . . . . . . . 54 Result of an expression. . . . . . . . . . . . . . . . . . . . . . 107 ridt.eps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Ring accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Rounded rectangles, around examples . . . . . . . . . 87 Rows, of a multitable . . . . . . . . . . . . . . . . . . . . . . . . 94 Running an Info formatter. . . . . . . . . . . . . . . . . . . . 22 Running Info-validate . . . . . . . . . . . . . . . . . . . . 209 Running macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 Running makeinfo in Emacs. . . . . . . . . . . . . . . . . 157

S

Sample @include file. . . . . . . . . . . . . . . . . . . . . . . . 197 Sample function definition . . . . . . . . . . . . . . . . . . . 126 Sample Texinfo file, no comments . . . . . . . . . . . . 191 Sample Texinfo file, with comments . . . . . . . . . . . 11 Sample Texinfo files. . . . . . . . . . . . . . . . . . . . . . . . . 191 Sample texts, GNU . . . . . . . . . . . . . . . . . . . . . . . . . 192 Scaled points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Schwab, Andreas . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Scribe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Sea surges. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Section structure of a file, showing it . . . . . . . . . . 16 Sections, raising and lowering . . . . . . . . . . . . . . . . . 49 Sentence ending punctuation . . . . . . . . . . . . . . . . 102 Sentence non-ending punctuation . . . . . . . . . . . . 102 ‘Separate’ footnote style. . . . . . . . . . . . . . . . . . . . . 110 SGML-tools output format . . . . . . . . . . . . . . . . . . . . 4 Sharp S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Shell formatting with tex and texindex . . . . . . 142 Shell printing, on MS-DOS/MS-Windows . . . . . 144 Shell, format, print from . . . . . . . . . . . . . . . . . . . . 145 Shell, running makeinfo in . . . . . . . . . . . . . . . . . . 157 Short nodes for menus . . . . . . . . . . . . . . . . . . . . . . . 57 Short table of contents . . . . . . . . . . . . . . . . . . . . . . . 42 Showing the section structure of a file . . . . . . . . . 16 Showing the structure of a file . . . . . . . . . . . . . . . 207 shrubbery . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 Site-wide Texinfo configuration file . . . . . . . . . . . 148 Size of printed book . . . . . . . . . . . . . . . . . . . . . . . . 150 slanted typewriter font, for @kbd . . . . . . . . . . . . . . 73 Small book size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Small caps font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 Small examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 Software copying permissions . . . . . . . . . . . . . . . . . 40 Sorting indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Source file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Space, inserting horizontal. . . . . . . . . . . . . . . . . . . 102 Space, inserting vertical . . . . . . . . . . . . . . . . . . . . . 114 Spaces in macros . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 Spaces, in menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Spacing, inserting. . . . . . . . . . . . . . . . . . . . . . . . . . . 101 Special characters, commands to insert . . . . . . . 101 Special insertions . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 Special typesetting commands . . . . . . . . . . . . . . . 104 Specifying index entries . . . . . . . . . . . . . . . . . . . . . . 97 Splitting an Info file manually . . . . . . . . . . . . . . . 210 Splitting of output files. . . . . . . . . . . . . . . . . . . . . . 155 ß . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 Stallman, Richard M. . . . . . . . . . . . . . . . . . . . . . . . . 13 Start of header line . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Starting chapters . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Structure of a file, showing it . . . . . . . . . . . . . . . . . 16 Structure, catching mistakes in . . . . . . . . . . . . . . 204 Structuring of chapters . . . . . . . . . . . . . . . . . . . . . . . 44 Subsection-like commands . . . . . . . . . . . . . . . . . . . . 48 Subsub commands . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Suggestions for Texinfo, making. . . . . . . . . . . . . . . . 3 Summary of document . . . . . . . . . . . . . . . . . . . . . . . 38 Syntactic conventions . . . . . . . . . . . . . . . . . . . . . . . . . 8 Syntactic tokens, indicating. . . . . . . . . . . . . . . . . . . 72 Syntax, of @-commands . . . . . . . . . . . . . . . . . . . . . 213 Syntax, optional & repeated arguments. . . . . . . 117

T

Table of contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Table of contents, after title page . . . . . . . . . . . . . 43 Tables and lists, making . . . . . . . . . . . . . . . . . . . . . . 89 Tables with indexes . . . . . . . . . . . . . . . . . . . . . . . . . . 93 Tables, making multi-column . . . . . . . . . . . . . . . . . 94 Tables, making two-column . . . . . . . . . . . . . . . . . . . 92 Tabs; don’t use! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Tag table, making automatically . . . . . . . . . . . . . 159 Tag table, making manually . . . . . . . . . . . . . . . . . 210 Targets for cross-references, arbitrary. . . . . . . . . . 56 Template for a definition . . . . . . . . . . . . . . . . . . . . 116 TEX commands, using ordinary . . . . . . . . . . . . . . 130 TEX index sorting . . . . . . . . . . . . . . . . . . . . . . . . . . 142 TEX input initialization . . . . . . . . . . . . . . . . . . . . . 148 TEX, how to obtain . . . . . . . . . . . . . . . . . . . . . . . . . 214 texi2dvi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 texi2dvi (shell script) . . . . . . . . . . . . . . . . . . . . . . 144 texi2roff, unsupported software . . . . . . . . . . . . . . 6 texindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Texinfo commands, defining new . . . . . . . . . . . . . 137 Texinfo file beginning . . . . . . . . . . . . . . . . . . . . . . . . 26 Texinfo file ending . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Texinfo file header . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 Texinfo file minimum. . . . . . . . . . . . . . . . . . . . . . . . . . 9 Texinfo file section structure, showing it . . . . . . . 16 Texinfo history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Texinfo mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Texinfo overview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Texinfo printed book characteristics . . . . . . . . . . . . 6 Texinfo requires @setfilename . . . . . . . . . . . . . . . . 28 Texinfo, introduction to . . . . . . . . . . . . . . . . . . . . . . . 3 texinfo.cnf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 texinfo.cnf installation . . . . . . . . . . . . . . . . . . . . 148 texinfo.tex, installing . . . . . . . . . . . . . . . . . . . . . 148 TEXINPUTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 TEXINPUTS environment variable . . . . . . . . . . . . . 148 Text width and height . . . . . . . . . . . . . . . . . . . . . . 150 Text, conditionally visible . . . . . . . . . . . . . . . . . . . 129 Text, marking up . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 Thin space between number, dimension. . . . . . . 103 Tie-after accent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Tied space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 Tilde accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Tips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 Title page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Title page, bastard. . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Title page, for plain text . . . . . . . . . . . . . . . . . . . . . 31 Titlepage end starts headings . . . . . . . . . . . . . . . . . 35 Top node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Top node example . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Top node is first . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 ‘Top’ node naming for references . . . . . . . . . . . . . . 67 Tree structuring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Two ‘First’ Lines for @deffn . . . . . . . . . . . . . . . . . 118 Two letter names for indices . . . . . . . . . . . . . . . . . . 98 Two named items for @table . . . . . . . . . . . . . . . . . 93 Two part menu entry . . . . . . . . . . . . . . . . . . . . . . . . 58 ‘txi-cc.tex’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 Typesetting commands for dots, etc. . . . . . . . . . 104

U

Ugly black rectangles in hardcopy . . . . . . . . . . . . 149 Umlaut accent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Unbreakable space . . . . . . . . . . . . . . . . . . . . . . . . . . 114 Uncluttered menu entry . . . . . . . . . . . . . . . . . . . . . . 58 Undefining macros . . . . . . . . . . . . . . . . . . . . . . . . . . 138 Underbar accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Underdot accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Uniform resource locator, indicating . . . . . . . . . . . 78 Uniform resource locator, referring to . . . . . . . . . 69 Unique nodename requirement . . . . . . . . . . . . . . . . 54 Unprocessed text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Unsplit file creation . . . . . . . . . . . . . . . . . . . . . . . . . 209 Up node of Top node. . . . . . . . . . . . . . . . . . . . . . . . . 54 Updating nodes and menus . . . . . . . . . . . . . . . . . . . 17 Updating requirements . . . . . . . . . . . . . . . . . . . . . . . 20 URI syntax for Info . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 URL, indicating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 URL, referring to . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Usage tips. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 user input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 User options, marking. . . . . . . . . . . . . . . . . . . . . . . 121 User-defined Texinfo commands. . . . . . . . . . . . . . 137 Using Texinfo in general. . . . . . . . . . . . . . . . . . . . . . . 3

V

Validating a large file . . . . . . . . . . . . . . . . . . . . . . . 209 Validation of pointers . . . . . . . . . . . . . . . . . . . . . . . 156 Value of an expression, indicating . . . . . . . . . . . . 107 Verbatim environment . . . . . . . . . . . . . . . . . . . . . . . 83 Verbatim in-line text . . . . . . . . . . . . . . . . . . . . . . . . . 75 Verbatim, include file . . . . . . . . . . . . . . . . . . . . . . . . 84 version number, finding . . . . . . . . . . . . . . . . . . . . . 166 Vertically holding text together . . . . . . . . . . . . . . 115 Visibility of conditional text . . . . . . . . . . . . . . . . . 129

W

Weinberg, Zack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Weisshaus, Melissa. . . . . . . . . . . . . . . . . . . . . . . . . . . 13 White space, excessive . . . . . . . . . . . . . . . . . . . . . . 113 Whitespace in macros . . . . . . . . . . . . . . . . . . . . . . . 137 Whitespace, inserting . . . . . . . . . . . . . . . . . . . . . . . 102 Width of images . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 Width of text area . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Widths, defining multitable column . . . . . . . . . . . 94 Wildcards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 Words and phrases, marking them . . . . . . . . . . . . 71 Writing a menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Writing an @node line . . . . . . . . . . . . . . . . . . . . . . . . 53 Writing index entries. . . . . . . . . . . . . . . . . . . . . . . . . 97

X

XPM image format. . . . . . . . . . . . . . . . . . . . . . . . . . . 111

Y

Years, in copyright line . . . . . . . . . . . . . . . . . . . . . . . 30

Z

Zaretskii, Eli . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Zuhn, David D.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13